A Complete Bibliography of Publications in Journal of Numerical Linear Algebra with Applications and

Numerical Linear Algebra with Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2023 Version 1.71

Title word cross-reference

(1, 1) [Cao08, Krz11]. (2, 2) [Li00]. (m, k) [MN00]. (q) [Jia96]. + [LJM14]. −Δu = λu [EFG^{+}18]. 0.822 [Ano09]. 16 [KM09]. 2 [AM96, BF19, BV13, DHBV21, Mar94, NBKS99, NSCTPW22, QB15, ZVO14, vKVW00]. 2 × 2 [AB10, AB13, Cao13, Kol05]. 3 [GKY97, KK16, LPW06, NBKS99, PM97, PR96, SY18b, mMP99, vKVW00]. 4 [MR14, SY18b]. A [CC07]. A − BX ± X ∗ B∗ [LT08]. A − XB [Den09]. Aαx = b [AL21]. α [Tre13, XCG16]. AXA^{∗} = B [Tia13]. AXB + CY D = E [yPxP06, WTZD10].

AXB = C [fLyHZ11, Miy15]. β [DP23]. H [Gra08, LOY08]. H^{2} [Bör17]. K [Mar95]. D [BLLA11]. f [LMM^{+}23]. f(A)x = b [AL21]. Γ [DP23]. GMRES(k) [KY95]. H [AMM04, BK21, BCGM09, Chu04, KPV08, KC17, Leb02, LP16, Sun06, ZSCX10, DMM^{+}08, FS21, Pul09]. H(div) [BO18]. H^{1} [AMM04]. H_{1} [LPW06]. H_{∞} [Özb13, TV20]. hp [DMM^{+}08]. IDR(s) [CvG11]. ILU [CGK94, KOV17]. k [BO08, VVM05a]. L [Aih20]. λ [FLPW01]. LDL^{T} [LSS18]. l_{p} [Dax94]. LU [KNY00, KOV17, Le 23, DHS95, Saa94]. M [BNT94, Sau95, Bea94, BCC98, HHLL16, IP13, JZ11, Kra02, LSL01, WQZ09, XZS10, ZJ06, vN00, CSB20]. P_{k} [RSCTP20]. R [DN12]. R^{3} [ST23]. H [HK02]. O(N) [Sac05]. Ω [CGS20]. P [LHL07a, Peñ09, AEHV15, Beu03, BB06, GKY97, LZ09, LO13, LH17, Pul09]. p × p × 2(p ≥ 2) [KJ12]. Q [Cha12, DBLP16, DOP21]. Q_{2} − Q_{1} [PT17]. QMR [FH94]. QR [ADP96, Cha12, FG02, Le 23, MVLB23, AG95, CH94]. QZ [CX23a]. R [DW15, BKM^{+}12]. s [CGY22, CK10]. S/P [Bea94, BNT94]. S_{n} [Lee12]. SSOR [JO94]. t [BSMN22, Lun20, RU22, ZSKA18]. tr(f(A)) [CS18]. UTU + UTR + RTU [Kap98]. uT (A)v [GR04]. V [BLZ08, Lai97, Lot23, NN10, Not98]. φ [MPR20]. X [fLyHZ11]. Z [CHCS22, HCD15, HHQ13, LQY13, XC13, ZZLX20].

-based [Le 23]. -circulants [Tre13, XCG16]. -conforming [AMM04, LPW06]. -cycle [BLZ08, Lai97, NN10, Not98]. -cycles [Lot23]. -decomposition [Kap98]. -divergences [LMM^{+}23]. -dominated [AMM04]. -eigenpairs [CHCS22, LP16, ZZLX20]. -eigenvalue [WQZ09]. -eigenvalues [HCD15, HHQ13, LQY13, XC13]. -elliptic [ZSCX10]. -factor [Cha12]. -factorization [KNY00]. -factors [Bea94]. -FEM [BB06]. -form [BO08]. -function [Lun20, XZS10]. -hierarchical [LO13, Pul09]. -independent [FS21]. -integers [DOP21]. -linear [DN12]. -matrices [BNT94, BCC98, BCGM09, KC17, Kra02, LSL01, LHL07a, Peñ09, Sun06, vN00]. -matrix [FLPW01, Sau95, Bör17, Gra08, HK02, HHLL16, IP13, JZ11, LOY08, ZJ06]. -monotonicity [Mar95]. -multisplittings [BCC98]. -optimization [Chu04]. -partitionings [GKY97]. -policy [BLLA11]. -product [BSMN22, RU22, ZSKA18]. -refinement [DMM^{+}08]. -self-adjoint [Leb02]. -stable [CGS20]. -step [CGY22, CK10, Li00]. -tensor [WCW20].

-th [AEHV15]. -Toeplitz [BF19].

-transformed [MPR20]. -version [Beu03]. -weighted [DBLP16].

0-521-48296-8 [Nab97].

14 [SB12]. 1st [NL09].

2010 [NL09]. 2017 [Den18]. 2D [BCV03]. 2nd [Kap02].

3-D [BG02]. 3-tensors [ED22]. 3D

[MM02, NH98].

4th [Web10a]. 4th-order [Web10a].

60th [Vas03].

70th [CLR13, Vas05]. 7th [BFG^{+}18].

80th [SGP14].

’97 [Axe98]. 98 [Axe99].

abs [Bos19, SCD94]. abs-normal [Bos19]. ABS-type [SCD94]. absorbing [Cas11, RV12]. abstract [NV08a]. accelerate [HY22]. Accelerated [IKAA22, BEH^{+}17, Ema12, HW18, LLPC23, PRPI09, Wan18b, YYN12]. Accelerating [ACGH21, MS22, PH19, KKPS18]. Acceleration [DE06, BGN07, DH18, EY23, rFS09, IIFM23, LZ22, MYD20, Pas19, Ris19, WM12]. Accuracy [LL97, BS01, DOP21, DP23, SWKW98, YM22]. Accurate [BP13, DPP16, DOP19, MP18a, MPR20, MPR22, KR06, LVW01, MPR23, Van00, Yan23, Ye20]. achieving [SWKW98]. acoustic [GM17, mM04]. acoustics [CCvG06]. active [TV20]. activity [MC04]. adapting [HY22]. Adaptive [MMM06, MM11, RR12, BLE97, BGM^{+}12, BE98, CSB20, CH21, DV19, DHR^{+}04, Fer96, GKL18, JYH17, LM06, LWS^{+}23, MMC12, MRK22, MWZ06, Mit10, MW21, RSR10, SWKW98, SY18b, TV20, Ver00, ZSCX10]. adaptively [YYN12]. adaptivity [vVW23]. addition [BH07]. additive [BK11, CL96, CZ02, DS08, FHM21, GH23, KV92, KLM15, MW21, NV08a, NWZ17, SWW21, VY14, XZS10]. ADI [Dam08, MP16]. ADI-preconditioned [Dam08]. adjoint [Leb02, MM11]. adjustments [FLR03]. admissible [VL11]. ADN [Lee19]. Advanced [VZ08]. Advances [MM18]. advection [BCV03, CCK06, DFF^{+}21a]. advection-diffusion [BCV03]. advection-dominated [CCK06]. aerodynamic [LW04]. affine [DGC19]. agglomeration [IV04, KV06, LV08, LV12]. aggregation [BMM^{+}08, BVV12, BDM^{+}14, CG15, DV19, GHT09, GHJV16, HST22, KWS^{+}18, MM98, NN11, NY03, OS10, Pul08, PM11, Sch12]. aggregation-based [CG15, DV19, NN11]. aggregation/disaggregation [MM98]. aggressive [Yan10]. AILU [GN00]. AINV [KKNY01]. AINV-type [KKNY01]. Algebra [NLA94, Ano09, SB12, JNL92, BDRS12, BNR18, BM05a, CSCTP05, Dat01, GGV13, GY08, Mar00, MV05, Özb13, PDV05]. Algebraic [Ada04, AN94, BD21, BBS12, BO08, FM18, GL95a, Kra06, LB21, Lee21a, LOS04, NN11, NFD10, Not05b, Not10, Pfl99, RBV08, Sim03, Web10b, Web10a, XM17, AB12, BGX06, BKY10, BF11a, BDV06, BCZ12, BVV12, BKM^{+}12, BDM^{+}14, CG15, CH21, DFNY08, Don10, Ema12, GB11, GHJV16, GMOS06, Het07, HM14, HHLL16, IP13, KH23, KS22, Kra02, Kuz92, KP10, LSS03, LB08, LS15, LCHH18, Liv04b, Lu05, LJM14, MMC12, MO14, MMPR10, MM95, MBW97, MC08, MSF21, Miy17, NL16, Not98, Not02b, OST10a, PM97, PT17, RS02, SS02, Sei10, Sha99, SH19, SY18b, TC10, VY14, XSZ09, XZS15, ZCW11, vN00]. algorithm [ARSO14, Amb15, AB12, ADO23, AMMR17, AG95, BCK05, BPS95, BCB14, BFdP13, BD15, BLP01, CD11, CX23a, CX23b, CC03, CP12, DW15, DDM23, DGC19, DDL^{+}21, ER96, FG02, FO95, Gan99, GM17, GP18, HK21, HVCY21, HNR^{+}18, Het07, HHLL16, HW22, JR94, JZ11, Jou94, Kap99, Kau07, KNY00, KMC16, LW21, LC21, LZ23, Liv04b, LYL15, MPV06, MCV01, MLV05, MVV08, MP13, MP16, MCLM20, MM18, MC04, MR14, NG15, NLZ11, OC04, PR16, RK18, RY08, RSR10, Roh92, SW96, ST17a, Shi04, SS97, SWKW98, Sto92, SHT11, TGKR10, VJM16, VVM05b, Van00, Vla00, WDS09, WM12, WL08, WW07, WtFW15, Yan23, YCY17, ZQ12, ZZ15, ZZLX20, ZWQA18]. algorithmic [DIPR19]. Algorithms [BVD^{+}18, GL96, AH02, AMP99, BH04, BT15, Bun92, CL96, CS96, Cao04, CGY22, CQ10, CJT03, DMY03, DFZ05, DKVB15, Du19, FLM09, FP95a, FH94, HJR97, HCGM23, HR05, HM16, HL21, KN14, KR14, Kub92, Lai97, LW98, LLLJ16, LZ22, Mar98, Mat96, MP18b, MK20, Pfl99, RS07, Sac05, SLK16, SHvBW21, Sha98, SX15, SCD94, SST18, SSB15, ŚLA^{+}21, US19, VP95, WWX10, XCGL10, XXW19, XZS10, YZ13, ZJ06, vGSZ15]. aligned [YXZ13]. alignment [YZ13]. All-at-once [DW21]. Almost [ACR^{+}00, Bos19, AW11, AMP99, BL20, EFG^{+}18]. almost-isotropic [BL20]. along [MM95]. Alternately [BGX06]. alternating [AG19, Bai12, DH18, DL23, EY23, LZZ20, Liu22, MS22, MVLB23, MYD20, ORU23, Wan18a, XJ12, ZN18, ZS08]. Alternative [GS99]. Alternatives [Sid97]. AMG [LOS04, BBM^{+}06, DV19, GX14, HVX16, KV06, MMM06, TT15, Vas02, Web18, XM17]. AMG-shifted [TT15]. AMGe [LV08]. AMLI [Beu03, Mar98]. among [Par92]. amplitude [TH19]. AMPS [YPC20]. analyse [AN13, HS13]. analyses [PM97]. Analysis [BEH^{+}17, BLP01, CCvG06, CG15, GR23, MSS07, Mat96, SPD05, SP06, Sha98, YZ13, YXZ13, Zhu14, AK23, Axe15, BPS15, BSC20, Bat95, BL22, BW17b, BBG13, BVV12, Cas11, CDDSC12, CTP09, CLC11, CL13, CMSW19, CLTW11, CV13, CDW06, DFHM20, Don10, DFF^{+}21b, DFF^{+}18, EFG^{+}18, EM11, FHM21, FM15, GMSCS20, GZ16, GCLG18, GX14, HJR97, HM18, HM20, He21, HHvR04, HW22, KO18, Lee10, Lee21b, LV04, LT09, LC21, LB08, LH17, MO11, MO14, MM98, MM02, NN11, NLZ11, NSCTPW22, Not10, PV99, PY22, Pas19, Pfl99, RSCTP20, RR12, Saa00b, ST17a, SHvBW21, Sha99, The98, WCZ15, WW08b, WW11, WF15, WX21, ZN22, ZY19, ZBCN23, mMvdV02, vRH05]. analytic [CLQY23, GN00, IT05]. analytical [SSB04]. Analyzing [RV12]. Anderson [LLPC23, LZ22, Pas19]. Anderson-type [Pas19]. angle [DMY03, Lee12]. angles [GH06]. anisotropic [BCZ12, CG15, GHT09, Höm06, KW99, KT08, KLM14, KNP03, RNV21, Sch12, XZS15, YXZ13]. anti [MMMM09, Per06, XHZ03]. anti-persymmetric [XHZ03]. anti-reflective [Per06]. anti-triangular [MMMM09]. antibandwidth [SH14]. Any [VL11]. AOR [WCW20]. Appl [SB12]. applicability [DOR19]. Application [CC03, Ibr02, LD08, MBW97, AM96, ABK15, BGS21, BGW05, BCC98, BDRZ21, Car97, CD11, DH18, DQW15, DCT18, EAA19, GKK04, GMTV16, IKAA22, KMM18, KK23a, Lam12, LY15, LQY13, NR11, SLK16, SHvBW21, Vas02, Wan18b, BG02, CPSM06, Leb02]. Applications [NLA94, LX08, Ada04, ACR^{+}00, AHJ20, JNL92, ABNP15, BNR18, BKP02, BSC20, BF96, BVD^{+}18, BFM12, CC07, CCS10, CEQN07, CNP96, CCLN05, CCLQ18, CNY05, DHW16, DFF^{+}21b, DKVB15, FBSC21, FJ05, FH94, GCLG18, HS21a, HPS15, Hua12, JNS19, KCC16, Kub92, LB17, LHW11, LW23, LNQ13, LT08, LWW09, LT11, LT13, LWS^{+}23, LPS15, MRK22, MV05, NPR13, NR14a, PN18, PRR^{+}16, SKR08, WWC^{+}15, XM17, ZZ15, NL09, Ano09]. applied [AGRR21, BCK05, CH05, GORR16, GH23, LMM00, LD07, MO11, Mit10, RU22, ZCW11]. approach [AMM04, AN13, BGM^{+}21, CCX23, CCLQ18, CJL08, DY04, DIPR19, DGRR11, DS02, FLPW01, GL21, GHW06, HKKP07, HG00, KV92, KNX01, KBF15, Laz16, LVD02, MZHB17, MM97, MC08, NWZ17, RT02, SP18, Sir19, Ste99]. approaches [KKPS18, KNY99, MMC12, MFFJ18, Mav01, NH98]. appropriate [KV96]. approximants [BLW08]. Approximate [AHJ20, Bea94, BPS00, HDIS18, LPSV18, MGF^{+}02, PPv95, ZS08, AW11, AK16, BPSH13, BSI17, CN21, DK23, Doh07, DS10, Gus03, Huc98, ISZ09, JZ09, JK17, KKNY01, KNY99, KM92, LS04, LB17, LPS15, NY03, SSB19, Sol14, VW97, WW20]. Approximated [NR17]. Approximating [DE98, VS17, AFSCSU14, CGS20, NSCTPW22, SS97]. Approximation [AEHV14, Vab20, AH02, BSC20, BE09, BF11a, BCV03, BMS17, BMS18, CCE^{+}18, CCX23, DW15, DK15, DK95, EFG^{+}18, ED22, FMPS13, HK02, HPS15, HKL21, ITS07, KKS19, KJ12, KT08, KLM15, KV15, LPS16, LV12, LZQ12, MO16, OST10b, PN18, PW12, SLV04, SLV06, Vab23, WN18, XG10, XHZ03]. approximations [CYZ99, DLVZ06, ESS23, FY01, HJR97, KKRS21, KN07, LW21, LO15, Mor07, Mor09, Per06, RSCTP15, US19]. arbitrarily [MK23]. arbitrary [BW17a, HR05]. arbitrary-degree [BW17a]. architectures [FO95]. arising [AN03b, BGM09, BFPS10, BMP11, BRT07, CZ15, DLSvL20, FP15, Gem00, HKKP07, HM14, LLV19, MZHB17, Mar16, MSV13, Miy17, MST16, PM97, Sei10, SMSW00, TC10, ZN18]. arithmetic [DK95, GKV12, TR21]. arithmetics [BB16].

ARMS [SS02]. Arnoldi [BHHJ13, GGV13, HLLL13, KR14, MP15, PRR^{+}16, VJM16, WW07, WtFW15, YYN12]. arrow [BFG95, GNQ15]. Arrowhead [Zha92]. assignment [CQX11, LC13, LW04, LW05]. assimilation [DLSvL20, TDH^{+}18, TDL^{+}22]. associated [CCG00, DKM^{+}22, IP13, MO94, Yan23]. Asymptotic [BGP97, BMS18, CGK05, FBSC21, Tre05, Lam12]. Asymptotical [DS02]. Asymptotics [BG23, TS20]. asynchronous [EGMS20, MW21, Sch99]. atmospheric [BNP15]. atomic [LO15]. Augmentation [Cao08]. Augmented [BR07, TT15, CS97, EG16, HW18, LD07, MG08, OZ22, Szu14, YPC20, Zít05]. Austin [Lee10]. Automated [SV11]. Auxiliary [KLM15, BC12, KS22, KLS23, KPV08]. averse [ADO23]. aware [DHR^{+}04]. away [IV04]. Axelsson [Cao13, Vas05]. axisymmetric [CP06].

B [Nab97, EFG^{+}18, MDMS23]. B-spline [EFG^{+}18, MDMS23]. background [LNY15]. Backward [AK23, CTP09, CMSW19, GL95a, DO18, EM11, LC07, LZ12, Peñ03, Sun05, WKS95, WW20, YDH11]. balance [GSS01]. balanced [Lot07]. Balancing [PY03, BPS13, LT09, MD03, NV08a, WLBH12]. Ball [MPR23]. BAMG [BKM^{+}12]. Banach [LZY11]. band [ESC20, VP95, dF20]. banded [BCR11, CSCTP05, CGK05, ESC18, FLM09, GSS01, Kau07, Lot07, MS14, SK21]. Barrier [Gar01, Mar95]. Barzilai [HD07]. basal [AMR18]. based [AMR18, AB12, AMMR17, AMMP06, Bai10, BZ13, BZ17, BEV22, BMAA16, BG05a, BBM^{+}06, BCZ12, BC12, BMM^{+}08, BLW08, CR20, CW97, CG15, CLNY15, Cho03, DV19, DMM^{+}08, DOR21, Don10, DKVB15, FP05, Fer96, GKL18, GN00, GB11, GZ16, GNQ15, GHW06, GKY97, HJ18, HM03, He23, Het07, HL16, HKLP19, HM16, HFG^{+}22, IV04, IIFM23, JK17, Kap98, KY95, KXZ03, KN14, KS22, KNY00, KWS^{+}18, KRW08, KLM15, Lam12, LO13, Le 23, LJ04, LNY15, LXS16, LZ22, LWZ22, LZ23, LM06, MMM06, MMPR10, MP18b, Mez20, MCLM20, NN11, Naz95, NA97, NV08b, QPS23, Reu96, RR12, RMM22, SW96, SPD05, SH14, TW20, TH19, Uhl23, UMO09, Vab23, WH94, WTWG14, XSZ09, Xie11, Xie21, XXCB20, wX15, XM17, ZSKA18, ZMO10, AL21]. bases [CV03, MP18a, MPR22, MPR23, MYZ16]. basic [BR99, BB96, MLV05]. basis [BGW05, BCHT04, CDDSC12, Gan05, KR14, LO13, Sid97, Sir19, VW97, Ver00]. Bayesian [SCP20]. BCCB [LJ04]. BDDC [Doh07, ŠBS15]. be [Ano09, PM97]. BE-FE [PM97]. becomes [NL09]. been [Ano09]. behavior [Jou94, Kem12, VL11]. behaviour [NSCTP05]. Bellman [ESS23]. BEM [FS09, HPPS03, HMS99]. Bénard [KABH17]. Beresford [Bun95]. Bernoulli [AB12]. Bernstein [BGW05, MPR22, YM22]. Besov [Dah02]. best [BDK^{+}15, FMPS13, KJ12]. better [Alb06, BG05b]. between [CH21, Li00, MC09, Tre05]. Beyond [ZN20]. BFGS [BDdSM18, DH18, YHAG20]. BFGS-like [BDdSM18]. Bi [Aih20, pLL07, LZQ12]. Bi-CGstab [Aih20]. bi-quadratic [LZQ12]. bi-symmetric [pLL07]. biconjugate [CL23]. bidiagonal [BP13, MPR20]. bidiagonalization [DDKR23, HDA19]. bidomain [MNCT07]. biharmonic [AY11, LLW09, Osw95]. bilinear [ABBP10, EAA19, GA18, JL09, STZ12]. Binary [BMP11]. Bingham [HG00]. biomechanical [LV99]. biomechanics [Axe99, NBKS99]. Biomedical [LD08, NL09]. Biot [HKLP19]. birthday [CLR13, LPQ06, SGP14, Vas03, Vas05]. Biswa [CLR13]. bisymmetric [yPyHZ04]. black [NA97, DM10, YW12]. Blaheta [AB13, Cao13]. BLAS [OOO11]. Blended [HK02, BM05a]. Block [Bai12, BHL^{+}22, CNZ17, CK10, FP15, FGNW14, GKY97, HK12, KABH17, MPS96, PS00, RS10, SLV06, AGRR21, ACR^{+}00, ACGH21, AB10, AN13, AB13, BPS15, BCR14, Bas00, BL22, Bot13, BHHJ13, CCS19, Cao08, Cao13, CNY05, CV03, CB21, DJW^{+}21, DHS95, DGM^{+}16, DFF^{+}21b, DL23, ES07, FJP12, FJP16, FS09, Gro00, HM18, Hem96, HS05, HDH19, ISZ09, KK02, KN07, KP00, KNY00, Kol05, KC17, KLMP21, Krz11, Lam12, LO13, Le 23, LPS15, MSS07, MR14, NZ14, Poi00, ST19, Ste95, SHJC18, Tre13, VVM05a, Van00, WCZ15, Wan18b, Wan18a, WH94, XCG16, YNP04, YZCQ23, Zho18, ZBCN23, Zou23, SP18]. Block-diagonal [PS00, BCR14, FS09]. block-Lanczos [Zho18]. block-preconditioned [DJW^{+}21]. block-preconditioner [ES07]. Block-row [SLV06]. block-semiseparable [VVM05a]. block-structured [HM18]. block-Toeplitz [CNY05, DFF^{+}21b]. Block-triangular [RS10]. blocking [NO04]. blocks [Cao08, JS96, KKMM12]. Blockwise [EY23]. blur [NWZ17]. Boltzmann [Lee12]. Book [DHBV21, Nab97]. Boolean [WWC^{+}15]. bootstrap [BKM^{+}12, MMPR10]. bordered [HS05, VP95]. bordering [KNY00]. Borwein [HD07]. Bott [LWW09]. bound [DD07, GX14, HVX16, LCN13, Mar94, SB03, WW20, YDH11]. bound-constrained [DD07]. boundary [BBP03, BWN05, Che15, HW19, IV04, Lee21a, MSB18, Per06, PR95, RR12, Rja98, RT99, ZYL13]. boundary-value [Lee21a, MSB18]. bounded [DGC19, EGMS20, LY15]. Bounding [AW11, Buc11, DHSW11, IK00]. Bounds [FSS18, Kol05, BC10, BF11b, CGM11, CSYS14, DS08, Du19, LS05, LW15, MO14, PPv95, Peñ09, PL21, TDL^{+}22, VR23, WKS95, WL03, YL08, YLH11, ZLLH23, Zik08, DHBV21]. Box [DM10, GH11, YW12]. box-shaped [GH11]. break [HM96]. Brinkman [He23]. Broadband [RSR10]. Broyden [DEM18, USS21]. Brualdi [Nab97]. BSSOR [GKY97]. Buckley [IK00]. building [PGT14]. Bunyakowski [AALS01]. BVM [LJ04]. BVM-based [LJ04].

C.B.S [Bla03]. C.B.S. [AM96, Mar94].

Cache [DHR^{+}04, MWZ06].

cache-oblivious [MWZ06]. CAGD [BGW05]. Calculating [RMM19]. calculation [AFS14, MK94, vKVW00]. calculations [LPW06]. Call [Ano08, LD08]. Cambridge [Nab97]. canonical [ADMS22, BVD^{+}18, DW15, EY23, MYD20]. Caputo [DA21]. Carlo [AK16, BEH^{+}17, RNV21]. carry [NN10]. Cartesian [KVC12]. cartoon [WSN19]. case [EZ96, Not10, PDV05, RSCTP20, Sha99]. Cauchy [AALS01, FG02, FSS18, SB03]. Cauchy-Bunyakowski-Schwarz [AALS01]. cavities [AG99]. cavity [WDS09]. Cayley [DHR20, HSY18]. Cell [CN21, ELV94, SRGL13, RSR10]. Cell-by-cell [CN21]. cell-centered [ELV94, SRGL13]. centered [ELV94, SRGL13]. centroidal [DE06]. centrosymmetric [HM03, fLyHZ11]. certain [BDS94, EG16, FSS18]. CFD [Tur00]. CG [Aih20, Bla02, KPT14, MB21]. CGNE [Egg07]. CGstab [Aih20]. chain [BH16, BF11b, FH94, NW15]. Chains [Ben11, BK11, BDS94, BCC98, Cas11, DMTY11, KNX01, LLV19, MPS96, NX03, Sid11]. Change [Gan05]. Changing [Mee01]. channel [PDV05]. chaos [Lee16]. chaotic [BW17b]. Characteristic [CCK06, HC20, ZYFG11]. Characteristic-mixed [CCK06]. characteristics [AC23]. Characterization [BGS21]. characterizations [ES09b]. Chasing [Zha92]. ChebFilterCG [ST17a, ST19]. ChebStaBlkCG [ST19]. Chebyshev [KR14, Li00, Ma22, PRPI09, PSK08, Wan18b]. Chebyshev-like [Ma22, PRPI09]. chemical [DO18, DK15]. Cholesky [EM95, FP95a, JO94, Kap02, Le 23, RTN03, Sau95, ZHJL12]. choosing [GNR14]. circuit [BvdV00]. Circulant [CC92, JLW05, CNY05, HN05, NR12, SPD05, VR23, WRW18, YNP04]. circulant-plus-diagonal [HN05]. circulants [GGV13, Tre13, XCG16]. class [CNT07, CCLQ18, CDDZ19, DEM18, DL23, DN12, GBB22, HES15, HL16, HZZC23, HM16, HXM19, HDH19, IK00, LT09, MP18a, Pul08, SPD05, SP06, SCD94, Wu15, Yan23, YLH11]. classes [BSI17, rFS09, Peñ09]. classic [MM97]. classical [Zou23]. classification [GMOS06, NLZ11]. CLC [Web18]. climbing [SH14]. CLJP [Alb06]. closed [EFG^{+}18, GH23]. closed-form [GH23]. closure [EJK01]. clustered [CP12]. clustering [CNZ17]. clusters [DHBV21, KBF15]. CNM [LD08, WW08a]. Coarse [GMOS06, TR21, AO07, AG19, BHL^{+}22, CRV14, DFF^{+}21a, KV96, LV12, NV08a, VSG09]. Coarse-arithmetic [TR21]. coarse-fine-mesh [AG19]. coarse-grid [AO07, DFF^{+}21a]. Coarsening [Liv04a, BBM^{+}06, DM10, GMOS06, IV04, Mar98, US19, Wan00, XM17, Yan10, YW12, ZMO10]. code [Bra02]. coefficient [DHR^{+}04, GVT03, HST22, Sau95]. coefficients [BKP02, BMMR18, CC20, KK23b, RBV08, SKKS22, Wan00, Zhu08, Zhu14]. Coffey [DPP16]. Collapsible [LD08]. collapsing [BB01]. collisions [LO15]. collocation [CDDSC12, FP15, MP18a, MPR20, MDMS23, PSK08]. column [KV15]. columns [How18]. Combination [Not02a, PW13, Shi02]. combined [KRW08, SLV13, ŠBS15]. coming [FBSC21]. Comment [Cao13, ESC20, dF20, AB13]. Comments [WTZD10, NT04].

Communication

[Lai97, Yon96, AMMR17, VY14]. Communications [LD08, NL09]. Commuting [VZ14, JMPR18]. Compact [BMM20, DEM18, DO18, DGP19]. compactly [FP15]. comparative [LR08, RS18, ST23]. comparing [MMC12]. Comparison [CGK94, CH21, Li00, PGT14, SY18b, SSB15, AG99, BB96, CP99, FLR03, FP95b, GLOW04, GLJ19, KP00, MC09, NV08a, Not05b]. comparisons [BT15]. compatible [CBE18, Liv04a]. compensated [AK94]. complement [BCGM09, CN21, HKKP07, KW99, KNX01, KLM15, LXS16, LW03, NG15, PW12, Rak99, SGP14, WW08b]. complement-based [LXS16]. complementarity [AS19, AW11, Bai10, BZ13, BZ17, CK14, DJ09, HL16, HM16, LZ22, LZ23, Mez20, XZS10, wX15]. Complementary [ZM08]. complements [BG05a, DHBV21, Kra06, MW16, NX03, WTWG14]. complete [JL09]. Completely [GL95b, BN21]. completion [BZ23, EHM95, HS18, JNS19, Laz16, SNZ20, SVV22, TYF23, WLC21, YZCQ23, YB23]. complex [AK00, CV13, GH06, HES15, HKH^{+}06, IK00, KR11, KH07, MZHB17, Not05a, SS97, Wan18a, Wu15, XQ09]. complexities [Alb06]. complexity [CCX23, DFZ05, GHJV16]. Compliant [LD08]. component [BF11b, HW22, LC21, MM02, NH06, PY22, ZN22]. component-wise [BF11b]. components [BDGL09, LB17]. componentwise [Dia09, DXW12, Lam12]. Composite [ALM18, DJW^{+}21, Fer96, MW21, RSR10, RR12]. Composite-based [RR12]. Composite-grid [ALM18]. Compress [KLMP21]. Compress-and-restart [KLMP21]. compressed [BT15]. compression [Bör17, Ibr02, LW21]. compressive [ZZ15]. Computation [BSMN22, EJK01, Mai06, Özb13, AT00, BB16, BV00, BEG18, Chu04, CLQY23, GL21, Huc98, MVK04, MM11, Miy17, MGF^{+}02, NX03, Sid97, Sir19, WLBH12, XM17]. Computational [BGM11, HJ18, CCvG06, DFF^{+}18, Ema12, GS97, Ian16, Mar00, SS07]. Computations [MPV06, Axe98, AC11, BP13, CRZT20, DPP16, DOP19, Kho96, MP18a, MPR22, MPR23, OST10b, QvGvW^{+}21]. Computed [GL95a]. computer [CZ15, DK95, GL02]. computers [JO94, MM97, Mez20, TSPSO06]. Computing [BDGL09, Dax04, GMS18, KKRS21, LCHH18, LMM^{+}23, Lor14, MRT98, NW15, YM22, YHAG20, vNR07, BL22, BGW05, BP22, CS18, CCLQ18, CJL08, CfX05, CC20, CHCS22, Dem21, DE06, FM99, HCGM23, HVCY21, KK23a, KNX01, KBF15, KMC16, KR06, LH17, LZY11, LP16, MM98, MVV08, MP16, MK20, Pul16, RT02, SLK16, SST18, SHT11, TS12, WQZ09, WW07, Yan18, YYN12, ZQ12, ZZLX20, MMMM09]. concept [Mey94]. concerning [BM05a]. Condition [BC10, CLTW11, MDB21, YDH11, ADT19, BB06, BT92, BG05b, CCG00, CDW06, DW07, Dia09, DXW12, DWWQ13, EHM95, EG16, LX08, LH08, LLW09, Pul08, TDL^{+}22, ZLLH23]. condition-number [ADT19]. conditioned [MM09, NCV05, Spi21, Ye20]. conditioning [BDGL09, LHW11, TDH^{+}18, YHAG20]. conditions [Per06, Szy94, XHZ03, Zít00, Zít05]. conduction [AJ94]. cones [LZ23]. conforming [AMM04, BMN05, KLS23, KM99, LPW06]. conic [Naz95]. conjugate [AM95, BL22, BGP97, BMSS09, BB96, CNT07, CGY22, Cha07, CL23, DMY03, DW15, DR03, GSTPT21, Hac92, HZZC23, IIFM23, Kap94, Kap02, MO94, Mey94, Not02a, PR95, WD08, Wei94, YBZ19]. Connection [MC09]. connectivity [CLQY23]. conquer [KNX01, LLLJ16, SK21]. Conservative [AIT05a, HKLP19, DKM^{+}22]. conserving [ABM17, KLS23]. Consistency [FLR03]. Consistent [Rie09, DBG06]. consistently [Bea94]. constant [AM96, Liv14, Mar94]. constrained [Ada04, AN03b, BD21, BVD^{+}18, DD07, DR03, ER96, GW00, HHM10, KV06, Lin12, LWC16, LWS^{+}23, LV98, NBKS99, PW12, PSW14, Pen08, RS10, SKKS22, SY18a, Sto92, SW12, Vla00, XJ12, PPS20]. Constraint [SL10, Ber12, BDdSM18, Cao09, DLSvL20, DGC19, fLyHZ11, pLL07, LW07, MRT02, yPyHZ04, WBL14]. constraint-preconditioned [Ber12, WBL14]. constraints [ADMS22, BPS13, Dob99, Lay05, LZQ12, MD03, MS07, dCSRS19, SW12, VFdV13]. Constructing [Uhl23, BFdP13, KKNY01, NY03]. construction [BC09, WWC^{+}15]. constructions [YNP04]. constructive [BW17a]. contact [Ada04, Hla99, IV04, NO04, ZVO14]. Continuation [DF01, HKL21, BP22, CWS97, CC03]. continuous [Cas11, LZZ20, SSB15]. continuous-time [Cas11]. continuously [Vos09]. contour [CZS22, HFG^{+}22, KKPS18]. contrast [AY11, GKK19]. contribution [WF15]. control [BLP08, BFPS10, BO13, DMS17, Dat01, GTZ18, HW19, KK13, LP22, LC13, LW05, MSS07, MP13, NV23, PSW14, QvGvW^{+}21, ROA13, SKKS22, SY18a, SW12, VFdV13, ZHJL12]. controlled [FJP16]. controller [CSB20]. controllers [Özb13]. convection [BR99, FY01, HP97, HK12, KABH17, KXZ03, PH19, RSCTP15, XG10, ZYFG11, vRH05]. convection-diffusion [BR99, FY01, KXZ03, PH19, ZYFG11, vRH05]. Convergence [BL22, BBG13, BH16, CL96, CP99, DFHM20, HNR^{+}18, HW21, IIFM23, KKO20, KO18, LT09, LB08, MD03, MM98, NH98, Pas19, Pul16, ST17a, Sch99, WCZ15, ZSCX10, Zho18, ZBCN23, Zít05, AC23, AJ94, BPS15, BS01, BGP97, BR99, BMS18, BMSS09, BLZ08, BVV12, CZ02, Che02, CJT03, CK14, DDM23, DFF^{+}21b, DS08, Du19, EN17, FVZ05, FS21, GR99, GD11, GX14, HVX16, HW18, JK09, Jou94, Kap94, Kap05, KPV06, Li00, Lin12, LW16, MRT96, MC08, ORU23, PS95, PRPI09, Pul08, RV12, SLV13, Szy94, VL11, ZW10, ZQ12, Zik08, Zít00, vdE02]. convergent [BSI17, CQ10, GT09, Sol14, ZZLX20]. convex [BGM^{+}21, Car97, GSTPT21, Laz16, LMV04, ODH21, Shi02, Shi04]. coordinate [ACGH21, BW19, TR21, YZCQ23]. copositivity [CCLQ18]. core [BH04, JYZ17, Mor07]. core-functions [Mor07]. corner [BLZ08]. corrected [BKM^{+}12, MZ15]. correction [CS02, CRV14, GS99, NV08a, NFD10]. corrections [LXS16, QXB09]. corrector [HM14]. corrector-type [HM14]. correlated [MK23, OZB^{+}18]. correlation [DBLP16, LW16]. corresponding [AT00]. Corrigendum [HS14, HS21b]. corrupted [NWZ17]. cosine [ROA13]. counts [DPS16]. Coupled [LNP12, GLOW04, HMS99, KWS^{+}18, LPV01, TSPSO06, WRW18]. coupling [FS09, HPPS03]. couplings [Yot01]. covolume [CCK06]. CP [KKS19, MVLB23]. crack [CKW02, LLW09]. Crank [LP22]. criteria [Bir15, Peñ07, Sot13]. critically [HHLL16]. Cross [OST10b, MO16]. Crout [May05, May07]. Crouzeix [KMS08, SSB04, Zhu14]. Crouzeix-Velte [SSB04]. cubic [HLLW05]. cubically [ZZLX20]. curl [BK21, CP06, KPV08, ZSCX10]. current [Bai12]. curvature [KRW08]. curvature-based [KRW08]. curvilinear [PSK08]. CutFEM [GR23]. cutting [AC23, YPC20]. cycle [BLZ08, GT09, Lai97, NN10, Not98, VL11]. cycle-convergence [VL11]. cycles [Lot23, NV08b, TGKR10, VL11, ZM08]. cyclic [MR14, NG15]. cyclically [GH11]. cylindrical [HG00]. Czech [FM99]. Czech-US [FM99].

D [DHBV21, GKY97, AM96, BV13, BG02, KK16, LPW06, Mar94, NBKS99, NSCTPW22, PM97, PR96, QB15, SY18b, ZVO14, mMP99, vKVW00]. DAE [ABK15]. damped [BC09, CMSW19]. damping [BTT13, MW21, TV20]. Darcy [He23]. Data [CCX23, GA18, Bau08, BF11a, BFdP13, BH04, CLNY15, DLSvL20, DQW15, KKS19, KK23a, LC21, NLZ11, PDV05, Rie09, TDH^{+}18, TDL^{+}22]. Data-driven [CCX23, GA18]. data-sparse [Bau08, BF11a]. Datta [CLR13]. Davidson [FJP16, GS99, HLLW05, MSV13, Not02a, Zho06, vNR07, vdE02]. DCT [CSCTP05]. DD [AB13, Cao13, AB10]. deblurring [BDRZ21, CFAM16, Don05, LNP12]. decay [FSS18]. decision [Buc11, CEQN07]. decomposable [Uhl23]. Decomposition [CGK94, AN03a, AN07, ADMS22, ADO23, AMMR17, AFK02, AG19, BP13, BW17a, Bla94, Bla02, BVD^{+}18, BPS13, BO18, BIA18, CS96, Car97, CGM01, CL13, CLNY15, CJT03, DH18, EM95, EY23, FLP00, FRR16, FGNW14, GVT03, GB15, GTI16, Gus03, HLM92, HDIS18, HC05, Ibr02, JK18, JM10, KV92, KKPS18, Kap98, Kap02, Kem12, KMMR10, Kho96, KN14, KNP03, LR95, LV99, LT09, LHW11, LXS16, LT11, LT13, Liu22, LMM00, MS22, MPR20, MD03, MK20, MM02, MVLB23, MM18, MYD20, MSF21, NR14b, PY03, PH19, SHvBW21, Sau95, SNZ20, TSPSO06, WQ07, WSN19, YL08, ZSKA18, Zhu08]. Decompositions [ZN20, BF96, BLW08, LS06, SSB04]. deconvolution [MLV05]. Decoupling [LVW01, HDIS18]. Dedicated [Bun95, SGP14, CLR13]. Dedication [NN15]. deep [PLMV23]. defect [NFD10]. defective [AFS14]. defects [KK16]. deficient [DE98, GS97]. definite [ARMW14, AIT05a, AV94, Bai16, Bai18a, BMAA16, BT03, BMM20, DJW^{+}21, DJ09, Ema12, Kap98, KH07, Kol05, LHL07b, MVV08, yPES07, SB12, WW08b]. definiteness [PW13]. definition [Lun20, VVM05c]. Deflated [CS97, MYZ16, SHJC18, MN00]. DEFLATED-GMRES [MN00]. deflation [NV08a, SLV13, ZBCN23]. degenerate [BMM06, Sto92]. degree [BW17a, DS10, Gus04b, HVX16]. delay [DOR19, DGRR11, JLW05, LC13, MSV13]. delay-differential [MSV13]. denoising [LNP12, ZZ15]. denoising/deblurring [LNP12]. dense [CDGmM04, DS10, GTY97, How18, KN07, KBF15, Ver00]. density [LMM^{+}23, NY03, OST10b]. dependency [RV12]. dependent [BEG18, CNT07, CRV14, GS05, HG00, KPT14, LP22, Mai06, MV13, RBV08, Sha98, Xie21, ZYFG11, vKVW00]. depending [Vos09]. derivative [DA21, IKAA22, KKRS21, LY15, LWS^{+}23, USS21]. derivative-free [IKAA22, LWS^{+}23, USS21]. derivatives [AT00, Xie11]. derived [BDV06]. deriving [Mey94]. descent [ACGH21, BW19, De 13, Liu22, NZ14, Shi02, Shi04, TR21, YZCQ23, ZBCN23]. design [AG99, BCK05, MC08, SMSW00]. designing [RS07]. designs [LW05]. determinantal [CC07]. determinants [MP15]. determining [WW20]. developments [SS07]. deviation [CCvG06, DDM23]. device [GMR05]. DFT [Not05a]. Diagonal [BLP17, SZ99, ACR^{+}00, BCR14, EW13, EM11, Fas05, FS09, HN05, HS05, KKMM12, MCV01, Par03, PS00, TS12, ZZ15]. diagonal-plus-semiseparable [Fas05]. diagonal-plus-Toeplitz [BLP17]. diagonalization [CCS19, MCLM20, WZZ18]. Diagonally [AK94, Yon96, MRT98, RT02]. diameter [Par03]. difference [AJ94, FY01, Fer96, Gem00, PR11, PL21, SCD94, Web10a, ZZ21]. different [DOR19, Tre05]. differentiable [Est09]. differential [AHJ20, BCR11, BCR14, BSC20, BD21, Bot13, DOR19, GBB22, HJ18, JLW05, KKO20, LH08, LHW11, LW03, MW11, MRK22, MSV13, MM11, PSK08, Rak99, RBV08, SW12, TC10, ZCW11, Zhu14]. differential-algebraic [ZCW11]. differentiation [DO18]. difficult [HST22]. diffusion [ALM18, BLP17, Bai18b, BL20, BEV22, BCV03, BR99, CCK06, CG15, DJW^{+}21, DHS23, DOR21, DA21, FY01, Gan99, GBB22, GKK19, KXZ03, KWS^{+}18, KRW08, KP10, Lee16, LCHH18, LPS15, Mav01, OC04, PH19, QPS23, RNV21, RSCTP15, Sch12, WBWM04, WZZ18, XG10, YXZ13, ZYFG11, vRH05]. diffusion- [KRW08]. diffusion-dominated [GBB22]. diffusion-wave [DA21]. digraphs [THC09]. dimension [BTT13, CLNY15, KCS11, VS17, vGSZ15]. dimensional [AALS01, CGPV13, CLNY15, DY04, DLSvL20, KK23b, KT08, LMM^{+}23, LS22, NLZ11, Özb13, Rja98, SKKS22, XSZ09, ZZ21]. dimensionality [LW21, PY22, YZ13]. dimensions [BO18, DHNR18, ŠBS15, XZS15, YZ13]. Direct [Dam08, GT19, JZ11, SH19, ZJ06, BLP01, CNY05, CS95, DOR21, ES09a, GMR05, HS05, MRT02, SW96, SST18, TW20, TSPSO06]. directed [FM18]. direction [BB96, DBG06, LZZ20, XJ12, ZN18]. Directional [Bör17]. directions [DS13b, ZS08]. Dirichlet [Rja98]. disaggregation [MM98, Pul08, PM11]. discontinuous [ABM17, BKP02, BBS12, DLVZ06, DFF^{+}18, EWY03, HHvR04, KT08, Wan00, WBWM04, vRH05]. discrepancy [BC02]. discrete [AGRR21, BCV03, BDRZ21, CLTW11, DGB^{+}13, DNR12, DHNR18, GORR16, Han13, HDA19, JK18, KM92, NR14b, PSK08, SSB04, Web10a]. discrete-difference [Web10a]. discretization [ABM17, BCR11, BS01, CGM11, DP03, GMSCS20, GTZ18, HHvR04, HK12, Lay05, LPV01, LOY08, LP22, SY18b, UMO09, Zhu14]. discretizations [AT15, BCR14, BK21, BBS12, CBE18, DMMR23, DKM^{+}22, EGF11, GHO15, HM20, HKLP19, KOV17, Lee12, Lee16, LOS04, MW11, Osw95, PT17, RS02, SRGL13, SSB15, XS11, XZS15]. discretized [Bai18b, BL20, CN21, GS07, KS04, MNCT07, MRK22, vRH05]. discriminant [NLZ11, WF15]. disks [Peñ07]. disordered [Sac05]. Displacement [Bla94, WN05, Bla02, HC20, KM99]. displaying [EJK01]. dissipative [BGS21]. dissipative-Hamiltonian [BGS21]. Distance [DFNY08, AFS14, LCHH18, NR11]. Distance-two [DFNY08]. distances [LMM^{+}23]. Distortion [BG02]. Distributed [GL18, FO95, JO94, MW16]. distribution [AFSCSU14, Ber12, BF11b, Cao09, DHSW11, GR05, MV19, SJBH14, WBL14]. Distributive [GGLO08, GLOW04]. div [AMM04, CP06, GGLO08]. divergence [MRT02]. divergences [LMM^{+}23]. divide [KNX01, LLLJ16, SK21]. divide-and-conquer [LLLJ16, SK21]. division [Kub92]. does [NN10]. Domain [BIA18, CGK94, Car97, HLM92, KKPS18, KNP03, RVW98, Zhu08, AFK02, AG19, BPS13, BO18, CS96, CGM01, FLP00, GVT03, Gus03, HKKP07, JM10, Kho96, LR95, LV99, LT09, LXS16, LMM00, MD03, MZHB17, MSF21, PY03, PR11, RT99, WLC21]. domains [Dah02, DS02, EGMS20, HKH^{+}06, KM92]. Dominant [Yon96, MRT98, RT02, ZQLX13]. dominated [AMM04, CCK06, GBB22, HP97, RSCTP15]. dominating [GGLO08]. double [AL21, DL23, QB15]. double-layer [QB15]. double-preconditioning [AL21]. doubling [GB11, HHLL16, LYL15, MP13, PR16]. doubly [GHR98]. Downwind [HP97]. DP [DHBV21]. DQGMRES [SW96]. Dr [KVW10]. Drazin [BNS20, WL03]. DRIC [Not94]. driven [CCX23, GA18]. drivings [PM97]. dual [DH04, FLP00, GH01, HP04, Saa94, Sto92, WSN19]. dual-dual [GH01]. dual-primal [FLP00]. Duffin [LWW09]. Dykstra [ER96]. dynamic [Not94]. Dynamical [Bat95, ESS23, BBJ17, BW17b]. dynamically [MN00]. dynamics [Ema12, HW19].

ECLES [dCSRS19]. eddy [Bai12]. edge [Dah02, KS22, RS02, ZSCX10]. editing [dCSRS19]. Editorial [Axe96, Axe99, Axe03, Axe04, BNR18, Bun95, KK23a, Lan97, NT03, Saa00a, Yav04, Mar00, NT04]. effect [BS01, LW04]. Effective [LH08, LLW09, HFG^{+}22]. Effectiveness [XXCB20]. Effects [CJT03]. Efficiency [DMM^{+}08, CNT07, KNY99, LH17, Tur00]. Efficiency-based [DMM^{+}08]. Efficient [BV00, BCV03, BEG18, DPS16, FJP12, Gem00, HPS15, HCGM23, Huc98, LV99, Poi00, SCP20, VP95, WWX10, mMP99, vVW23, BDS94, CP12, DJW^{+}21, DGC19, DDL^{+}21, EGF11, GM17, HS13, KBF15, KR14, LR08, LM22, MPR22, OOO11, yPxP06, RGG07, TSPSO06, WTZD10, XZS15, Zha18]. eigCG [ARSO14]. eigendata [BC09, YBZ19]. eigenfrequencies [BTT13]. eigenpair [MPV06]. Eigenpairs [ESC20, dF20, AK23, CHCS22, DK95, LP16, Xie11, ZZLX20]. eigenparameter [Vos09]. eigenproblem [BGP97, FT98, Not02a, XHZ03]. eigenproblems [Bas00, BPS00, BFG95, DS13b, FLPW01, FJP12, KCS11, Ney02, SGSM15, TY10, Vos09, XCG16, vdE02]. eigensolution [Mar16]. eigensolver [BMM^{+}08, HFG^{+}22, ZBCN23]. eigensolvers [BM17, GKL18]. eigenspaces [Zít05]. Eigenvalue [AN06, AB13, Cao13, KY95, LV04, Peñ09, AS19, AFSCSU14, AG99, AB10, Bai95, Ber12, CQX11, CR16, CR20, CCvG06, CS02, CCLQ18, CWwS18, CMSW19, CZS22, CX22, DL97, DHR20, DPS16, Dia09, DGP19, EKS02, GP18, HKST12, HS08, HMMP19, HLLL13, HLLW05, Jia17, KKPS18, KH23, LLL97, LLK14, LLLJ16, fLWyL^{+}21, Liv04b, LYL15, LS22, MMMM09, MVV08, Mee01, MSV13, MP15, MZ98, PPv95, RMM19, RMM22, SLK16, SJBH14, Sim03, Sot13, VJM16, WQZ09, WBL14, Xie21, XX22, YBZ19, YCY17, YLH11, ZQ12, ZQLX13, ZQW13]. Eigenvalues [ESC18, AT00, BB16, BL22, BWN05, BG23, CSYS14, EFG^{+}18, HCD15, HVCY21, HHQ13, Kol05, KCV09, KVC12, LS05, LQY13, Mai06, MM11, MV19, PL21, SHT11, TS20, XC13, YM22]. Eigenvector [NR19, LW98]. eigenvectors [AT00, CCS19, ESC18, HVCY21, Mai06, Pul16]. Einstein [HXM19, HL21]. elastic [GT19, Höm06, NSCTPW22]. elasticity [AM96, AALS01, Axe99, BKY10, BLE97, Bla94, BC12, BIA18, GLGR10, GL98, GL02, GL13, HNR^{+}18, KK02, KS04, KS22, KMM19, Mar94, Mar98, Pad99, Rja98, XSZ09, XS11, XZS15]. elastoplastic [MBW97]. elastoplasticity [MM97]. electrical [MC04]. electrodynamics [KMMR10]. electromagnetic [WDS09]. electromagnetism [CDG00, CDGmM04]. electron [OST10b]. Element [LV12, RSCTP20, AK99, AMM04, BBP03, BMN05, BC12, CYZ99, CKW02, CGL05, DMM^{+}08, Dob99, EGF11, EWY03, GLGR10, GTZ18, HH06, HM20, HS13, HK12, HC20, IV04, KMM18, KMMR10, KR11, KS04, KV06, Kra06, KLM14, Lai97, LV08, LR95, LMM00, LPW06, MW16, MSB18, NSCTPW22, PY03, PS00, PR95, PL21, RS02, Rja98, RSCTP15, SGP14, SY18b, SSB15, The98, Vas92, VL96, Vas02, WBWM04, XSZ09, XS11, ZYFG11, ZSCX10]. elements [BB00, GL13, HHvR04, KKO20, Lee10, Osw95, Pul09, RS02, ZHJL12, vVW23]. elimination [GIK02, Gro00, IK00, Peñ03, Reu96]. Elliptic [CGK94, AV94, BBP03, BBS12, BCZ12, CC92, CW97, CS02, CGL05, CEL^{+}96, DHS23, DMMR23, DLVZ06, Dob99, DHR^{+}04, DP03, ELV94, EWY03, GN00, GTZ18, HKST12, KKO20, KW99, KK23b, KR06, KT08, KMS08, KLM14, KM92, LPV01, Lee19, Lee21a, LW03, MRT02, MSS07, MM11, Ney02, Rak99, RT99, SKKS22, SY18a, Sta96, Vab20, VL96, Wan00, ZSCX10, Zhu08, Zhu14]. Embedded [GNR14]. embedding [FLPW01, QV21, RVW98]. EMC [Ver00]. enables [MC08]. enclosure [Miy15, OOO11]. encountered [BMMR18]. energetic [Lee12]. Energy [VSG09, BBM^{+}06, KV06, Lee12, MD03, SWY07]. energy-based [BBM^{+}06]. Energy-minimizing [VSG09]. Engine [RSR10]. Engineering [LD08, NL09, WW08a, CEQN07, Mar16, Ano08]. Enhanced [KH23, RNV21]. enriched [HDA19]. entries [EW13, FSS18, Par03]. envelope [BPS95]. Environment [ADP96, CEQN07, TT10]. environmental [MS07]. epidemic [GCLG18]. equalities [CPSM06]. equality [DR03, LV98]. equation [AY11, AB12, AJ94, BD21, BMP11, Bot13, CKW02, CD11, Cha07, CGM11, Dah02, DO18, DA21, DK15, ESS23, FZwCW17, GH23, KP10, Lee12, Lee16, LB08, LS15, LLV19, Lu05, Miy15, Osw95, yPxP06, yPES07, RV12, SY18b, SW12, Tia13, TH19, WTZD10, Zhu14, vRH05]. Equations [BFG^{+}18, GL95a, ARMW14, ABM17, AB12, Axe99, AC11, BPS15, BGX06, BCR11, BCR14, BLP17, Bai18b, BL20, BKP02, BEV22, BSC20, Bau08, BMAA16, Ben08, BLP08, BES14, BR99, BG05a, BMMR18, BG00, BHHJ13, BCZ12, BFM12, CLR01, CN21, Che02, CH03, CQ10, CL23, CSZ21, CB21, Cor04, DJW^{+}21, Dam08, DSV18, DW21, DIPR19, DOR19, DOR21, DBG06, DMMR23, DXW12, DLVZ06, DFF^{+}18, FHM21, Gan99, GB11, GBB22, Gem00, GS99, Gra08, GS07, GD11, HS21a, HM18, HFW01, HNR^{+}18, HES15, HM14, HHLL16, HXM19, HL21, IP13, IKAA22, JMPR18, JLW05, JL09, JO94, KKO20, KW99, KK23b, KXZ03, KLM^{+}06, KS04, KWS^{+}18, KOV17, KPT14, KS15, KLMP21, LR08, Lee10, LB21, LH08, LLW09, LHW11, LGS12, LXX17, LSC21, LWZ22, LZZ20, LWS^{+}23, Liv14, LW03, LPS15, LPSV18, LMM00]. equations [LRGO17, MV13, MNCT07, MW11, Mar94, MZHB17, MRK22, MM09, MCV01, MDMS23, MSV13, MM11, Miy17, NFD10, NQ96, Ols99, PM97, PR95, PR16, PT17, QPS23, Rak99, RBV08, RSCTP15, SL19, Śmi19, SCD94, Ste99, Szy94, TC10, TSPSO06, Tyr05, USS21, Var08, VHM^{+}22, WRW18, WCW20, Web10a, WZZ18, XSZ09, YDH11, YXZ13, ZCW11, ZZ15, ZN18, ZZ21, ZSWX13, Zhu08, vVW23]. equidistantly [Rie09]. equilateral [RSCTP15]. equilibrium [DHSW11, GT19]. equispaced [FP05]. Equivalence [Szy94]. equivalent [MZHB17]. Errata [SB12]. Erratum [BN12]. Error [GL95a, OOO16, AM96, AK23, AW11, CGM11, CS18, CMSW19, HJR97, IIFM23, LO13, MMN^{+}10, Ney02, Pul09, WW11]. Error-free [OOO16]. errors [LC07, LZ12, Sun05]. essentially [ZQLX13]. Estimate [AM96, CS18, ES05]. estimates [AN06, AB10, AB13, BB06, CL96, Cao13, DLSvL20, FVZ05, LZ12, MST16, Pul09, Zho18]. Estimating [FOV21, LW98, NR22]. Estimation [BNP15, GR04, ADT19, Baz08, BT92, DPS16, DXW12, LX08, LZ23, NG15, Ney02, SZ11, ZM20, ZG22]. estimations [CD11, MDB21]. estimator [MVK04]. estimators [AM96, MMN^{+}10]. Euclidean [YHAG20]. Euler [Cor04, LH17, NFD10]. European [Rag14]. Evaluating [BB01]. evaluation [FÇ23, HK21]. evaluation-interpolation [HK21]. evaluations [KS10]. even [BNS20, Not05a, XC13]. even-order [BNS20]. evolution [BBG13, vVW23]. Ewing [LPQ06]. Exact [KV15, Bot13, DK95, Pul16]. expanded [HC20]. expansion [DS02, GTI16, MS07, RR12, ROA13, SLK16]. expansions [Tre05]. expectations [FOV21]. experience [BGM11]. Experimental [RR12]. experiments [ABK97, GL02]. Explicit [Lam12, VR23]. exploiting [VJM16]. exploits [NL16]. Exploring [AMR18]. Exponential [PDV05, BV00, BCV03, DQW15, LLS12, Mor07, PS11, Rag14, Śmi19, VS17, WtFW15]. exponentially [TS20]. expressions [LT08, Not05a]. extended [BL22, DPP16, Du19, KS10, VEV23, WX21, ZHZ10]. Extending [ARSO14]. Extension [BKP02, BCB14, DHR20, HW22]. extensions [Sun06]. exterior [GH01]. extracted [SPD05, SP06]. extraction [LNY15]. Extrapolation [CRZT20]. Extremal [Jia17, LT08, Vla00, Zho16]. extreme [BL22, BHL^{+}22, HCD15, HHQ13].

F.E.M. [AM96]. Faber [Nov03]. factor [Ano09, Cha12, DM10, GIK02, GH23, HW18, IK00, KM09]. factored [KKNY01]. factoring [BG05a, Kau07]. Factorization [ADP96, BN21, ACGH21, BT03, Bla94, CCG00, CGK05, Cha12, DHS95, DCT18, FG02, GN00, KNY00, KM92, Le 23, LSS18, MW16, OS01, RTN03, Saa94, SK01, ST17b, XXCB20, XQ09, ZHJL12]. factorizations [AMMP06, Bea94, CCS10, CH94, CV03, GNQ15, HCGM23, KOV17, LW15, MS14, RS18, mMvdV02, mM04]. Factorized [KNY99, NY03]. factors [Bea94, BF11a, WL08]. families [AABHV18]. family [AEHV14, AEHV15, GGZ12, LZ09, LWC16, LPW06, MG08, Sot13, vV94]. Fast [BO13, Cao04, CWWZ22, DMTY11, DQW15, FGT11, FP05, FS09, KK16, KK23b, LLS12, LP22, LO15, LPS15, MS14, MCV01, MLV05, Miy15, Miy17, RS07, STZ12, SSB19, XCGL10, ZG22, vKVW00, BB16, CGJ21, DPP16, DS10, Fer96, HW22, JR94, Kho96, Lee10, LLLJ16, MRT02, MVV08, MW21, Rak99, RSR10, Sol14, SKR08, SK21, US19, WF15, Yan23, ZWQA18, RR12, TW20]. fast-adaptive [RSR10]. faster [Kap99]. fault [NO04]. fault-zone [NO04]. FDFD [PR11]. FE [GKY97, PM97]. feasible [AW11]. FEAST [GP18, YCY17]. feedback [DGRR11, LW05]. FEM [AB10, AB13, Beu03, BB06, Cao13, FS09, GM17, HPPS03, HMS99, KM99, Mar94]. FEM-BEM [HPPS03]. FEM/BEM [HMS99]. FETI [DH04, DKVB15]. FFT [ZVO14]. fictitious [HKKP07, RT99]. field [KMMR10, Uhl23]. fields [HPS15, OZB^{+}18]. filter [RGG07]. filtering [AN03a, AN07, BPSH13, FGNW14, LNY15]. filtering-based [LNY15]. filters [RS07]. Finding [EW13, HHQ13, DK23, PRPI09, Roh92]. fine [AG19]. finer [Vöm12]. finer-grain [Vöm12]. finger [ISZ09]. Finite [DDM23, Dob99, KMMR10, RSCTP20, AK99, AMM04, BBP03, BB00, BMN05, BC12, CYZ99, CKW02, CGL05, DJW^{+}21, DMS17, DMM^{+}08, EGF11, EWY03, FY01, Fer96, GLGR10, GCLG18, GTZ18, GL13, HH06, HM20, HK12, HC20, KMM18, KR11, Kra06, KLM14, Lai97, LR95, Lee10, LMM00, LPW06, MW16, MSB18, NSCTPW22, Osw95, PY03, PS00, PR11, Pul09, PL21, RS02, RSCTP15, SGP14, SY18b, SSB15, The98, Vas92, VL96, WBWM04, XSZ09, XS11, ZYFG11, ZSCX10, vVW23]. finite-difference [PR11]. finite-element [HM20, SY18b]. FIR [RS07]. First [KLM^{+}06, BBJ17, BGM^{+}12, GHR98, Hem96, KNX01, LY15, MMN^{+}10]. first- [BBJ17]. First-order [KLM^{+}06, BGM^{+}12, Hem96, MMN^{+}10]. FISTA [CNXY20]. fit [BDK^{+}15]. fitting [DQW15, PDV05]. fixed [BG05a, Bir15, CRZT20, HW21, KO18, LLPC23]. fixed-point [Bir15, CRZT20, HW21, KO18, LLPC23]. Fletcher [USS21, YBZ19]. Flexible [HFG^{+}22, ZHJL12, vGSZ15, Zou23]. flow [BLLA11, HG00, HK12, KR11, KRW08, Lay05, LV04, Mar00, MRT96, ŠBS15, Tur00, Web10b, Web10a, Yot01, vKVW00, LD08]. fluid [BLLA11, Ema12, HG00, HW19, Mar00, MRT96, OZ22, SV11, Web10b, Web10a]. fluid-solid [SV11]. fluidity [AMR18]. fluidity-based [AMR18]. flux [HKLP19]. flux-based [HKLP19]. Fokker [ZZ21]. following [Uhl23]. FOM [GR99]. Form [Zha92, AB10, AB13, BCB14, BO08, BWN05, BBG13, Bos19, Cao13, EFG^{+}18, GS07, GH23, GNQ15, Han13, KKNY01, LGS12, MMMM09, vNR07]. formal [Tre05]. format [BG13, BMAA16, CL23, Gra08, GL18, MRK22]. formats [DO18, DK15, HKST12]. forms [Bra02, HS05, LPS16]. formula [HS21a, MS14]. formulas [BWN05]. formulation [CQX11, GH01, ŠBS15, Ypm95]. formulations [MZHB17, PS00, Sim03]. FOSLS [MMN^{+}10, AMR18]. four [DLSvL20]. four-dimensional [DLSvL20]. Fourier [TW20, CVY21, CV13, Don10, FHM21, HM18, HM20, He21, HHvR04, MO11, ROA13, TSPSO06]. fourth [UMO09, WQZ09]. fourth-order [UMO09, WQZ09]. Foz2006 [GY08]. FP [BCB14]. fractional [BLP17, Bai18b, BL20, BEV22, DJW^{+}21, DHS23, DOR19, DOR21, DA21, DKM^{+}22, HLM^{+}18, LPS15, LPSV18, MDMS23, QPS23, SKKS22, Vab20, WRW18, WZZ18, ZZ21]. fractional-order [DOR19]. framework [BD15, FOV21]. Fréchet [KKRS21]. Fredholm [MM09]. free [ABBP10, AD11, BK21, Bos19, CH21, GTY97, IKAA22, LWS^{+}23, MP16, Not02b, OOO16, RSR10, Sim03, TT10, USS21, YNP04, ZYL13]. free-space [RSR10]. frequency [AN07, Bör17, EKS02, MZHB17, MC09, MN00, PR11]. frequency-domain [PR11]. friction [Hla99]. frictional [ZVO14]. Friedrichs [RSCTP20]. Frobenius [CDG00, DW07, ES09b, MGF^{+}02]. Frobenius-norm [CDG00]. frontal [RS01, Sco99]. frozen [AABHV18]. FSAI [FJP12, FJP16, MFFJ18]. full [BMS17, BMS18, DEM18, LW23, MWZ06, SKR08, TGKR10]. fully [KWS^{+}18, MC04]. function [CDDSC12, GGZ12, KKRS21, KKS19, KS10, LZ09, Lun20, Par03, PSW14, SP18, SST18, Tre05, XZS10]. function-generated [KKS19]. functional [KN14]. functionals [AMM04]. functions [BEG18, Bos19, CKW02, CLC11, CJL08, DOR21, DP23, DK95, Est09, FBSC21, FSS18, LMM^{+}23, Lun20, MN05, Mor07, Mor09, MP14, Naz95, Vab23, Xie11]. fundamental [ZLLH23, ZYL13]. Further [BNS20, MMN^{+}10, Saa00b]. fuzzy [CEQN07].

Galerkin [ABM17, BBS12, CGM11, DLVZ06, DFF^{+}18, HHvR04, KT08, LPV01, NSCTP05, SGP14, WTWG14, vRH05]. games [AD12]. gauge [KMMR10]. Gauss [Du19, GL21, HP97, KLN99, LO13, Peñ03, Sun06, ZG22]. Gaussian [GIK02, IK00, Reu96]. Gay [Adi08]. GCV [FRR16]. General [JK09, AN13, BCB14, BCGM09, BDR17, CS96, CCS19, CCX23, CHCS22, FÇ23, Kap98, KS15, Lor14, MP18a, NR19, dCSRS19, SZ99, SS02, Uhl23, ZW10, ZXS20]. general-form [BCB14]. Generalization [CNP96, Zít00, Don10, LWW09]. Generalizations [SSB04]. Generalized [Amb15, Che15, KKR14, NR12, AC23, AM95, BSMN22, Bla02, BC12, BMM^{+}08, CC07, Cao09, CD11, CL13, CX23a, CV03, CBE18, DL97, DHR20, Dam08, DIPR19, FT98, FM15, GIK02, GW00, HS21a, HLLL13, JMPR18, KV92, KH23, KCV09, KVC12, LR08, LW21, fLWyL^{+}21, LZY11, LT13, Mai06, MP15, MP13, MSB18, RY08, SLK16, SHvBW21, SX15, ŚLA^{+}21, WW08b, Wei94, YCY17, Zha18, Zho06, vNR07, RSCTP20]. Generalizing [BT92]. generated [KKS19, KK23b, Tre05]. Generating [Ste99, Est09, OZB^{+}18, SP18, Vöm12]. generation [BG02, Gar01, Gar04, LM06, MS07]. Genocchi [DOR19]. geodesics [SVV22]. geometric [BS10, CCX23, Cho03, CH21, CBE18, Gar04, HS11, HS14, HS21b, Ian16, LJM14, XSZ09, YHAG20, ZMO10]. geometric-based [XSZ09]. geometries [HKH^{+}06, PSK08]. Gerschgorin [LHLS07, Peñ07]. Gerschgorin-type [LHLS07]. Gersgorin [KCV09, KMC16]. Gersgorin-type [KCV09]. GES [BMM^{+}08]. GES-SA [BMM^{+}08]. gigaflops [Tur00]. given [BFdP13, CC20]. Givens [MCLM20]. Givens-like [MCLM20]. GKB [BCB14]. Global [CGM11, BS10, FRR16, GD11, LSJ18]. Globalization [NQ96]. Globally [CQ10]. GMRES [BR07, BE98, CSB20, CZ02, De 13, DS08, DN12, GR99, JYH17, Jou94, LSJ18, LM22, MYZ16, MN00, Sid11, Sim99, SWKW98, SHJC18, VL11, WZ94, ZM08, Zít00, Zít05, vV94]. GMRES-type [BR07]. GMRESR [vV94]. Goldfarb [USS21]. Golub [AGRR21, DDKR23, FRR16, GORR16, RU22]. GPBi [Aih20].

GPBi-CG [Aih20]. GPBi-CGstab [Aih20]. GPCG [Bla02]. GPCG-generalized [Bla02]. GPU [HCGM23]. grad [GGLO08]. grade [IT05]. graded [BLZ08, BCS09]. Gradient [LWZ22, AM95, BL22, BGP97, BN21, BMSS09, CNT07, CGY22, Cha07, CL23, CHCS22, DMY03, DW15, DR03, Hac92, HZZC23, IIFM23, Kap94, Kap02, LCZZ21, Liu22, MO94, Mey94, ODH21, PR95, SZ11, WD08, Wei94, YBZ19, ZM20]. gradient-like [Mey94]. gradient-type [LCZZ21]. gradients [GSTPT21, Not02a]. grain [Vöm12]. Gram [Dax04, LBG13, LL97, MPR23, ŚLA^{+}21, Van00, WL08, Zou23]. graph [FM18, KXZ03, QV21]. graphs [CNZ17, EJK01, LCHH18, VZ14]. Greedy [ZL22, BW19, BT15, HVCY21]. Grid [GVT03, ALM18, Alb06, AO07, BG02, BHL^{+}22, CGPV13, CSCTP05, CG15, CRV14, DFF^{+}21a, Don10, ELV94, FVZ05, Fer96, GKK04, Gar04, GMOS06, GHO15, HVX16, HC20, KV96, Lee18, MC08, NV08a, NN10, NH98, Not10, RSR10, RR12, ZSWX13]. grids [BH04, Bla03, ELV94, Gar01, GLGR10, LPW06, Mit10, OCYM08, YXZ13, ZMO10]. group [WN05]. growth [GIK02, IK00, KM09, WL08]. GSOR [HES15]. guaranteed [PL21]. guarantees [BZ23]. Guest [Mar00].

H [DHBV21, WCW20, Cha07, HMS99]. h-optimally [Cha07]. h-p} [HMS99]. H-TFETI-DP [DHBV21]. Hadamard [KM09]. Hamilton [ESS23]. Hamiltonian [AIT05a, AIT05b, BGS21]. Hammerstein [EAA19]. hand [ARSO14, ARMW14, SHJC18]. handy [Adi08]. Hankel [BB16, DQW15, KN07, OS01, SLV06, SB03]. Hankel-like [OS01]. Hanson [DDM23]. hardback [Nab97]. Harmonic [HS08, MZ98, Bai12, GR99, GS07, Kho96, LGS12, Vöm10, ZSWX13]. having [BG23]. heart [MC04]. heat [AJ94, SY18b]. heavily [CMSW19]. Helmholtz [CGPV13, CV13, CRV14, KMMR10, Liv04b, Liv14, OS10, RV12, TW20, TH19, TT15, UMO09]. help [GKV12]. Hermitian [LT13, SB12, BGN07, Bai16, Bai18a, BL20, BSC20, CPS01, CSYS14, DPRV19, DBG06, Fas05, HM03, HSCTP05, Kol05, KKR14, LHL07b, LC05, Mee01, NCV05, SLK16, WD08, Wu15, ZW10, ZM20, vdE02]. Hermitian-type [LT13]. Hessenberg [CGK05, Gem00, Ste95]. Hessian [BD21, TDL^{+}22]. heterogeneous [BBS12, CGPV13, GM17, KP10, KNP03, NH06]. heuristics [SH14]. Hierarchical [BH04, GBB22, MRK22, SGP14, BH07, BM13, CV03, EGF11, GL18, LO13, OZB^{+}18, Pul09, VW97, GL18]. Hierarchical-matrix [GBB22]. hierarchically [XCGL10, Xia12]. hierarchies [Alb06, DHR^{+}04, EJK01]. hierarchy [CCE^{+}18]. High [DOP21, DP23, DMMR23, Kap98, SST18, AY11, AEHV14, AEHV15, AABHV18, ABK15, BK21, Bör17, BSI17, GM17, GKK19, GKY97, Lam12, LW21, LMM^{+}23, NLZ11, NY03, SWKW98, SSB15, TSPSO06, YM22]. high-contrast [AY11, GKK19]. high-dimensional [LMM^{+}23, NLZ11]. high-frequency [Bör17]. High-order [DMMR23, AEHV14, AEHV15, AABHV18, ABK15, BK21, BSI17, GM17, GKY97, Lam12, SSB15, TSPSO06]. High-performance [SST18]. high-quality [NY03]. Higham [GIK02]. higher [BH16, GHW06, GL13, HM20, LLV19, MCLM20, WQ07, XSZ09, XS11]. higher-order [BH16, GHW06, HM20, LLV19, MCLM20, WQ07, XSZ09]. highly [BKP02, GVT03, MYZ16, Wan00]. hill [SH14]. HILUCSI [CGJ21]. homogenization [KKO20]. homotopic [CCvG06]. horizon [DMS17]. horizontal [LZ22, Mez20]. HOSVD [ZN20].

Householder [Dax04, LL97]. hp [Mit10]. hp-adaptive [Mit10]. hp-multigrid [Mit10]. HSS [Bai09, Bai18a, Bai18b, DGM^{+}16, GD11, Yan18, Zha18]. HSS-like [Bai09]. Hurwitz [KSB13]. Hybrid [HDA19, BH04, CNY05, DMMR23, FÇ23, Lai97, LWS^{+}23, LJM14, ODH21, RTN03, RK18, ŠBS15, Yan04]. hybridized [GT09]. hydrodynamics [XM17]. hyper [CH05]. hyper-power [CH05]. hyperbolic [BBG13, DFHM20, JO01, KMM18]. hyperelastic [RGM17]. hyperellipsoids [BDK^{+}15]. hypergraph [LQY13, XC13]. hypergraphs [CCLQ18, CLQY23]. hyperspectral [BNP15, LNP12].

IBLU [BLW08]. ice [AMR18]. ice-sheet [AMR18]. identification [EAA19, LNP12, ZYL13]. identify [GB15]. II [ELV94, GL02]. III [CSCTP05, GKY97, GL13]. IJNMBE [NL09]. ILDLT [Bas00]. Ill [LHW11, AGRR21, BDRZ21, CLTW11, DNR12, DHNR18, Est09, GORR16, HDA19, NR14b, NCV05, RU22, Spi21, Ye20]. ill-conditioned [NCV05, Spi21, Ye20]. Ill-conditioning [LHW11]. ill-posed [AGRR21, BDRZ21, CLTW11, DNR12, DHNR18, Est09, GORR16, HDA19, NR14b, RU22]. ILU [AMMP06, CGJ21, May05, May07, SZ99]. ILUCP [May05]. ILUT [Bas00, Saa94]. ILUT/ILDLT [Bas00]. image [BC02, BDRZ21, CFAM16, CNXY20, CNSY05, Don05, GHW06, HHM10, Höm06, JNS19, LW21, PN18, Per06, RGM17, SKR08]. images [BNT94, NWZ17]. imaging [BNP15]. IMMB [Axe99]. Impact [SWW21, Ano09]. Implementation [AK99, BISC14, BM05a, DMY03, MM18, WF15, YHAG20]. Implicit [FP95a, BGX06, Bai12, BM05a, BD15, Che15, DL23, HL16, ISZ09, LVW01, LZZ20, MC04, PBN05, VVM05b, Wan18a, ZZ21, ZS08, ZBCN23, mMvdV02]. Imposing [Szu14]. Improved [ARMW14, Cor04, JO94, LW15, BVV12, CGPV13, HZZC23, LV12, LC21, Sun06]. improvement [WCZ15, WL03]. Improvements [BB06]. improves [HVX16]. Improving [BKY10, CSB20, DV19, GKL18, GKV12, ST17b]. impulse [LCZZ21]. including [CDDZ19]. inclusion [LHLS07, LLK14, THC09]. inclusions [GKK19]. Incomplete [FP21, Jia96, BT03, Bla94, CCS10, GNQ15, Gro00, JO94, Kap02, KNY00, RTN03, Reu96, Saa94, SW96, Sau95, ST17b, VS17, XXCB20, ZHJL12, mMvdV02, mM04, GKY97]. incompressible [BKP02, DFF^{+}18, HW19, HK12, KOV17, LV04, Ols99, OZ22, Tur00, Web10b, Web10a, vKVW00]. incorporated [TYF23]. increasing [DMY03, HVX16, NR22]. increasing-angle [DMY03]. Incremental [CCS10, BT92]. indefinite [BRT07, CL96, CK01, CS95, CRV14, GM17, GMTV16, Krz11, LT09, Liv14, PS00, ST17b, SL10, TT15, Vas92]. Indefinitely [DR03, LV98]. Independence [He21, DS08]. independent [CCS19, CJL08, FS21, KPV06]. index [RNV21]. indirect [BLP01]. individual [PL21]. induced [Lay05, vGSZ15]. industry [mM04]. inequalities [AM96, CPSM06, DKVB15]. inequality [AALS01, Bla03, DGRR11, DH04, DR03, EM95, Mar94]. Inertia [CR20, DCT18, KC17]. Inertia-based [CR20]. inertia-revealing [DCT18]. Inexact [ABK97, DDKR23, HD07, Sid11, XX22, Bir15, CQ10, Dax19, FK15, GB11, GP18, HLM92, HW18, KK02, KPV06, LLL97, LZZ20, LV98, MB21, Sim03, WtFW15]. infimum [Chu04]. infinite [BMMR18, Özb13, VJM16]. Information [Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano14b, Ano14f, Ano15f, Ano15a, Ano15b, Ano15c, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18f, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano14a, Ano14c, Ano14d, BF96, FJ05, TYF23, Ano14e, Ano15d, Ano16f]. information-incorporated [TYF23]. initial [Nov03, PBN05, VL11]. initializing [BMM^{+}08]. inner [DDKR23, FJP16, Gus04a, Mey94, MGF^{+}02, Xia12]. Innovative [BDRS12]. inpainting [JNS19]. integer [CP12]. integers [DOP21]. integrable [SHT11]. integral [AFSCSU14, CZS22, HSY18, MM09]. integrals [LO15]. integration [ABK15, FS21, HFG^{+}22, KKPS18, LLS12, MC09]. integration-based [HFG^{+}22]. integrator [SL19]. integrators [Ber01, LJ04, Mor07, Rag14]. integro [GBB22]. integro-differential [GBB22]. intensity [GKV12]. inter [MC08]. inter-grid [MC08]. interaction [SV11]. interchanges [EM11]. interdisciplinary [BNR18]. interest [FOV21]. Interface [Wan00, JM10, XM17, Yot01, ZYL13]. interface-based [XM17]. Interior [LMV04, PPS20, BMM06, BGM^{+}21, BCS09, BPS13, HP04, MST16]. Interior-point [LMV04, PPS20]. internal [HKH^{+}06]. International [NL09]. Interpolating [MN05]. interpolation [BKY10, DFNY08, Gan05, HK21, HM03, KKS19, KV06, KV15, LY15, MMPR10, Pul16, Rie09, TV20, Vla00, Web10b, Yan10]. Interpreting [CPSM06]. interval [DPS16, Jia17, KSB13, Roh92, SH19, YLH11]. intervals [Jia17, LHLS07, THC09].

Introducing [MS07]. invariance [JYZ17]. invariant [AG95, DF01, MK94, MP16, YL08]. Inverse [LC05, NR14a, SP18, Tre13, AS19, AEHV14, BF11a, BNS20, BM13, BPS00, BFG95, BFM12, BSI17, CC07, CWwS18, DL97, DHR20, DW07, DWWQ13, EW13, EKS02, Egg07, EHM95, FGT11, FP21, FK15, Han13, ISZ09, JZ09, JK17, JK18, KKNY01, Kho96, KNY99, KKMM12, LLL97, pLL07, LWW09, LZY11, MB21, Ma22, MV13, MP16, MGF^{+}02, NY03, yPyHZ04, SCP20, Sol14, Sot13, TS12, WL03, XHZ03, XCG16, YBZ19, ZN18, Zho06, Ney05]. inverse-free [MP16]. inverses [BSMN22, Cor04, DK23, FSS18, Gus03, Huc98, LXW13, VR23, WN05]. inversion [BO13, KK02, LPS15, LPSV18]. inversions [Dax19]. invert [MP14, PS11, WtFW15, Sim03]. invertibility [Den09]. investigation [KS10]. involving [DA21, DWWQ13, HL21, PPS20]. IOM [Jia96]. ion [LO15, TC10]. ion-atomic [LO15]. IPARS [LVW01]. IRAM [Xie11]. IRAM-based [Xie11]. Irreversible [BL03]. ISBN [Nab97]. isogeometric [CBE18, EFG^{+}18, GMSCS20]. isolation [EKS02]. isometric [Gar01, Gar02]. isospectrally [VW15]. isotropic [BL20]. Issue [Ano08, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano15f, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano18f, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, LD08, CLR13, Dat01, Fal06, VW01, Vas05, Axe99]. issues [BM05a]. Iterated [BDR17, AN03a]. iterates [DS13b]. iteration [AS19, AT15, AN94, BGX06, Bai10, Bai12, BZ13, BLP17, Bai18a, Bai18b, BL20, BM13, CH05, Che15, CX22, DL23, Egg07, FK15, GB11, GH01, HMS99, HL16, HW21, HFG^{+}22, KO18, Kra02, KKR14, LLL97, LLPC23, Lam12, LS15, LZZ20, MM18, PS95, Pas19, Spi21, Wan18b, Wan18a, wX15, YHS18, Zho06, ZS08, Ney05]. iterations [BGN07, BG05a, FJP16, GGZ12, HN05, Kap05, KLN99, LZ09, Lin12, Lu05, NZ14, Saa00b, Sch99, ZM20, vdE02]. Iterative [AT00, BF11b, CGK94, DBG06, GMR05, LPV01, MO16, MSB18, NZ14, PM97, AEHV14, AEHV15, AK00, ABNP15, BEH^{+}17, BM17, Ber01, BR99, BN21, BDRZ21, CR16, CH05, CK01, CK10, DA21, ELV94, FM99, GKK19, GTY97, Gus97, HG00, HES15, HY22, HM14, HW22, LR08, Lee10, LP22, LSL01, LWZ22, LZY11, LW16, LCZZ21, LJM14, MM98, NO04, OC22, Ols99, yPxP06, PR96, PR11, PH19, Pul08, PM11, SH19, Śmi19, Sol14, Sun06, Szy94, WDS09, WCW20, WTZD10, WW11, WX21, ZW10, Axe99]. IV [KNY99]. Ivo [SGP14].

J [NN15]. Jacobi [ESS23, BFdP13, BFG95, FJP16, GS99, HLLW05, MSV13, Not02a, Sch99, Zho06, vNR07, vdE02]. Jacobi-Newton-iterations [Sch99]. Jacobian [BS01]. January [NL09]. joint [CCS19, MCLM20]. Jordan [EJK01, GH06, Peñ03]. Journal [JNL92, NL09]. Jr. [KVW10]. jump [DSV18, VFdV13, Zhu08, Zhu14]. jumping [KK23b]. jumps [LLS12]. justifications [Gar04].

Kaczmarz [Du19, YMS^{+}23, ZL22]. Kahan [RU22, AGRR21, Bun95, DDKR23, FRR16, GORR16]. Kalman [BPSH13]. Karhunen [SLK16]. Keller [RSCTP20]. kernel [CCX23, HK02, MN05]. kernels [NWZ17]. kind [AK19, MM09]. kinds [DOR19]. KKT [BGM09, BDdSM18, MST16]. known [CC20, EFG^{+}18]. Kronecker [BW17a, Che15, DWWQ13, EJK01, KN07, LPS16, LS04, Per06, XG10]. Krukier [JK09]. Krylov [HS14, HS21b, OC04, AGG^{+}16, AFSCSU14, BPSH13, BMAA16, Bot13, BD15, CS97, CQ10, CK10, Dam08, DHS23, Dax19, DGP19, DK95, Ema12, EN17, Fas05, GP18, GLJ19, HJ18, HS11, HXM19, IT05, JMPR18, KKRS21, KS10, KLMP21, Mor07, MP14, NV08b, PPv95, Rag14, RLG12, RV12, SCP20, Sid97, SS07, VS17, WW20, XX22, Yot01]. Krylov-accelerated [Ema12]. Krylov-based [HJ18, NV08b]. Kutta [Che15, FS21].

L [Nab97, CZ02, DH18, ZMO10]. L-BFGS [DH18]. L-shaped [ZMO10]. L. [JK09]. Lagrange [Cor04]. Lagrangian [EG16, MG08, MP16, OZ22]. Lagrangian-type [EG16]. Lagrangians [LD07]. Laguerre [DOP19]. Lamé [BKP02]. LAMG [FM18]. Lanczos [ARSO14, Aih20, AGRR21, BB16, BBJ17, CGY22, CS18, CWS97, CC03, FG02, FJ05, GORR16, Lam12, LW98, Mee01, Mor09, PV99, PS11, Par92, Sim03, Zho18]. Lanczos-type [Aih20, CWS97, FG02]. Laplace [QB15, SLV13]. Laplacian [CV13, FM18, HM20, TT15, UMO09, XC13, DHBV21]. Laplacians [BO08]. Large [Ben08, Jia96, VW01, AHJ20, AG99, ADT19, Axe98, BW19, BBJ17, Bar02, BCB14, BLP08, BES14, BV00, BG00, BG05b, BHHJ13, CLR01, CRS05, CGJ21, DMY03, Dax94, DNR12, DGM^{+}16, DGP19, DHBV21, DR03, EW13, ED22, FBSC21, FJP12, GLJ19, GTY97, Gra08, GR04, HJ18, JZ09, JK17, KBF15, LLL97, Lee16, LV98, Mar16, MZ98, RK18, SCP20, Sid11, Sir19, VS17, WDS09, Xie11, vGSZ15].

Large-Scale [VW01, Ben08, BBJ17, Bar02, BCB14, BLP08, BES14, CGJ21, DMY03, DGP19, GLJ19, Gra08, GR04, HJ18, Sir19]. large-size [FJP12]. largest [LW98, WQZ09, ZQ12, ZQW13]. latency [RTN03]. lattices [KK16]. Lawson [DDM23]. layer [QB15, RV12]. layered [BDM^{+}14]. layout [SWW21]. Lazarov [Vas03]. LD [GCLG18]. LD-QBD [GCLG18]. LDL [mM04]. leading [CC20]. learning [SZ11]. Least [CYZ99, pLL07, PY22, Tia13, AB00, AK99, BDGL09, BW19, Bar02, BMM06, BGM09, BGM11, BGM^{+}12, CNP96, CTP09, CP12, CP06, Dax94, DE98, DH18, DW07, DWWQ13, DDL^{+}21, EAA19, ES07, ES09a, ER96, EY23, FB95, GW00, GR05, KMM18, KLM^{+}06, LVD02, LZ12, LW17, LL97, MS22, MMN^{+}10, MVK04, MLV05, MDB21, MVLB23, MYD20, Miy15, Pen08, Ren98, RLG12, Sto92, TDH^{+}18, Vab23, WKS95, WWC^{+}15, XXW19, ZHZ10, ZY19]. least-rank [Tia13]. Least-squares [CYZ99, pLL07, Tia13, AK99, BDGL09, BW19, Bar02, BMM06, BGM09, BGM11, CTP09, CP06, DW07, ES07, ES09a, ER96, KMM18, LVD02, Pen08, TDH^{+}18, ZHZ10]. Left [WD08]. lemma [Gus04a, Mar95]. length [BDK^{+}15]. Level [SH14, CGM01, CS02, CRV14, DLVZ06, EN17, GVT03, HH06, HM20, HHvR04, KM99, KWS^{+}18, KV96, NCV05, OC04, SZ99, SP06, SV11, TSMM21, VSG09, XZS10, YXZ13, Zik08, vRH05]. Level-based [SH14]. level-dependent [CRV14]. Levinson [Bun92]. life [KVW10]. like [Bai09, Bai16, BMM06, BDdSM18, GL21, Lee10, Ma22, Mey94, MCLM20, OS01, PRPI09, mMP99]. likelihood [ES05, NG15]. limit [LY15]. Limited [GMTV16]. limiting [DS13b]. line [BDK^{+}15, DMY03, MM95]. Linear [NLA94, Ano09, BVD^{+}18, ITS07, Jia96, Nab97, ZQ12, ARSO14, ARMW14, Ada04, AGG^{+}16, AW11, ADMS22, ACR^{+}00, AIT05b, JNL92, AL21, AMP99, AK00, AN03b, BDGL09, BPS15, Bai10, BCR11, BZ13, BCR14, BLP17, BZ17, Bai18a, BW19, BKY10, BG13, Bas00, BLE97, BLP08, BFPS10, BEH^{+}17, BDdSM18, BMM20, BGM^{+}21, Ber01, BWN05, Bla02, BMS18, BvdV00, Bos19, Bot13, BC12, BFM12, BM05a, BDRZ21, BIA18, BSI17, CCX23, CS09, CS11, CDGmM04, CPSM06, CSCTP05, CGL05, CC03, CK01, CK14, DGB^{+}13, DMS17, DSV18, DW21, Dat01, DDG99, DFF^{+}21a, DGRR11, DW07, DWWQ13, DNR12, DGM^{+}16, DFF^{+}21b, DJ09, DL23, DN12, FZwCW17, FÇ23, FGT11, FP15, FM18, FS09, Gem00, GLJ19, GM11, GSS01, GY08, GTY97, GS05, GW00, GL98, GL02, GL13, HS21a, HLM^{+}18]. linear [HHvR04, HNR^{+}18, HES15, HY22, HSCTP05, IIFM23, JZ09, JK17, JYH17, Jou94, JO94, KMM18, KK02, KPV06, KS04, KBF15, Kra02, KS15, KKR14, KMM19, LX08, LHL07b, LT09, LC13, LZ23, LL97, LV98, LMV04, Mar00, MCV01, MV05, Mav01, MP13, Mey94, Mez20, MV19, MC04, Naz95, NQ96, NLZ11, Nov03, OC04, Özb13, Pad99, PBN05, PM97, PGT14, RK18, RGG07, dCSRS19, RT99, SZ99, SS02, SCP20, SB12, SS07, SMSW00, Spi21, Sto92, Sun05, SL10, SHJC18, Szu14, TYZL23, TT10, TR21, VFdV13, VW01, WKS95, WD08, WM12, Wan18b, Wu15, WF15, XSZ09, XS11, XJ12, XZS15, wX15, YDH11, Yan23, Ye20, ZW10, ZL22, ZXS20, vGSZ15]. linear-constrained [XJ12]. linear-quadratic [BLP08]. Linearization [LZ12, KABH17]. linearizations [CMSW19, KR14]. linearized [BGX06, NFD10, SL19]. linearly [Bla94, CCS19, LVD02, NSCTPW22, Sto92]. lines [ZLLH23]. Lipschitzian [DS02]. load [WLBH12]. Local [BZ23, CGM01, CV13, ELV94, HM18, MO11, ORU23, BS01, Don10, FHM21, He21, Kra06, MMN^{+}10, MM95, Pul08, dCSRS19].

Localization [KVC12]. localizations [KCV09, KCC16]. localized [HVCY21]. Locally [RSCTP20, BB00, BL22, KR11, MK23, ZZLX20]. location [LC21]. locations [BB97]. Loève [SLK16]. logarithm [Lor14]. logarithmic [DHW16]. Long [Kem12, KK16, Yan10]. long-range [KK16, Yan10]. Long-time [Kem12]. look [LYL15]. loosely [TSPSO06]. Low [AN07, Bau08, BF96, CH94, DFZ05, ŚLA^{+}21, VHM^{+}22, WN18, WLC21, YZCQ23, AT15, AMMR17, BE09, BHL^{+}22, CCX23, CWWZ22, DPRV19, DBLP16, ESS23, ED22, Gra08, HS18, HC05, JMPR18, KKRS21, KO18, KPT14, KS15, Laz16, LXS16, LW21, LO15, MRK22, NL16, NY03, ORU23, QXB09, SLV04, SLV06, Tyr92, ZXS20, ZG22]. low-communication [AMMR17]. Low-complexity [DFZ05]. low-density [NY03]. Low-order [VHM^{+}22]. Low-rank [BF96, CH94, WN18, WLC21, AT15, BE09, BHL^{+}22, CCX23, CWWZ22, ESS23, ED22, Gra08, HS18, HC05, JMPR18, KKRS21, KO18, KPT14, KS15, Laz16, LXS16, LO15, MRK22, NL16, ORU23, QXB09, ZXS20, ZG22]. Lower [ZLLH23, Alb06, SPD05, SP06]. LQ [BG00]. LQ-Schur [BG00]. LSQR [RY08]. LTI [ZS08]. LU [CCS10, LW15]. Lucas [DOR21]. Luré [PR16]. Lyapunov [BLP08, CSZ21, Dam08, DSV18, KPT14].

M [KVW10]. maintaining [Par92]. Making [LSJ18, CEQN07]. manifold [KO18]. manifolds [MK94, SZ11, SVV22]. manipulations [HK21]. Manteuffel [Lee10]. manufacturing [CNY05]. mapping [BG02]. mappings [BGS21, Gar02]. maps [MK94]. Marek [SGP14]. Markov [AD11, BLLA11, Ben11, BK11, BL03, BDS94, BH16, BCC98, Buc11, BF11b, Cas11, DSV18, DHSW11, DMTY11, FH94, KNX01, LLV19, MPS96, NX03, NW15, Sid11, VFdV13].

Markov-modulated [BLLA11]. Markovian [BMP11]. mass [ABM17, EKS02, KLS23]. mass-conserving [ABM17]. master [DO18, DK15]. matching [BCZ12, DGC19, KXZ03]. matchings [HS15]. material [LNP12]. materials [BDM^{+}14, PR11]. Mathematical [SWY07]. mathematician [Voe92]. Matlab [Bra02]. Matrices [DKM^{+}22, Yon96, AFSCSU14, AIT05a, ADT19, AN94, AN06, AB10, AN13, AB13, Axe15, Bai16, BB16, BPS95, BP13, BNT94, BH07, BF11a, BF19, BM13, BT03, BV00, Ber12, BWN05, BG05a, BG23, BFG95, BN21, BG05b, BFM12, BCC98, BCGM09, BM05b, BM06, CS96, CCX23, Cao08, Cao09, Cao13, CDDSC12, CCLN05, CGK05, CfX05, CDDZ19, DLSvL20, DPP16, DOP19, DOP21, DP23, Dem21, Dia09, DS10, Don10, DNR12, DS13a, Dos99, DCT18, DHNR18, ESC18, ESC20, ES09b, Est09, EG16, FLR03, FG02, Fas05, FP95a, FBSC21, FP21, FSS18, GIK02, GS97, GR04, HH06, HLM^{+}18, HR05, HS15, How18, Hua12, HC05, Ian16, IK00, JR94, Jia17, Kau07, KN07, KS22, Kol05, KC17, Kra02, Kra06, Le 23, Leb02, LVD02, LSL01, LS05, LS06, LHL07a]. matrices [LW21, pLL07, Mai06, MP18a, MPR20, MPR22, MPR23, MM98, Mar16, MM09, Mat96, MW16, MDMS23, MCC^{+}12, MN05, MYZ16, NSCTPW22, NR11, NPR13, NR19, OS01, Peñ09, yPyHZ04, Poi00, RSCTP20, RMM19, Sei10, SJBH14, SS97, SB03, Sol14, SST18, Sun06, SK21, TS20, Tre05, Uhl23, VVM05a, VP95, VVM05b, VVM05c, VW15, Vas92, VR23, WBL14, XCGL10, XHZ03, XM17, YM22, YPC20, YLH11, ZLLH23, ZHZ10, Zho16, dF20, vN00, Nab97]. Matrix [AB00, AG95, AC11, BK21, BFG^{+}18, Bun92, GTY97, Not05a, YNP04, Zha92, AK23, AFS14, AH02, AEHV15, AD11, Bai10, BSC20, BBJ17, BE09, BFdP13, BB01, Ben08, BGW05, BG05a, BMMR18, BEG18, BG00, Bör17, Bos19, BHHJ13, CCE^{+}18, CCS19, CCG00, CH03, CLC11, CSYS14, CGS20, CH21, DPRV19, DBG06, DGRR11, DGM^{+}16, DK95, DBLP16, EW13, EM95, EHM95, ER96, FLPW01, GBB22, GMS18, GHR98, GGZ12, Gra08, HJ18, HK02, HK21, HM03, HS21a, HVCY21, HL16, HM16, HHLL16, IP13, Ibr02, JNS19, JZ11, KV92, KKRS21, Kap98, Kap99, KNX01, KH07, KS10, KO18, KM09, KR14, KPT14, KS15, KLMP21, LZ09, Laz16, LOY08, fLyHZ11, fLWyL^{+}21, LZ22, LZ23, pLL07, LH17, LT08, LT11, Lor14, LPS15, MVV08, MSS07, MRT98, Miy15, Mor09, MP14]. matrix [OOO11, OOO16, PS11, yPxP06, yPES07, QvGvW^{+}21, Rja98, Roh92, ST23, Sau95, Sha98, Ste99, SHT11, TS12, TT10, THC09, Tia13, TY10, US19, Vas02, VS17, WW08b, WTZD10, WtFW15, WF15, XJ12, Xie11, XQ09, wX15, YDH11, Yan23, YHAG20, ZJ06, ZN18]. matrix-dependent [Sha98]. Matrix-free [BK21, GTY97, YNP04, AD11, Bos19, CH21, TT10]. matrix-sequences [BSC20]. matrix-valued [DGM^{+}16, Xie11]. max [BDK^{+}15]. max-length-vector [BDK^{+}15]. maximal [LW16, RMM19]. maximization [DDM23, SH14]. Maximum [BCHT04, Gar02, CCLQ18, ES05, NG15]. Maximum-weight-basis [BCHT04]. Maxwell [GS07, LGS12, MV13, MZHB17, ZSWX13]. McCormick [Lee10]. mean [Ian16, KNX01, YHAG20]. means [MS14, RNV21]. measure [BG02]. measures [Buc11, OST10a]. mechanical [LV99]. mechanics [Ada04, Axe99, GMTV16]. mechanism [DH18]. mechanisms [MYD20]. Medal [Ano08]. media [BKP02, CGPV13, GM17, KP10, NH06, ŠBS15, WWX10, Yot01]. Median [LNY15]. Memory [KR14, FO95, GMTV16, JO94]. Memory-efficient [KR14]. memoryless [USS21]. meromorphic [BEG18]. Mesh [KPV06, AG19, BC10, BGM^{+}12, DJW^{+}21, DHR^{+}04, DS08, HST22, KPV08, ŠBS15, YPC20]. Mesh-independent [KPV06]. meshes [BB00, BLZ08, BCS09, CH21, HMS99, KR11, KV96, Mav01, OZB^{+}18, RSCTP15, SRGL13, XZS15]. meshfree [CN21, LOY08, LOS04]. Meshing [HKH^{+}06]. Method [Jia96, AC23, Aih20, ABBP10, AK99, AN94, AM95, AFK02, AG19, BC09, BG13, BB16, BBJ17, BMM06, BES14, BL22, BS01, BGM^{+}21, Bla02, Bot13, BHHJ13, BMSS09, BCZ12, BC12, BCS09, BPS13, BP22, BDRZ21, CKW02, CZ02, CNT07, CQX11, Cha07, CGL05, CH05, CG15, CS18, CZS22, CNY05, Cho03, CK01, CBE18, CP06, CHCS22, CK14, DL97, DMY03, DHR20, DHS23, DDKR23, Dax94, Dax19, DOR19, DOR21, Dem21, DA21, DGM^{+}16, DJ09, DGP19, DS13b, DR03, EKS02, ES09a, EWY03, FLP00, Fer96, GBB22, GHT09, GS99, GT09, GT19, GD11, Hac92, HK21, HCD15, HKKP07, HS18, HES15, Höm06, HD07, HHQ13, HC20, HLLL13, HW18, HSY18, IKAA22, JM10, Kap94, Kem12, KY95, KKNY01, KK16, KW99, KXZ03, KPV06, KR11, KS10, KS22]. method [KLS23, Kra02, KT08, KLM15, KPT14, KM92, LV08, LPV01, Li00, LT09, fLWyL^{+}21, LW23, LB08, LS15, LH17, LW17, LCZZ21, Liv14, LJM14, LP16, LPSV18, LS22, LMM00, LV98, LMV04, MZ15, MB21, Ma22, MO94, MM98, MRT96, Mee01, MSV13, MP15, MWZ06, MBW97, Mit10, MP14, MN00, NQ96, NR14b, Not94, ODH21, PS11, PS95, PY22, yPxP06, PR95, PR96, PR11, PT17, Rak99, RU22, RS01, RS02, RV12, Reu96, dCSRS19, RT99, ROA13, RMM22, SKKS22, Sha99, Sim03, Śmi19, Sun06, SHJC18, SK21, TS12, Uhl23, USS21, WD08, WQZ09, WCZ15, Wan18b, Wan18a, WBWM04, WSN19, WTZD10, Wu15, WX21, XSZ09, XJ12, XZS15, Xie11, Xie21, XQ09, XX22, YHS18, Yan18, YBZ19, YYN12, YXZ13, YZCQ23, YHAG20, ZYFG11, ZN18, ZLLH23, ZYL13, Zít05, ZM20, ZMO10, ZG22, vNR07, vRH05]. Methods [Ano08, CGK94, Den18, LD08, NL09, QACT18, VW01, WW08a, ARMW14, AM96, Ada04, AD12, AGRR21, AEHV14, AEHV15, AABHV18, AMMP06, AK94, AV94, Axe98, Axe99, AK00, AN03b, ABNP15, ABK15, Axe15, BR07, BGX06, Bai09, Bai10, Bai12, BDRS12, BZ13, BCR14, BLP17, BZ17, BNR18, Bai18a, Bai18b, BW19, BL20, BP13, BLE97, Baz08, BMAA16, BGM11, BK11, BEH^{+}17, BGP97, BR99, BGW05, BDV06, Bra21, BCS09, BO18, BB96, BM05a, BSI17, BHL^{+}22, CEQN07, CS09, CS11, CGM01, CS02, CSCTP05, CEL^{+}96, Che02, CCK06, Che15, CNZ17, CWwS18, CL23, CWS97, CK10, CRZT20, Dam08, DMM^{+}08, DMTY11, Den12, Den14, DBG06, Dob99, DFF^{+}21b, DL23, DFF^{+}18, EZ96, EGMS20, EN17, EM11, ELV94, Fal06, Fal10, rFS09, FM99, FM15, FP95b, GB11, GMSCS20, GLGR10, GORR16, GLJ19, GZ16]. methods [GVT03, GMR05, GGV13, GKK19, GMOS06, Gus97, GL95b, HJ18, HS11, HS14, HS21b, HMMP19, HL16, HY22, HP04, How18, HKL21, HZZC23, HXM19, HDH19, HLLW05, IV04, JMPR18, JS96, KMM18, KKRS21, KMMR10, KK23a, KP00, KCS11, KLM14, KS15, KLMP21, KKR14, Lee10, Lee12, Li00, LSL01, LHL07b, LLW09, LNY15, LXX17, LWZ22, LZY11, LW16, LLV19, LWS^{+}23, LMM00, MMC12, MMMM09, Mar00, MG08, MDMS23, Mez20, MPS96, MZ98, MST16, NBKS99, NSCTP05, Not05b, Not10, ORU23, PBN05, PPS20, PY03, PRPI09, Pul08, PM11, Rag14, SCP20, SRGL13, SB12, ST23, SK01, SWY07, Sei10, Sid11, SS07, SH19, SGP14, Sta96, SY18b, Szy94, TYZL23, VSG09, VZ08, WCW20, Wei94, Wie99, wX15, ZW10, ZL22, ZSCX10, ZSWX13, Zho18, Zik08, vV94, Fal08, GL02]. Meyer [WSN19]. MILU [WH94]. Mindlin [CYZ99]. minimal [BGX06, CfX05, JR94, KMC16, MRT96, RMM19, SW96, Sta96, ŚLA^{+}21]. minimal/maximal [RMM19]. Minimization [EHM95, CDG00, Car97, DMY03, DFZ05, Het07, KV06, MD03, NZ14, XJ12, Yan18]. Minimizing [CvG11, GSTPT21, AMM04, VSG09]. Minimum [GH01, DE98, DBG06, DS10, Gus03, HMS99, Kap05, Miy15, Saa00b]. minmax [Vos09]. MINRES [SHZ20, KK13]. mirror [BCK05]. miscible [HC20]. Mixed [CGY22, DXW12, KMM18, AB10, AB13, BBG13, Cao13, CEL^{+}96, CCK06, GH01, GTZ18, GT09, GS07, HC20, Lai97, LPV01, LGS12, LW17, OC22, PY03, PS00, PT17, RVW98, SWW21, ŠBS15, VL96, WBWM04, Web10b, YZ13, ZY19]. mixed-hybrid [ŠBS15]. mixed-order [Web10b]. mode [STZ12]. Model [Lay05, Sha99, AMR18, BLLA11, BBJ17, FLPW01, GA18, Gus98, KNP03, Lee18, MV13, WSN19, XG10, ZS08]. model-order [MV13]. modeling [FH94, WWX10]. modelling [Gar04, GMR05, LO15, NH06, SWY07]. Models [CEQN07, Bai12, BL03, BV13, Buc11, DHSW11, GM17, GCLG18, GB15, HKLP19, LNP12, PGT14, QXB09, TC10]. modern [MM97]. Modifiable [BE09]. modification [CSYS14, ZG22]. Modified [LHL07b, wX15, Bea94, CS95, DJ09, Kap02, KPV06, NR14b, Sun06, WL08, ZZ15, SB12]. Modifying [Alb06]. Modular [BC02]. modulated [BLLA11]. Modulus [Bai10, BZ13, BZ17, HL16, Mez20, DJ09, HM16, LZ22, LZ23, wX15, YHS18]. Modulus-based [Bai10, BZ13, BZ17, HL16, Mez20, HM16, LZ22, LZ23, wX15]. moment [AK16, GHR98, VFdV13]. Moments [BFM12, HFG^{+}22]. Momentum [MYD20]. Monotone [IV04, IKAA22, LWS^{+}23, USS21, ZZ15]. monotonic [LD07]. monotonicity [Mar95]. Monte [AK16, BEH^{+}17, RNV21]. Moore [DW07, DWWQ13, KKMM12, LXW13].

Moreau [PSW14]. Morrison [HS21a]. mortar [DP03, PY03]. motivation [MM18]. Motzkin [ZL22]. MRRR [MPV06]. MSMAOR [CK14]. Multi [NH06, TYF23, BCK05, CS02, CLNY15, Lee12, PDV05, RNV21, SZ99, SV11, TC10, XM17, ZHJL12, vGSZ15]. multi-channel [PDV05]. multi-dimensional [CLNY15]. multi-energetic [Lee12]. multi-index [RNV21]. multi-ion [TC10]. multi-level [CS02, SZ99, SV11]. multi-mirror [BCK05]. multi-parameters [ZHJL12]. Multi-scale [NH06, XM17]. multi-shift [vGSZ15]. Multi-view [TYF23]. Multicolor [ZXS20]. multidimensional [BBKY06, LO15]. Multifrontal [ADP96, NL16]. Multigrid [AD12, BB00, BW17b, Bra21, BCS09, BO18, BBKY06, Den12, Den14, DFF^{+}21b, Fal08, Fal10, GLGR10, KRW08, Lee18, Lee19, Mav01, SRGL13, Wie99, WTWG14, ZVO14, Ada04, ALM18, AY11, AK19, BZ17, BKY10, BD21, BLE97, BBS12, BO08, BH04, BISC14, BMS17, BDV06, BLZ08, BMM^{+}08, BVV12, BKM^{+}12, BDM^{+}14, BS10, BHL^{+}22, Cho03, CH21, CBE18, DY04, DFNY08, DFF^{+}21a, DMMR23, Don05, Don10, DKM^{+}22, DHR^{+}04, EZ96, Ema12, Fal06, FM18, FM15, FS21, GM17, GLOW04, GGLO08, GHT09, GKV12, GT09, Gra08, GHJV16, GH23, GMOS06, HBH10, HM18, HM20, He23, HNR^{+}18, Het07, Höm06, IV04, KXZ03, KR11, KS22, KR06, KLM15, Lee12, Lee16, LB21, Lee21a, LOS04, LCHH18, Liv04b, Liv14, Lot23, LJM14, LD07, LRGO17, MO11, MMC12, MO14, MMPR10, MWZ06]. multigrid [MBW97, MC08, Mit10, MSF21, MW21, NN11, NFD10, NSCTP05, NSCTPW22, Not05b, NV08b, OST10a, Pfl99, PT17, RS02, RV12, Reu96, RNV21, RBV08, RGM17, Sei10, Sha98, SY18b, SKR08, SSSF23, SSB15, TGKR10, TC10, TY10, TH19, UMO09, VZ08, VY14, Wan00, Web10b, Web10a, WZZ18, XSZ09, XZS15, YW12, Zhu14, ZMO10, vRH05, DM10, Den18]. multigrid-based [UMO09]. Multigrid-in-time [BW17b]. multigrid-reduction-in-time [FS21]. multigroup [KWS^{+}18]. Multilevel [AT15, CEL^{+}96, CV03, LSC21, MFFJ18, Osw95, PLMV23, QV21, Sta96, AM96, AMM04, AN94, AV94, BMN05, BCZ12, CL96, CGJ21, DMTY11, DGM^{+}16, FOV21, Kra02, Kra06, KT08, KMS08, KLM14, KP10, Lai97, LSS03, LM06, MM95, May07, Not98, Not02b, Not05b, Pad99, QvGvW^{+}21, SS02, Sha99, SLV13, The98, US19, XCG16, Yot01, vN00]. Multilevel-in-width [PLMV23]. multilinear [BP22, CRZT20, HW21, LLPC23, LPS16, LLNV17, MP18b, PDV05]. multiparameter [RMM22]. multiparameter-eigenvalue [RMM22]. multiphysics [Yot01]. multiple [ARSO14, ARMW14, CNZ17, HKLP19, Mai06, RNV21, SHJC18]. multiple-network [HKLP19]. multiplication [Kap99, OOO11, OOO16, WF15]. multiplicative [CL96, LSC21]. multiplicity [CC20]. multipliers [BGM^{+}21, ZN18]. multiprecision [BB16]. Multiprocessor [ADP96]. Multiscale [HPPS03, BIA18, FP15, VSG09, WWX10]. multisecant [rFS09]. multisensors [CNSY05]. Multisplitting [RLG12, AMP99, BZ13, CS09, CS11, JS96, LSL01, Mez20, Ren98]. multisplittings [BCC98, CP99, FP95b]. Multistage [OC22]. multistep [BWN05]. multivariate [HDIS18, LZQ12, MVK04]. multiwavelet [DOR21].

Nath [CLR13]. Navier

[AB12, CA99, DFF^{+}18, HFW01, KOV17, LMM00, Ols99, PT17, QvGvW^{+}21]. near [CNY05, Ver00]. near-circulant-block [CNY05]. near-singularity [Ver00]. nearby [AFS14]. nearest [CGS20, DBLP16, GMS18, GHR98, MRT98, NW15, ST23]. nearly [BKP02, HFW01, NA97, RSCTP15]. Nearness [BF19]. Necessary [Pul08]. negative [BMM06, CfX05, PR11, Yan23]. Nested [Bla03, GNQ15, MO16, MM18, vV94]. Nesterov [HY22, MYD20]. network [CHCS22, HKLP19, NR22, PLMV23]. networks [GB15, Lee18, WWC^{+}15]. Neumann [KMM19, RT99]. neural [CHCS22, PLMV23]. neutral [ZCW11]. neutron [Cha07, CGM11, KWS^{+}18]. Newton [ABBP10, AABHV18, AMMP06, ABK97, AFK02, BC09, BMM06, BMM20, CQ10, CWwS18, DL97, DEM18, DS13b, GB11, GKK04, GD11, HP04, KPV06, LB08, Lu05, LS22, LV98, MB21, NQ96, OC04, Sch99, Vla00, Yot01, ZZ15, Zho06, ZG22]. Newton-like [BMM06]. Newton-type [ABBP10, AABHV18, CWwS18, LS22, Vla00]. Nicolson [LP22]. NLA [Axe10, Vas05]. nnCANDELINC [ADMS22]. nodal [BDV06]. nodes [FP05]. noise [LCZZ21, MK23, NWZ17]. noisy [BC09, NWZ17]. Non [AMP99, BSC20, VW01, Bai16, BMM06, Bla02, BMN05, CL96, Cao04, Car97, CGM01, CPS01, CGL05, CK01, CfX05, DS02, EZ96, FP05, GB11, GM11, GVT03, HKKP07, HSCTP05, KPV06, KM99, Kra02, LVD02, LHL07b, Lu05, LMM00, LV98, LMV04, Mav01, MZ98, MC04, NQ96, OC04, RT99, SB12, Sei10, WD08, vN00, Bai18a]. non-conforming [BMN05, KM99]. non-convex [LMV04]. non-equispaced [FP05]. Non-Hermitian [BSC20, SB12, Bai16, CPS01, HSCTP05, LHL07b, WD08, Bai18a]. Non-linear [VW01, Bla02, CGL05, KPV06, Kra02, LV98, LMV04, Mav01, MC04, NQ96, OC04, RT99]. non-linearly [LVD02]. non-Lipschitzian [DS02]. non-negative [BMM06, CfX05]. non-overlapping [CGM01, GVT03, LMM00]. non-smooth [Car97]. Non-stationary [AMP99, LMM00]. non-symmetric [Bla02, CL96, Cao04, CK01, EZ96, GB11, GM11, HKKP07, Lu05, MZ98, Sei10, vN00]. nonaligned [YXZ13]. Nonconvex [YB23, HZZC23, Laz16]. Nonequivalence [FLPW01]. Nonlinear [Gra08, Vab23, AMMP06, AC11, BRT07, De 13, DGRR11, rFS09, GD11, HM16, IKAA22, LB21, LZ22, LWS^{+}23, MV13, MSV13, Naz95, yPES07, SGSM15, SCD94, USS21, VJM16, Vos09, WRW18, Xie21, XZS10, ZZ15]. Nonlinearly [DH18, DW15]. nonmatching [OZB^{+}18]. Nonnegative [ADMS22, ACGH21, BGX06, BGM09, BGM11, CQZ13, DDL^{+}21, HKL21, Sot13, Vab23, WWC^{+}15, ZQ12, ZQLX13, ZQW13, ZWQA18]. nonnormal [MYZ16]. Nonnormality [Baz08]. Nonoverlapping [CB21, BO18]. nonpositive [Hua12]. nonrestarted [Zho18]. Nonsingular [DIPR19]. nonsingularity [Peñ07]. nonsmooth [Che02, CQ10, Śmi19]. nonsquare [fLWyL^{+}21]. nonstationary [BDRZ21]. Nonsymmetric [CGK94, YW12, ARSO14, AHJ20, Bai95, BGX06, Ema12, GLJ19, HM14, IP13, Jou94, LW07, LB08, Mey94, MCLM20, Not10, SJBH14, SX15, Sta96, SL10, VEV23, Vas92, WTWG14]. nonzero [ZHJL12]. norm [CDG00, Dax94, DE98, DBG06, DHW16, EM95, EHM95, Gar02, Miy15, XJ12, YL08, Yan18]. Normal [Gus04b, SZ11, Bos19, FSS18, LS05]. normality [NR11]. norms [GZ16, SB03]. normwise [DW07, FT98]. notch [RS07]. Note [LZY11, CNT07, Cao09, CK14, DS10, DS08, DN12, FT98, GM11, GX14, JO01, KH07, Lai97, LXW13, LW07, LC07, Lot07, Ney05, SB03, Sun05, SHT11, VVM05c, Vöm10, Vöm12, Wan18a, WBL14]. notion [DGM^{+}16]. novel [DOR21, NPR13, SP06, BNR18]. NS [FM18]. NS-LAMG [FM18]. nuclear [XJ12]. null [How18, ITS07, RS18, WF15]. null-space [ITS07, RS18]. nullspace [Sim03]. nullspace-free [Sim03]. number [ADT19, BB06, BC10, EHM95, EG16, LH08, LLW09, RV12, TDL^{+}22, TGKR10, ZHJL12, ZLLH23]. numbers [BG05b, CCG00, CLTW11, CDW06, DW07, Dia09, DXW12, DWWQ13, Liv14, MDB21, YDH11]. Numer [SB12]. Numerical [AGG^{+}16, NLA94, Ano08, Ano09, BLP08, Ben11, CH03, CSZ21, CA99, DMS17, DSV18, FZwCW17, GS05, HHM10, HJR97, HL21, KKO20, fLyHZ11, LD08, MK94, MMMM09, MV05, NBKS99, NSCTP05, NL09, WW08a, JNL92, Bai95, BDRS12, BNR18, BKP02, BSC20, Bat95, BGM11, Ber01, BDS94, CQX11, CJW06, Cor04, CJT03, Dat01, DS02, GY08, HPS15, KK23a, LJ04, LH08, LHW11, LGS12, Lin12, MM09, MP13, MM18, OCYM08, Ols99, Özb13, Spi21, SHT11, Tur00, Mar00]. NURBS [GMSCS20].

Objective [Ris19]. Oblique [Han13, YCY17]. oblivious [MWZ06]. observations [CZ02]. observer [CLR01, CD11]. obstacle [JZ11, ZJ06]. occasion [CLR13, LPQ06, SGP14, Vas03, Vas05]. occur [CC03]. occurring [AG99]. oceanography [Rak99]. odd [Not05a]. ODE [AL21]. ODE-based [AL21]. Odir [CK01]. off [EW13]. off-diagonal [EW13]. once [DW21]. One [OC04, CSYS14, EGMS20, FMPS13, O’H14, Pul08]. One-level [OC04]. one-way [EGMS20]. open [Gar04, RR12]. OpenMG [BISC14]. operations [STZ12]. Operator [Gus97, Gus98, Gus03, MMPR10, Alb06, BV00, BCV03, BFM12, Den09, GN00, GH11, Liv04b, MP15, SKKS22, Tyr05, Vab20]. Operator-based [MMPR10]. operators [AFSCSU14, ABBP10, AEHV14, BKY10, DFF^{+}21a, Don10, DKM^{+}22, GGLO08, GVT03, Kho96, MC08, PSK08, Yan10]. optical [BCK05, KRW08]. Optimal [Bai09, BTT13, ELV94, FS21, GHO15, HLM^{+}18, LHLS07, Lot23, LD07, MM95, Not98, WKS95, BLP08, BL22, BFPS10, BMN05, CDDZ19, DH04, EG16, GTZ18, GR23, HFW01, HW18, KK13, Lai97, LP22, LZ23, MNCT07, MSS07, MP13, NV23, NA97, PSW14, RGG07, RSCTP15, SKKS22, SY18a]. optimality [NN10]. optimally [Cha07]. optimization [ADO23, AN03b, BD21, BDK^{+}15, BZ23, CWWZ22, Chu04, De 13, DD07, Gar02, GY08, HHM10, HP04, HZZC23, HW22, KCS11, Laz16, LZQ12, Lin12, LMV04, MV13, NBKS99, ORU23, PW12, PPS20, RS10, Ris19, SW12, TV20, WCZ15, WN18, YB23]. optimize [MC08]. optimized [EGMS20, OOO11]. Optimizing [DFF^{+}21a, TGKR10]. option [LLS12, Rag14]. order [ABBP10, AEHV14, AEHV15, AABHV18, ABK15, BCR11, BCR14, BK21, BBJ17, BNS20, BH16, BGM^{+}12, BSI17, CEL^{+}96, DOR19, DOR21, DMMR23, DLVZ06, ELV94, GM17, GA18, GTI16, GHW06, GKY97, GL13, HM20, Hem96, JM10, Kap02, KLM^{+}06, KPV06, KM09, Lam12, Lee19, LY15, LZ23, LLV19, Lun20, MV13, MMN^{+}10, MNCT07, MCLM20, RS01, SSB15, TSPSO06, UMO09, VHM^{+}22, WQ07, WQZ09, Web10b, Web10a, XSZ09, XS11, YB23, ZN20]. order-reducible [BCR14]. ordered [Bea94]. Ordering [HS05, HS15, Sco99]. orderings [DS10, NA97]. ordinary [BCR11, BCR14, Bot13, ZCW11]. oriented [TC10]. Orthogonal [FB95, VVM05a, AM95, BF96, DBG06, Kem12, MO94, MK23, PN18]. orthogonality [Par92]. Orthogonalization [Jia96, LBG13, LW23, LL97, SW96, VS17]. orthogonalizations [Dax04]. orthogonalizing [Mat96]. Orthotropic [GL96]. oscillators [MV19]. Oseen [CBE18, HBH10, KLM^{+}06, Ols99]. outer [Cor04, DDKR23, Xia12]. output [LW05]. outs [LPW06]. ovals [KVC12]. over-penalized [BPS13]. overall [BS01]. overlap [KK02, mMvdV02]. Overlapping [CS96, GNQ15, CGM01, Gan99, GVT03, JS96, KP00, LMM00, MO11]. overrelaxation [BGN07, Gus03, ORU23]. Owe [Cao13, Vas05].

p [SP06, HMS99]. p-level [SP06]. Padé [BLW08, GGZ12, LZ09]. PageRank [BP22, CRZT20, HW21, LLPC23, LLNV17, MP18b, WW07, YYN12]. pairs [CLC11, GMS18]. pairwise [FLR03, MS22]. palindromic [LYL15, MMMM09]. panel [PR96]. Papers [Ano08, LD08]. parabolic [AT15, DHS23, JM10, KK13, LSC21, vVW23]. Parallel [AO07, AMMP06, Bas00, BLE97, BGM^{+}12, BS10, CR16, FJP16, GL21, GR05, GL96, KR11, Lee16, LSL01, LGS12, MSF21, NO04, RT99, The98, Voe92, WH94, ZYFG11, AGG^{+}16, ACR^{+}00, AMMR17, AMP99, BPS00, BMS17, BMS18, BvdV00, CS09, CS11, CJT03, DFNY08, DFHM20, FJP12, FM99, GMR05, GSS01, GMOS06, GL98, GL02, GL13, Hac92, HS05, JO94, KK02, Kuz92, LVW01, LSS03, LWC16, MW16, MM97, MBW97, Mez20, MC04, MR14, Pad99, PR95, PR96, Rak99, RK18, Ren98, SL19, Sid97, TSPSO06, Van00, WLBH12, mMvdV02, mM04]. parallel-in-time [DFHM20]. parallelism [Vöm12]. parallelizable [GL95b]. Parameter [ZM20, AK99, BEG18, GNR14, GS05, He23, HMMP19, HKLP19, KPT14, LZ23, MSV13, Not02b, Xie21, Yan18]. parameter-dependent [BEG18, GS05, KPT14, Xie21]. parameter-free [Not02b]. parameter-robust [He23, HKLP19]. parameterized [CCvG06, DHR20, HW18, fLWyL^{+}21, RMM19, TS20, VEV23]. parameters [Bai09, BNP15, GHO15, HW18, Mai06, dCSRS19, Yan04, Yan18, ZHJL12]. parametric [SH19]. parametrization [Hua12]. Parareal [DFF^{+}21a, FS21, GL21].

Parareal-like [GL21]. ParaStieltjes [GL21]. Parlett [Bun95, EM95]. pARMS [LSS03]. Part [GL98, GL02, GL13]. Partial [LW04, LW05, BSC20, BD21, BGP97, CQX11, GBB22, KKO20, LH08, LHW11, LW03, MW11, MRK22, MM11, Not02a, Rak99, RBV08, SW12, TC10, YBZ19, Zhu14, vNR07]. partially [DD07, WQZ09]. particle [Sei10]. particular [ESC20, dF20]. partition [BDV06]. partitioned [AB10, AB13, Cao13, Poi00]. partitioning [CJT03, ED22]. partitionings [GKY97]. parts [Bai18a]. passage [KNX01]. past [Axe10]. path [Uhl23]. pathology [PM11]. pattern [CDG00, ISZ09]. PDE [BDM^{+}14, GHW06, Lin12, OZB^{+}18, PW12, PPS20, RS10]. PDE-based [GHW06]. PDE-constrained [PPS20, Lin12, PW12, RS10]. PDEs [AT15, AMMR17, CGJ21, Hem96, Höm06, LP22, MO11, VSG09, VZ08, ZMO10]. Peaceman [LR95]. PEERS [KS04]. penalized [BPS13, Dos99]. penalties [MG08]. penalty [BCS09, BPS13, BDR17, DH04, Lai97, PSW14]. pencil [LW05]. pencils [BB01, fLWyL^{+}21]. Penrose [DW07, DWWQ13, KKMM12, LXW13]. pentadiagonal [TS20]. Performance [BT15, Sei10, mM04, Alb06, BE98, MO14, MSF21, SST18]. periodic [CX22, KK13, Var08, WZZ18]. periodicity [BDS94]. permanents [WLBH12]. permittivity [PR11]. permutation [May07]. Perron [Dem21, ES09b, KNX01, LCN13, MP18b, NX03]. Perron-based [MP18b]. perspective [BMS17, OST10a]. persymmetric [XHZ03]. Perturbation [Cas11, CLC11, GCLG18, GW00, WW08b, YL08, ZY19, CTP09, Cha12, CLTW11, FT98, JLW05, LS05, LS06, LCN13, LW15, MS22, O’H14, WKS95, WW20, WL03, Xie21, YDH11]. perturbation-based [Xie21]. perturbations [AIT05a, AIT05b, BGS21, BSC20, LXW13, NR19]. perturbed [Sau95]. Petrov [CGM11]. phase [DY04, HS13, HHLL16, NH06, SY18a]. phylogenetic [BL03]. physics [TC10]. physics-oriented [TC10]. Physiology [PM11]. Piecewise [HM96, Bos19]. piezoelectric [CN21]. pinch [LPW06]. pinch-outs [LPW06]. pipes [HG00]. pivoted [HC05]. pivoting [BM05b, BM06, EM11, LSS18, May05, May07]. placement [Dod11, He21]. planar [GLGR10]. Planck [ZZ21]. plane [BLE97, Ypm95]. planewise [mMP99]. planewise-like [mMP99]. plants [Özb13]. plasticity [ABK97, Car97, HJR97, Wie99]. plate [AY11, CYZ99]. player [AD12]. Plemmons [NN15]. plus [BLP17, DPRV19, Fas05, HN05, KN07, MCV01]. point [AN06, Axe15, Bai09, Bai12, BMM06, Ber12, BGM^{+}21, BG05a, Bir15, Cao04, Cao08, Cao09, CJZ11, CH03, CGJ21, CRZT20, DLSvL20, DL23, EG16, HP04, HD07, HDH19, HW21, KP00, KO18, KKR14, Krz11, KKMM12, LLPC23, LOY08, LOS04, LW07, LMV04, LSS18, MZ15, PPS20, PW13, RS18, SJBH14, SX15, TH19, VEV23, VL96, Wan18a, Web18, WBL14, Zha18, MST16]. point-proximal [BGM^{+}21]. point-type [Cao08]. points [HM96]. Poisson [CKW02, CJL08, Dah02, GH23, RSR10, TSPSO06]. polar [CCG00, LS06, RT02, YL08]. Pole [Dod11, LC13, LW04, LW05]. poles [Mee01]. policy [BLLA11]. pollution [LC21]. polyadic [BVD^{+}18]. polyhedral [Dah02]. polynomial [CCS19, CR16, CZS22, Gan05, GKV12, HM96, HS08, HVX16, Lee16, LW98, LM22, Lot23, WCZ15]. polynomials [AK23, BB97, BGW05, BG05a, HDIS18, KR14, MO94, MN05, Nov03]. population [DHSW11]. poroelasticity [GLOW04, HKLP19, LRGO17]. porous [NH06, ŠBS15, WWX10, Yot01]. posed [AGRR21, BDRZ21, CLTW11, DNR12, DHNR18, Est09, GORR16, HDA19, NR14b, RU22]. positive [ARMW14, AIT05a, AV94, Bai16, Bai18a, BP13, BMAA16, BT03, BMM20, BN21, CS09, CS11, DJW^{+}21, DPP16, DJ09, Kap98, Kol05, LHL07b, MVV08, PS11, yPES07, PW13, SB12, WW08b]. positive-definite [DJW^{+}21, DJ09, Kol05, LHL07b, MVV08, SB12]. positivity [KSB13, MPR23]. possible [VL11]. Post [KLN99]. Post-processing [KLN99]. posterior [FOV21]. posteriori [AM96, BLP01, CS18, OOO16, Pul09, Ney02]. potential [Kho96, MRT96, Shi02, Shi04]. potential-reduction [Shi04]. potentials [KK16]. £1000 [Ano08]. Power [AS19, BG23, CEQN07, CH05, DS13b, GGV13, JZ09, Lee18, LP16, Vab20, WW07]. powerflow [LB21]. powers [HLM^{+}18]. practical [DGB^{+}13, Kap99, WQZ09, WM12]. Prandtl [Wie99]. Prandtl-Reuss [Wie99]. precision [BHL^{+}22, CGY22, GSTPT21, OC22, SWW21]. Preconditioned [Axe98, CGK94, CL23, DGM^{+}16, GKK19, HMS99, HES15, WCW20, AN06, BM13, BL22, Ber12, Ber01, BWN05, BB06, Bla02, BHHJ13, BDRZ21, BE98, CZ02, Cao09, DJW^{+}21, Dam08, DW15, DH18, DS08, DR03, GLJ19, IIFM23, KK13, KPT14, LD07, LV98, PR95, PR96, PL21, RV12, SJBH14, SHZ20, TDL^{+}22, WBL14, ZZ21, ZBCN23]. Preconditioner [TT10, BPS15, BT03, Beu03, BC12, BPS13, CGPV13, CJZ11, CNP96, CJW06, CS95, CV13, Doh07, ES07, EGF11, GN00, GTZ18, HFW01, ISZ09, KS04, KLS23, KWS^{+}18, KV96, Kuz92, KP10, LS04, LSC21, May05, May07, MC09, NL16, OZ22, SPD05, SP06, SLV13, SGP14, UMO09, VEV23, Xia12, XS11, XM17, Zha18, ZXS20, Zhu14, vN00]. Preconditioners [BEV22, CPS01, Est09, GS07, NV23, PSW14, AY11, AN13, Bai16, BM17, BDdSM18, BMM20, Bla02, BMN05, BCHT04, BIA18, BSI17, Cao08, CDG00, CDGmM04, CGM01, CC92, CW97, CEL^{+}96, CDDZ19, DDG99, DP03, FP15, FK15, FS09, GMTV16, GNQ15, GR23, HLM92, HH06, Hem96, HKLP19, HK12, JLW05, KABH17, KY95, KKNY01, KK23b, KP00, Krz11, LVW01, LOY08, Lee16, LJ04, LXS16, LC05, LW07, LWC16, Mar16, MSS07, NV08a, NR12, Osw95, PW12, PS00, QB15, RS10, RSCTP15, RVW98, SZ99, ST17b, The98, TT15, Tyr92, Tyr05, WRW18, XG10, YNP04, Yan18, ZCW11, ZHJL12, Zhu08, mMP99]. Preconditioning [ABM17, AN03b, AB10, ABNP15, ABK15, CFAM16, Egg07, Gro00, HW19, HSCTP05, MW11, Pul09, QvGvW^{+}21, SMSW00, SW12, Vas92, VL96, WDS09, WBWM04, Ye20, AFSCSU14, AT15, AL21, AK94, AV94, AFK02, Axe15, BCR11, BCR14, BD21, BK21, Bas00, BGM09, BPS00, BGM^{+}21, Bla94, CN21, CDDSC12, De 13, DLVZ06, DD07, Dos99, DKVB15, FBSC21, FJP12, FJP16, GM11, Gus03, GL95b, HPPS03, JZ09, JK17, Kap94, Kap98, KK02, Kap02, KM99, KPV08, KOV17, Kra02, Kra06, KMS08, LV04, LM22, LW03, MFFJ18, MM95, MM02, NO04, NR11, NA97, Not98, Not02b, NCV05, PPS20, PW13, Poi00, QPS23, SP18, SL10, TSMM21, Vas02, VHM^{+}22, WH94, XXCB20, AB13, Cao13]. preconditionings [GKY97, KNY99, NY03]. prediction [BS10, PGT14]. predictive [FM15]. predictor [BB97, HM14]. Preface [Axe02, AK10, Cve09, Dat01, NT04]. Prefiltration [NY03]. Preordering [LSS18]. presentation [EJK01]. preserving [HHLL16, LW23, PR16, Wan00, WRW18]. Press [Nab97, Amb15]. pressure [Lay05, LWC16, vKVW00]. Prestructuring [How18]. Price [Nab97]. pricing [LLS12, Rag14]. Primal [HP04, RT02, FLP00, WSN19]. Primal-dual [HP04, WSN19]. primitive [Dem21]. principal [GH06, HW22, LB17, LC21, PY22]. principle [BC02, Vos09]. principles [Gar04]. priori [HM96]. PRISM [Axe98]. Prize [Ano08]. probabilistic [WWC^{+}15]. probabilities [NX03]. probability [BH16, LCN13, LMM^{+}23, MM98]. probing [TS12]. problem [AS19, AH02, AK99, ABK15, Bai95, BDK^{+}15, BFPS10, CCS19, CZ15, Car97, CPSM06, CGL05, CG15, CFAM16, CMSW19, CZS22, CH21, CJT03, DL97, DHR20, DMS17, DWWQ13, Dod11, DBLP16, DDL^{+}21, ES07, ES09a, ER96, GKK04, GT19, Gus98, HBH10, Hla99, HS08, HC20, IV04, KABH17, KPV06, KH07, KMM19, KNP03, LLLJ16, fLWyL^{+}21, pLL07, LYL15, LD07, MV13, MRT96, MLV05, Mee01, MP15, MDB21, Ols99, OC04, yPyHZ04, Ren98, RSR10, Rja98, RT99, ST23, Sau95, SH14, Sim03, Sot13, TDL^{+}22, TW20, VFdV13, Vla00, WKS95, Xie21, XZS10, YHS18, ZJ06, ZYFG11, ZYL13, ZVO14]. Problems [CGK94, GL96, Ada04, AB00, AW11, AGRR21, AIT05b, AHJ20, AG99, AV94, Axe98, AN03b, BBP03, Bai09, Bai10, Bai12, BZ13, BZ17, BW19, BKY10, BKP02, Bar02, BLE97, BBS12, BMM06, BGM09, BGM11, BLP08, BCV03, Bla94, BC02, BBG13, BMS18, BvdV00, BRT07, Bör17, BO13, BDM^{+}14, BDRZ21, BIA18, CL96, CNT07, CQX11, CGPV13, CRS05, CR16, CR20, CEQN07, Cao04, CJZ11, CCvG06, CC92, CNP96, CW97, CS02, CTP09, CEL^{+}96, CCK06, CWwS18, CGJ21, CX22, CWS97, CC03, CLTW11, CP12, CBE18, CV13, CRV14, CK14, DHS23, Dax94, DE98, DW07, Dia09, DNR12, DJ09, DGP19, DHR^{+}04, DP03, DR03, DHNR18, Egg07, EAA19, EGF11, ELV94, EWY03, FY01, FGT11, Gar04, GGLO08, GH01, GORR16, GHT09, GVT03, GGZ12, GTZ18, GKK19, GMTV16, GL98, GL02, GL13]. problems [HJ18, HP97, HKST12, HJR97, Han13, HDA19, He23, HW19, HMMP19, HS13, HL16, HY22, HD07, HST22, HLLL13, HM16, HDH19, HLLW05, JZ11, JK18, JM10, KKPS18, KH23, KMM18, KK02, KR11, KP00, KK13, KR06, KT08, KMS08, KLM14, Krz11, KM92, LLL97, LR95, Lay05, LPV01, LV99, Lee21a, LZ23, LW07, Lin12, LZ12, LW16, LW17, Liv04b, LL97, LS22, LV98, MZ15, MB21, Ma22, MMMM09, MS07, Mar00, Mar98, MRT02, Mar16, MSS07, Mav01, MSV13, MP13, MM97, MBW97, Mez20, MM02, MSB18, MZ98, NV23, NR14a, NR14b, Nov03, OS10, Pad99, PBN05, PSW14, PPS20, Pen08, PH19, PL21, RU22, RR12, RNV21, ROA13, RMM22, SLK16, SCP20, SKKS22, SX15, Shi02, Shi04, SY18a, SV11, Sta96, Sto92, TDH^{+}18, Tre13, TT15, VEV23, VJM16]. problems [VL96, Ver00, Wan00, Wan18a, Web18, WWC^{+}15, XG10, XZS15, XXW19, wX15, XCG16, XX22, YBZ19, YCY17, ZZ15, ZN18, Zha18, ZHZ10, ZY19, ZSCX10, mMP99, mM04, VW01]. Procedure [IDVV96, GL21, JZ09, JK17, LR95]. process [PRR^{+}16]. processes [AD11, BMMR18, BL03, Buc11, DGB^{+}13, GCLG18, NH06]. processing [Dat01, KLN99, SKR08]. Procrustes [CZ15, KH07, XCG16]. producing [SH19]. product [Aih20, BW17a, BSMN22, Che15, DQW15, DK15, FZwCW17, Gus04a, HXM19, HL21, KN07, LS04, MGF^{+}02, Per06, RU22, XG10, ZSKA18]. products [BB01, DWWQ13, LPS16, Mat96, Mey94]. Professor [SGP14]. profile [HR05]. program [CCLQ18]. programming [BDdSM18, BGM^{+}21, BRT07, HHQ13, LV98, Naz95, ODH21, RGG07, Shi02, Shi04]. Progress [Bai95]. project [TYZL23]. Projected [HKKP07, BN21, KO18, Shi04]. projected-steepest-descent [Shi04]. projection [BG13, Baz08, BG00, FB95, GKL18, HCD15, ITS07, LWS^{+}23, MZ98, RT02, RMM19, YCY17]. projection-based [GKL18]. projection-type [Baz08]. projections [Dax04, Han13, LCHH18, VZ14, WTWG14]. Projector [DD07]. prolongator [KV15]. prolongators [BDV06]. proof [Adi08]. propagation [BO13, mM04]. proper [Kem12]. Properties [CLQY23, PSK08, Wei94, Yon96, ZHZ10, BDS94, Bun92, CGK05, CJW06, CDDZ19, LV12, MDMS23, NPR13]. property [DMY03, EZ96, ES09b, NL16, YLH11]. proposal [NCV05]. proposals [NSCTPW22]. proving [BBP03]. proximal [BGM^{+}21]. pseudo [BFdP13, ZLLH23, mMvdV02]. pseudo-Jacobi [BFdP13]. pseudo-overlap [mMvdV02]. Pseudoeigenvector [MYZ16]. Pseudospectra [KCC16, VW15, NR17, Sir19]. PSF [BNP15]. published [Ano09]. pure [KM99]. purely [BF11a]. Python [BISC14].

QBD [BMMR18, GCLG18]. QLP [HC05]. QR [CGK05, CX23b, Fas05, LW15, VVM05b]. QTT [VR23]. quadratic [BLP08, BDdSM18, BGM^{+}21, BG05a, BMP11, BMMR18, CQX11, CR20, CCvG06, CMSW19, DMS17, DD07, DR03, EGF11, GA18, HLLL13, KLM14, LC13, LW05, LZQ12, LYL15, MP13, ODH21, QXB09, Ste99, XZS15]. quadratic-bilinear [GA18]. quadratics [GSTPT21]. quadrature [GL21]. quality [BC10, Kap98, NY03]. quantification [Lee21b, SCP20]. quantity [FOV21]. Quantized [KKS19]. Quantum [CVY21, KMMR10]. Quasi [RSCTP15, BMM20, DEM18, Gar01, Gar02, HMS99, LY15, MN05, SW96, YM22, ZZ15, Bai18a]. Quasi-HSS [Bai18a]. quasi-isometric [Gar01, Gar02]. quasi-kernel [MN05]. quasi-minimal [SW96]. quasi-Newton [BMM20, DEM18, ZZ15]. Quasi-optimal [RSCTP15]. quasi-rational [YM22]. quasi-uniform [HMS99]. quasiseparable [BEG18]. quaternion [JNS19, LW23]. question [JK09]. queueing [BLLA11]. quotient [CX22, CX23a, CX23b, CHCS22, FK15, Het07, NZ14, PS95, Zho06]. quotient-gradient [CHCS22].

R [Nab97]. Rachford [LR95]. radial [CDDSC12, ZLLH23]. radial-lines [ZLLH23]. radiation [OC04, WBWM04, XM17]. radii [CfX05, ZWQA18]. Radim [Cao13]. radiosity [Leb02]. radix [MR14]. radix- [MR14]. random [HPS15, LW98, LCHH18, OZB^{+}18, WF15]. Randomized [SLK16, SHvBW21, XXW19, YMS^{+}23, BW19, CWWZ22, Du19, LW21, WX21, ZSKA18]. randomly [KK23b]. range [AMMR17, CJW06, KK16, MM18, Yan10, ZW10]. range-Hermitian [ZW10]. Rank [GS97, Kub92, AT15, BE09, Bau08, BF96, BHL^{+}22, CCX23, CH94, CWWZ22, CSYS14, DE98, DW15, DPRV19, DBLP16, ESS23, ES05, ED22, FMPS13, Gra08, HCGM23, HS18, HR05, HKL21, HC05, JMPR18, KKRS21, KO18, KJ12, KPT14, KS15, Laz16, LXS16, LW21, LO15, MRK22, NL16, O’H14, ORU23, QXB09, SPD05, SP06, SLV04, SLV06, Tia13, Tyr92, VVM05a, VJM16, VR23, WQ07, WN18, WLC21, YZCQ23, ZXS20, ZG22]. rank- [DW15]. rank-1 [HKL21, KJ12, WQ07]. Rank-deficient [GS97, DE98]. rank-exploiting [VJM16]. rank-one [CSYS14, O’H14]. rank-revealing [HCGM23]. ranks [LT08, STZ12]. Rapid [LO13]. rarely [BG05b]. rate [BS01, CJT03, KKO20, MRT96, RV12, Zik08]. rates [Li00]. Rational [Fas05, Mor09, Rag14, BBJ17, DHS23, DGP19, HK21, Mee01, Mor07, PRR^{+}16, Tre05, XX22, YM22]. Raviart [KMS08, LV12, Zhu14]. ray [Liv04b]. Rayleigh [CX22, CX23a, CX23b, CHCS22, FK15, Het07, HS08, KABH17, NZ14, PS95, Zho06]. Raytcho [Vas03]. RBFs [FP15]. RD [Mor07]. RD-rational [Mor07]. reaction [DOR21, Gan99, TC10]. reaction-diffusion [DOR21]. Real [AK00, YPC20, BF19, Bra02, CHV05, GHR98, MZHB17, MSV13, MV19, Sot13, vNR07]. real-equivalent [MZHB17]. Real-time [YPC20]. realizability [Sot13]. realizable [CfX05]. realization [Baz08, PR96]. reciprocals [Vöm10]. reconstruction [CNSY05, PN18]. Recovering [MK23]. recovery [AGG^{+}16, BZ23, DDM23]. rectangular [BS01, Le 23, LS06, Osw95, Pul09]. Recursive [FLM09, HSY18, NV08b, LSS03, Not05a, NA97, SS02]. Recycling [OZ22, RLG12, SGSM15]. red [NA97]. red-black [NA97]. reduced [ES05, GH11, KN14, Sir19, VW15]. reduced-rank [ES05]. reducible [BCR14, ZWQA18]. Reducing [GHJV16, VY14, Zha92]. reduction [AGRR21, AK94, BBJ17, BPS95, BTT13, DFF^{+}21a, FS21, GORR16, GA18, HNR^{+}18, KCS11, Lay05, LO13, Lee18, LW21, MMM06, MV13, MR14, PV99, PY22, Shi02, Shi04, SSSF23, VP95, YZ13, ZS08, vGSZ15]. reduction-based [MMM06]. reduction-in-time [DFF^{+}21a]. reductions [KNX01]. Reeves [YBZ19]. refined [BB00, HS08, KR11]. Refinement [GL95a, BS01, BGM^{+}12, CR16, DMM^{+}08, DHR^{+}04, ELV94, MMN^{+}10, MM95, Mit10, OC22, WW11]. Refining [Peñ07]. reflective [Per06]. regenerative [AD11]. region [HS18, fLWyL^{+}21]. Regions [PS95, Naz95]. registration [GHW06, HHM10, Höm06, RGM17]. Regression [TSMM21, ES05, PY22, PLMV23]. regular [CLC11, FG02, FT98]. regularity [Dah02]. Regularization [BGM09, DHNR18, IDVV96, BCB14, BDR17, CRS05, CLTW11, Don05, DNR12, FRR16, GNR14, LHW11, Spi21, WLC21]. regularized [BL20, ES07, ES09a, FGT11, MLV05, RLG12]. regularizer [KRW08].

Regularizing [CDDZ19]. Reissner [CYZ99]. related [AK94, DSV18, DP23, DKM^{+}22, ESC18, GGZ12, Li00, MPR22, Mor09]. relations [Tia13]. relationships [JYZ17, Tre05]. relative [DOP21, DP23, YM22]. Relaxation [BKM^{+}12, LLV19, DDKR23, Dax94, FHM21, FP95b, Gan99, HM18, He23, LZQ12, Liv04a, PBN05, SX15, SSSF23, Yan04]. Relaxation-corrected [BKM^{+}12]. Reliable [Ber01, Hla99]. remarks [LS06, Mar95]. removal [LCZZ21]. Reorthogonalization [DKVB15, Van00, Zou23]. Reorthogonalization-based [DKVB15]. Repairing [Ver00]. repeated [AT00]. repetitive [DGB^{+}13]. representation [CC07, DEM18, VVM05c, Vöm12]. representations [RMM22, VR23]. representative [KKO20]. representing [MO16]. reservoir [LVW01, LWC16]. residual [AC23, AM95, GH01, Gus03, HMS99, JR94, JK17, Kap05, LWC16, MO94, MRT96, SW96, Saa00b, Sta96, ŚLA^{+}21]. resilient [AGG^{+}16]. resolution [CNSY05, JK09, TR21]. resonant [AG99]. Respectively [Bai18b]. Response [AB13]. restart [KLMP21, MYD20, MN00]. restarted [Dax19, Jou94, MP14, Sim99, VL11, ZM08, Zho18, Zít00, Zít05]. restarting [BD15, SHJC18]. restoration [BC02, CNXY20, Per06]. restoring [NWZ17]. Restricted [BK11]. result [FP95b]. resultant [BGW05]. results [BF19, BNS20, BMSS09, DFF^{+}18, Kap94, MMN^{+}10, MM18, NH98]. retinex [YHS18]. retraction [Kau07]. retrieval [BF96, FJ05]. Reuss [Wie99]. revealing [CH94, DCT18, HCGM23]. reversible [NW15]. Reversing [RS01]. Review [DHBV21, Nab97, MO14]. revisited [CVY21, XXCB20]. Revisiting [AD11]. reward [Buc11]. rewards [Par92]. RIC [Not94]. Riccati [AHJ20, BGX06, BLP08, GB11, GL95a, Gra08, HM14, HHLL16, IP13, LB08, LS15, Lu05, Miy17, Var08]. Richard [LPQ06]. Richardson [Pas19]. Ridge [TSMM21]. Riemannian [BZ23, FJ05, HS18, HW22, MB21, YBZ19, YHAG20]. Riesz [DKM^{+}22, MDMS23]. Right [SHZ20, ARSO14, ARMW14, Lin12, SHJC18]. right-hand [ARSO14, ARMW14, SHJC18]. rigorous [LW15]. ring [MK20]. rising [KNY99]. risk [ADO23]. Ritz [GR99, HS08, Vöm10]. RLSL [BLP01]. Robert [NN15]. Robust [AY11, BMN05, JNS19, KSB13, KW99, KLM14, MMC12, Not02b, SNZ20, ZN22, AMM04, BT03, CDG00, CGJ21, GTZ18, He23, HKLP19, KKNY01, Lee10, NV23, SZ99, Xia12, XS11, vN00]. robustness [NR22, ST17b, XXCB20]. root [AEHV15, Dem21, LZ09, LH17, Mor09, PRPI09]. root-finding [PRPI09]. roots [CC20, MO94]. rotated [CG15]. rotating [SL19]. rotation [ST23]. rotations [MCLM20, Ypm95]. rounded [BH07]. roundoff [WW11]. row [Dax94, May07, RS01, SLV06, Sco99, WX21, ZHZ10]. row-by-row [RS01]. rows [DS10, How18]. RSCG [FO95]. rules [GL21]. Runge [Che15, FS21]. Rybicki [Amb15].

SA [BMM^{+}08, GX14, HVX16]. SA-AMG [HVX16]. saddle [AN06, Axe15, Bai09, Bai12, Ber12, Cao04, Cao08, Cao09, CJZ11, CH03, CGJ21, DLSvL20, DL23, EG16, HD07, HDH19, KP00, KKR14, Krz11, KKMM12, LOY08, LOS04, LW07, LSS18, MZ15, PW13, RS18, SJBH14, SX15, VEV23, VL96, Wan18a, Web18, WBL14, Zha18]. saddle-point [Bai09, Bai12, CGJ21, EG16, HDH19, KKR14, KKMM12, LOY08, LSS18, VL96, Wan18a, Web18]. Said [MPR23]. same [GHR98]. sample [DXW12]. sampled [Rie09]. sampler [OZB^{+}18]. Sampling [US19, AFSCSU14, FGT11, IIFM23]. SANs [LS04]. SAXPY [Ypm95]. Scalable [DH04, OZB^{+}18, FOV21, FLP00, Liv14, MW16]. Scale [VW01, Axe98, BBJ17, Bar02, BCB14, Ben08, BLP08, BES14, BHL^{+}22, CGJ21, DMY03, DGP19, GLJ19, Gra08, GR04, HJ18, Lee16, NH06, Sir19, XM17]. Scaled [Yan18, Bai18b, CTP09]. scaling [BBKY06, CZS22, GHO15, HS15, USS21]. scattering [FGT11, MV13, WDS09]. Scheme [Zha92, BS01, BMS17, BMS18, CRV14, GB11, GSS01, GMOS06, HY22, KV15, LLS12, Poi00, Pul16, RR12, ZZ21]. schemes [AIT05b, AJ94, Bir15, DE06, Gus03, HM18, HM14, KABH17, OCYM08]. Schmidt [Dax04, LBG13, LL97, ŚLA^{+}21, Van00, WL08, Zou23]. Schoenmakers [DPP16]. Schrödinger [CJL08, WRW18]. Schur [BG00, DHBV21, BCK05, BG05a, Bra02, BCGM09, BD15, Bun92, CN21, HKKP07, KSB13, KW99, Kra06, KLM15, LXS16, LW03, MMMM09, MW16, NG15, PW12, Rak99, SGP14, TSPSO06, WW08b, WTWG14, vNR07]. Schwarz [AB13, Cao13, AALS01, AB10, AG19, BK11, CZ02, DS08, EGMS20, KP00, KWS^{+}18, LSC21, OC04, SWW21, VSG09, XZS10]. science [KK23a]. scientific [Axe98, KK23a]. searches [DMY03]. Second [JM10, VFdV13, AK19, BBJ17, CEL^{+}96, DLVZ06, GTI16, KPV06, Lee19, LZ23, LM06, MM09]. second-generation [LM06]. Second-order [JM10, BBJ17, DLVZ06, GTI16, Lee19, LZ23]. sector [LZ09]. seed [ARMW14]. segmentation [LNP12]. Segrè [SVV22]. Seidel [Du19, HP97, KLN99, LO13, Sun06]. select [Alb06]. Selectable [YMS^{+}23]. selected [BTT13]. selection [AO07, CDG00, MO16, MYD20]. selective [NO04]. self [Leb02, MWZ06, MM11]. self-adaptive [MWZ06]. self-adjoint [MM11]. Selfadjoint [AV94]. Semi [Mar98, TV20, CH05, Ema12, GLGR10, KH07, LJM14, LD07, MCV01, MC08, Par92, WW08b, Xia12]. Semi-active [TV20]. semi-algebraic [MC08]. Semi-coarsening [Mar98]. semi-definite [Ema12, KH07, WW08b]. semi-iterative [CH05, LJM14]. semi-monotonic [LD07]. semi-orthogonality [Par92]. semi-separable [MCV01, Xia12]. semi-structured [GLGR10]. semicoarsened [RNV21]. semiconductor [GMR05]. Semiconvergence [CS11, WX21]. Semidefinite [LZQ12, CS09, CS11, CCLQ18, HHQ13, PS11, TR21]. semidiscrete [GB15]. semilinear [ZZ21]. Semilocal [GD11]. semiorthogonal [HLLL13]. semiseparability [VVM05a]. semiseparable [Fas05, QvGvW^{+}21, VVM05a, VVM05b, VVM05c, XCGL10]. sensing [BT15, ZZ15]. sensitive [LC21]. Sensitivity [CL13, GL95a, PV99, BW17b, NR19, SHvBW21]. separable [MCV01, Xia12]. separation [CCE^{+}18]. sequence [IIFM23]. sequences [AFSCSU14, BSC20, BDdSM18, Not05a, TT10]. sequential [ACR^{+}00, HCD15, HHQ13]. sequentially [QvGvW^{+}21]. serendipity [HH06]. series [GZ16, ROA13]. Set [YMS^{+}23, BDK^{+}15, MO16]. sets [KMC16, LLK14]. Several [Wu15, DHNR18, Pul16]. Shader [Nab97]. shallow [SL19]. Shamanskii [LB08]. Shanno [USS21]. shape [HP04]. shaped [GH11, ZMO10]. shapes [AG95]. shared [JO94]. sheet [AMR18]. shell [MBW97, The98]. Sherman [HS21a]. Shift [PS11, BBP03, IP13, MC09, MP14, VEV23, WtFW15, vGSZ15, Sim03]. shift-and-invert [MP14, WtFW15, Sim03]. Shift-invert [PS11]. shift-splitting [VEV23]. Shifted [LP16, CV13, JR94, JYH17, SLV13, TT15, UMO09]. shifts [CX23a, CX23b, SHJC18]. Short [Lai97, SHT11, Yon96]. shrinkage [HW22]. shrinkage-thresholding [HW22]. side [TYF23]. sided [CX22, CX23a, CX23b, DJW^{+}21, FK15, JZ11, PL21, WW20, ZJ06, Zik08]. sides [ARSO14, ARMW14, SHJC18]. Sign [Nab97, CLC11, GM17, GGZ12, SST18]. sign-indefinite [GM17]. Sign-Solvable [Nab97]. signal [Dat01, HM03, ZG22]. signless [XC13]. Signorini [Hla99, IV04]. similarity [VVM05a]. similarly [Tre05]. Simple [CGJ21, LV04, KNY99]. simpler [JYH17, LSJ18, WZ94]. simplified [BM06, ZVO14]. simulating [MC04]. simulation [BFPS10, BvdV00, BO13, PR11]. simulations [AK16, KR11, LWC16, NO04, YPC20]. simulator [LVW01]. Simultaneous [DK15, Peñ03, AT15, GM11, LT11]. sinc [BCR11, BCR14, NSCTP05]. Sine [CW97, QPS23]. single [BHL^{+}22, PDV05]. single-channel [PDV05]. singly [HS05]. Singular [AFSCSU14, BCC98, CKW02, Cao08, CWWZ22, CL13, CSZ21, Dod11, EN17, FP95a, FH94, GTI16, HS11, HS14, HS21b, HHLL16, JLW05, JK18, KR06, Krz11, KKMM12, KMM19, LSL01, LHLS07, LHW11, LT13, MB21, Ma22, MPS96, NR14b, PH19, Roh92, SHvBW21, Sau95, SS97, SNZ20, SHZ20, Szy94, THC09, Tre05, ZW10, ZSKA18]. Singular-value [AFSCSU14]. singularities [BLZ08, CKW02, Dah02, LLW09]. singularity [BG23, Ver00]. sixtieth [LPQ06]. size [BMMR18, FJP12, KKO20, MYD20]. skeleton [Zou23]. Sketch [TYZL23]. Sketch-and-project [TYZL23]. skew [BGN07, KKR14, LHL07b, SB12, Wu15, ZM20, Bai18a]. skew-Hermitian [SB12, BGN07, KKR14, LHL07b, Wu15, ZM20, Bai18a]. slicing [CR20]. sliding [AMR18]. small [DXW12, KV96]. smallest [MVV08, MM11]. SMASH [CCE^{+}18]. Smith [BES14]. smooth [Car97, HKKP07, The98]. Smoothed [BDM^{+}14, CDW06, HST22, OS10, Sch12, BMM^{+}08, BVV12, GHT09, KWS^{+}18]. smoother [GH23, LRGO17, ZVO14]. smoothers [AK19, BO18, CB21, GGLO08, GKV12, HBH10, Lot23, LJM14, MO11, SWW21, Yan04]. smoothing [BC09, EZ96, GLOW04, GH23, HP97, TC10]. smoothness [Cho03]. SNAP [ITS07]. Sobolev [AFK02]. social [GB15]. software [Voe92]. solid [Ada04, SV11]. Solution [Bar02, BFPS10, Ben11, JL09, ST23, ACR^{+}00, AD11, Axe98, Axe99, BDGL09, Bai95, BKP02, Bau08, BMM06, BLP08, BS01, BPS00, BMP11, BEG18, BRT07, BDS94, Bot13, BVD^{+}18, BFM12, CGPV13, CLR01, Che15, CSZ21, CA99, Cor04, DMS17, DSV18, DW21, DO18, DFHM20, DBLP16, FZwCW17, FJP12, Gem00, GTY97, Gra08, GT19, GS05, GL98, GL02, GL13, HJR97, HG00, Hla99, HC20, ITS07, JZ11, JO94, KK23b, KRW08, LX08, LPV01, LV99, LGS12, Lin12, LL97, Lot07, MS14, MZHB17, MP13, MM97, MBW97, Miy15, MSB18, Ols99, yPES07, PH19, RU22, Ren98, SGSM15, Sim03, Spi21, Ste95, TSPSO06, WWC^{+}15, ZN18, ZYL13, VW01]. Solutions [GL95a, Pen08, AW11, AHJ20, BGX06, CH03, DE98, DBG06, ESS23, HM96, KR06, fLyHZ11, pLL07, Miy17, PPv95, SH19, Tia13, Ye20, ZLLH23]. solvability [XHZ03]. Solvable [Nab97]. solve [BG13, DV19, HXM19, KBF15, Liv04b, MZHB17, ZJ06]. solved [CZS22]. solver [BvdV00, Bos19, CHV05, DJW^{+}21, GKK04, IIFM23, KK13, KR06, LP22, LSS03, LM06, MNCT07, MRT02, Ols99, OZ22, Pad99, PR11, RTN03, Rak99, RGG07, RGM17, SS02, Sol14, SKR08, TW20, TH19, Vab23, Yot01]. solvers [AGG^{+}16, AG99, ABK97, Ber01, BC02, BO13, BHL^{+}22, DKM^{+}22, FÇ23, FS09, HLM92, HLM^{+}18, HS05, KKPS18, LR08, Lee16, Mey94, MSB18, NO04, Sch12, Sco99, Web18]. solves [Cha07, GP18]. Solving [BG05a, CCS19, EAA19, Nov03, WZZ18, AH02, AL21, AMMR17, AK99, AK00, Bai18b, BW19, BL20, BSI17, Cao04, CQ10, CWwS18, CL23, CC03, CNY05, DA21, DN12, DDL^{+}21, EM11, FH94, HKKP07, HM14, HL21, JLW05, Jou94, KS15, KKMM12, KM92, LT09, LWZ22, LLV19, Liv14, MZ15, MLV05, Mez20, NQ96, PM97, yPxP06, QACT18, RSR10, SL19, Shi02, Śmi19, Sto92, SHJC18, TT10, USS21, Var08, Vla00, WTZD10, Yan23, ZL22, mMP99, mM04, vGSZ15]. Some [BFG95, BM05a, CGK94, CZ02, HM14, LS06, Mar95, Sun06, Ber01, BB06, CDW06, DP23, DS10, ESC20, GL02, LV08, LHL07a, Peñ09, XZS15, dF20]. SOR [Che02]. sorting [Bra02]. source [TH19]. Space [Lee12, AT15, AMM04, AFK02, BPSH13, BMS17, BMS18, BV13, BC12, DHNR18, GB15, How18, ITS07, KV92, KLS23, KLM15, LSC21, QPS23, RS18, RSR10, SY18b, WRW18, ZZ21, vVW23]. Space-angle-energy [Lee12]. space-time [LSC21, vVW23]. spaces [GH06, LV12, LZY11, LPW06, VSG09]. Sparse [CDG00, CDGmM04, Vas02, WWC^{+}15, AB00, AK23, ADT19, BPS95, Bas00, Bau08, BF11a, BEH^{+}17, BPS00, BV00, BG00, CS96, DDM23, DCT18, DR03, EW13, ED22, FÇ23, FJP12, FSS18, GHO15, Gus03, HLM^{+}18, HS15, How18, HS05, HW22, Huc98, ISZ09, JZ09, JK17, KKNY01, KKS19, KNY99, LLL97, LV98, LSS18, Mey94, NLZ11, NY03, NH98, RTN03, RK18, RS18, SZ99, SS02, VS17, WLBH12, WN18, WLC21, XM17, ZXS20, vGSZ15]. sparsity [PPS20, Poi00]. spatial [BLP17, Bai18b, BL20]. spatially [OZB^{+}18]. SPD [HLM^{+}18, Mar16]. Special [Ano08, CLR13, Fal06, LD08, VW01, Vas05, Ben08, Dat01, DP23, ES07, Mey94, Axe99]. specially [SHT11]. specified [AK23, fLyHZ11]. spectra [DHBV21, FBSC21]. Spectral [ADT19, CDDSC12, DLSvL20, ED22, Lee21b, MST16, NSCTPW22, RSCTP20, SGSM15, mMvdV02, BEV22, BSC20, BPS95, BFdP13, BM17, CQZ13, CNZ17, CfX05, DGC19, GMSCS20, LQY13, LNQ13, MS14, MDMS23, MC09, Par03, SK01, SK21, ZWQA18, DFF^{+}18]. spectrum [CR20, Cao09, Lor14]. Speed [LY15]. sphere [ALM18, SL19]. spheres [WCZ15]. Spline [LPS16, EFG^{+}18, MDMS23]. splines [LY15]. Split [HR05]. Splitting [HN05, LXX17, BGN07, Bai10, Bai12, BLP17, BL20, CJZ11, Che15, DL23, Gan99, HL16, HM16, KKR14, LHL07b, LZ22, LZ23, SB12, VEV23, Wan18b, Wan18a, WCW20, Wu15, wX15, ZM20]. spring [EKS02]. spring-mass [EKS02]. SQP [AH02]. Square [DNR12, TY10, Mor09]. squared [BES14, CL23]. squares [AB00, AK99, BDGL09, BW19, Bar02, BMM06, BGM09, BGM11, BGM^{+}12, CYZ99, CNP96, CTP09, CP12, CP06, DE98, DH18, DW07, DWWQ13, DDL^{+}21, EAA19, ES07, ES09a, ER96, EY23, FB95, GW00, GR05, KMM18, KLM^{+}06, LVD02, pLL07, LZ12, LW17, LL97, MS22, MMN^{+}10, MVK04, MLV05, MDB21, MVLB23, MYD20, Miy15, PY22, Pen08, Ren98, RLG12, Sto92, TDH^{+}18, Tia13, Vab23, WKS95, WWC^{+}15, XXW19, ZHZ10, ZY19]. squares-total [ZY19]. SSOR [Bai16, GKY97, WH94]. SSOR-like [Bai16]. Stability [CJW06, DHS95, OCYM08, BV13, DGB^{+}13, DS13a, EM11, KSB13, Lee10, NX03, Peñ03, Sau95, ST17b]. stabilization [AB12, AG19, DGB^{+}13, DGRR11, Lay05]. Stabilized [BH07, MW21, Cao04, CL23, EWY03, KOV17, LMM00, RGM17]. Stabilizing [VW97]. Stable [OS01, ABK15, CGS20, Gem00, GMS18, LXW13, LSJ18, MCV01, ZG22]. Stably [CC20]. stage [AMMP06, BM17, JS96, MPS96]. Staggered [DFF^{+}18, OCYM08]. standard [Han13, LPV01]. standard-form [Han13]. standpoint [Voe92]. start [LW98]. State [DGRR11, BV13, BF11b, CD11, DK15, KV92, LCHH18, PSW14]. state-constrained [PSW14]. state-space [BV13, KV92]. state-time [DK15]. static [LNY15]. stationary [AMP99, BH16, LMM00, MM98, NX03, RBV08, ZW10]. statistical [DXW12, LX08, LT08]. Steady [HG00, BF11b]. steady-state [BF11b]. Steepest [De 13, NZ14, Shi02, Shi04, ZBCN23]. Stein [BES14, CL23]. stents [GT19]. step [AV94, CGY22, CWwS18, CK10, Li00, Ma22, MYD20, PBN05]. stepping [Lam12]. steps [BN21, Fas05, Shi02]. Stewart [HC05]. Stiefel [CZ15]. Stieltjes [AN94, FSS18, GL21]. stiffness [DKVB15]. stochastic [AD12, BMMR18, BDM^{+}14, DMS17, GHR98, KKO20, Lee16, MB21, MM98, RBV08, ROA13, SGP14, TY10, YBZ19]. Stokes [ABM17, AB12, AK99, BKP02, CA99, CB21, CH21, CBE18, DFF^{+}18, FHM21, HM18, He23, HFW01, KOV17, LR08, Lee10, LMM00, LD07, Ols99, PT17, QvGvW^{+}21, VHM^{+}22]. Stokes-like [Lee10]. storage [SWW21]. Strang [ZCW11, CNP96, NR12]. Strang-type [ZCW11, NR12]. strategies [AGG^{+}16, BE98, CDG00, DMM^{+}08, DMMR23, GTY97, HSCTP05, Kap94, PM97, PGT14, SGSM15, SMSW00, WW20]. strategy [BBM^{+}06, BM05b, BM06, DDKR23, GP18, Sco99, SY18a, WLBH12]. strength [OST10a]. Strengthened [AALS01, AM96, Bla03, Mar94]. stress [KLS23, MM02]. stress-yielding [KLS23]. stretch [TY10]. stretched [HST22, KM92, ZMO10]. stretching [AB00]. strictly [ODH21]. Strong [DGB^{+}13, Bai18a, DK23, DS13a, MK23]. strongly [ABK15, KW99]. structural [GMTV16, NR22]. Structure [LW23, BS01, FZwCW17, FBSC21, Hem96, HHLL16, PR16, RU22, Rja98, WRW18, WN05]. structure-preserving [HHLL16, PR16]. Structured [BGW05, BG05b, CCE^{+}18, CCLN05, CNXY20, MCC^{+}12, SLV04, Tyr05, CCLQ18, DDG99, Dia09, GLGR10, Gem00, HM18, KK23b, LVD02, LYL15, MMC12, MVK04, MLV05, MP13, NR11, NR17, NR19, Poi00, Sun05, SHT11, Tre05, XXCB20]. structures [BCK05, BH04, EJK01, NSCTPW22]. structuring [SV11]. Studies [Zho06]. study [KKO20, LR08, RS18]. sub [CZ15, LPS15, SV11]. sub-diffusion [LPS15]. sub-Stiefel [CZ15]. sub-structuring [SV11]. subclasses [LHL07a]. subdiffusion [DOR21, LPSV18]. subdivision [EGMS20]. subdomain [HLM92]. subgraph [BCZ12]. submatrix [KK02, fLyHZ11, pLL07, yPyHZ04]. suboptimal [HS15]. subsets [MPV06]. Subspace [CS02, DDG99, HMMP19, RMM22, BMAA16, Bot13, CS97, Dam08, DK95, GLJ19, GZ16, GTI16, HCD15, HS11, HS14, HS21b, HXM19, HL21, HFG^{+}22, IP13, KKRS21, KS10, KLMP21, LS15, NR14a, RLG12, SCP20, Sid97, SS07, XX22, ZS08]. subspace-based [GZ16]. Subspace-by-subspace [DDG99]. subspaces [BDK^{+}15, DF01, IT05, MP16, PPv95, VS17, WW20]. substructuring [GMR05, KH23]. Subtracting [KJ12]. successive [BGN07, Gus03, WQ07]. successive-overrelaxation [BGN07]. sufficient [Pul08]. suitable [HK21]. sum [AD12]. Summation [LC21, FP05, KK16]. Super [CNSY05]. Super-resolution [CNSY05]. Superconvergence [FY01]. superfast [CHV05]. Superlinear [Kap05]. superlinearly [CQ10]. superoptimal [CJW06, CDDZ19]. supersymmetric [HCD15, WQ07]. supervised [PY22]. supply [CPSM06]. supported [FP15]. supports [Pul09]. Surfaces [LD08]. surgical [YPC20]. surveillance [LNY15]. survey [CQZ13, ST23, SK01]. SVD [FJ05, GL18, SSB19, XQ09, ZN20]. sweeping [BPS15]. switching [CSB20, MN00]. Sylvester [Bau08, BMAA16, BHHJ13, CLR01, CD11, DIPR19, DXW12, HJ18, JMPR18, KLMP21, LZZ20, MP15]. Sylvester-observer [CLR01, CD11]. symbol [BEV22, BG23, DGM^{+}16]. symbols [ESC18]. Symmetric [AIT05b, Liu22, PN18, QXB09, Zha92, ARMW14, AG95, AK00, BGP97, BV00, Ber12, BMM20, Bla02, BCS09, BPS13, BM05b, BM06, CL96, CRS05, CR20, Cao04, CS09, CS11, CDGmM04, CK01, CHV05, CS95, DS10, DJ09, EW13, EZ96, EN17, GB11, GM11, GMTV16, HKKP07, HR05, HVCY21, HES15, HS15, IK00, Jia17, KKPS18, KH23, Kap98, Kau07, LOY08, LQY13, LLLJ16, LWZ22, pLL07, Lu05, MVV08, MZ98, MST16, NSCTP05, Not02a, O’H14, PS11, PS00, RT02, RS18, RMM19, ST17b, Sei10, SS97, Sot13, VVM05b, WQZ09, WBL14, Wu15, XHZ03, XQ09, YZCQ23, YLH11, ZZLX20, ZQW13, vN00]. symmetrization [FBSC21, GM11]. symmetrizing [Tyr92]. symmetry [Pen08, Szu14]. symmetry-constrained [Pen08]. symplectic [DS13a]. synchronization [CvG11, ŚLA^{+}21]. Synchronous [EGMS20, BZ13, Mez20]. synthesis [RGG07]. system [AALS01, BC09, Baz08, BB06, BvdV00, BGM^{+}12, CJL08, GLOW04, GP18, HES15, ITS07, KLM^{+}06, KRW08, LW04, MMN^{+}10, SB12, SCD94, USS21, ZS08]. systematic [GLOW04]. Systems [Jia96, Nab97, ARSO14, AM96, Ada04, ACR^{+}00, AL21, AMP99, AMMP06, AK00, AN03b, BPS15, BGS21, BLP17, Bai18a, BG13, BBJ17, Bas00, Bat95, BGM09, BFPS10, BEH^{+}17, BDdSM18, BMM20, BMN05, BW17b, BEG18, BRT07, Bot13, BVD^{+}18, BSI17, CS09, CS11, CDGmM04, CD11, CPSM06, CPS01, CSCTP05, CC03, CNY05, CK01, CA99, CHV05, CS95, CP06, DSV18, DDG99, DFHM20, DIPR19, DGRR11, Dob99, Dod11, DGM^{+}16, DFF^{+}21b, DL23, DN12, EKS02, Ema12, EN17, EM11, FP15, FLM09, FM18, FH94, Gem00, GLJ19, GM11, GSS01, GTY97, GA18, GKY97, GS05, GD11, HLM^{+}18, HKKP07, HS11, HN05, HW18, HSCTP05, IIFM23, JZ09, JK17, JYH17, JL09, Jou94, KBF15, KM99, KKR14, KKMM12, Lai97, LX08, LOY08, Lee19, Lee21a, LOS04, LJ04, LHL07b, LT09, LC13, LZ22, LC05]. systems [LC07, LW03, Lot07, LSS18, MO11, MPR20, MS14, MW11, MZHB17, MCV01, Mey94, MPS96, MST16, NSCTP05, NCV05, PM97, PW13, QACT18, RK18, RS18, RVW98, SZ99, SS02, Sac05, SPD05, SP06, SP18, SS07, SH19, SMSW00, Spi21, Ste95, SHZ20, Sun05, SL10, SHJC18, Szu14, TYZL23, TT10, TR21, TC10, VFdV13, VZ08, WD08, WM12, Wan18b, WTWG14, Wu15, Yan23, Ye20, ZW10, ZL22, ZXS20, vGSZ15, HS14, HS21b].

t [mM04, ZN20]. T-SVD [ZN20].

tangential [AN03a, AN07]. technique [CN21, HM03, IP13, NY03, WZZ18, XXCB20]. techniques [ACR^{+}00, BB00, Bla94, CDDSC12, CS97, CFAM16, Dat01, ELV94, GKL18, GNR14, HK02, HS05, LM06, QPS23, SZ99, Ver00, YHAG20, BFG^{+}18]. technology [RSCTP20]. template [LB17]. Tensor [ADO23, BFG^{+}18, BZ23, DHW16, JYZ17, KK23a, SKKS22, SVV22, VR23, ADMS22, AT15, ACGH21, AK16, BG13, BMAA16, BH16, CWWZ22, CLNY15, CL23, DW15, DH18, DQW15, DK15, DDL^{+}21, DS13b, EAA19, ED22, EY23, FZwCW17, HKST12, HS18, HDIS18, HXM19, HL21, KK16, KKS19, KK23b, KN14, LQY13, LCN13, LXX17, LWZ22, LLV19, Liu22, LP16, Lun20, MS22, MK20, MYD20, OST10b, RU22, RMM22, STZ12, SNZ20, TYZL23, TYF23, WQZ09, WCW20, WLC21, XC13, YZCQ23, YB23, ZQ12, ZQLX13, ZSKA18, ZN22, ZQW13]. tensor-structured [KK23b]. tensor-train [RMM22]. tensors [BW17a, BNS20, BSMN22, CQZ13, CCLQ18, CHCS22, DK23, ED22, FMPS13, HCD15, HKL21, HHQ13, KJ12, LLK14, LNQ13, Lun20, MCLM20, MCC^{+}12, MK23, MVLB23, O’H14, PN18, WQ07, WN18, ZN20, ZZLX20, ZWQA18]. term [BDR17, Lai97, WM12]. Termination [Bir15]. terms [PPS20]. tessellations [DE06]. test [BC09, CCLQ18]. tetrahedral [Bla03]. texture [WSN19]. TFETI [DHBV21]. th [AEHV15, LZ09, LH17]. their [BGS21, BKP02, CEQN07, KK23a, KCC16, Kub92, LY15, LHW11, LWS^{+}23, MDMS23, MDB21, Tia13, Vöm10, Xie11]. theorem [Adi08]. theorems [BBP03, BKP02, CP99]. Theoretical [MO14, Gar04, MM18, Not05b, WF15]. theories [BDRS12, BNR18]. theory [AHJ20, ABK97, ABNP15, CCvG06, CQZ13, FT98, GW00, GL98, HM14, JL09, LQY13, LNQ13, Miy17, Pul16, VW97]. thermal [HK12]. thermoacoustics [SGSM15]. thin [The98]. third [ABBP10, BCR11, BCR14, Lun20, YB23, ZN20]. third-order [ABBP10, BCR11, BCR14, Lun20, YB23, ZN20]. Thomas [LV12]. Three [LWS^{+}23, AALS01, BO18, BB96, CGPV13, DM10, HMMP19, HW18, HDH19, Ibr02, KK23b, KT08, Rja98, SKKS22, XZS15, YW12]. three-by-three [HDH19]. three-dimensional [AALS01, CGPV13, KK23b, KT08, Rja98, SKKS22]. three-parameter [HMMP19]. three-way [Ibr02]. threshold [Saa94, SZ99]. thresholding [CWWZ22, HW22, LM06]. Tight [Du19, OOO11]. Tikhonov [BCB14, BDR17, CRS05, CLTW11, Don05, FRR16, GNR14, LHW11, RU22]. time [AT15, ABK15, Bai12, BW17b, BMS17, BMS18, Bot13, CNT07, Cas11, CLNY15, CJL08, DV19, DFHM20, DFF^{+}21a, DGRR11, DK15, FS21, GZ16, GS07, HG00, HNR^{+}18, Kem12, KK13, Lam12, LLS12, LP22, LGS12, LC13, LSC21, LPSV18, MV13, MC09, RBV08, SL19, SY18b, SSSF23, TH19, WZZ18, YPC20, ZYFG11, ZZ21, ZSWX13, vKVW00, vVW23]. time-delay [DGRR11, LC13]. time-dependent [CNT07, LP22, MV13, RBV08, ZYFG11]. time-exact [Bot13]. time-fractional [LPSV18]. time-harmonic [Bai12, GS07, LGS12, ZSWX13]. time-independent [CJL08]. time-periodic [KK13, WZZ18]. time-space [ZZ21]. times [KVW10]. tire [SMSW00]. Toeplitz [BF19, AH02, BLP17, BG05a, BG23, BG05b, CNP96, CPS01, CGK05, CNY05, CHV05, CS95, Don10, DGM^{+}16, DFF^{+}21b, ESC18, ESC20, Est09, FLM09, FBSC21, HR05, Hem96, HSCTP05, KN07, LC05, LC07, Lot07, LPS15, MS14, MVV08, NR11, NPR13, NR19, NCV05, PS11, RSCTP20, SP18, WtFW15, dF20]. Toeplitz-block [SP18]. Toeplitz-plus-Hankel [KN07]. Toeplitz-type [NR19]. tolerant [RTN03]. tool [FM15, GS97]. tools [BBP03]. topology [HP04, KS22, Vas02]. Total [CLNY15, MPR23, CTP09, FB95, GR05, LVD02, LW17, MVK04, MLV05, MDB21, XXW19, ZZ15, ZY19]. totally [BP13, Hua12]. Trace [KCS11, BFM12]. tracking [LB17]. train [ADO23, BZ23, RMM22, ZN22]. training [PLMV23]. transfer [Don10, GVT03, KV92]. transfers [WTWG14]. transform [CVY21, CW97, DHR20, QPS23, YB23, TW20]. transformation [FLPW01, HSY18, LL97, MC09, OOO16, VVM05a]. transformations [CHV05, Dax04, Han13, JO01]. transformed [MPR20, SNZ20, WLC21]. transforming [Lin12]. transforms [FP05]. transient [KWS^{+}18]. transition [BH16, LCN13]. translation [KY95]. transmission [GH01]. transport [AHJ20, Cha07, CGM11, HM14, Miy17, TC10]. travel [TH19]. treatment [JM10, MM09]. tree [Vöm12]. Trees [BMP11]. Treffitz [LLW09]. triangle [RSCTP15]. Triangular [Zho16, BNT94, BF11a, FP95a, GLGR10, KABH17, KKR14, LPS15, MMMM09, Mit10, RS10, SRGL13, SST18]. triangulations [RSCTP20]. Tridiagonal [NPR13, Zha92, BPS15, BF19, BM05b, BM06, EM11, Jia17, LLLJ16, NR19, XQ09, YLH11]. trigonometric [CHV05, FP05]. trigonometry [Gus97, Gus98, Gus03]. trilinear [BG02]. triplet [LT11]. triplets [SS97]. tropical [CZS22]. Truncated [GKK04, KS15, GTI16, LHW11, MDB21, NR14b, SSB19]. truncation [STZ12, Zho16]. Trust [Naz95, HS18, fLWyL^{+}21]. trust-region [HS18, fLWyL^{+}21]. TTRISK [ADO23]. Tubes [LD08]. Tucker [GL18, JYZ17, YZCQ23]. Tuned [FK15, Mar16]. tuning [FLPW01]. tunnel [PM97]. twisted [XQ09]. Two [BM17, CSCTP05, CWwS18, CX22, DLVZ06, ES09b, rFS09, HH06, HM20, JS96, KM99, KV96, Ma22, PBN05, PL21, TSMM21, XZS10, Yon96, Zha92, ZSWX13, Zik08, vRH05, AM96, AD12, AABHV18, AMMP06, AN13, BSI17, CGPV13, CGM01, CG15, CX23a, CX23b, DJW^{+}21, DY04, DFNY08, DBG06, DKM^{+}22, EN17, ELV94, FVZ05, FK15, FH94, GVT03, HK21, HHvR04, HVX16, HC20, HHLL16, Ian16, JZ11, KWS^{+}18, LS22, MCV01, MSV13, MP15, MPS96, NN10, NH06, Not10, NCV05, SY18a, VSG09, WM12, Wan18b, WW20, XSZ09, Yan18, YXZ13, ZJ06, ZZ21]. two-by-two [AN13, MP15, Wan18b]. two-component [NH06]. two-dimensional [DY04, LS22, XSZ09, ZZ21]. Two-grid [CSCTP05, ZSWX13, CGPV13, CG15, ELV94, FVZ05, HVX16, HC20, NN10, Not10]. Two-level [DLVZ06, HH06, HM20, KM99, KV96, TSMM21, XZS10, vRH05, CGM01, EN17, GVT03, HHvR04, KWS^{+}18, NCV05, VSG09, YXZ13, Zik08]. two-parameter [Yan18]. two-phase [HHLL16, NH06, SY18a]. two-player [AD12]. two-real-parameter [MSV13]. Two-sided [CX22, PL21, Zik08, CX23a, CX23b, DJW^{+}21, FK15, JZ11, WW20, ZJ06]. Two-stage [BM17, JS96, AMMP06, MPS96]. Two-step [CWwS18, Ma22, PBN05]. two-term [WM12]. two-variable [HK21]. Two-Way [Zha92, MCV01]. type [Aih20, ABBP10, AABHV18, BR07, Baz08, Ben08, Cao08, CWwS18, CWS97, EG16, FG02, GKK04, GH23, HM14, KKNY01, KCV09, LHLS07, LT13, LCZZ21, LS22, MP15, NR12, NR19, Pas19, SCD94, Vla00, ZCW11]. typical [XZS15].

UK£30.00 [Nab97]. Ulm [Ma22]. unbalanced [FLM09]. uncertain [DGB^{+}13, DGRR11]. uncertainty [Lee21b, NV23, SCP20]. unconstrained [HZZC23, Ris19]. underdetermined [QACT18]. Unified [Axe15, DHS23]. Uniform [BLZ08, Lee10, CLQY23, HMS99, XC13]. uniformization [Sid11]. unifying [Aih20]. unilevel [Tre13]. uniqueness [LLNV17]. unit [WCZ15]. unitarily [YL08]. unitary [DPRV19, JR94, Lor14, Mat96, MCLM20]. unity [BDV06]. University [Nab97]. unreduced [MST16]. unsteady [OC04]. unstructured [Cho03, KV96, Mav01]. Unsymmetric [Jia96, EM11, GR04, HS05, MS14]. untangling [GKK04]. unwrapping [DY04]. update [ZZ15]. updates [BDdSM18, DEM18, TT10, Tyr92, Zho06]. Updating [LB17]. Upper [Mar94, BNT94, Du19, GX14, ZLLH23]. US$49.95 [Nab97]. Use [HKST12, Bla02, BDS94, FH94, HS15, Yan04]. used [KV15]. users [GB15]. Using [BBP03, GB15, Kap02, AW11, AMR18, AFK02, BB16, BP13, BTT13, BC02, BN21, Buc11, CSB20, CKW02, CS18, CNSY05, CHV05, DK15, ED22, FS21, GL21, GP18, HDIS18, ISZ09, KKRS21, KMMR10, KK23b, KRW08, Kra06, Laz16, MGF^{+}02, MW21, NG15, NX03, OZB^{+}18, OOO11, Özb13, PDV05, PH19, Pul09, QvGvW^{+}21, RTN03, RMM19, Sim03, SNZ20, SVV22, Uhl23, VY14, VS17, YZCQ23, ZLLH23, ZM20, vNR07, vVW23]. usual [BG05b]. Uzawa [Cao04, HW18, HDH19, LRGO17, MZ15, SX15].

validation [CH03, OOO16]. value [AFSCSU14, BBP03, BWN05, CWWZ22, CL13, Che15, JK18, Lee21a, LHW11, LT13, MB21, Ma22, MSB18, NR14b, Nov03, PBN05, PH19, RT99, SHvBW21, SNZ20, ZSKA18]. valued [AK00, DGM^{+}16, MZHB17, Xie11]. values [FP95a, GR99, LHLS07, THC09, Tre05, Uhl23, Vöm10]. Vandermonde [YM22, ZLLH23]. Vandermonde-wise [ZLLH23]. Vanka [FHM21, GH23, He23]. Vanka-based [He23]. Vanka-type [GH23]. Variable [AV94, CH21, DOR21, DHR^{+}04, GVT03, GSTPT21, GR05, HK21, RS07, SKKS22, SX15, VS17]. variable-order [DOR21]. Variable-step [AV94]. variant [JYH17, ST19, Sim99, YBZ19, Zha18]. variants [VY14, Wu15, WX21]. variation [CLNY15, HST22, ZZ15]. Variational [Gar04, DLSvL20, DH04, DKVB15, Gar01, NWZ17, TDH^{+}18, TDL^{+}22]. variations [HLLL13]. various [BE98]. vector [BDK^{+}15, BL22, BMP11, BH16, DO18, DQW15, GH06, GB15, IIFM23, KV15, LCN13]. vectors [AW11, FP95a, LLNV17, MM98, Par92]. velocity [CD11]. Velte [SSB04]. verified [KBF15, Miy17]. version [Beu03, HMS99, LSS03, Not94, PR95]. versions [LV08, NV08a]. versus [GMR05, Kra02, Tur00, YHAG20]. very [KBF15]. via [BSMN22, BZ23, CCS19, Dax04, DGM^{+}16, FRR16, GKL18, How18, HHQ13, HXM19, KSB13, Kho96, KNX01, LCHH18, MVLB23, NY03, RSCTP20, WN18, WZZ18]. vibrating [BC09, CD11, LW04]. video [LNY15]. view [HS11, HS14, HS21b, TYF23]. viewpoint [Ian16]. viscoelastic [MV19]. viscosity [CH21]. vision [CZ15]. volume [CGL05, DJW^{+}21, KKO20]. Voronoi [DE06].

water [CPSM06, SL19]. wave

[BO13, CJL08, DW21, DA21, Liv04b, Liv14, RV12, mM04]. wave-ray [Liv04b]. waveform [FP95b, Gan99, PBN05]. wavelet [DOR19, ISZ09, LM06, YB23]. wavelets [VW97, vVW23]. Way [Zha92, EGMS20, Ibr02, MCV01]. weak [DLSvL20]. weakly [BPS13, HM16, ZWQA18]. weather [BS10]. weight [BCHT04]. Weighted [GZ16, SSSF23, Bar02, DBLP16, HDIS18, JK18, MPR20, SLV04, SLV06, YL08]. weighting [LW17]. Well [MM09]. Well-conditioned [MM09]. Weyl [DA21]. where [Sau95]. width [PLMV23]. Wilkinson [DCT18]. William [Bun95]. wind [BFPS10]. wise [BF11b, ZLLH23]. within [BS01, NFD10]. without [EM11, HM96, LSS18, Van00]. Woodbury [HS21a, MS14]. Workshop [BFG^{+}18, FM99]. Wronskian [MPR22].

years [Axe10, LBG13]. yielding [KLS23]. Yosida [PSW14]. Young [KVW10].

zero [AD12, BB97]. zero-sum [AD12]. zeros [MN05]. Zienkiewicz [Ano08]. ZNN [Uhl23]. zone [NO04].

References

Amat:2018:THO

[AABHV18] S. Amat, I. Argyros, S. Busquier, and M. A. Hernández-Verón. On two high-order families of frozen Newton-type methods. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Achchab:2001:SCB

[AALS01] B. Achchab, O. Axelsson, L. Laayouni, and A. Souissi. Strengthened Cauchy-Bunyakowski-Schwarz inequality for a threedimensional elasticity system. Numerical Linear Algebra with Applications, 8 (3):191–205, April/May 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76506672/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76506672&PLACEBO=IE.pdf.

Adlers:2000:MSS

[AB00] Mikael Adlers and Åke Björck.

Matrix stretching for sparse least squares problems. Numerical Linear Algebra with Applications, 7(2):51–65, March 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71008527/
START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=71008527&PLACEBO=IE.pdf.

Axelsson:2010:PMP

[AB10] Owe Axelsson and Radim Blaheta. Preconditioning of matrices partitioned in 2 × 2 block form: eigenvalue estimates and Schwarz DD for mixed FEM. Numerical Linear Algebra with Applications, 17(5): 787–810, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See comment [Cao13, AB13].

Amodei:2012:SAN

[AB12] L. Amodei and J.-M. Buchot. A stabilization algorithm of the Navier–Stokes equations based on algebraic Bernoulli equation. Numerical Linear Algebra with Applications, 19(4): 700–727, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2013:RCP

[AB13] Owe Axelsson and Radim Blaheta. Response to Comment on ‘Preconditioning of matrices partitioned in 2 × 2 block form: Eigenvalue estimates and Schwarz DD for mixed FEM’ by Owe Axelsson, Radim Blaheta. Numerical Linear Algebra with Applications, 20(3): 536–539, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [AB10, Cao13].

Amat:2010:TON

[ABBP10] S. Amat, C. Bermúdez, S. Busquier, and S. Plaza. On a third-order Newton-type method free of bilinear operators. Numerical Linear Algebra with Applications, 17(4): 639–653, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1997:INS

[ABK97] Owe Axelsson, Radim Blaheta, and Roman Kohut. Inexact Newton solvers in plasticity: theory and experiments. Numerical Linear Algebra with Applications, 4 (3):133–152, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15031;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15031&PLACEBO=IE.pdf.

Axelsson:2015:PMH

[ABK15] Owe Axelsson, Radim Blaheta, and Roman Kohut. Preconditioning methods for high-order strongly stable time integration methods with an application for a DAE problem. Numerical Linear Algebra with Applications, 22(6):930–949, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Adler:2017:PMC

[ABM17] James H. Adler, Thomas R. Benson, and Scott P. MacLachlan. Preconditioning a mass-conserving discontinuous Galerkin discretization of the Stokes equations. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2015:PIM

[ABNP15] O. Axelsson, R. Blaheta, M. Neytcheva, and I. Pultarovà. Preconditioning of iterative methods — theory and applications. Numerical Linear Algebra with Applications, 22(6):901–902, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2011:MCN

[AC11] Owe Axelsson and Xiaojun Chen. Matrix computations and nonlinear equations. Numerical Linear Algebra with Applications, 18(2): 175–176, March 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Abe:2023:GRC

[AC23] T. Abe and A. T. Chronopoulos. The generalized residual cutting method and its convergence characteristics. Numerical Linear Algebra with Applications, 30(6):e2517:1– e2517:??, December 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ang:2021:ABC

[ACGH21] Andersen Man Shun Ang, Jeremy E. Cohen, Nicolas Gillis, and Le Thi Khanh Hien. Accelerating block coordinate descent for nonnegative tensor factorization. Numerical Linear Algebra with Applications, 28(5):e2373:1–e2373:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amodio:2000:ABD

[ACR^{+}00] P. Amodio, J. R. Cash, G. Roussos, R. W. Wright, G. Fairweather, I. Gladwell, G. L. Kraut, and M. Paprzycki. Almost block diagonal linear systems: sequential and parallel solution techniques, and applications. Numerical Linear Algebra with Applications, 7(5): 275–317, July/August 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72508407/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72508407&PLACEBO=IE.pdf.

Amparore:2011:RMF

[AD11] Elvio Gilberto Amparore and Susanna Donatelli. Revisiting the matrix-free solution of Markov regenerative processes. Numerical Linear Algebra with Applications, 18(6):1067–1083, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Akian:2012:MMT

[AD12] Marianne Akian and Sylvie Detournay. Multigrid methods for two-player zero-sum stochastic games. Numerical Linear Algebra with Applications, 19(2): 313–342, March 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Adams:2004:AMM

[Ada04] Mark F. Adams. Algebraic multigrid methods for constrained linear systems with applications to contact problems in solid mechanics. Numerical Linear Algebra with Applications, 11(2–3):141–153, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Adib:2008:HPG

[Adi08] Majid Adib. A handy proof of Gay’s theorem. Numerical Linear Algebra with Applications, 15(7):637–642, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Alexandrov:2022:NCT

[ADMS22] Boian Alexandrov, Derek F. DeSantis, Gianmarco Manzini, and Erik W. Skau. Nonnegative canonical tensor decomposition with linear constraints: nnCANDELINC. Numerical Linear Algebra with Applications, 29(6):e2443:1–e2443:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Antil:2023:TTT

[ADO23] Harbir Antil, Sergey Dolgov, and Akwum Onwunta. TTRISK: Tensor train decomposition algorithm for risk averse optimization. Numerical Linear Algebra with Applications, 30(3):e2481:1–e2481:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amestoy:1996:MFM

[ADP96] P. R. Amestoy, I. S. Duff, and C. Puglisi. Multifrontal QR factorization in a multiprocessor environment. Numerical Linear Algebra with Applications, 3(4): 275–300, July/August 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000995.

Avron:2019:SCN

[ADT19] Haim Avron, Alex Druinsky, and Sivan Toledo. Spectral condition-number estimation of large sparse matrices. Numerical Linear Algebra with Applications, 26(3): e2235:1–e2235:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amat:2014:AIO

[AEHV14] S. Amat, J. A. Ezquerro, and M. A. Hernández-Verón. Approximation of inverse operators by a new family of high-order iterative methods. Numerical Linear Algebra with Applications, 21(5): 629–644, October 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amat:2015:NFH

[AEHV15] S. Amat, J. A. Ezquerro, and M. A. Hernández-Verón. On a new family of high-order iterative methods for the matrix pth root. Numerical Linear Algebra with Applications, 22(4): 585–595, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2002:SSP

[AFK02] O. Axelsson, I. Faragó, and J. Karátson. Sobolev space preconditioning for Newton’s method using domain decomposition. Numerical Linear Algebra with Applications, 9(6–7):585–598, September/ November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Akinola:2014:CDN

[AFS14] R. O. Akinola, M. A. Freitag, and A. Spence. The calculation of the distance to a nearby defective matrix. Numerical Linear Algebra with Applications, 21(3):403–414, May 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Al-Fhaid:2014:SVE

[AFSCSU14] A. S. Al-Fhaid, S. Serra-Capizzano, D. Sesana, and M. Zaka Ullah. Singular-value (and eigenvalue) distribution and Krylov preconditioning of sequences of sampling matrices approximating integral operators. Numerical Linear Algebra with Applications, 21(6): 722–743, December 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arbenz:1995:MSI

[AG95] Peter Arbenz and Gene H.

Golub. Matrix shapes invariant under the symmetric QR algorithm. Numerical Linear Algebra with Applications, 2(2):87– 93, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arbenz:1999:CSL

[AG99] Peter Arbenz and Roman Geus. A comparison of solvers for large eigenvalue problems occurring in the design of resonant cavities. Numerical Linear Algebra with Applications, 6(1):3–16, January/February 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002983;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002983&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part I (Milovy, 1997).

Axelsson:2019:CFM

[AG19] Owe Axelsson and Ivar Gustafsson. A coarse-fine-mesh stabilization for an alternating Schwarz domain decomposition method. Numerical Linear Algebra with Applications, 26(3): e2236:1–e2236:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Agullo:2016:NRS

[AGG^{+}16] Emmanuel Agullo, Luc Giraud, Abdou Guermouche, Jean Roman, and Mawussi Zounon. Numerical recovery strategies for parallel resilient Krylov linear solvers. Numerical Linear Algebra with Applications, 23(5):888–905, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Alqahtani:2021:BLB

[AGRR21] Abdulaziz Alqahtani, Silvia Gazzola, Lothar Reichel, and Giuseppe Rodriguez. On the block Lanczos and block Golub–Kahan reduction methods applied to discrete ill-posed problems. Numerical Linear Algebra with Applications, 28 (5):e2376:1–e2376:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Al-Homidan:2002:SAS

[AH02] Suliman S. Al-Homidan. SQP algorithms for solving Toeplitz matrix approximation problem. Numerical Linear Algebra with Applications, 9(8): 619–627, December 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Angelova:2020:ASL

[AHJ20] V. Angelova, M. Hached, and K. Jbilou. Approximate solutions to large nonsymmetric differential Riccati problems with applications to transport theory. Numerical Linear Algebra with Applications, 27 (1):e2272:1–e2272:??, January 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Aihara:2020:GCL

[Aih20]  Kensuke Aihara.  GPBi-

CGstab(L): A Lanczos-type product method unifying bi-CGstab(L) and GPBi-CG. Numerical Linear Algebra with Applications, 27(3):e2298:1– e2298:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amodio:2005:CPP

[AIT05a] P. Amodio, F. Iavernaro, and D. Trigiante. Conservative perturbations of positive definite Hamiltonian matrices. Numerical Linear Algebra with Applications, 12(2–3):117–125, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Amodio:2005:SSH

[AIT05b] P. Amodio, F. Iavernaro, and D. Trigiante. Symmetric schemes and Hamiltonian perturbations of linear Hamiltonian problems. Numerical Linear Algebra with Applications, 12(2–3):171–179, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Andreev:1994:CDS

[AJ94] Andrey Andreev and Hussain Juboury. On the convergence of difference schemes for a heat conduction equation. Numerical Linear Algebra with Applications, 1(3):237–245, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1994:DCR

[AK94] O. Axelsson and L. Kolotilina. Diagonally compensated reduction and related preconditioning methods. Numerical Linear Algebra with Applications, 1(2):155–177, 1994. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arushanian:1999:ILS

[AK99] I. O. Arushanian and G. M.

Kobelkov. Implementation of a least-squares finite element method for solving the Stokes problem with a parameter. Numerical Linear Algebra with Applications, 6(7): 587–597, October/November 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/68502742/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=68502742&PLACEBO=IE.pdf.

Axelsson:2000:RVI

[AK00] Owe Axelsson and Andrey Kucherov. Real valued iterative methods for solving complex symmetric linear systems. Numerical Linear Algebra with Applications, 7(4):197–218, May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72507874/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72507874&PLACEBO=IE.pdf.

Axelsson:2010:P

[AK10] Owe Axelsson and David R.

Kincaid. Preface. Numerical Linear Algebra with Applications, 17(5):741–742, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arismendi:2016:MCA

[AK16] Juan Carlos Arismendi and Herbert Kimura. Monte Carlo approximate tensor moment simulations. Numerical Linear Algebra with Applications, 23(5):825–847, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Aksoylu:2019:SMS

[AK19] Burak Aksoylu and Adem Kaya. On smoothers for multigrid of the second kind. Numerical Linear Algebra with Applications, 26(6):e2267:1– e2267:??, December 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ahmad:2023:BEA

[AK23] Sk Safique Ahmad and Prince Kanhya. Backward error analysis of specified eigenpairs for sparse matrix polynomials. Numerical Linear Algebra with Applications, 30(4):e2476:1– e2476:??, August 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Antoine:2021:OBD

[AL21] Xavier Antoine and Emmanuel Lorin. ODE-based doublepreconditioning for solving linear systems Aαx = b and f(A)x = b. Numerical Linear Algebra with Applications, 28 (6):e2399:1–e2399:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Alber:2006:MCS

[Alb06] David M. Alber. Modifying CLJP to select grid hierarchies with lower operator complexities and better performance. Numerical Linear Algebra with Applications, 13(2–3):87–104, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Adler:2018:CGM

[ALM18] James H. Adler, Ilya Lashuk, and Scott P. MacLachlan. Composite-grid multigrid for diffusion on the sphere. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1995:GCG

[AM95] Owe Axelsson and M. Makarov.

On a generalized conjugate gradient orthogonal residual method. Numerical Linear Algebra with Applications, 2 (5):467–479, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Special issue dedicated to David M. Young, Jr.

Achchab:1996:ECT

[AM96] B. Achchab and J. F. Maître. Estimate of the constant in two strengthened C.B.S. inequalities for F.E.M. systems of 2D elasticity: application to multilevel methods and a posteriori error estimators. Numerical Linear Algebra with Applications, 3(2): 147–159, March/April 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15000981.

Ambikasaran:2015:GRP

[Amb15] Sivaram Ambikasaran. Generalized Rybicki Press algorithm. Numerical Linear Algebra with Applications, 22(6):1102–1114, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Austin:2004:RMA

[AMM04] Travis M. Austin, Thomas A. Manteuffel, and Steve Mc-Cormick. A robust multilevel approach for minimizing H(div)-dominated functionals in an H^{1}-conforming finite element space. Numerical Linear Algebra with Applications, 11(2–3):115–140, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arnal:2006:PNT

[AMMP06] J. Arnal, H. Migallón, V. Migallón, and J. Penadés. Parallel Newton two-stage methods based on ILU factorizations for nonlinear systems. Numerical Linear Algebra with Applications, 13(7):553–572, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Appelhans:2017:LCP

[AMMR17] David J. Appelhans, Tom Manteuffel, Steve McCormick, and John Ruge. A lowcommunication, parallel algorithm for solving PDEs based on range decomposition. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Arnal:1999:NSP

[AMP99] Josep Arnal, Violeta Migallón, and José Penadés. Nonstationary parallel multisplitting algorithms for almost linear systems. Numerical Linear Algebra with Applications, 6(2):79–92, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002990;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002990&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Allen:2018:EBS

[AMR18] Jeffery Allen, Tom Manteuffel, and Harihar Rajaram. Exploring basal sliding with a fluidity-based, ice-sheet model using FOSLS. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1994:AMI

[AN94] O. Axelsson and M. Neytcheva. Algebraic multilevel iteration method for Stieltjes matrices. Numerical Linear Algebra with Applications, 1(3):213– 236, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Achdou:2003:ITF

[AN03a] Y. Achdou and F. Nataf. An iterated tangential filtering decomposition. Numerical Linear Algebra with Applications, 10(5–6):511–539, July/ September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2003:PML

[AN03b] Owe Axelsson and Maya Neytcheva. Preconditioning methods for linear systems arising in constrained optimization problems. Numerical Linear Algebra with Applications, 10(1–2):3–31, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2006:EEP

[AN06] Owe Axelsson and Maya Neytcheva. Eigenvalue estimates for preconditioned saddle point matrices. Numerical Linear Algebra with Applications, 13(4):339–360, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Achdou:2007:LFT

[AN07] Y. Achdou and F. Nataf. Low frequency tangential filtering decomposition. Numerical Linear Algebra with Applications, 14(2):129–147, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2013:GAA

[AN13] Owe Axelsson and Maya Neytcheva. A general approach to analyse preconditioners for two-by-two block matrices. Numerical Linear Algebra with Applications, 20(5): 723–742, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2008:CPS

[Ano08] Anonymous. Call for papers: Special issue on ‘The Zienkiewicz Medal and £1000 Prize Numerical Methods in Engineering’. Numerical Linear Algebra with Applications, 15(6):573, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2009:NLA

[Ano09] Anonymous. Numerical Linear Algebra with Applications impact factor for 2008 has been published to be 0.822. Numerical Linear Algebra with Applications, 16(9):i, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IIa

[Ano12a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 19(1):1–3, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IIb

[Ano12b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 19(2): i–iii, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IIc

[Ano12c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 19(3):i–iii, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IId

[Ano12d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 19(4): i–iii, August 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IIe

[Ano12e] Anonymous. Issue information.

Numerical Linear Algebra with Applications, 19(5):i–iii, October 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2012:IIf

[Ano12f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 19(6):1– 3, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2013:IIa

[Ano13a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 20(3):i–iii, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2013:IIb

[Ano13b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 20(4): i–iii, August 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2013:IIc

[Ano13c] Anonymous. Issue information.

Numerical Linear Algebra with Applications, 20(5):i–iii, October 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2013:IId

[Ano13d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 20(6):i– iii, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IIa

[Ano14a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 21(1):i–iii, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IIb

[Ano14b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 21(2): i–iii, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IIc

[Ano14c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 21(3):i–iii, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IId

[Ano14d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 21(4): i–iii, August 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IIe

[Ano14e] Anonymous. Issue Information. Numerical Linear Algebra with Applications, 21(5):i– iii, October 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2014:IIf

[Ano14f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 21(6):i– iii, December 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:IIa

[Ano15a] Anonymous. Issue information.

Numerical Linear Algebra with Applications, 22(1):i–iii, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:IIb

[Ano15b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 22(2): i–iii, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:IIc

[Ano15c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 22(3):i–iii, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:IId

[Ano15d] Anonymous. Issue Information. Numerical Linear Algebra with Applications, 22(4): i–iii, August 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:IIe

[Ano15e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 22(5):i–iii, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2015:II

[Ano15f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 22(6):i– iii, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:IIa

[Ano16a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 23(1):1–3, January 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:IIb

[Ano16b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 23(2): 205–207, March 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:IIc

[Ano16c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 23(3): 391–393, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:IId

[Ano16d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 23(4): 585–587, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:IIe

[Ano16e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 23(5): 773–775, October 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2016:II

[Ano16f] Anonymous. Issue Information. Numerical Linear Algebra with Applications, 23(6): 969–971, December 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IIa

[Ano17a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(1): ??, January 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IIb

[Ano17b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(2): ??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IIc

[Ano17c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IId

[Ano17d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(4): ??, August 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IIe

[Ano17e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2017:IIf

[Ano17f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 25(1): ??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 25(2): ??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information.

Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2018:II

[Ano18f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IIa

[Ano19a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26(1): ??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IIb

[Ano19b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26(2): e2193:1–e2193:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IIc

[Ano19c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26(3): e2194:1–e2194:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IId

[Ano19d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26 (4):e2195:1–e2195:??, August 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IIe

[Ano19e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26 (5):e2196:1–e2196:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2019:IIf

[Ano19f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 26 (6):e2197:1–e2197:??, December 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27 (1):e2247:1–e2247:??, January 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27(2): e2248:1–e2248:??, March 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27(3): e2249:1–e2249:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27 (4):e2250:1–e2250:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27 (5):e2251:1–e2251:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 27 (6):e2252:1–e2252:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IIa

[Ano21a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28 (1):e2307:1–e2307:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IIb

[Ano21b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28(2): e2308:1–e2308:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IIc

[Ano21c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28(3): e2309:1–e2309:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IId

[Ano21d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28 (4):e2310:1–e2310:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IIe

[Ano21e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28 (5):e2311:1–e2311:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2021:IIf

[Ano21f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 28 (6):e2312:1–e2312:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29 (1):e2392:1–e2392:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29(2): e2393:1–e2393:??, March 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29(3): e2394:1–e2394:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29 (4):e2395:1–e2395:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29 (5):e2396:1–e2396:??, October 2022. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2022:IIf

[Ano22f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 29 (6):e2397:1–e2397:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30 (1):e2449:1–e2449:??, January 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30(2): e2450:1–e2450:??, March 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30(3): e2451:1–e2451:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IId

[Ano23d] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30 (4):e2452:1–e2452:??, August 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IIe

[Ano23e] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30 (5):e2453:1–e2453:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:2023:IIf

[Ano23f] Anonymous. Issue information. Numerical Linear Algebra with Applications, 30 (6):e2454:1–e2454:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Alber:2007:PCG

[AO07] David M. Alber and Luke N. Olson. Parallel coarse-grid selection. Numerical Linear Algebra with Applications, 14(8): 611–643, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Abdel-Rehim:2014:ISM

[ARMW14] Abdou M. Abdel-Rehim, Ronald B. Morgan, and Walter Wilcox. Improved seed methods for symmetric positive definite linear equations with multiple right-hand sides. Numerical Linear Algebra with Applications, 21(3):453–471, May 2014. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Abdel-Rehim:2014:EEA

[ARSO14] A. M. Abdel-Rehim, Andreas Stathopoulos, and Kostas Orginos. Extending the eigCG algorithm to nonsymmetric Lanczos for linear systems with multiple right-hand sides. Numerical Linear Algebra with Applications, 21(4): 473–493, August 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Abdi:2019:PII

[AS19] Fatemeh Abdi and Fatemeh Shakeri. Power iteration and inverse power iteration for eigenvalue complementarity problem. Numerical Linear Algebra with Applications, 26 (4):e2244:1–e2244:??, August 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Andrew:2000:ICD

[AT00] Alan L. Andrew and Roger C. E. Tan. Iterative computation of derivatives of repeated eigenvalues and the corresponding eigenvectors. Numerical Linear Algebra with Applications, 7(4):151–167, May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72507873/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=72507873&PLACEBO=IE.pdf.

Andreev:2015:MPL

[AT15] Roman Andreev and Christine Tobler. Multilevel preconditioning and low-rank tensor iteration for space–time simultaneous discretizations of parabolic PDEs. Numerical Linear Algebra with Applications, 22(2):317–337, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1994:VSM

[AV94] O. Axelsson and P. S. Vassilevski. Variable-step multilevel preconditioning methods. I. selfadjoint and positive definite elliptic problems. Numerical Linear Algebra with Applications, 1(1):75–101, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Alefeld:2011:BEA

[AW11] Götz Alefeld and Zhengyu Wang. Bounding the error for approximate solutions of almost linear complementarity problems using feasible vectors. Numerical Linear Algebra with Applications, 18(2): 177–187, March 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:1996:E

[Axe96]  Owe Axelsson.  Editorial.

Numerical Linear Algebra with Applications, 3(5):349– 350, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001003.

Axelsson:1998:PSM

[Axe98]  Owe Axelsson.  Precondi-

tioned solution methods for large scale problems in scientific computations PRISM ’97. Numerical Linear Algebra with Applications, 5(5): 319, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62000049;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62000049&
PLACEBO=IE.pdf. Special Is-

sue: PRISM 97.

Axelsson:1999:ESI

[Axe99] Owe Axelsson. Editorial: Special Issue: Iterative solution methods for the elasticity equations in mechanics and biomechanics — IMMB 98. Numerical Linear Algebra with Applications, 6(6):409–410, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.

                                        com/cgi-bin/abstract/67501475/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=67501475&PLACEBO=IE.pdf.

Axelsson:2002:P

[Axe02] Owe Axelsson. Preface. Numerical Linear Algebra with Applications, 9(6–7):399–400, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2003:E

[Axe03] Owe Axelsson. Editorial. Numerical Linear Algebra with Applications, 10(7):561, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2004:E

[Axe04] Owe Axelsson. Editorial. Numerical Linear Algebra with Applications, 11(5–6):411–412, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2010:PYN

[Axe10] Owe Axelsson. The past 16 years of NLA. Numerical Linear Algebra with Applications, 17(1):1, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Axelsson:2015:UAP

[Axe15] Owe Axelsson. Unified analysis of preconditioning methods for saddle point matrices. Numerical Linear Algebra with Applications, 22(2): 233–253, March 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Aksoylu:2011:RMP

[AY11] Burak Aksoylu and Zuhal Yeter. Robust multigrid preconditioners for the highcontrast biharmonic plate equation. Numerical Linear Algebra with Applications, 18(4): 733–750, August 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:1995:PNS

[Bai95] Zhaojun Bai. Progress in the numerical solution of the nonsymmetric eigenvalue problem. Numerical Linear Algebra with Applications, 2(3):219– 234, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2009:OPH

[Bai09] Zhong-Zhi Bai. Optimal parameters in the HSS-like methods for saddle-point problems. Numerical Linear Algebra with Applications, 16(6): 447–479, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2010:MBM

[Bai10] Zhong-Zhi Bai. Modulus-based matrix splitting iteration methods for linear complementarity problems. Numerical Linear Algebra with Applications, 17(6):917–933, December 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2012:BAS

[Bai12] Zhong-Zhi Bai. Block alternating splitting implicit iteration methods for saddle-point problems from time-harmonic eddy current models. Numerical Linear Algebra with Applications, 19(6):914–936, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2016:SLP

[Bai16] Zhong-Zhi Bai. On SSORlike preconditioners for non-Hermitian positive definite matrices. Numerical Linear Algebra with Applications, 23(1): 37–60, January 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2018:QHI

[Bai18a] Zhong-Zhi Bai. Quasi-HSS iteration methods for non-Hermitian positive definite linear systems of strong skew-Hermitian parts. Numerical Linear Algebra with Applications, 25(4):??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2018:RSH

[Bai18b] Zhong-Zhi Bai. Respectively scaled HSS iteration methods for solving discretized spatial fractional diffusion equations. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Baryamureeba:2002:SLS

[Bar02] Venansius Baryamureeba. Solution of large-scale weighted least-squares problems. Numerical Linear Algebra with Applications, 9(2):93–106, March 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/89013910/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=89013910&PLACEBO=IE.pdf.

Basermann:2000:PBI

[Bas00] Achim Basermann. Parallel block ILUT/ILDLT preconditioning for sparse eigenproblems and sparse linear systems. Numerical Linear Algebra with Applications, 7(7–8): 635–648, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505472/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505472&PLACEBO=IE.pdf.

Batterson:1995:DAN

[Bat95] Steve Batterson. Dynamical analysis of numerical systems. Numerical Linear Algebra with Applications, 2 (3):297–310, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Baur:2008:LRS

[Bau08] U. Baur. Low rank solution of data-sparse Sylvester equations. Numerical Linear Algebra with Applications, 15(9): 837–851, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bazan:2008:NEP

[Baz08] F. S. V. Bazán. Nonnormality estimation in projectiontype system realization methods. Numerical Linear Algebra with Applications, 15(10): 969–983, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Broyden:1996:CTB

[BB96] C. G. Broyden and M. A. Boschetti. A comparison of three basic conjugate direction methods. Numerical Linear Algebra with Applications, 3(6): 473–489, November/December 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15001008.

Bazan:1997:ZLP

[BB97] Fermin S. V. Bazán and Licio H. Bezerra. On zero locations of predictor polynomials. Numerical Linear Algebra with Applications, 4(6): 459–468, November/December 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15049;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15049&PLACEBO=IE.pdf.

Becker:2000:MTF

[BB00] R. Becker and M. Braack.

Multigrid techniques for finite elements on locally refined meshes. Numerical Linear Algebra with Applications, 7(6):363–379, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516700/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516700&PLACEBO=IE.pdf.

Benner:2001:EPM

[BB01] Peter Benner and Ralph Byers. Evaluating products of matrix pencils and collapsing matrix products. Numerical Linear Algebra with Applications, 8(6–7):357–380, September/November 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007289/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=85007289&PLACEBO=IE.pdf.

Beuchler:2006:ISC

[BB06] Sven Beuchler and Dietrich Braess. Improvements for some condition number estimates for preconditioned system in p-FEM. Numerical Linear Algebra with Applications, 13(7): 573–588, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bangay:2016:FLM

[BB16] Shaun Bangay and Gleb Beliakov. On the fast Lanczos method for computation of eigenvalues of Hankel matrices using multiprecision arithmetics. Numerical Linear Algebra with Applications, 23(3): 485–500, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Boffi:2013:CAH

[BBG13] Daniele Boffi, Annalisa Buffa, and Lucia Gastaldi. Convergence analysis for hyperbolic evolution problems in mixed form. Numerical Linear Algebra with Applications, 20(4): 541–556, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barkouki:2017:MRL

[BBJ17] H. Barkouki, A. H. Bentbib, and K. Jbilou. A matrix rational Lanczos method for model reduction in large-scale first- and second-order dynamical systems. Numerical Linear Algebra with Applications, 24(1):??, January 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bronstein:2006:MMS

[BBKY06] M. M. Bronstein, A. M.

Bronstein, R. Kimmel, and I. Yavneh. Multigrid multidimensional scaling. Numerical Linear Algebra with Applications, 13(2–3):149–171, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brannick:2006:EBA

[BBM^{+}06] J. Brannick, M. Brezina, S. MacLachlan, T. Manteuffel, S. McCormick, and J. Ruge. An energy-based AMG coarsening strategy. Numerical Linear Algebra with Applications, 13(2–3):133–148, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bacuta:2003:UFE

[BBP03] Constantin Bacuta, James H.

Bramble, and Joseph E. Pasciak. Using finite element tools in proving shift theorems for elliptic boundary value problems. Numerical Linear Algebra with Applications, 10(1–2): 33–64, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bastian:2012:AMD

[BBS12] Peter Bastian, Markus Blatt, and Robert Scheichl. Algebraic multigrid for discontinuous Galerkin discretizations of heterogeneous elliptic problems. Numerical Linear Algebra with Applications, 19(2): 367–388, March 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blomgren:2002:MSI

[BC02] Peter Blomgren and Tony F. Chan. Modular solvers for image restoration problems using the discrepancy principle. Numerical Linear Algebra with Applications, 9(5):347– 358, July/August 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2009:SNM

[BC09] Zheng-Jian Bai and Wai-Ki Ching. A smoothing Newton’s method for the construction of a damped vibrating system from noisy test eigendata. Numerical Linear Algebra with Applications, 16(2): 109–128, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Branets:2010:CNB

[BC10] L. Branets and G. F. Carey.

Condition number bounds and mesh quality. Numerical Linear Algebra with Applications, 17(5):855–869, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brannick:2012:ASP

[BC12] James Brannick and Durkbin Cho. An auxiliary space preconditioner for linear elasticity based on the generalized finite element method. Numerical Linear Algebra with Applications, 19(3):471–484, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bazan:2014:EGF

[BCB14] Fermín S. Viloche Bazán, Maria C. C. Cunha, and Leonardo S. Borges. Extension of GKB–FP algorithm to largescale general-form Tikhonov regularization. Numerical Linear Algebra with Applications, 21(3):316–339, May 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bru:1998:MSM

[BCC98] Rafael Bru, Rafael Cantó, and Joan-Josep Climent. On Mmultisplittings of singular Mmatrices with application to Markov chains. Numerical Linear Algebra with Applications, 5(4):299–311, July/August 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
10005755;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=10005755&
PLACEBO=IE.pdf.

Bru:2009:SCG

[BCGM09] R. Bru, C. Corral, I. Giménez, and J. Mas. Schur complement of general H-matrices. Numerical Linear Algebra with Applications, 16(11-12):935– 947, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Boman:2004:MWB

[BCHT04] Erik G. Boman, Doron Chen, Bruce Hendrickson, and Sivan Toledo. Maximum-weight-basis preconditioners. Numerical Linear Algebra with Applications, 11(8–9):695–721, October/November 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bae:2005:SAA

[BCK05] Jinho Bae, Joohwan Chun, and Thomas Kailath. The Schur algorithm applied to the design of optical multi-mirror structures. Numerical Linear Algebra with Applications, 12(2– 3):283–292, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2011:SDB

[BCR11] Zhong-Zhi Bai, Raymond H.

Chan, and Zhi-Ru Ren. On sinc discretization and banded preconditioning for linear thirdorder ordinary differential equations. Numerical Linear Algebra with Applications, 18 (3):471–497, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2014:ORS

[BCR14] Zhong-Zhi Bai, Raymond H.

Chan, and Zhi-Ru Ren. On order-reducible sinc discretizations and block-diagonal preconditioning methods for linear third-order ordinary differential equations. Numerical Linear Algebra with Applications, 21(1):108–135, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brenner:2009:MMS

[BCS09] S. C. Brenner, J. Cui, and L.-Y. Sung. Multigrid methods for the symmetric interior penalty method on graded meshes. Numerical Linear Algebra with Applications, 16(6): 481–501, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2003:EAE

[BCV03] Luca Bergamaschi, Marco Caliari, and Marco Vianello. Efficient approximation of the exponential operator for discrete 2D advection-diffusion problems. Numerical Linear Algebra with Applications, 10 (3):271–289, April/May 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brannick:2012:AMM

[BCZ12] James Brannick, Yao Chen, and Ludmil Zikatanov. An algebraic multilevel method for anisotropic elliptic equations based on subgraph matching. Numerical Linear Algebra with Applications, 19(2): 279–295, March 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bujanovic:2015:NFI

[BD15] Zvonimir Bujanović and Zlatko Drmač. A new framework for implicit restarting of the Krylov–Schur algorithm. Numerical Linear Algebra with Applications, 22(2): 220–232, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barker:2021:AMP

[BD21] Andrew T. Barker and Andrei Draganescu. Algebraic multigrid preconditioning of the Hessian in optimization constrained by a partial differential equation. Numerical Linear Algebra with Applications, 28(1):e2333:1–e2333:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2018:BLU

[BDdSM18] L. Bergamaschi, V. De Simone, D. di Serafino, and A. Martínez. BFGS-like updates of constraint preconditioners for sequences of KKT linear systems in quadratic programming. Numerical Linear Algebra with Applications, 25 (5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Baboulin:2009:CCC

[BDGL09] Marc Baboulin, Jack Dongarra, Serge Gratton, and Julien Langou. Computing the conditioning of the components of a linear least-squares solution. Numerical Linear Algebra with Applications, 16(7): 517–533, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bates:2015:MLV

[BDK^{+}15] Daniel J. Bates, Brent R.

Davis, Michael Kirby, Justin Marks, and Chris Peterson. The max-length-vector line of best fit to a set of vector subspaces and an optimization problem over a set of hyperellipsoids. Numerical Linear Algebra with Applications, 22 (3):453–464, May 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brezina:2014:SAA

[BDM^{+}14] Marian Brezina, Alireza Doostan, Tom Manteuffel, Steve Mc-Cormick, and John Ruge. Smoothed aggregation algebraic multigrid for stochastic PDE problems with layered materials. Numerical Linear Algebra with Applications, 21 (2):239–255, March 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buccini:2017:ITR

[BDR17] Alessandro Buccini, Marco Donatelli, and Lothar Reichel. Iterated Tikhonov regularization with a general penalty term. Numerical Linear Algebra with Applications, 24(4):??, August 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2012:IMT

[BDRS12] Zhong-Zhi Bai, Iain S. Duff, Lothar Reichel, and Zhong-Ci Shi. Innovative methods and theories in numerical algebra. Numerical Linear Algebra with Applications, 19(6): 893–895, December 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buccini:2021:NNP

[BDRZ21] Alessandro Buccini, Marco Donatelli, Lothar Reichel, and Wei-Hong Zhang. A new nonstationary preconditioned iterative method for linear discrete ill-posed problems with application to image deblurring. Numerical Linear Algebra with Applications, 28(2): e2353:1–e2353:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bonhoure:1994:UPP

[BDS94] François Bonhoure, Yves Dallery, and William J. Stewart. On the use of periodicity properties for the efficient numerical solution of certain Markov chains. Numerical Linear Algebra with Applications, 1(3):265–286, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Boonen:2006:AMM

[BDV06] Tim Boonen, Geoffrey Deliége, and Stefan Vandewalle. On algebraic multigrid methods derived from partition of unity nodal prolongators. Numerical Linear Algebra with Applications, 13(2–3):105–131, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Burrage:1998:PVA

[BE98] Kevin Burrage and Jocelyne Erhel. On the performance of various adaptive preconditioned GMRES strategies. Numerical Linear Algebra with Applications, 5(2): 101–121, March/April 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5965;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5965&PLACEBO=IE.pdf.

Barlow:2009:MLR

[BE09] Jesse L. Barlow and Hasan Erbay. Modifiable lowrank approximation to a matrix. Numerical Linear Algebra with Applications, 16(10): 833–860, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Beauwens:1994:AFM

[Bea94] R. Beauwens. Approximate factorizations with modified S/P consistently ordered Mfactors. Numerical Linear Algebra with Applications, 1 (1):3–17, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Boito:2018:ESP

[BEG18] P. Boito, Y. Eidelman, and L. Gemignani. Efficient solution of parameter-dependent quasiseparable systems and computation of meromorphic matrix functions. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benzi:2017:AMC

[BEH^{+}17] Michele Benzi, Thomas M.

Evans, Steven P. Hamilton, Massimiliano Lupo Pasini, and Stuart R. Slattery. Analysis of Monte Carlo accelerated iterative methods for sparse linear systems. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2008:LSM

[Ben08]  Peter Benner.  Large-scale

matrix equations of special type. Numerical Linear Algebra with Applications, 15(9): 747–754, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benzi:2011:NSM

[Ben11] Michele Benzi. Numerical solution of Markov chains. Numerical Linear Algebra with Applications, 18(6):897–900, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bertaccini:2001:RPI

[Ber01] D. Bertaccini. Reliable preconditioned iterative linear solvers for some numerical integrators. Numerical Linear Algebra with Applications, 8(2):111–125, March 2001.

CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501760/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501760&PLACEBO=IE.pdf.

Bergamaschi:2012:EDC

[Ber12]  Luca Bergamaschi.  On

eigenvalue distribution of constraint-preconditioned symmetric saddle point matrices. Numerical Linear Algebra with Applications, 19(4): 754–772, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2014:SSM

[BES14] Peter Benner, Grece El Khoury, and Miloud Sadkane. On the squared Smith method for large-scale Stein equations. Numerical Linear Algebra with Applications, 21(5): 645–665, October 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Beuchler:2003:APV

[Beu03] Sven Beuchler. AMLI preconditioner for the p-version of the FEM. Numerical Linear Algebra with Applications, 10(8): 721–732, December 2003. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barakitis:2022:PFD

[BEV22] Nikos Barakitis, Sven-Erik Ekström, and Paris Vassalos. Preconditioners for fractional diffusion equations based on the spectral symbol. Numerical Linear Algebra with Applications, 29(5):e2441:1– e2441:??, October 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Berry:1996:LRO

[BF96] Michael W. Berry and Ricardo D. Fierro. Low-rank orthogonal decompositions for information retrieval applications. Numerical Linear Algebra with Applications, 3(4): 301–327, July/August 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000993.

Bebendorf:2011:PAD

[BF11a] M. Bebendorf and T. Fischer.

On the purely algebraic datasparse approximation of the inverse and the triangular factors of sparse matrices. Numerical Linear Algebra with Applications, 18(1):105–122, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Busic:2011:ICW

[BF11b] Ana Bušić and Jean-Michel Fourneau. Iterative componentwise bounds for the steadystate distribution of a Markov chain. Numerical Linear Algebra with Applications, 18(6): 1031–1049, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bebiano:2019:NRR

[BF19] Natália Bebiano and Susana Furtado. Nearness results for real tridiagonal 2-toeplitz matrices. Numerical Linear Algebra with Applications, 26 (5):e2257:1–e2257:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bebiano:2013:ACP

[BFdP13] N. Bebiano, S. Furtado, and J. da Providência. An algorithm for constructing a pseudo-Jacobi matrix from given spectral data. Numerical Linear Algebra with Applications, 20(2):185–197, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Borges:1995:SIE

[BFG95] Carlos F. Borges, Ruggero Frezza, and William B. Gragg. Some inverse eigenproblems for Jacobi and arrow matrices. Numerical Linear Algebra with Applications, 2(3):195– 203, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2018:WME

[BFG^{+}18] Peter Benner, Heike Faßbender, Lars Grasedyck, Daniel Kressner, Beatrice Meini, and Valeria Simoncini. 7th Workshop on Matrix Equations and Tensor Techniques. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brezinski:2012:MLO

[BFM12] Claude Brezinski, Paraskevi Fika, and Marilena Mitrouli. Moments of a linear operator, with applications to the trace of the inverse of matrices and the solution of equations. Numerical Linear Algebra with Applications, 19(6): 937–953, December 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benzi:2010:SLS

[BFPS10] M. Benzi, L. Ferragut, M. Pennacchio, and V. Simoncini. Solution of linear systems from an optimal control problem arising in wind simulation. Numerical Linear Algebra with Applications, 17(6):895–915, December 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Boley:2000:LSP

[BG00] Daniel Boley and Todd Goehring.

LQ-schur projection on large sparse matrix equations. Numerical Linear Algebra with Applications, 7(7–8):491–503, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505470/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505470&PLACEBO=IE.pdf.

Branets:2002:DMT

[BG02] L. V. Branets and V. A. Garanzha. Distortion measure of trilinear mapping. Application to 3-D grid generation. Numerical Linear Algebra with Applications, 9(6–7):511–526, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bini:2005:SQM

[BG05a] Dario A. Bini and Luca Gemignani. Solving quadratic matrix equations and factoring polynomials: new fixed point iterations based on Schur complements of Toeplitz matrices. Numerical Linear Algebra with Applications, 12(2–3):181–189, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bottcher:2005:SCN

[BG05b] A. Böttcher and S. M. Grudsky. Structured condition numbers of large Toeplitz matrices are rarely better than usual condition numbers. Numerical Linear Algebra with Applications, 12(2–3):95–102, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ballani:2013:PMS

[BG13] Jonas Ballani and Lars Grasedyck.

A projection method to solve linear systems in tensor format. Numerical Linear Algebra with Applications, 20(1):27–43, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bogoya:2023:AET

[BG23] Manuel Bogoya and Sergei M.

Grudsky. Asymptotics for the eigenvalues of Toeplitz matrices with a symbol having a power singularity. Numerical Linear Algebra with Applications, 30(5):e2496:1– e2496:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bellavia:2009:RPK

[BGM09] Stefania Bellavia, Jacek Gondzio, and Benedetta Morini. Regularization and preconditioning of KKT systems arising in nonnegative least-squares problems. Numerical Linear Algebra with Applications, 16 (1):39–61, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bellavia:2011:CEN

[BGM11] Stefania Bellavia, Jacek Gondzio, and Benedetta Morini. Computational experience with numerical methods for nonnegative least-squares problems. Numerical Linear Algebra with Applications, 18(3):363–385, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brezina:2012:PAM

[BGM^{+}12] M. Brezina, J. Garcia, T. Manteuffel, S. McCormick, J. Ruge, and L. Tang. Parallel adaptive mesh refinement for first-order system least squares. Numerical Linear Algebra with Applications, 19(2):343–366, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2021:NPA

[BGM^{+}21] Luca Bergamaschi, Jacek Gondzio, Ángeles Martínez, John W. Pearson, and Spyridon Pougkakiotis. A new preconditioning approach for an interior point-proximal method of multipliers for linear and convex quadratic programming. Numerical Linear Algebra with Applications, 28 (4):e2361:1–e2361:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2007:SOA

[BGN07] Zhong-Zhi Bai, Gene H. Golub, and Michael K. Ng. On successive-overrelaxation acceleration of the Hermitian and skew-Hermitian splitting iterations. Numerical Linear Algebra with Applications, 14(4): 319–335, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See erratum [BN12].

Bergamaschi:1997:ACC

[BGP97] Luca Bergamaschi, Giuseppe Gambolati, and Giorgio Pini. Asymptotic convergence of conjugate gradient methods for the partial symmetric eigenproblem. Numerical Linear Algebra with Applications, 4 (2):69–84, March/April 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001016;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=15001016&
PLACEBO=IE.pdf.

Baghel:2021:CDM

[BGS21] Mohit Kumar Baghel, Nicolas Gillis, and Punit Sharma. Characterization of the dissipative mappings and their application to perturbations of dissipative-Hamiltonian systems. Numerical Linear Algebra with Applications, 28 (6):e2402:1–e2402:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bini:2005:SMM

[BGW05] Dario A. Bini, Luca Gemignani, and Joab R. Winkler. Structured matrix methods for CAGD: an application to computing the resultant of polynomials in the Bernstein basis. Numerical Linear Algebra with Applications, 12(8): 685–698, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2006:ALI

[BGX06] Zhong-Zhi Bai, Xiao-Xia Guo, and Shu-Fang Xu. Alternately linearized implicit iteration methods for the minimal nonnegative solutions of the nonsymmetric algebraic Riccati equations. Numerical Linear Algebra with Applications, 13(8):655–674, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergen:2004:HHG

[BH04] Benjamin Karl Bergen and Frank Hülsemann. Hierarchical hybrid grids: data structures and core algorithms for multigrid. Numerical Linear Algebra with Applications, 11(2– 3):279–291, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bebendorf:2007:SRA

[BH07] M. Bebendorf and W. Hackbusch. Stabilized rounded addition of hierarchical matrices. Numerical Linear Algebra with Applications, 14(5): 407–423, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bozorgmanesh:2016:CTP

[BH16] Hassan Bozorgmanesh and Masoud Hajarian. Convergence of a transition probability tensor of a higher-order Markov chain to the stationary probability vector. Numerical Linear Algebra with Applications, 23(6): 972–988, December 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bouhamidi:2013:PBA

[BHHJ13] A. Bouhamidi, M. Hached, M. Heyouni, and K. Jbilou. A preconditioned block Arnoldi method for large Sylvester matrix equations. Numerical Linear Algebra with Applications, 20(2):208–219, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buttari:2022:BLR

[BHL^{+}22] Alfredo Buttari, Markus Huber, Philippe Leleux, Theo Mary, Ulrich Rüde, and Barbara Wohlmuth. Block lowrank single precision coarse grid solvers for extreme scale multigrid methods. Numerical Linear Algebra with Applications, 29(1):e2407:1– e2407:??, January 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buck:2018:DDP

[BIA18] Marco Buck, Oleg Iliev, and Heiko Andrä. Domain decomposition preconditioners for multiscale problems in linear elasticity. Numerical Linear Algebra with Applications, 25 (5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Birken:2015:TCI

[Bir15] Philipp Birken. Termination criteria for inexact fixed-point schemes. Numerical Linear Algebra with Applications, 22(4): 702–716, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bertalan:2014:ONM

[BISC14] Tom S. Bertalan, Akand W. Islam, Roger B. Sidje, and Eric S. Carlson. OpenMG: a new multigrid implementation in Python. Numerical Linear Algebra with Applications, 21(5):685–700, October 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benzi:2011:RAS

[BK11] Michele Benzi and Verena Kuhlemann. Restricted additive Schwarz methods for Markov chains. Numerical Linear Algebra with Applications, 18(6):1011–1029, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barker:2021:MFP

[BK21] Andrew T. Barker and Tzanio Kolev. Matrix-free preconditioning for high-order H (curl) discretizations. Numerical Linear Algebra with Applications, 28(2):e2348:1–e2348:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brezina:2012:RCB

[BKM^{+}12] M. Brezina, C. Ketelsen, T. Manteuffel, S. McCormick, M. Park, and J. Ruge. Relaxation-corrected bootstrap algebraic multigrid (rBAMG). Numerical Linear Algebra with Applications, 19(2):178– 193, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bakhvalov:2002:ETS

[BKP02] N. S. Bakhvalov, A. V.

Knyazev, and R. R. Parashkevov. Extension theorems for Stokes and Lamé equations for nearly incompressible media and their applications to numerical solution of problems with highly discontinuous coefficients. Numerical Linear Algebra with Applications, 9(2):115–139, March 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/89013912/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=89013912&PLACEBO=IE.pdf.

Baker:2010:IAM

[BKY10] A. H. Baker, Tz. V. Kolev, and U. M. Yang. Improving algebraic multigrid interpolation operators for linear elasticity problems. Numerical Linear Algebra with Applications, 17 (2-3):495–517, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bohl:2003:IMP

[BL03] Erich Bohl and Peter Lancaster. Irreversible Markov processes for phylogenetic models. Numerical Linear Algebra with Applications, 10(7): 577–593, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2020:RHS

[BL20] Zhong-Zhi Bai and Kang-Ya Lu. On regularized Hermitian splitting iteration methods for solving discretized almost-isotropic spatial fractional diffusion equations. Numerical Linear Algebra with Applications, 27(1):e2274:1– e2274:??, January 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2022:CAV

[BL22] Peter Benner and Xin Liang.

Convergence analysis of vector extended locally optimal block preconditioned extended conjugate gradient method for computing extreme eigenvalues. Numerical Linear Algebra with Applications, 29 (6):e2445:1–e2445:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blaheta:1994:DDI

[Bla94] Radim Blaheta. Displacement decomposition—incomplete factorization preconditioning techniques for linearly elasticity problems. Numerical Linear Algebra with Applications, 1 (2):107–128, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blaheta:2002:GGP

[Bla02]  Radim Blaheta.  GPCG-

generalized preconditioned CG method and its use with nonlinear and non-symmetric displacement decomposition preconditioners. Numerical Linear Algebra with Applications, 9(6–7):527–550, September/ November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blaheta:2003:NTG

[Bla03] Radim Blaheta. Nested tetrahedral grids and strengthened C.B.S. inequality. Numerical Linear Algebra with Applications, 10(7):619–637, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bastian:1997:PAM

[BLE97] Peter Bastian, Stefan Lang, and Knut Eckstein. Parallel adaptive multigrid methods in plane linear elasticity problems. Numerical Linear Algebra with Applications, 4 (3):153–176, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15029;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15029&PLACEBO=IE.pdf.

Baek:2011:MMF

[BLLA11] Jung Woo Baek, Ho Woo Lee, Se Won Lee, and Soohan Ahn. A Markov-modulated fluid flow queueing model under D-policy. Numerical Linear Algebra with Applications, 18 (6):993–1010, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bunch:2001:ADI

[BLP01] J. R. Bunch, R. C. Le Borne, and I. K. Proudler. Analysis of the direct and indirect a posteriori RLSL algorithm. Numerical Linear Algebra with Applications, 8(6–7):453–466, September/November 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007284/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85007284&PLACEBO=IE.pdf.

Benner:2008:NSL

[BLP08] Peter Benner, Jing-Rebecca Li, and Thilo Penzl. Numerical solution of large-scale Lyapunov equations, Riccati equations, and linear-quadratic optimal control problems. Numerical Linear Algebra with Applications, 15(9):755–777, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2017:DTS

[BLP17] Zhong-Zhi Bai, Kang-Ya Lu, and Jian-Yu Pan. Diagonal and Toeplitz splitting iteration methods for diagonalplus-Toeplitz linear systems from spatial fractional diffusion equations. Numerical Linear Algebra with Applications, 24 (4):??, August 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buzdin:2008:IDB

[BLW08] A. Buzdin, D. Logashenko, and G. Wittum. IBLU decompositions based on Padé approximants. Numerical Linear Algebra with Applications, 15(8): 717–746, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brannick:2008:UCM

[BLZ08] James J. Brannick, Hengguang Li, and Ludmil T. Zikatanov. Uniform convergence of the multigrid V -cycle on graded meshes for corner singularities. Numerical Linear Algebra with Applications, 15(2-3): 291–306, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brugnano:2005:SLA

[BM05a] Luigi Brugnano and Cecilia Magherini. Some linear algebra issues concerning the implementation of blended implicit methods. Numerical Linear Algebra with Applications, 12(2–3):305–314, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bunch:2005:PSS

[BM05b] James R. Bunch and Roummel F. Marcia. A pivoting strategy for symmetric tridiagonal matrices. Numerical Linear Algebra with Applications, 12(9):911–922, November 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bunch:2006:SPS

[BM06] James R. Bunch and Roummel F. Marcia. A simplified pivoting strategy for symmetric tridiagonal matrices. Numerical Linear Algebra with Applications, 13(10): 865–867, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2013:PII

[BM13] Peter Benner and Thomas Mach. The preconditioned inverse iteration for hierarchical matrices. Numerical Linear Algebra with Applications, 20(1): 150–166, January 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2017:TSS

[BM17] Luca Bergamaschi and Ángeles Martínez. Two-stage spectral preconditioners for iterative eigensolvers. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Beik:2016:KSM

[BMAA16] Fatemeh Panjeh Ali Beik, Farid Saberi Movahed, and Salman Ahmadi-Asl. On the Krylov subspace methods based on tensor format for positive definite Sylvester tensor equations. Numerical Linear Algebra with Applications, 23 (3):444–466, May 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bellavia:2006:IPN

[BMM06] Stefania Bellavia, Maria Macconi, and Benedetta Morini. An interior point Newton-like method for non-negative leastsquares problems with degenerate solution. Numerical Linear Algebra with Applications, 13 (10):825–846, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brezina:2008:GEB

[BMM^{+}08] M. Brezina, T. Manteuffel, S. McCormick, J. Ruge, G. Sanders, and P. Vassilevski. A generalized eigensolver based on smoothed aggregation (GES-SA) for initializing smoothed aggregation (SA) multigrid. Numerical Linear Algebra with Applications, 15(2-3):249–269, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2020:CQN

[BMM20] Luca Bergamaschi, José Marín, and Ángeles Martínez. Compact quasi-Newton preconditioners for symmetric positive definite linear systems. Numerical Linear Algebra with Applications, 27(6):e2322:1– e2322:??, December 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bini:2018:QME

[BMMR18] D. A. Bini, S. Massei, B. Meini, and L. Robol. On quadratic matrix equations with infinite size coefficients encountered in QBD stochastic processes. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blaheta:2005:ROM

[BMN05] R. Blaheta, S. Margenov, and M. Neytcheva. Robust optimal multilevel preconditioners for non-conforming finite element systems. Numerical Linear Algebra with Applications, 12(5–6):495–514, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bini:2011:SQV

[BMP11] Dario A. Bini, Beatrice Meini, and Federico Poloni. On the solution of a quadratic vector equation arising in Markovian Binary Trees. Numerical Linear Algebra with Applications, 18(6):981–991, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bolten:2017:MPP

[BMS17] Matthias Bolten, Dieter Moser, and Robert Speck. A multigrid perspective on the parallel full approximation scheme in space and time. Numerical Linear Algebra with Applications, 24 (6):??, December 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bolten:2018:ACP

[BMS18] Matthias Bolten, Dieter Moser, and Robert Speck. Asymptotic convergence of the parallel full approximation scheme in space and time for linear problems. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bouyouli:2009:NRC

[BMSS09] R. Bouyouli, G. Meurant, L. Smoch, and H. Sadok. New results on the convergence of the conjugate gradient method. Numerical Linear Algebra with Applications, 16 (3):223–236, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2012:E

[BN12] Zhong-Zhi Bai and Michael K. Ng. Erratum. Numerical Linear Algebra with Applications, 19(5):891, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [BGN07].

Bot:2021:FCP

[BN21] Radu Ioan Bot and Dang-Khoa Nguyen. Factorization of completely positive matrices using iterative projected gradient steps. Numerical Linear Algebra with Applications, 28 (6):e2391:1–e2391:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Berisha:2015:EAP

[BNP15] Sebastian Berisha, James G.

Nagy, and Robert J. Plemmons. Estimation of atmospheric PSF parameters for hyperspectral imaging. Numerical Linear Algebra with Applications, 22(5):795–813, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2018:ENM

[BNR18] Zhong-Zhi Bai, Maya G.

Neytcheva, and Lothar Reichel. Editorial: Novel methods and theories in numerical algebra with interdisciplinary applications. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Behera:2020:FRD

[BNS20] Ratikanta Behera, Ashish Kumar Nandi, and Jajati Keshari Sahoo. Further results on the Drazin inverse of evenorder tensors. Numerical Linear Algebra with Applications, 27(5):e2317:1–e2317:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Beauwens:1994:IUT

[BNT94] R. Beauwens, Y. Notay, and B. Tombuyses. S/P images of upper triangular M-matrices. Numerical Linear Algebra with Applications, 1(1):19–31, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bell:2008:AMF

[BO08] Nathan Bell and Luke N. Olson. Algebraic multigrid for k-form Laplacians. Numerical Linear Algebra with Applications, 15(2-3):165–185, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Borzi:2013:FSS

[BO13] Alfio Borzi and Cornelis W. Oosterlee. Fast solvers for simulation, inversion, and control of wave propagation problems. Numerical Linear Algebra with Applications, 20(4): 539–540, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brenner:2018:MMT

[BO18] Susanne C. Brenner and Duk-Soon Oh. Multigrid methods for H(div) in three dimensions with nonoverlapping domain decomposition smoothers. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Borm:2017:DMC

[Bör17] Steffen Börm. Directional H^{2}- matrix compression for highfrequency problems. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bosse:2019:AMF

[Bos19]  Torsten Bosse.  (almost)

matrix-free solver for piecewise linear functions in absnormal form. Numerical Linear Algebra with Applications, 26(5):e2258:1–e2258:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Botchev:2013:BKS

[Bot13] M. A. Botchev. A block Krylov subspace time-exact solution method for linear ordinary differential equation systems. Numerical Linear Algebra with Applications, 20(4): 557–574, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barreras:2013:ACM

[BP13] A. Barreras and J. M. Peña.

Accurate computations of matrices with bidiagonal decomposition using methods for totally positive matrices. Numerical Linear Algebra with Applications, 20(3):413–424, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bucci:2022:CMC

[BP22] Alberto Bucci and Federico Poloni. A continuation method for computing the multilinear PageRank. Numerical Linear Algebra with Applications, 29 (4):e2432:1–e2432:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barnard:1995:SAE

[BPS95] Stephen T. Barnard, Alex Pothen, and Horst Simon. A spectral algorithm for envelope reduction of sparse matrices. Numerical Linear Algebra with Applications, 2(4):317– 334, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2000:AIP

[BPS00] Luca Bergamaschi, Giorgio Pini, and Flavio Sartoretto. Approximate inverse preconditioning in the parallel solution of sparse eigenproblems. Numerical Linear Algebra with Applications, 7 (3):99–116, April/May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72001232/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72001232&PLACEBO=IE.pdf.

Brenner:2013:BDD

[BPS13] Susanne C. Brenner, Eun-Hee Park, and Li-Yeng Sung. A balancing domain decomposition by constraints preconditioner for a weakly overpenalized symmetric interior penalty method. Numerical Linear Algebra with Applications, 20(3):472–491, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bagci:2015:CAS

[BPS15] Hakan Bağcı, Joseph E. Pasciak, and Kostyantyn Y. Sirenko. A convergence analysis for a sweeping preconditioner for block tridiagonal systems of linear equations. Numerical Linear Algebra with Applications, 22(2): 371–392, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bardsley:2013:KSA

[BPSH13] Johnathan M. Bardsley, Albert Parker, Antti Solonen, and Marylesa Howard. Krylov space approximate Kalman filtering. Numerical Linear Algebra with Applications, 20(2): 171–184, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bey:1999:CBI

[BR99] Jürgen Bey and Arnold Reusken. On the convergence of basic iterative methods for convectiondiffusion equations. Numerical Linear Algebra with Applications, 6(5):329–352, July/August 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/65500099/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=65500099&PLACEBO=IE.pdf.

Baglama:2007:AGT

[BR07] James Baglama and Lothar Reichel. Augmented GMREStype methods. Numerical Linear Algebra with Applications, 14(4):337–350, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Brandts:2002:MCS

[Bra02]  J. H. Brandts.  Matlab

code for sorting real Schur forms. Numerical Linear Algebra with Applications, 9 (3):249–261, April/May 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/90512120/
                                        START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=90512120&PLACEBO=IE.pdf.

Brannick:2021:MM

[Bra21] James Brannick. Multigrid methods. Numerical Linear Algebra with Applications, 28(3): e2359:1–e2359:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bonettini:2007:SIS

[BRT07] Silvia Bonettini, Valeria Ruggiero, and Federica Tinti. On the solution of indefinite systems arising in nonlinear programming problems. Numerical Linear Algebra with Applications, 14(10):807–831, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bennett:2001:EOD

[BS01] Beth Anne V. Bennett and Mitchell D. Smooke. The effect of overall discretization scheme on Jacobian structure, convergence rate, and solution accuracy within the local rectangular refinement method. Numerical Linear Algebra with Applications, 8(8):513–536, December 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88010580/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88010580&PLACEBO=IE.pdf.

Buckeridge:2010:PGM

[BS10] Sean Buckeridge and Robert Scheichl. Parallel geometric multigrid for global weather prediction. Numerical Linear Algebra with Applications, 17 (2-3):325–342, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Barbarino:2020:NHP

[BSC20] Giovanni Barbarino and Stefano Serra-Capizzano. Non-Hermitian perturbations of Hermitian matrix-sequences and applications to the spectral analysis of the numerical approximation of partial differential equations. Numerical Linear Algebra with Applications, 27(3):e2286:1–e2286:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buranay:2017:TCH

[BSI17] Suzan C. Buranay, Dervis Subasi, and Ovgu C. Iyikal. On the two classes of high-order convergent methods of approximate inverse preconditioners for solving linear systems. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Behera:2022:CGI

[BSMN22] Ratikanta Behera, Jajati Keshari Sahoo, Ram N. Mohapatra, and M. Zuhair Nashed.

Computation of generalized inverses of tensors via t-product. Numerical Linear Algebra with Applications, 29(2):e2416:1– e2416:??, March 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bischof:1992:GIC

[BT92] Christian H. Bischof and Ping Tak Peter Tang. Generalizing incremental condition estimation. Journal of Numerical Linear Algebra with Applications, 1(2):149–163, 1992. CODEN ???? ISSN 0129-3281.

Benzi:2003:RIF

[BT03] Michele Benzi and Miroslav Tůma. A robust incomplete factorization preconditioner for positive definite matrices. Numerical Linear Algebra with Applications, 10(5–6):385–400, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blanchard:2015:PCG

[BT15] Jeffrey D. Blanchard and Jared Tanner. Performance comparisons of greedy algorithms in compressed sensing. Numerical Linear Algebra with Applications, 22(2):254–282, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Benner:2013:ODS

[BTT13] Peter Benner, Zoran Tomljanović, and Ninoslav Truhar.

Optimal damping of selected eigenfrequencies using dimension reduction. Numerical Linear Algebra with Applications, 20(1):1–17, January 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Buchholz:2011:BRM

[Buc11] Peter Buchholz. Bounding reward measures of Markov models using the Markov decision processes. Numerical Linear Algebra with Applications, 18 (6):919–930, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bunch:1992:MPL

[Bun92]  James R. Bunch.  Matrix

properties of the Levinson and Schur algorithms. Journal of Numerical Linear Algebra with Applications, 1(2):183– 198, 1992. CODEN ???? ISSN 0129-3281.

Bunch:1995:EDB

[Bun95] James R. Bunch. Editorial: Dedicated to Beresford Parlett and William Kahan. Numerical Linear Algebra with Applications, 2(2):85, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bergamaschi:2000:ECE

[BV00] Luca Bergamaschi and Marco Vianello. Efficient computation of the exponential operator for large, sparse, symmetric matrices. Numerical Linear Algebra with Applications, 7(1):27–45, January/February 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71001528/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71001528&PLACEBO=IE.pdf.

Bouagada:2013:SSS

[BV13] Djillali Bouagada and Paul Van Dooren. On the stability of 2D state-space models. Numerical Linear Algebra with Applications, 20(2): 198–207, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bousse:2018:LSC

[BVD^{+}18] M. Boussé, N. Vervliet, I. Domanov, O. Debals, and L. De Lathauwer. Linear systems with a canonical polyadic decomposition constrained solution: Algorithms and applications. Numerical Linear Algebra with Applications, 25(6): ??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bomhof:2000:PLS

[BvdV00] C. W. Bomhof and H. A.

van der Vorst. A parallel linear system solver for circuit simulation problems. Numerical Linear Algebra with Applications, 7(7–8):649–665, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505475/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505475&PLACEBO=IE.pdf.

Brezina:2012:ICA

[BVV12] Marian Brezina, Petr Vaněk, and Panayot S. Vassilevski. An improved convergence analysis of smoothed aggregation algebraic multigrid. Numerical Linear Algebra with Applications, 19(3):441–469, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Batselier:2017:CAD

[BW17a] Kim Batselier and Ngai Wong. A constructive arbitrary-degree Kronecker product decomposition of tensors. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Blonigan:2017:MTS

[BW17b] Patrick Blonigan and Qiqi Wang. Multigrid-in-time for sensitivity analysis of chaotic dynamical systems. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2019:GRC

[BW19] Zhong-Zhi Bai and Wen-Ting Wu. On greedy randomized coordinate descent methods for solving large linear least-squares problems. Numerical Linear Algebra with Applications, 26(4):e2237:1– e2237:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bertaccini:2005:EPM

[BWN05] Daniele Bertaccini, Youwei Wen, and Michael K. Ng. The eigenvalues of preconditioned matrices for linear multistep formulas in boundary value form. Numerical Linear Algebra with Applications, 12(2– 3):315–325, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2013:MBS

[BZ13] Zhong-Zhi Bai and Li-Li Zhang. Modulus-based synchronous multisplitting iteration methods for linear complementarity problems. Numerical Linear Algebra with Applications, 20(3):425–439, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bai:2017:MBM

[BZ17] Zhong-Zhi Bai and Li-Li Zhang. Modulus-based multigrid methods for linear complementarity problems. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Budzinskiy:2023:TTC

[BZ23] Stanislav Budzinskiy and Nikolai Zamarashkin. Tensor train completion: Local recovery guarantees via Riemannian optimization. Numerical Linear Algebra with Applications, 30 (6):e2520:1–e2520:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cihlar:1999:NSN

[CA99] Jan Cihlář and Philippe Angot. Numerical solution of Navier–Stokes systems. Numerical Linear Algebra with Applications, 6(1):17–27, January/February 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002984;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002984&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part I (Milovy, 1997).

Cao:2004:FUA

[Cao04] Zhi-Hao Cao. Fast Uzawa algorithms for solving nonsymmetric stabilized saddle point problems. Numerical Linear Algebra with Applications, 11(1):1–24, February 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2008:ABP

[Cao08] Zhi-Hao Cao. Augmentation block preconditioners for saddle point-type matrices with singular (1, 1) blocks. Numerical Linear Algebra with Applications, 15(6):515–533, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2009:NSD

[Cao09] Zhi-Hao Cao. A note on spectrum distribution of constraint preconditioned generalized saddle point matrices. Numerical Linear Algebra with Applications, 16(6): 503–516, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2013:CPM

[Cao13] Zhi-Hao Cao. Comment on ‘Preconditioning of matrices partitioned in 2 × 2 block form: Eigenvalue estimates and Schwarz DD for mixed FEM’ by Owe Axelsson, Radim Blaheta. Numerical Linear Algebra with Applications, 20(3):

533–535, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [AB10, AB13].

Carstensen:1997:DDN

[Car97]  Carsten Carstensen.  Do-

main decomposition for a nonsmooth convex minimization problem and its application to plasticity. Numerical Linear Algebra with Applications, 4 (3):177–190, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15027;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15027&PLACEBO=IE.pdf.

Caswell:2011:PAC

[Cas11] Hal Caswell. Perturbation analysis of continuous-time absorbing Markov chains. Numerical Linear Algebra with Applications, 18(6):901–917, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Claus:2021:NBS

[CB21] Lisa Claus and Matthias Bolten. Nonoverlapping block smoothers for the Stokes equations. Numerical Linear Algebra with Applications, 28 (6):e2389:1–e2389:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Coley:2018:GMM

[CBE18] Christopher Coley, Joseph Benzaken, and John A. Evans. A geometric multigrid method for isogeometric compatible discretizations of the generalized Stokes and Oseen problems. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1992:CPE

[CC92] Raymond H. Chan and Tony F.

Chan. Circulant preconditioners for elliptic problems. Journal of Numerical Linear Algebra with Applications, 1(1):77– 101, 1992. CODEN ???? ISSN 0129-3281.

Chien:2003:ALA

[CC03] C.-S. Chien and S.-L. Chang.

Application of the Lanczos algorithm for solving the linear systems that occur in continuation problems. Numerical Linear Algebra with Applications, 10(4):335–355, June 2003. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2007:DRG

[CC07] Jing Cai and Guoliang Chen.

On determinantal representation for the generalized inverse A and its applications. Numerical Linear Algebra with Applications, 14(3):169–182, ???? 2007. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Clark:2020:SCM

[CC20] Gregory J. Clark and Joshua N. Cooper. Stably computing the multiplicity of known roots given leading coefficients. Numerical Linear Algebra with Applications, 27(2):e2275:1– e2275:??, March 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2018:SSM

[CCE^{+}18] Difeng Cai, Edmond Chow, Lucas Erlandson, Yousef Saad, and Yuanzhe Xi. SMASH: Structured matrix approximation by separation and hierarchy. Numerical Linear Algebra with Applications, 25(6): ??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chaitin-Chatelin:2000:CNA

[CCG00] F. Chaitin-Chatelin and S. Gratton. On the condition numbers associated with the polar factorization of a matrix. Numerical Linear Algebra with Applications, 7(5): 337–354, July/August 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72508408/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72508408&PLACEBO=IE.pdf.

Chen:2006:CMC

[CCK06] Zhangxin Chen, So-Hsiang Chou, and Do Young Kwak. Characteristic-mixed covolume methods for advectiondominated diffusion problems. Numerical Linear Algebra with Applications, 13(9): 677–697, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:2005:SMA

[CCLN05] Raymond Chan, Wai-Ki Ching, Franklin Luk, and Michael Ng. Structured matrices with applications. Numerical Linear Algebra with Applications, 12(2– 3):87, March/April 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2018:SPA

[CCLQ18] Haibin Chen, Yannan Chen, Guoyin Li, and Liqun Qi. A semidefinite program approach for computing the maximum eigenvalue of a class of structured tensors and its applications in hypergraphs and copositivity test. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Calgaro:2010:IIL

[CCS10] Caterina Calgaro, Jean-Paul Chehab, and Yousef Saad. Incremental incomplete LU factorizations with applications. Numerical Linear Algebra with Applications, 17(5): 811–837, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2019:SGJ

[CCS19] Yunfeng Cai, Guanghui Cheng, and Decai Shi. Solving the general joint block diagonalization problem via linearly independent eigenvectors of a matrix polynomial. Numerical Linear Algebra with Applications, 26 (4):e2238:1–e2238:??, August 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chaitin-Chatelin:2006:APQ

[CCvG06] F. Chaitin-Chatelin and M. B. van Gijzen. Analysis of parameterized quadratic eigenvalue problems in computational acoustics with homotopic deviation theory. Numerical Linear Algebra with Applications, 13(6):487–512, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2023:DDL

[CCX23] Difeng Cai, Edmond Chow, and Yuanzhe Xi. Data-driven linear complexity low-rank approximation of general kernel matrices: a geometric approach. Numerical Linear Algebra with Applications, 30 (6):e2519:1–e2519:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Carvalho:2011:NAG

[CD11] João B. Carvalho and Biswa N.

Datta. A new algorithm for generalized Sylvester-observer equation and its application to state and velocity estimations in vibrating systems. Numerical Linear Algebra with Applications, 18(4):719–732, August 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cavoretto:2012:SAP

[CDDSC12] R. Cavoretto, A. De Rossi, M. Donatelli, and S. Serra-Capizzano. Spectral analysis and preconditioning techniques for radial basis function collocation matrices. Numerical Linear Algebra with Applications, 19(1):31–52, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cipolla:2019:RPC

[CDDZ19] Stefano Cipolla, Carmine Di Fiore, Fabio Durastante, and Paolo Zellini. Regularizing properties of a class of matrices including the optimal and the superoptimal preconditioners. Numerical Linear Algebra with Applications, 26(2): e2225:1–e2225:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Carpentieri:2000:SPS

[CDG00] B. Carpentieri, I. S. Duff, and L. Giraud. Sparse pattern selection strategies for robust Frobenius-norm minimization preconditioners in electromagnetism. Numerical Linear Algebra with Applications, 7(7–8):667–685, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505479/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505479&PLACEBO=IE.pdf.

Carpentieri:2004:SSP

[CDGmM04] B. Carpentieri, I. S. Duff, L. Giraud, and M. Magolu monga Made. Sparse symmetric preconditioners for dense linear systems in electromagnetism. Numerical Linear Algebra with Applications, 11(8–9):753–771, October/November 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cucker:2006:SAS

[CDW06] F. Cucker, H. Diao, and Y. Wei. Smoothed analysis of some condition numbers. Numerical Linear Algebra with Applications, 13(1): 71–84, February 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:1996:MPM

[CEL^{+}96] Zhangxin Chen, Richard E.

Ewing, Raytcho D. Lazarov, Serguei Maliassov, and Yuri A. Kuznetsov. Multilevel preconditioners for mixed methods for second order elliptic problems. Numerical Linear Algebra with Applications, 3(5): 427–453, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001001.

Canha:2007:MMD

[CEQN07] L. Canha, P. Ekel, J. Queiroz, and F. Schuffner Neto. Models and methods of decision making in fuzzy environment and their applications to power engineering problems. Numerical Linear Algebra with Applications, 14(4):369–390, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2016:PTI

[CFAM16] Ke Chen, Faisal Fairag, and Adel Al-Mahdi. Preconditioning techniques for an image deblurring problem. Numerical Linear Algebra with Applications, 23(3):570–584, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chu:2005:CMR

[CfX05] Moody T. Chu and Shu fang Xu. On computing minimal realizable spectral radii of nonnegative matrices. Numerical Linear Algebra with Applications, 12(1):77–86, February 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2015:AAB

[CG15] Meng-Huo Chen and Anne Greenbaum. Analysis of an aggregation-based algebraic two-grid method for a rotated anisotropic diffusion problem. Numerical Linear Algebra with Applications, 22(4): 681–701, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2021:HSR

[CGJ21] Qiao Chen, Aditi Ghai, and Xiangmin Jiao. HILUCSI: Simple, robust, and fast multilevel ILU for large-scale saddlepoint problems from PDEs. Numerical Linear Algebra with Applications, 28(6):e2400:1– e2400:??, December 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:1994:CSD

[CGK94] Xiao-Chuan Cai, William D. Gropp, and David E. Keyes. A comparison of some domain decomposition and ILU preconditioned iterative methods for nonsymmetric elliptic problems. Numerical Linear Algebra with Applications, 1 (5):477–504, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chang:2005:APQ

[CGK05] Xiao-Wen Chang, Martin J.

Gander, and Samir Karaa. Asymptotic properties of the QR factorization of banded Hessenberg–Toeplitz matrices. Numerical Linear Algebra with Applications, 12(7):659–682, September 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chatzipantelidis:2005:FVE

[CGL05] P. Chatzipantelidis, V. Ginting, and R. D. Lazarov. A finite volume element method for a non-linear elliptic problem. Numerical Linear Algebra with Applications, 12(5–6):515–546, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Carvalho:2001:LPT

[CGM01] L. M. Carvalho, L. Giraud, and G. Meurant. Local preconditioners for two-level nonoverlapping domain decomposition methods. Numerical Linear Algebra with Applications, 8(4):207–227, June 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76509568/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=76509568&PLACEBO=IE.pdf.

Chang:2011:GEB

[CGM11] Britton Chang, Anne Greenbaum, and Eric Machorro. Global error bounds for the Petrov–Galerkin discretization of the neutron transport equation. Numerical Linear Algebra with Applications, 18(1): 141–154, January 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Calandra:2013:ITG

[CGPV13] Henri Calandra, Serge Gratton, Xavier Pinel, and Xavier Vasseur. An improved two-grid preconditioner for the solution of three-dimensional Helmholtz problems in heterogeneous media. Numerical Linear Algebra with Applications, 20(4): 663–688, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Choudhary:2020:ANS

[CGS20] Neelam Choudhary, Nicolas Gillis, and Punit Sharma. On approximating the nearest Ωstable matrix. Numerical Linear Algebra with Applications, 27(3):e2282:1–e2282:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Carson:2022:MPS

[CGY22] Erin Carson, Tomás Gergelits, and Ichitaro Yamazaki. Mixed precision s-step Lanczos and conjugate gradient algorithms. Numerical Linear Algebra with Applications, 29(3):e2425:1– e2425:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1994:LRR

[CH94] Tony F. Chan and Per Christian Hansen. Low-rank revealing QR factorizations. Numerical Linear Algebra with Applications, 1(1):33–44, 1994. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2003:NVS

[CH03] Xiaojun Chen and Kouji Hashimoto. Numerical validation of solutions of saddle point matrix equations. Numerical Linear Algebra with Applications, 10(7):661–672, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2005:SIM

[CH05] Xuzhou Chen and Robert E.

Hartwig. The semi-iterative method applied to the hyperpower iteration. Numerical Linear Algebra with Applications, 12(9):895–910, November 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Clevenger:2021:CBA

[CH21] Thomas C. Clevenger and Timo Heister. Comparison between algebraic and matrixfree geometric multigrid for a Stokes problem on adaptive meshes with variable viscosity. Numerical Linear Algebra with Applications, 28(5):e2375:1– e2375:??, October 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chang:2007:CGM

[Cha07] B. Chang. The conjugate gradient method solves the neutron transport equation hoptimally. Numerical Linear Algebra with Applications, 14 (10):751–769, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chang:2012:PFF

[Cha12] X.-W. Chang. On the perturbation of the Q-factor of the QR factorization. Numerical Linear Algebra with Applications, 19(3):607–619, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cui:2022:RQG

[CHCS22] Lu-Bin Cui, Qing Hu, Ying Chen, and Yi-Sheng Song. A Rayleigh quotient-gradient neural network method for computing Z-eigenpairs of general tensors. Numerical Linear Algebra with Applications, 29(3):e2420:1–e2420:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2002:CSM

[Che02] Xiaojun Chen. On convergence of SOR methods for nonsmooth equations. Numerical Linear Algebra with Applications, 9 (1):81–92, January/February 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013648/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013648&PLACEBO=IE.pdf.

Chen:2015:GKP

[Che15] Hao Chen. Generalized Kronecker product splitting iteration for the solution of implicit Runge–Kutta and boundary value methods. Numerical Linear Algebra with Applications, 22(2):357–370, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chow:2003:UMM

[Cho03] Edmond Chow. An unstructured multigrid method based on geometric smoothness. Numerical Linear Algebra with Applications, 10(5–6):401–421, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chu:2004:CIO

[Chu04] Delin Chu. On the computation of the infimum in H-optimization. Numerical Linear Algebra with Applications, 11(7):619–648, September 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Codevico:2005:SSR

[CHV05] G. Codevico, G. Heinig, and M. Van Barel. A superfast solver for real symmetric Toeplitz systems using real trigonometric transformations. Numerical Linear Algebra with Applications, 12(8): 699–713, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chien:2008:TIA

[CJL08] C.-S. Chien, B.-W. Jeng, and Z.-C. Li. A time-independent approach for computing wave functions of the Schrödinger– Poisson system. Numerical Linear Algebra with Applications, 15(1):55–82, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cullum:2003:EPD

[CJT03] Jane K. Cullum, Keith Johnson, and Miroslav Tůma. Effects of problem decomposition (partitioning) on the rate of convergence of parallel numerical algorithms. Numerical Linear Algebra with Applications, 10(5–6):445–465, July/ September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cheng:2006:SPS

[CJW06] Che-Man Cheng, Xiao-Qing Jin, and Yi-Min Wei. Stability properties of superoptimal preconditioner from numerical range. Numerical Linear Algebra with Applications, 13(7): 513–521, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2011:SPS

[CJZ11] Yang Cao, Mei-Qun Jiang, and Ying-Long Zheng. A splitting preconditioner for saddle point problems. Numerical Linear Algebra with Applications, 18(5):875–895, October 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chronopoulos:2001:OIM

[CK01] Anthony T. Chronopoulos and David Kincaid. On the Odir iterative method for non-symmetric indefinite linear systems. Numerical Linear Algebra with Applications, 8(2):71–82, March 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501757/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=76501757&PLACEBO=IE.pdf.

Chronopoulos:2010:BSK

[CK10] Anthony T. Chronopoulos and Andrey B. Kucherov. Block s-step Krylov iterative methods. Numerical Linear Algebra with Applications, 17(1): 3–15, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cvetkovic:2014:NCM

[CK14] Ljiljana Cvetković and Vladimir Kostić. A note on the convergence of the MSMAOR method for linear complementarity problems. Numerical Linear Algebra with Applications, 21(4):534–539, August 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2002:FEM

[CKW02] Zhiqiang Cai, Seokchan Kim, and Gyungsoo Woo. A finite element method using singular functions for the Poisson equation: crack singularities. Numerical Linear Algebra with Applications, 9(6–7):445–455, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:1996:CEM

[CL96] Zhiqiang Cai and Chen-Yao G. Lai. Convergence estimates of multilevel additive and multiplicative algorithms for nonsymmetric and indefinite problems. Numerical Linear Algebra with Applications, 3 (3):205–220, May/June 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000990.

Chen:2013:SAG

[CL13] Xiao Shan Chen and Wen Li.

Sensitivity analysis for the generalized singular value decomposition. Numerical Linear Algebra with Applications, 20(1): 138–149, January 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2023:PTF

[CL23] Yuhan Chen and Chenliang Li.

Preconditioned tensor format conjugate gradient squared and biconjugate gradient stabilized methods for solving Stein tensor equations. Numerical Linear Algebra with Applications, 30(5):e2502:1–e2502:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2011:PAS

[CLC11] Xiao Shan Chen, Wen Li, and Wai-Ki Ching. Perturbation analysis for the sign functions of regular matrix pairs. Numerical Linear Algebra with Applications, 18(2):189–203, March 2011. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-

1506 (electronic).

Chen:2015:TVB

[CLNY15] Chuan Chen, Xutao Li, Michael K. Ng, and Xiaoming Yuan. Total variation based tensor decomposition for multidimensional data with time dimension. Numerical Linear Algebra with Applications, 22(6): 999–1019, December 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cui:2023:ACU

[CLQY23] Chunfeng Cui, Ziyan Luo, Liqun Qi, and Hong Yan. The analytic connectivity in uniform hypergraphs: Properties and computation. Numerical Linear Algebra with Applications, 30(2):e2468:1– e2468:??, March 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Calvetti:2001:SLS

[CLR01] D. Calvetti, B. Lewis, and L. Reichel. On the solution of large Sylvester-observer equations. Numerical Linear Algebra with Applications, 8(6–7):435–451, September/November 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/85007287/
                                        START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=85007287&PLACEBO=IE.pdf.

Chu:2013:SID

[CLR13] E. K-W. Chu, W.-W. Lin, and L. Reichel. Special issue dedicated to Biswa Nath Datta on the occasion of his 70th birthday. Numerical Linear Algebra with Applications, 20(2): 167–170, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chu:2011:CNP

[CLTW11] Delin Chu, Lijing Lin, Roger C. E. Tan, and Yimin Wei. Condition numbers and perturbation analysis for the Tikhonov regularization of discrete ill-posed problems. Numerical Linear Algebra with Applications, 18(1):87–103, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2019:BEA

[CMSW19] Hongjia Chen, Jie Meng, Tetsuya Sakurai, and Xiang Wang. Backward error analysis for linearizations in heavily damped quadratic eigenvalue problem. Numerical Linear Algebra with Applications, 26(4):e2253:1– e2253:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2021:CCA

[CN21] Yang Cao and Maya Neytcheva. Cell-by-cell approximate Schur complement technique in preconditioning of meshfree discretized piezoelectric equations. Numerical Linear Algebra with Applications, 28 (4):e2362:1–e2362:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1996:GSP

[CNP96] Raymond H. Chan, Michael K. Ng, and Robert J. Plemmons. Generalization of Strang’s preconditioner with applications to Toeplitz least squares problems. Numerical Linear Algebra with Applications, 3(1): 45–64, January/February 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000497.

Ching:2005:SRI

[CNSY05] Wai-Ki Ching, Michael K. Ng, Kenton N. Sze, and Andy C. Yau. Super-resolution image reconstruction using multisensors. Numerical Linear Algebra with Applications, 12(2– 3):271–281, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2007:NEC

[CNT07] Xing Cai, Bjørn Fredrik Nielsen, and Aslak Tveito. A note on the efficiency of the conjugate gradient method for a class of time-dependent problems. Numerical Linear Algebra with Applications, 14(5): 459–467, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2020:SFI

[CNXY20] Zixuan Chen, James G. Nagy, Yuanzhe Xi, and Bo Yu. Structured FISTA for image restoration. Numerical Linear Algebra with Applications, 27(2): e2278:1–e2278:??, March 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ching:2005:DMS

[CNY05] Wai-Ki Ching, Michael K. Ng, and Wai-On Yuen. A direct method for solving block-Toeplitz with near-circulant-block systems with applications to hybrid manufacturing systems. Numerical Linear Algebra with Applications, 12 (10):957–966, December 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2017:BSC

[CNZ17] Chuan Chen, Michael K. Ng, and Shuqin Zhang. Block spectral clustering methods for multiple graphs. Numerical Linear Algebra with Applications, 24(1):??, January 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Coroian:2004:IOI

[Cor04] Dan I. Coroian. Improved outer inverses for the numerical solution of Euler–Lagrange equations. Numerical Linear Algebra with Applications, 11 (10):853–866, December 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Climent:1999:CCT

[CP99] Joan-Josep Climent and Carmen Perea. Convergence and comparison theorems for multisplittings. Numerical Linear Algebra with Applications, 6(2):93–107, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002991;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002991&
                                        PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Copeland:2006:LSM

[CP06] Dylan M. Copeland and Joseph E. Pasciak. A leastsquares method for axisymmetric div–curl systems. Numerical Linear Algebra with Applications, 13(9):733–752, ???? 2006. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Chun:2012:EAC

[CP12] Joohwan Chun and Jaehyun Park. An efficient algorithm for clustered integer least squares problems. Numerical Linear Algebra with Applications, 19 (1):19–30, January 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:2001:PNH

[CPS01] Raymond H. Chan, Daniel Potts, and Gabriele Steidl. Preconditioners for non-Hermitian Toeplitz systems. Numerical Linear Algebra with Applications, 8(2):83–98, March 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501758/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501758&PLACEBO=IE.pdf.

Castillo:2006:ILS

[CPSM06] E. Castillo, R. E. Pruneda, C. Solares, and R. Mínguez. Interpreting linear systems of equalities and inequalities. Application to the water supply problem. Numerical Linear Algebra with Applications, 13(5): 361–397, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2010:GSC

[CQ10] Jinhai Chen and Liqun Qi. Globally and superlinearly convergent inexact Newton– Krylov algorithms for solving nonsmooth equations. Numerical Linear Algebra with Applications, 17(1):155–174, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:2011:FNM

[CQX11] Yun-Feng Cai, Jiang Qian, and Shu-Fang Xu. The formulation and numerical method for partial quadratic eigenvalue assignment problems. Numerical Linear Algebra with Applications, 18(4):637–652, August 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chang:2013:SST

[CQZ13] Kungching Chang, Liqun Qi, and Tan Zhang. A survey on the spectral theory of nonnegative tensors. Numerical Linear Algebra with Applications, 20(6):891–912, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Campos:2016:PIR

[CR16] Carmen Campos and Jose E. Roman. Parallel iterative refinement in polynomial eigenvalue problems. Numerical Linear Algebra with Applications, 23(4):730–745, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Campos:2020:IBS

[CR20] Carmen Campos and Jose E.

Roman. Inertia-based spectrum slicing for symmetric quadratic eigenvalue problems. Numerical Linear Algebra with Applications, 27(4):e2293:1– e2293:??, August 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Calvetti:2005:TRL

[CRS05] D. Calvetti, L. Reichel, and A. Shuibi. Tikhonov regularization of large symmetric problems. Numerical Linear Algebra with Applications, 12(2–3):127–139, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cools:2014:NLD

[CRV14] Siegfried Cools, Bram Reps, and Wim Vanroose. A new level-dependent coarse grid correction scheme for indefinite Helmholtz problems. Numerical Linear Algebra with Applications, 21(4):513–533, August 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cipolla:2020:EMF

[CRZT20] Stefano Cipolla, Michela Redivo-Zaglia, and Francesco Tudisco. Extrapolation methods for fixed-point multilinear PageRank computations. Numerical Linear Algebra with Applications, 27(2):e2280:1– e2280:??, March 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Concus:1995:MDP

[CS95] Paul Concus and Paul Saylor. A modified direct preconditioner for indefinite symmetric Toeplitz systems. Numerical Linear Algebra with Applications, 2(5):415–429, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Special issue dedicated to David M. Young, Jr.

Cai:1996:ODD

[CS96] Xiao-Chuan Cai and Yousef Saad. Overlapping domain decomposition algorithms for general sparse matrices. Numerical Linear Algebra with Applications, 3 (3):221–237, May/June 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000988.

Chapman:1997:DAK

[CS97] Andrew Chapman and Yousef Saad. Deflated and augmented Krylov subspace techniques. Numerical Linear Algebra with Applications, 4(1): 43–66, January/February 1997.

CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001011;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=15001011&
PLACEBO=IE.pdf.

Chan:2002:SCM

[CS02] Tony F. Chan and Ilya Sharapov. Subspace correction multi-level methods for elliptic eigenvalue problems. Numerical Linear Algebra with Applications, 9(1): 1–20, January/February 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013647/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013647&PLACEBO=IE.pdf.

Cao:2009:PMM

[CS09] Guangxi Cao and Yongzhong Song. On parallel multisplitting methods for symmetric positive semidefinite linear systems. Numerical Linear Algebra with Applications, 16(4): 301–318, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cao:2011:SPM

[CS11] Guangxi Cao and Yongzhong Song. Semiconvergence of parallel multisplitting methods for symmetric positive semidefinite linear systems. Numerical Linear Algebra with Applications, 18(3):317–324, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2018:PEE

[CS18] Jie Chen and Yousef Saad. A posteriori error estimate for computing tr(f(A)) by using the Lanczos method. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cabral:2020:IGU

[CSB20] Juan C. Cabral, Christian E. Schaerer, and Amit Bhaya. Improving GMRES(m) using an adaptive switching controller. Numerical Linear Algebra with Applications, 27 (5):e2305:1–e2305:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:2005:TGM

[CSCTP05] R. H. Chan, S. Serra-Capizzano, and C. Tablino-Possio. Twogrid methods for banded linear systems from DCT III algebra. Numerical Linear Algebra with Applications, 12(2–3):241–249, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cheng:2014:BER

[CSYS14] Guanghui Cheng, Zhida Song, Jianfeng Yang, and Jia Si. The bounds of the eigenvalues for rank-one modification of Hermitian matrix. Numerical Linear Algebra with Applications, 21(1):98–107, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chu:2021:NSS

[CSZ21] Eric K.-W. Chu, Daniel B.

Szyld, and Jieyong Zhou. Numerical solution of singular Lyapunov equations. Numerical Linear Algebra with Applications, 28(5):e2381:1– e2381:??, October 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chang:2009:BPA

[CTP09] X.-W. Chang and D. Titley-Peloquin. Backward perturbation analysis for scaled total least-squares problems. Numerical Linear Algebra with Applications, 16(8):627–648, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chow:2003:MBF

[CV03] Edmond Chow and Panayot S.

Vassilevski. Multilevel block factorizations in generalized hierarchical bases. Numerical Linear Algebra with Applications, 10(1–2):105–127, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cools:2013:LFA

[CV13] Siegfried Cools and Wim Vanroose. Local Fourier analysis of the complex shifted Laplacian preconditioner for Helmholtz problems. Numerical Linear Algebra with Applications, 20 (4):575–597, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cvetkovic:2009:P

[Cve09] Ljiljana Cvetković. Preface. Numerical Linear Algebra with Applications, 16(11-12):861– 862, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Collignon:2011:MS

[CvG11] Tijmen P. Collignon and Martin B. van Gijzen. Minimizing synchronization in IDR(s). Numerical Linear Algebra with Applications, 18(5): 805–825, October 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Camps:2021:QFT

[CVY21] Daan Camps, Roel Van Beeumen, and Chao Yang. Quantum Fourier transform revisited. Numerical Linear Algebra with Applications, 28 (1):e2331:1–e2331:??, January 2021. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Chan:1997:STB

[CW97] Raymond H. Chan and C. K.

Wong. Sine transform based preconditioners for elliptic problems. Numerical Linear Algebra with Applications, 4(5):351–368, September/October 1997. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15041;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15041&PLACEBO=IE.pdf.

Chien:1997:LTM

[CWS97] C.-S. Chien, Z.-L. Weng, and C.-L. Shen. Lanczos-type methods for continuation problems. Numerical Linear Algebra with Applications, 4(1): 23–41, January/February 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001012;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=15001012&
PLACEBO=IE.pdf.

Chen:2018:TSN

[CWwS18] Xiao Shan Chen, Chao Tao Wen, and Hai wei Sun. Twostep Newton-type methods for solving inverse eigenvalue problems. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Che:2022:FRT

[CWWZ22] Maolin Che, Xuezhong Wang, Yimin Wei, and Xile Zhao. Fast randomized tensor singular value thresholding for lowrank tensor optimization. Numerical Linear Algebra with Applications, 29(6):e2444:1– e2444:??, December 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2022:TSR

[CX22] Xiao Shan Chen and Hongguo Xu. Two-sided Rayleigh quotient iteration for periodic eigenvalue problems. Numerical Linear Algebra with Applications, 29(3):e2422:1– e2422:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2023:QATa

[CX23a] Xiao Shan Chen and Hongguo Xu. QZ algorithm with twosided generalized Rayleigh quotient shifts. Numerical Linear Algebra with Applications, 30 (4):e2475:1–e2475:??, August 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2023:QATb

[CX23b] Xiao Shan Chen and Hongguo Xu. QR algorithm with two-sided Rayleigh quotient shifts. Numerical Linear Algebra with Applications, 30 (5):e2487:1–e2487:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cai:1999:LSF

[CYZ99] Zhiqiang Cai, Xiu Ye, and Huilong Zhang. Least-squares finite element approximations for the Reissner–Mindlin plate. Numerical Linear Algebra with Applications, 6 (6):479–496, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/67501480/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501480&PLACEBO=IE.pdf.

Cai:2002:SOC

[CZ02] Xiao-Chuan Cai and Jun Zou.

Some observations on the l 2 convergence of the additive Schwarz preconditioned GM-RES method. Numerical Linear Algebra with Applications, 9(5):379–397, July/August 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Cardoso:2015:SSP

[CZ15] João R. Cardoso and Krystyna Ziȩtak. On a sub-Stiefel Procrustes problem arising in computer vision. Numerical Linear Algebra with Applications, 22(3):523–547, May 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Chen:2022:TSP

[CZS22] Hongjia Chen, Ke Zhang, and Tetsuya Sakurai. The tropical scaling for the polynomial eigenvalue problem solved by a contour integral method. Numerical Linear Algebra with Applications, 29(2):e2413:1– e2413:??, March 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Derakhshan:2021:IMS

[DA21] Mohammadhossein Derakhshan and Azim Aminataei. An iterative method for solving fractional diffusion-wave equation involving the Caputo– Weyl fractional derivative. Numerical Linear Algebra with Applications, 28(2):e2345:1– e2345:??, March 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dahlke:2002:BRE

[Dah02] Stephan Dahlke. Besov regularity of edge singularities for the Poisson equation in polyhedral domains. Numerical Linear Algebra with Applications, 9(6–7):457–466, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Damm:2008:DMA

[Dam08]  T. Damm.  Direct meth-

ods and ADI-preconditioned Krylov subspace methods for generalized Lyapunov equations. Numerical Linear Algebra with Applications, 15(9): 853–871, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Datta:2001:PSI

[Dat01] Biswa Nath Datta. Preface for the special issue on numerical linear algebra techniques for control and signal processing. Numerical Linear Algebra with Applications, 8(6–7):355–356, September/November 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007285/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85007285&PLACEBO=IE.pdf.

Dax:1994:RRM

[Dax94] Achiya Dax. A row relaxation method for large l_{p} least norm problems. Numerical Linear Algebra with Applications, 1 (3):247–263, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dax:2004:CPH

[Dax04] Achiya Dax. Computing projections via Householder transformations and Gram–Schmidt orthogonalizations. Numerical Linear Algebra with Applications, 11(7):675–692, September 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dax:2019:RKM

[Dax19]  Achiya Dax.  A restarted

Krylov method with inexact inversions. Numerical Linear Algebra with Applications, 26(1): ??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Deng:2006:IOD

[DBG06] Yuan-Bei Deng, Zhong-Zhi Bai, and Yong-Hua Gao. Iterative orthogonal direction methods for Hermitian minimum norm solutions of two consistent matrix equations. Numerical Linear Algebra with Applications, 13(10):801–823, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Duan:2016:LRS

[DBLP16] Xue-Feng Duan, Jian-Chao Bai, Jiao-Fen Li, and Jing-Jing Peng. On the low rank solution of the Q-weighted nearest correlation matrix problem. Numerical Linear Algebra with Applications, 23(2): 340–355, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rodrigues:2019:EGM

[dCSRS19] Elisa de Cássia Silva Rodrigues and Jorge Stolfi. ECLES: A general method for local editing of parameters with linear constraints. Numerical Linear Algebra with Applications, 26 (6):e2268:1–e2268:??, December 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Druinsky:2018:WIR

[DCT18] Alex Druinsky, Eyal Carlebach, and Sivan Toledo. Wilkinson’s inertia-revealing factorization and its application to sparse matrices. Numerical Linear Algebra with Applications, 25(2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Domoradova:2007:PPP

[DD07] Marta Domorádová and Zdeněk Dostál. Projector preconditioning for partially boundconstrained quadratic optimization. Numerical Linear Algebra with Applications, 14 (10):791–806, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dayde:1999:SSP

[DDG99] Michel J. Daydé, Jérôme P.

Décamps, and Nicholas I. M. Gould. Subspace-by-subspace preconditioners for structured linear systems. Numerical Linear Algebra with Applications, 6(3):213–234, April/May 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        63003630;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=63003630&
                                        PLACEBO=IE.pdf.

Darrigrand:2023:IIO

[DDKR23] Vincent Darrigrand, Andrei Dumitrasc, Carola Kruse, and Ulrich Rüde. Inexact inner– outer Golub–Kahan bidiagonalization method: a relaxation strategy. Numerical Linear Algebra with Applications, 30(5):e2484:1–e2484:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Duan:2021:EAS

[DDL^{+}21] Xue-Feng Duan, Shan-Qi Duan, Juan Li, Jiao fen Li, and Qing-Wen Wang. An efficient algorithm for solving the nonnegative tensor least squares problem. Numerical Linear Algebra with Applications, 28 (6):e2385:1–e2385:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dessole:2023:LHA

[DDM23] Monica Dessole, Marco Dell’Orto, and Fabio Marcuzzi. The Lawson–Hanson algorithm with deviation maximization: Finite convergence and sparse recovery. Numerical Linear Algebra with Applications, 30 (5):e2490:1–e2490:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2013:SDP

[De 13]  Hans De Sterck.  Steep-

est descent preconditioning for nonlinear GMRES optimization. Numerical Linear Algebra with Applications, 20(3): 453–471, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dax:1998:AMN

[DE98] Achiya Dax and Lars Eldén.

Approximating minimum norm solutions of rank-deficient least squares problems. Numerical Linear Algebra with Applications, 5(2):79–99, March/April 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5964;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5964&PLACEBO=IE.pdf.

Du:2006:ASC

[DE06] Qiang Du and Maria Emelianenko. Acceleration schemes for computing centroidal Voronoi tessellations. Numerical Linear Algebra with Applications, 13(2–3):173–192, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeGuchy:2018:CRF

[DEM18] Omar DeGuchy, Jennifer B. Erway, and Roummel F. Marcia. Compact representation of the full Broyden class of quasi-Newton updates. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dembele:2021:MCP

[Dem21] Doulaye Dembélé. A method for computing the Perron root for primitive matrices. Numerical Linear Algebra with Applications, 28(1):e2340:1– e2340:??, January 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Deng:2009:IO

[Den09] Chun Yuan Deng. On the invertibility of the operator A − XB. Numerical Linear Algebra with Applications, 16(10): 817–831, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dendy:2012:MM

[Den12] J. E. Dendy, Jr. Multigrid methods. Numerical Linear Algebra with Applications, 19(2): 175–177, March 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dendy:2014:MM

[Den14]  J. E. Dendy, Jr.  Multi-

grid methods 2013. Numerical Linear Algebra with Applications, 21(2):175–176, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dendy:2018:MM

[Den18] Joel E. Dendy, Jr. Multigrid Methods 2017. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dieci:2001:CIS

[DF01] L. Dieci and M. J. Friedman.

Continuation of invariant subspaces. Numerical Linear Algebra with Applications, 8(5): 317–327, July/August 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/82004029/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=82004029&PLACEBO=IE.pdf.

daFonseca:2020:ESP

[dF20] Carlos M. da Fonseca. Eigenpairs of some particular band Toeplitz matrices: A comment. Numerical Linear Algebra with Applications, 27 (1):e2270:1–e2270:??, January 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dumbser:2018:SDG

[DFF^{+}18] M. Dumbser, F. Fambri, I. Furci, M. Mazza, S. Serra-Capizzano, and M. Tavelli. Staggered discontinuous Galerkin methods for the incompressible Navier–Stokes equations: Spectral analysis and computational results. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2021:OMR

[DFF^{+}21a] Hans De Sterck, Robert D. Falgout, Stephanie Friedhoff, Oliver A. Krzysik, and Scott P. MacLachlan. Optimizing multigrid reduction-in-time and parareal coarse-grid operators for linear advection. Numerical Linear Algebra with Applications, 28(4):e2367:1– e2367:??, August 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2021:MMB

[DFF^{+}21b] Marco Donatelli, Paola Ferrari, Isabella Furci, Stefano Serra-Capizzano, and Debora Sesana. Multigrid methods for block-Toeplitz linear systems: convergence analysis and applications. Numerical Linear Algebra with Applications, 28 (4):e2356:1–e2356:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2020:CAP

[DFHM20] Hans De Sterck, Stephanie Friedhoff, Alexander J. M. Howse, and Scott P. MacLachlan. Convergence analysis for parallel-in-time solution of hyperbolic systems. Numerical Linear Algebra with Applications, 27(1):e2271:1– e2271:??, January 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2008:DTI

[DFNY08] Hans De Sterck, Robert D. Falgout, Joshua W. Nolting, and Ulrike Meier Yang. Distancetwo interpolation for parallel algebraic multigrid. Numerical Linear Algebra with Applications, 15(2-3):115–139, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DiFiore:2005:LCM

[DFZ05] Carmine Di Fiore, Stefano Fanelli, and Paolo Zellini. Lowcomplexity minimization algorithms. Numerical Linear Algebra with Applications, 12(8): 755–768, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dabkowski:2013:SPS

[DGB^{+}13] Pawel Dabkowski, Krzysztof Galkowski, Olivier Bachelier, Eric Rogers, Anton Kummert, and James Lam. Strong practical stability and stabilization of uncertain discrete linear repetitive processes. Numerical Linear Algebra with Applications, 20(2):220–233, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dou:2019:EAB

[DGC19] Yan Dou, Ming Gu, and Shivkumar Chandrasekaran. An efficient algorithm for bounded spectral matching with affine constraint. Numerical Linear Algebra with Applications, 26(5):e2264:1–e2264:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2016:PHM

[DGM^{+}16] Marco Donatelli, Carlo Garoni, Mariarosa Mazza, Stefano Serra-Capizzano, and Debora Sesana. Preconditioned HSS method for large multilevel block Toeplitz linear systems via the notion of matrix-valued symbol. Numerical Linear Algebra with Applications, 23(1): 83–119, January 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dopico:2019:CRK

[DGP19] Froilán M. Dopico and Javier González-Pizarro. A compact rational Krylov method for large-scale rational eigenvalue problems. Numerical Linear Algebra with Applications, 26 (1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dey:2011:SFS

[DGRR11] Rajeeb Dey, Sandip Ghosh, G. Ray, and A. Rakshit. State feedback stabilization of uncertain linear time-delay systems: a nonlinear matrix inequality approach. Numerical Linear Algebra with Applications, 18 (3):351–361, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dostal:2004:SFO

[DH04] Zdeněk Dostál and David Horák. Scalable FETI with optimal dual penalty for a variational inequality. Numerical Linear Algebra with Applications, 11(5–6):455–472, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2018:NPB

[DH18] Hans De Sterck and Alexander J. M. Howse. Nonlinearly preconditioned L-BFGS as an acceleration mechanism for alternating least squares with application to tensor decomposition. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dostal:2021:BRB

[DHBV21] Zdenek Dostál, David Horák, Tomás Brzobohatý, and Petr Vodstrcil. Book review: Bounds on the spectra of Schur complements of large H-TFETI-DP clusters for 2D Laplacian. Numerical Linear Algebra with Applications, 28(2):e2344:1–e2344:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dykes:2018:RMD

[DHNR18] Laura Dykes, Guangxin Huang, Silvia Noschese, and Lothar Reichel. Regularization matrices for discrete ill-posed problems in several space dimensions. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Douglas:2004:CAM

[DHR^{+}04] C. C. Douglas, J. Hu, J. Ray, D. T. Thorne, and R. S. Tuminaro. Cache aware multigrid for variable coefficient elliptic problems on adaptive mesh refinement hierarchies. Numerical Linear Algebra with Applications, 11(2–3):173–187, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dalvand:2020:ECT

[DHR20] Zeynab Dalvand, Masoud Hajarian, and Jose E. Roman.

An extension of the Cayley transform method for a parameterized generalized inverse eigenvalue problem. Numerical Linear Algebra with Applications, 27(6):e2327:1–e2327:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Demmel:1995:SBF

[DHS95] James W. Demmel, Nicholas J.

Higham, and Robert S. Schreiber. Stability of block LU factorization. Numerical Linear Algebra with Applications, 2(2):173– 190, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Danczul:2023:URK

[DHS23] Tobias Danczul, Clemens Hofreither, and Joachim Schöberl. A unified rational Krylov method for elliptic and parabolic fractional diffusion problems. Numerical Linear Algebra with Applications, 30(5):e2488:1– e2488:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dayar:2011:BED

[DHSW11] Tuğrul Dayar, Holger Hermanns, David Spieler, and Verena Wolf. Bounding the equilibrium distribution of Markov population models. Numerical Linear Algebra with Applications, 18(6):931–946, November 2011. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Ding:2016:TLN

[DHW16] Weiyang Ding, Zongyuan Hou, and Yimin Wei. Tensor logarithmic norm and its applications. Numerical Linear Algebra with Applications, 23(6): 989–1006, December 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Diao:2009:CCN

[Dia09] Huaian Diao. On componentwise condition numbers for eigenvalue problems with structured matrices. Numerical Linear Algebra with Applications, 16(2):87–107, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeTeran:2019:NSG

[DIPR19] Fernando De Terán, Bruno Iannazzo, Federico Poloni, and Leonardo Robol. Nonsingular systems of generalized Sylvester equations: An algorithmic approach. Numerical Linear Algebra with Applications, 26(5):e2261:1– e2261:??, October 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dong:2009:MMM

[DJ09] Jun-Liang Dong and Mei-Qun Jiang. A modified modulus method for symmetric positivedefinite linear complementarity problems. Numerical Linear Algebra with Applications, 16 (2):129–143, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dai:2021:EPD

[DJW^{+}21] Pingfei Dai, Jinhong Jia, Hong Wang, Qingbiao Wu, and Xiangcheng Zheng. An efficient positive-definite blockpreconditioned finite volume solver for two-sided fractional diffusion equations on composite mesh. Numerical Linear Algebra with Applications, 28 (5):e2372:1–e2372:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Druskin:1995:KSA

[DK95] Vladimir Druskin and Leonid Knizhnerman. Krylov subspace approximation of eigenpairs and matrix functions in exact and computer arithmetic. Numerical Linear Algebra with Applications, 2 (3):205–217, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dolgov:2015:SST

[DK15] Sergey Dolgov and Boris Khoromskij. Simultaneous state-time approximation of the chemical master equation using tensor product formats. Numerical Linear Algebra with Applications, 22(2): 197–219, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dehdezi:2023:FSA

[DK23] Eisa Khosravi Dehdezi and Saeed Karimi. On finding strong approximate inverses for tensors. Numerical Linear Algebra with Applications, 30 (1):e2460:1–e2460:??, January 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2022:MAT

[DKM^{+}22] Marco Donatelli, Rolf Krause, Mariarosa Mazza, Matteo Semplice, and Ken Trotti. Matrices associated to two conservative discretizations of Riesz fractional operators and related multigrid solvers. Numerical Linear Algebra with Applications, 29(5):e2436:1– e2436:??, October 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dostal:2015:RBS

[DKVB15] Z. Dostál, T. Kozubek, O. Vlach, and T. Brzobohatý. Reorthogonalizationbased stiffness preconditioning in FETI algorithms with applications to variational inequalities. Numerical Linear Algebra with Applications, 22(6): 987–998, December 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dai:1997:NMG

[DL97] Hua Dai and Peter Lancaster.

Newton’s method for a generalized inverse eigenvalue problem. Numerical Linear Algebra with Applications, 4(1): 1–21, January/February 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001013;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=15001013&
PLACEBO=IE.pdf.

Dou:2023:CBA

[DL23] Yan Dou and Zhao-Zheng Liang. A class of block alternating splitting implicit iteration methods for double saddle point linear systems. Numerical Linear Algebra with Applications, 30(1):e2455:1– e2455:??, January 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dauzickaite:2020:SES

[DLSvL20] Ieva Dauzickaite, Amos S. Lawless, Jennifer A. Scott, and Peter Jan van Leeuwen. Spectral estimates for saddle point matrices arising in weak constraint four-dimensional variational data assimilation. Numerical Linear Algebra with Applications, 27(5):e2313:1– e2313:??, October 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dobrev:2006:TLP

[DLVZ06] Veselin A. Dobrev, Raytcho D. Lazarov, Panayot S. Vassilevski, and Ludmil T. Zikatanov. Two-level preconditioning of discontinuous Galerkin approximations of second-order elliptic equations. Numerical Linear Algebra with Applications, 13(9):753–770, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dendy:2010:BBM

[DM10] J. E. Dendy, Jr. and J. D. Moulton. Black Box Multigrid with coarsening by a factor of three. Numerical Linear Algebra with Applications, 17(2-3):577–598, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2008:EBR

[DMM^{+}08] H. De Sterck, T. Manteuffel, S. McCormick, J. Nolting, J. Ruge, and L. Tang. Efficiency-based h- and hprefinement strategies for finite element methods. Numerical Linear Algebra with Applications, 15(2-3):89–114, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DiPietro:2023:HOM

[DMMR23] Daniele A. Di Pietro, Pierre Matalon, Paul Mycek, and Ulrich Rüde. High-order multigrid strategies for hybrid highorder discretizations of elliptic equations. Numerical Linear Algebra with Applications, 30 (1):e2456:1–e2456:??, January 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Damm:2017:NSF

[DMS17] Tobias Damm, Hermann Mena, and Tony Stillfjord. Numerical solution of the finite horizon stochastic linear quadratic control problem. Numerical Linear Algebra with Applications, 24(4):??, August 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2011:FMM

[DMTY11] Hans De Sterck, Killian Miller, Eran Treister, and Irad Yavneh. Fast multilevel methods for Markov chains. Numerical Linear Algebra with Applications, 18(6):961–980, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dai:2003:IAP

[DMY03] Yu-Hong Dai, José Mario Martínez, and Jin-Yun Yuan. An increasing-angle property of the conjugate gradient method and the implementation of large-scale minimization algorithms with line searches. Numerical Linear Algebra with Applications, 10(4): 323–334, June 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Du:2012:NLG

[DN12] Kui Du and Olavi Nevanlinna. A note on R-linear GMRES for solving a class of R-linear systems. Numerical Linear Algebra with Applications, 19(5): 880–884, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2012:SRM

[DNR12] Marco Donatelli, Arthur Neuman, and Lothar Reichel. Square regularization matrices for large linear discrete illposed problems. Numerical Linear Algebra with Applications, 19(6):896–913, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dayar:2018:CVF

[DO18] Tuğrul Dayar and M. Can Orhan. On compact vector formats in the solution of the chemical master equation with backward differentiation. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dobrowolski:1999:FEM

[Dob99]  Manfred Dobrowolski.  Fi-

nite element methods for elliptic systems with constraints. Numerical Linear Algebra with Applications, 6(2):115–124, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002993;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002993&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Dodig:2011:PPP

[Dod11] Marija Dodig. Pole placement problem for singular systems. Numerical Linear Algebra with Applications, 18(3): 283–297, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dohrmann:2007:ABP

[Doh07] C. R. Dohrmann. An approximate BDDC preconditioner. Numerical Linear Algebra with Applications, 14(2): 149–168, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2005:MID

[Don05] M. Donatelli. A multigrid for image deblurring with Tikhonov regularization. Numerical Linear Algebra with Applications, 12(8):715–729, October 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Donatelli:2010:AGL

[Don10] M. Donatelli. An algebraic generalization of local Fourier analysis for grid transfer operators in multigrid based on Toeplitz matrices. Numerical Linear Algebra with Applications, 17(2-3):179–197, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Delgado:2019:ACL

[DOP19] Jorge Delgado, Héctor Orera, and Juan Manuel Peña. Accurate computations with Laguerre matrices. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Delgado:2021:HRA

[DOP21] Jorge Delgado, Héctor Orera, and Juan M. Peña. High relative accuracy with matrices of q-integers. Numerical Linear Algebra with Applications, 28(5):e2383:1–e2383:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dehestani:2019:AGW

[DOR19] Haniye Dehestani, Yadollah Ordokhani, and Mohsen Razzaghi. On the applicability of Genocchi wavelet method for different kinds of fractionalorder differential equations with delay. Numerical Linear Algebra with Applications, 26(5):e2259:1–e2259:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dehestani:2021:NDM

[DOR21] Haniye Dehestani, Yadollah Ordokhani, and Mohsen Razzaghi. A novel direct method based on the Lucas multiwavelet functions for variable-order fractional reaction-diffusion and subdiffusion equations. Numerical Linear Algebra with Applications, 28(2):e2346:1–e2346:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dostal:1999:PPM

[Dos99]  Zdeněk Dostál.  On pre-

conditioning and penalized matrices. Numerical Linear Algebra with Applications, 6(2):109–114, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002992;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002992&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Dryja:2003:PMD

[DP03] M. Dryja and W. Proskurowski. On preconditioners for mortar discretization of elliptic problems. Numerical Linear Algebra with Applications, 10(1–2): 65–82, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Delgado:2023:HRA

[DP23] Jorge Delgado and Juan Manuel Peña. High relative accuracy with some special matrices related to Γ and β functions. Numerical Linear Algebra with Applications, 30(5):e2494:1– e2494:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Delgado:2016:AFC

[DPP16] Jorge Delgado, Guillermo Peña, and Juan Manuel Peña. Accurate and fast computations with positive extended Schoenmakers–Coffey matrices. Numerical Linear Algebra with Applications, 23(6):1023– 1031, December 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DelCorso:2019:WMU

[DPRV19] Gianna M. Del Corso, Federico Poloni, Leonardo Robol, and Raf Vandebril. When is a matrix unitary or Hermitian plus low rank? Numerical Linear Algebra with Applications, 26 (6):e2266:1–e2266:??, December 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DiNapoli:2016:EEE

[DPS16] Edoardo Di Napoli, Eric Polizzi, and Yousef Saad. Efficient estimation of eigenvalue counts in an interval. Numerical Linear Algebra with Applications, 23(4):674–692, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ding:2015:FHT

[DQW15] Weiyang Ding, Liqun Qi, and Yimin Wei. Fast Hankel tensor–vector product and its application to exponential data fitting. Numerical Linear Algebra with Applications, 22(5): 814–832, October 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Durazzi:2003:IPC

[DR03] C. Durazzi and V. Ruggiero. Indefinitely preconditioned conjugate gradient method for large sparse equality and inequality constrained quadratic problems. Numerical Linear Algebra with Applications, 10(8):673–688, December 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dimitrov:2002:AEN

[DS02] A. Dimitrov and E. Schnack. Asymptotical expansion in non-Lipschitzian domains — a numerical approach. Numerical Linear Algebra with Applications, 9(6–7):467–492, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Du:2008:NMI

[DS08] Xiuhong Du and Daniel B. Szyld. A note on the mesh independence of convergence bounds for additive Schwarz preconditioned GM-RES. Numerical Linear Algebra with Applications, 15(6): 547–557, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dollar:2010:NFA

[DS10] H. S. Dollar and J. A. Scott. A note on fast approximate minimum degree orderings for symmetric matrices with some dense rows. Numerical Linear Algebra with Applications, 17(1):43–55, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dosso:2013:SSS

[DS13a] M. Dosso and M. Sadkane. On the strong stability of symplectic matrices. Numerical Linear Algebra with Applications, 20 (2):234–249, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dupont:2013:PMT

[DS13b] Todd F. Dupont and L. Ridgway Scott. The power method for tensor eigenproblems and limiting directions of Newton iterates. Numerical Linear Algebra with Applications, 20(6): 956–971, December 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Damm:2018:NSL

[DSV18] Tobias Damm, Kazuhiro Sato, and Axel Vierling. Numerical solution of Lyapunov equations related to Markov jump linear systems. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Du:2019:TUB

[Du19] Kui Du. Tight upper bounds for the convergence of the randomized extended Kaczmarz and Gauss–Seidel algorithms. Numerical Linear Algebra with Applications, 26(3):e2233:1– e2233:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DAmbra:2019:IST

[DV19] Pasqua D’Ambra and Panayot S.

Vassilevski. Improving solve time of aggregation-based adaptive AMG. Numerical Linear Algebra with Applications, 26(6):e2269:1–e2269:??, December 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Diao:2007:FNC

[DW07] Huaian Diao and Yimin Wei. On Frobenius normwise condition numbers for Moore– Penrose inverse and linear least-squares problems. Numerical Linear Algebra with Applications, 14(8):603–610, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

DeSterck:2015:NPC

[DW15] Hans De Sterck and Manda Winlaw. A nonlinearly preconditioned conjugate gradient algorithm for rank-R canonical tensor approximation. Numerical Linear Algebra with Applications, 22(3):410–432, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Danieli:2021:AOS

[DW21] Federico Danieli and Andrew J. Wathen. All-at-once solution of linear wave equations. Numerical Linear Algebra with Applications, 28(6):e2386:1– e2386:??, December 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Diao:2013:CNM

[DWWQ13] Huaian Diao, Weiguo Wang, Yimin Wei, and Sanzheng Qiao. On condition numbers for Moore–Penrose inverse and linear least squares problem involving Kronecker products. Numerical Linear Algebra with Applications, 20(1):44–59, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Diao:2012:MCC

[DXW12] Huaian Diao, Hua Xiang, and Yimin Wei. Mixed, componentwise condition numbers and small sample statistical condition estimation of Sylvester equations. Numerical Linear Algebra with Applications, 19 (4):639–654, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dardyk:2004:MAT

[DY04] Gregory Dardyk and Irad Yavneh. A multigrid approach to two-dimensional phase unwrapping. Numerical Linear Algebra with Applications, 11(2–3):241–259, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elden:2019:SBT

[EAA19] Lars Eldén and Salman Ahmadi-Asl. Solving bilinear tensor least squares problems and application to Hammerstein identification. Numerical Linear Algebra with Applications, 26 (2):e2226:1–e2226:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elden:2022:SPL

[ED22] Lars Eldén and Maryam Dehghan. Spectral partitioning of large and sparse 3-tensors using low-rank tensor approximation. Numerical Linear Algebra with Applications, 29 (5):e2435:1–e2435:??, October 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ekstrom:2018:EBS

[EFG^{+}18] Sven-Erik Ekström, Isabella Furci, Carlo Garoni, Carla Manni, Stefano Serra-Capizzano, and Hendrik Speleers. Are the eigenvalues of the B-spline isogeometric analysis approximation of −Δu = λu known in almost closed form? Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Estrin:2016:TOC

[EG16] R. Estrin and C. Greif. Towards an optimal condition number of certain augmented Lagrangian-type saddle-point matrices. Numerical Linear Algebra with Applications, 23(4): 693–705, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elmaliki:2011:EHP

[EGF11] A. El maliki, R. Guénette, and M. Fortin. An efficient hierarchical preconditioner for quadratic discretizations of finite element problems. Numerical Linear Algebra with Applications, 18(5):789–803, October 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Egger:2007:PCI

[Egg07] H. Egger. Preconditioning CGNE iteration for inverse problems. Numerical Linear Algebra with Applications, 14 (3):183–196, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

ElHaddad:2020:SAO

[EGMS20] Mireille El Haddad, José C.

Garay, Frédéric Magoulès, and Daniel B. Szyld. Synchronous and asynchronous optimized Schwarz methods for one-way subdivision of bounded domains. Numerical Linear Algebra with Applications, 27(2): e2279:1–e2279:??, March 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elsner:1995:MNN

[EHM95] Ludwig Elsner, Chun Yang He, and Volker Mehrmann. Minimization of the norm, the norm of the inverse and the condition number of a matrix by completion. Numerical Linear Algebra with Applications, 2(2):155–171, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elmroth:2001:CPG

[EJK01] Erik Elmroth, Pedher Johansson, and Bo Kågström. Computation and presentation of graphs displaying closure hierarchies of Jordan and Kronecker structures. Numerical Linear Algebra with Applications, 8(6–7):381–399, September/November 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007286/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85007286&PLACEBO=IE.pdf.

Egana:2002:IEM

[EKS02] Juan C. Egaña, Nelson M. Kuhl, and Luis C. Santos. An inverse eigenvalue method for frequency isolation in springmass systems. Numerical Linear Algebra with Applications, 9(1):65–79, January/February 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013651/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013651&PLACEBO=IE.pdf.

Ewing:1994:LRT

[ELV94] R. E. Ewing, R. D. Lazarov, and P. S. Vassilevski. Local refinement techniques for elliptic problems on cell-centered grids. II. optimal order twogrid iterative methods. Numerical Linear Algebra with Applications, 1(4):337–368, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Edelman:1995:PMN

[EM95] Alan Edelman and Walter F.

Mascarenhas. On Parlett’s matrix norm inequality for the Cholesky decomposition. Numerical Linear Algebra with Applications, 2(3):243–250, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Erway:2011:BSA

[EM11] Jennifer B. Erway and Roummel F. Marcia. A backward stability analysis of diagonal pivoting methods for solving unsymmetric tridiagonal systems without interchanges. Numerical Linear Algebra with Applications, 18(1):41–54, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Emans:2012:KAA

[Ema12] Maximilian Emans. Krylovaccelerated algebraic multigrid for semi-definite and nonsymmetric systems in computational fluid dynamics. Numerical Linear Algebra with Applications, 19(2):210–231, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Erlangga:2017:CTL

[EN17] Yogi A. Erlangga and Reinhard Nabben. On the convergence of two-level Krylov methods for singular symmetric systems. Numerical Linear Algebra with Applications, 24(6): ??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Escalante:1996:DAC

[ER96] René Escalante and Marcos Raydan. Dykstra’s algorithm for a constrained least-squares matrix problem. Numerical Linear Algebra with Applications, 3(6): 459–471, November/December 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001009.

Elden:2005:MLE

[ES05] Lars Eldén and Berkant Savas. The maximum likelihood estimate in reduced-rank regression. Numerical Linear Algebra with Applications, 12(8): 731–741, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elfving:2007:BPS

[ES07] Tommy Elfving and Ingegerd Skoglund. A block-preconditioner for a special regularized leastsquares problem. Numerical Linear Algebra with Applications, 14(6):469–484, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elfving:2009:DMR

[ES09a] Tommy Elfving and Ingegerd Skoglund. A direct method for a regularized least-squares problem. Numerical Linear Algebra with Applications, 16(8): 649–675, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Elhashash:2009:TCM

[ES09b] Abed Elhashash and Daniel B.

Szyld. Two characterizations of matrices with the Perron– Frobenius property. Numerical Linear Algebra with Applications, 16(11-12):863–869, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ekstrom:2018:EEB

[ESC18] S.-E. Ekström and S. Serra-Capizzano. Eigenvalues and eigenvectors of banded Toeplitz matrices and the related symbols. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ekstrom:2020:ESP

[ESC20] Sven-Erik Ekström and Stefano Serra-Capizzano. Eigenpairs of some particular band Toeplitz matrices: A comment. Numerical Linear Algebra with Applications, 27 (1):e2273:1–e2273:??, January 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Eigel:2023:DLR

[ESS23] Martin Eigel, Reinhold Schneider, and David Sommer. Dynamical low-rank approximations of solutions to the Hamilton–Jacobi–Bellman equation. Numerical Linear Algebra with Applications, 30(3): e2463:1–e2463:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Estatico:2009:PIP

[Est09] C. Estatico. Preconditioners for ill-posed Toeplitz matrices with differentiable generating functions. Numerical Linear Algebra with Applications, 16 (3):237–257, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Eastwood:2013:FDE

[EW13] Shawn Eastwood and Justin W. L. Wan. Finding offdiagonal entries of the inverse of a large symmetric sparse matrix. Numerical Linear Algebra with Applications, 20(1):

74–92, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ewing:2003:SDF

[EWY03] Richard E. Ewing, Junping Wang, and Yongjun Yang. A stabilized discontinuous finite element method for elliptic problems. Numerical Linear Algebra with Applications, 10 (1–2):83–104, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Evans:2023:BAA

[EY23] David Evans and Nan Ye.

Blockwise acceleration of alternating least squares for canonical tensor decomposition. Numerical Linear Algebra with Applications, 30(6):e2516:1– e2516:??, December 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ecker:1996:SPM

[EZ96] Alois Ecker and Walter Zulehner.

On the smoothing property of multigrid methods in the non-symmetric case. Numerical Linear Algebra with Applications, 3(2):161–172, March/April 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000982.

Falgout:2006:SIM

[Fal06]  Robert D. Falgout.  Spe-

cial issue on multigrid methods. Numerical Linear Algebra with Applications, 13(2–3): 85, March/April 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Falgout:2008:MM

[Fal08] Robert D. Falgout. Multigrid Methods. Numerical Linear Algebra with Applications, 15(2-3):85–87, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Falgout:2010:MM

[Fal10] Robert D. Falgout. Multigrid methods. Numerical Linear Algebra with Applications, 17(2-3):175–178, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fasino:2005:RKM

[Fas05]  Dario Fasino.  Rational

Krylov matrices and QR steps on Hermitian diagonal-plus-semiseparable matrices. Numerical Linear Algebra with Applications, 12(8):743–754, October 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fierro:1995:OPT

[FB95] Ricardo D. Fierro and James R. Bunch. Orthogonal projection and total least squares. Numerical Linear Algebra with Applications, 2(2):135–153, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ferrari:2021:ASL

[FBSC21] Paola Ferrari, Nikos Barakitis, and Stefano Serra-Capizzano. Asymptotic spectra of large matrices coming from the symmetrization of Toeplitz structure functions and applications to preconditioning. Numerical Linear Algebra with Applications, 28(1):e2332:1– e2332:??, January 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Farea:2023:EGS

[FÇ23] Afrah Farea and M. Serdar Çelebi. On the evaluation of general sparse hybrid linear solvers. Numerical Linear Algebra with Applications, 30(2):e2469:1–e2469:??, March 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ferket:1996:FDB

[Fer96] P. J. J. Ferket. The finite difference based fast adaptive composite grid method. Numerical Linear Algebra with Applications, 3(5):391– 411, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15000999.

Fasino:2002:LTA

[FG02] Dario Fasino and Luca Gemignani. A Lanczos-type algorithm for the QR factorization of regular Cauchy matrices. Numerical Linear Algebra with Applications, 9(4):305–319, June 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/92012855/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=92012855&PLACEBO=IE.pdf.

Fezzani:2014:BFD

[FGNW14] Riadh Fezzani, Laura Grigori, Frédéric Nataf, and Ke Wang. Block filtering decomposition. Numerical Linear Algebra with Applications, 21(6):703–721, December 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fares:2011:FRL

[FGT11] M’Barek Fares, Serge Gratton, and Philippe L. Toint. Fast regularized linear sampling for inverse scattering problems. Numerical Linear Algebra with Applications, 18(1):55–68, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Freund:1994:UTA

[FH94] Roland W. Freund and Marlis Hochbruck. On the use of two QMR algorithms for solving singular systems and applications in Markov chain modeling. Numerical Linear Algebra with Applications, 1(4):403– 420, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Farrell:2021:LFA

[FHM21] Patrick E. Farrell, Yunhui He, and Scott P. MacLachlan. A local Fourier analysis of additive Vanka relaxation for the Stokes equations. Numerical Linear Algebra with Applications, 28(3):e2306:1–e2306:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fierro:2005:LRS

[FJ05] Ricardo D. Fierro and Eric P.

Jiang. Lanczos and the Riemannian SVD in information retrieval applications. Numerical Linear Algebra with Applications, 12(4):355–372, May 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ferronato:2012:EPS

[FJP12] Massimiliano Ferronato, Carlo Janna, and Giorgio Pini. Efficient parallel solution to large-size sparse eigenproblems with block FSAI preconditioning. Numerical Linear Algebra with Applications, 19(5): 797–815, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ferronato:2016:PJD

[FJP16] Massimiliano Ferronato, Carlo Janna, and Giorgio Pini. Parallel Jacobi–Davidson with block FSAI preconditioning and controlled inner iterations. Numerical Linear Algebra with Applications, 23(3):394–409, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Freitag:2015:TPI

[FK15] Melina A. Freitag and Patrick Kürschner. Tuned preconditioners for inexact two-sided inverse and Rayleigh quotient iteration. Numerical Linear Algebra with Applications, 22(1): 175–196, January 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Favati:2009:RAU

[FLM09] P. Favati, G. Lotti, and O. Menchi. Recursive algorithms for unbalanced banded Toeplitz systems. Numerical Linear Algebra with Applications, 16(7):561–587, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Farhat:2000:SDP

[FLP00] Charbel Farhat, Michael Lesoinne, and Kendall Pierson. A scalable dual-primal domain decomposition method. Numerical Linear Algebra with Applications, 7(7–8):687–714, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505483/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505483&PLACEBO=IE.pdf.

Ferng:2001:NTM

[FLPW01] William R. Ferng, Wen-Wei Lin, Daniel J. Pierce, and Chern-Shuh Wang. Nonequivalence transformation of λmatrix eigenproblems and model embedding approach to model tuning. Numerical Linear Algebra with Applications, 8(1):53–70, January/February 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501455/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501455&PLACEBO=IE.pdf.

Farkas:2003:CAP

[FLR03] András Farkas, Peter Lancaster, and Pál Rózsa. Consistency adjustments for pairwise comparison matrices. Numerical Linear Algebra with Applications, 10(8):689–700, December 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2021:TRM

[fLWyL^{+}21] Jiao fen Li, Kai Wang, Yue yuan Liu, Xue feng Duan, and Xue lin Zhou. A trustregion method for the parameterized generalized eigenvalue problem with nonsquare matrix pencils. Numerical Linear Algebra with Applications, 28 (4):e2363:1–e2363:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2011:NSC

[fLyHZ11] Jiao fen Li, Xi yan Hu, and Lei Zhang. Numerical solutions of AXB = C for centrosymmetric matrix X under a specified submatrix constraint. Numerical Linear Algebra with Applications, 18(5):857–873, October 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Freund:1999:CUW

[FM99] Ronold W. Freund and Ivo Marek. Czech-US workshop on iterative methods and parallel computing. Numerical Linear Algebra with Applications, 6 (1):1, January/February 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002988;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002988&
                                        PLACEBO=IE.pdf. Special Is-

sue: Czech-US Workshop on Iterative Methods and Parallel Computing.

Friedhoff:2015:GPA

[FM15] S. Friedhoff and S. MacLachlan. A generalized predictive analysis tool for multigrid methods. Numerical Linear Algebra with Applications, 22(4): 618–647, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fox:2018:AMD

[FM18] Alyson Fox and Thomas Manteuffel. Algebraic multigrid for directed graph Laplacian linear systems (NS-LAMG). Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Friedland:2013:BRO

[FMPS13] S. Friedland, V. Mehrmann, R. Pajarola, and S. K. Suter. On best rank one approximation of tensors. Numerical Linear Algebra with Applications, 20(6):942–955, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Freitag:1995:RAD

[FO95] Lori A. Freitag and James M.

Ortega. The RSCG algorithm on distributed memory architectures. Numerical Linear Algebra with Applications, 2(5):401–414, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Special issue dedicated to David M. Young, Jr.

Fairbanks:2021:EPQ

[FOV21] Hillary R. Fairbanks, Sarah Osborn, and Panayot S. Vassilevski. Estimating posterior quantity of interest expectations in a multilevel scalable framework. Numerical Linear Algebra with Applications, 28(3):e2352:1–e2352:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fernando:1995:ICA

[FP95a] K. Vince Fernando and Beresford N. Parlett. Implicit Cholesky algorithms for singular values and vectors of triangular matrices. Numerical Linear Algebra with Applications, 2(6):507–531, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Frommer:1995:CRM

[FP95b] Andreas Frommer and Bert Pohl. A comparison result for multisplittings and waveform relaxation methods. Numerical Linear Algebra with Applications, 2(4):335–346, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fenn:2005:FSB

[FP05] Markus Fenn and Daniel Potts.

Fast summation based on fast trigonometric transforms at non-equispaced nodes. Numerical Linear Algebra with Applications, 12(2–3):161–169, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Farrell:2015:BPL

[FP15] Patricio Farrell and Jennifer Pestana. Block preconditioners for linear systems arising from multiscale collocation with compactly supported RBFs. Numerical Linear Algebra with Applications, 22(4): 731–747, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Filelis-Papadopoulos:2021:IIM

[FP21] Christos K. Filelis-Papadopoulos.

Incomplete inverse matrices. Numerical Linear Algebra with Applications, 28(5):e2380:1– e2380:??, October 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fenu:2016:GTR

[FRR16] Caterina Fenu, Lothar Reichel, and Giuseppe Rodriguez. GCV for Tikhonov regularization via global Golub–Kahan decomposition. Numerical Linear Algebra with Applications, 23(3): 467–484, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Funken:2009:FSB

[FS09] Stefan A. Funken and Ernst P. Stephan. Fast solvers with block-diagonal preconditioners for linear FEM–BEM coupling. Numerical Linear Algebra with Applications, 16(5): 365–395, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Friedhoff:2021:OHI

[FS21] Stephanie Friedhoff and Ben S. Southworth. On “optimal” h-independent convergence of parareal and multigridreduction-in-time using Runge– Kutta time integration. Numerical Linear Algebra with Applications, 28(3):e2301:1– e2301:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Frommer:2018:BDE

[FSS18] Andreas Frommer, Claudia Schimmel, and Marcel Schweitzer. Bounds for the decay of the entries in inverses and Cauchy–Stieltjes functions of certain sparse, normal matrices. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fraysse:1998:NNP

[FT98] Valérie Frayssé and Vincent Toumazou. A note on the normwise perturbation theory for the regular generalized eigenproblem. Numerical Linear Algebra with Applications, 5(1):1–10, January/February 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5959;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5959&PLACEBO=IE.pdf.

Falgout:2005:TGC

[FVZ05] Robert D. Falgout, Panayot S.

Vassilevski, and Ludmil T. Zikatanov. On two-grid convergence estimates. Numerical Linear Algebra with Applications, 12(5–6):471–494, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Fang:2001:SFD

[FY01] Qing Fang and Tetsuro Yamamoto. Superconvergence of finite difference approximations for convection-diffusion problems. Numerical Linear Algebra with Applications, 8(2):99–110, March 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501759/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=76501759&PLACEBO=IE.pdf.

Fan:2017:NSL

[FZwCW17] Hung-Yuan Fan, Liping Zhang, Eric King wah Chu, and Yimin Wei. Numerical solution to a linear equation with tensor product structure. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gosea:2018:DDM

[GA18] Ion Victor Gosea and Athanasios C. Antoulas. Datadriven model order reduction of quadratic-bilinear systems. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gander:1999:WRA

[Gan99] Martin J. Gander. A waveform relaxation algorithm with overlapping splitting for reaction diffusion equations. Numerical Linear Algebra with Applications, 6(2):125–145, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002994;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002994&
                                        PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Gander:2005:CBP

[Gan05] W. Gander. Change of basis in polynomial interpolation. Numerical Linear Algebra with Applications, 12(8): 769–778, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Garanzha:2001:BVG

[Gar01] V. A. Garanzha. Barrier variational generation of quasiisometric grids. Numerical Linear Algebra with Applications, 8(5):329–353, July/August 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/82004030/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=82004030&PLACEBO=IE.pdf.

Garanzha:2002:MNO

[Gar02] V. A. Garanzha. Maximum norm optimization of quasiisometric mappings. Numerical Linear Algebra with Applications, 9(6–7):493–510, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Garanzha:2004:VPG

[Gar04] V. A. Garanzha. Variational principles in grid generation and geometric modelling: theoretical justifications and open problems. Numerical Linear Algebra with Applications, 11(5–6):535–563, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gao:2011:INM

[GB11] Yong-Hua Gao and Zhong-Zhi Bai. On inexact Newton methods based on doubling iteration scheme for non-symmetric algebraic Riccati equations. Numerical Linear Algebra with Applications, 18(3):325–341, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gosnell:2015:USD

[GB15] Denise K. Gosnell and Michael W. Berry. Using semidiscrete decomposition and vector space models to identify users of social networks. Numerical Linear Algebra with Applications, 22(5):833–844, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gathungu:2022:HMM

[GBB22] Duncan Gathungu, Mario Bebendorf, and Alfio Borzì. Hierarchical-matrix method for a class of diffusiondominated partial integrodifferential equations. Numerical Linear Algebra with Applications, 29(1):e2410:1– e2410:??, January 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gomez-Corral:2018:PAF

[GCLG18] A. Gómez-Corral and M. López-García. Perturbation analysis in finite LD-QBD processes and applications to epidemic models. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Guo:2011:SGC

[GD11] Xue-Ping Guo and Iain S.

Duff. Semilocal and global convergence of the Newton–HSS method for systems of nonlinear equations. Numerical Linear Algebra with Applications, 18(3):299–315, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gemignani:2000:ESS

[Gem00] Luca Gemignani. Efficient and stable solution of structured Hessenberg linear systems arising from difference equations. Numerical Linear Algebra with Applications, 7(5): 319–335, July/August 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72508406/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72508406&PLACEBO=IE.pdf.

Gaspar:2008:DSM

[GGLO08] F. J. Gaspar, J. L. Gracia, F. J. Lisbona, and C. W. Oosterlee. Distributive smoothers in multigrid for problems with dominating grad–div operators. Numerical Linear Algebra with Applications, 15(8): 661–683, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gleich:2013:PAM

[GGV13] David F. Gleich, Chen Greif, and James M. Varah. The power and Arnoldi methods in an algebra of circulants. Numerical Linear Algebra with Applications, 20(5): 809–831, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gomilko:2012:PFI

[GGZ12] Oleksandr Gomilko, Federico Greco, and Krystyna Ziȩtak. A Padé family of iterations for the matrix sign function and related problems. Numerical Linear Algebra with Applications, 19(3):585–605, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gatica:2001:MRI

[GH01] Gabriel N. Gatica and Norbert Heuer. Minimum residual iteration for a dual-dual mixed formulation of exterior transmission problems. Numerical Linear Algebra with Applications, 8(3):147–164, April/ May 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76506693/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76506693&PLACEBO=IE.pdf.

Galantai:2006:JPA

[GH06] A. Galántai and Cs. J.

Hegedűs. Jordan’s principal angles in complex vector spaces. Numerical Linear Algebra with Applications, 13(7): 589–598, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Greif:2011:BSC

[GH11] Chen Greif and Robert L.

Hocking. A box-shaped cyclically reduced operator. Numerical Linear Algebra with Applications, 18(4):671–693, August 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Greif:2023:CFM

[GH23] Chen Greif and Yunhui He. A closed-form multigrid smoothing factor for an additive Vanka-type smoother applied to the Poisson equation. Numerical Linear Algebra with Applications, 30(5):e2500:1– e2500:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gratton:2016:RCA

[GHJV16] Serge Gratton, Pascal Hénon, Pavel Jiránek, and Xavier Vasseur. Reducing complexity of algebraic multigrid by aggregation. Numerical Linear Algebra with Applications, 23 (3):501–518, May 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Griebel:2015:OSP

[GHO15] Michael Griebel, Alexander Hullmann, and Peter Oswald. Optimal scaling parameters for sparse grid discretizations. Numerical Linear Algebra with Applications, 22(1): 76–100, January 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Glunt:1998:NDS

[GHR98] William Glunt, Thomas L. Hayden, and Robert Reams. The nearest “doubly stochastic” matrix to a real matrix with the same first moment. Numerical Linear Algebra with Applications, 5(6): 475–482, November/December 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002967;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002967&
                                        PLACEBO=IE.pdf.

Gee:2009:NSA

[GHT09] Michael W. Gee, Jonathan J.

Hu, and Raymond S. Tuminaro. A new smoothed aggregation multigrid method for anisotropic problems. Numerical Linear Algebra with Applications, 16(1):19–37, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Grimm:2006:HOP

[GHW06] Volker Grimm, Stefan Henn, and Kristian Witsch. A higherorder PDE-based image registration approach. Numerical Linear Algebra with Applications, 13(5):399–417, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

George:2002:GFG

[GIK02] Alan George, Khakim D.

Ikramov, and Andrey B. Kucherov. On the growth factor in Gaussian elimination for generalized Higham matrices. Numerical Linear Algebra with Applications, 9(2):107–114, March 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/89013911/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=89013911&PLACEBO=IE.pdf.

Garanzha:2004:TNT

[GKK04] V. Garanzha, I. Kaporin, and I. Konshin. Truncated Newton type solver with application to grid untangling problem. Numerical Linear Algebra with Applications, 11(5–6):525–533, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gorb:2019:PIM

[GKK19] Yuliya Gorb, Vasiliy Kramarenko, and Yuri Kuznetsov. Preconditioned iterative methods for diffusion problems with high-contrast inclusions. Numerical Linear Algebra with Applications, 26(4):e2243:1– e2243:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Galgon:2018:IPB

[GKL18] Martin Galgon, Lukas Krämer, and Bruno Lang. Improving projection-based eigensolvers via adaptive techniques. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ghysels:2012:IAI

[GKV12] P. Ghysels, P. Klosiewicz, and W. Vanroose. Improving the arithmetic intensity of multigrid with the help of polynomial smoothers. Numerical Linear Algebra with Applications, 19(2):253–267, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gruzinov:1997:BSP

[GKY97] F. A. Gruzinov, L. Yu.

Kolotilina, and A. Yu. Yeremin. Block SSOR preconditionings for high-order 3D FE systems. III. Incomplete BSSOR preconditionings based on ppartitionings. Numerical Linear Algebra with Applications, 4(5):393–423, September/October 1997. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15043;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15043&PLACEBO=IE.pdf.

Ghavimi:1995:BES

[GL95a] Ali R. Ghavimi and Alan J.

Laub. Backward error, sensitivity, and refinement of computed solutions of algebraic Riccati equations. Numerical Linear Algebra with Applications, 2(1):29–49, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gustafsson:1995:CPP

[GL95b] Ivar Gustafsson and Gunhild Lindskog. Completely parallelizable preconditioning methods. Numerical Linear Algebra with Applications, 2(5):447– 465, 1995. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic). Special issue dedicated to David M. Young, Jr.

Gustafsson:1996:PAO

[GL96] Ivar Gustafsson and Gunhild Lindskog. Parallel algorithms for orthotropic problems. Numerical Linear Algebra with Applications, 3 (3):185–203, May/June 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000991.

Gustafsson:1998:PSL

[GL98] Ivar Gustafsson and Gunhild Lindskog. On parallel solution of linear elasticity problems: Part I: theory. Numerical Linear Algebra with Applications, 5(2): 123–139, March/April 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5966;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5966&PLACEBO=IE.pdf.

Gustafsson:2002:PSL

[GL02] I. Gustafsson and G. Lindskog. On parallel solution of linear elasticity problems. Part II: Methods and some computer experiments. Numerical Linear Algebra with Applications, 9(3):205–221, April/ May 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/90512112/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=90512112&PLACEBO=IE.pdf.

Gustafsson:2013:PSL

[GL13] Ivar Gustafsson and Gunhild Lindskog. On parallel solution of linear elasticity problems. Part III: higher order finite elements. Numerical Linear Algebra with Applications, 20(5): 869–887, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Grasedyck:2018:DHS

[GL18] Lars Grasedyck and Christian Löbbert. Distributed hierarchical SVD in the Hierarchical Tucker format. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gander:2021:PPC

[GL21] Martin J. Gander and Thibaut Lunet. ParaStieltjes: Parallel computation of Gauss quadrature rules using a Parareal-like approach for the Stieltjes procedure. Numerical Linear Algebra with Applications, 28(3): e2314:1–e2314:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gaspar:2010:MFE

[GLGR10] F. J. Gaspar, F. J. Lisbona, J. L. Gracia, and C. Rodrigo. Multigrid finite element methods on semi-structured triangular grids for planar elasticity. Numerical Linear Algebra with Applications, 17(2-3): 473–493, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ghai:2019:CPK

[GLJ19] Aditi Ghai, Cao Lu, and Xiangmin Jiao. A comparison of preconditioned Krylov subspace methods for large-scale nonsymmetric linear systems. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gaspar:2004:SCC

[GLOW04] F. J. Gaspar, F. J. Lisbona, C. W. Oosterlee, and R. Wienands. A systematic comparison of coupled and distributive smoothing in multigrid for the poroelasticity system. Numerical Linear Algebra with Applications, 11(2–3):93– 113, March/April 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ghoussoub:2011:NSP

[GM11] Nassif Ghoussoub and Amir Moradifam. A note on simultaneous preconditioning and symmetrization of nonsymmetric linear systems. Numerical Linear Algebra with Applications, 18(3):343–349, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ganesh:2017:EMA

[GM17] Mahadevan Ganesh and Charles Morgenstern. An efficient multigrid algorithm for heterogeneous acoustic media sign-indefinite high-order FEM models. Numerical Linear Algebra with Applications, 24 (3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Griebel:2006:CGC

[GMOS06] Michael Griebel, Bram Metsch, Daniel Oeltz, and Marc Alexander Schweitzer. Coarse grid classification: a parallel coarsening scheme for algebraic multigrid methods. Numerical Linear Algebra with Applications, 13(2–3):193–214, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Giraud:2005:IVD

[GMR05] L. Giraud, A. Marrocco, and J.-C. Rioual. Iterative versus direct parallel substructuring methods in semiconductor device modelling. Numerical Linear Algebra with Applications, 12(1):33–53, February 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gillis:2018:CNS

[GMS18] Nicolas Gillis, Volker Mehrmann, and Punit Sharma. Computing the nearest stable matrix pairs. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Garoni:2020:NID

[GMSCS20] Carlo Garoni, Carla Manni, Stefano Serra-Capizzano, and Hendrik Speleers. NURBS in isogeometric discretization methods: A spectral analysis. Numerical Linear Algebra with Applications, 27 (6):e2318:1–e2318:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gratton:2016:LMP

[GMTV16] Serge Gratton, Sylvain Mercier, Nicolas Tardieu, and Xavier Vasseur. Limited memory preconditioners for symmetric indefinite problems with application to structural mechanics. Numerical Linear Algebra with Applications, 23(5): 865–887, October 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gander:2000:APB

[GN00] Martin J. Gander and Frédéric Nataf. AILU: a preconditioner based on the analytic factorization of the elliptic operator. Numerical Linear Algebra with Applications, 7(7–8): 505–526, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505477/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505477&PLACEBO=IE.pdf.

Grigori:2015:OPB

[GNQ15] Laura Grigori, Frédéric Nataf, and Long Qu. Overlapping for preconditioners based on incomplete factorizations and nested arrow form. Numerical Linear Algebra with Applications, 22(1):48–75, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gazzola:2014:ETC

[GNR14] S. Gazzola, P. Novati, and M. R. Russo. Embedded techniques for choosing the parameter in Tikhonov regularization. Numerical Linear Algebra with Applications, 21(6): 796–812, December 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gazzola:2016:LGK

[GORR16] Silvia Gazzola, Enyinda Onunwor, Lothar Reichel, and Giuseppe Rodriguez. On the Lanczos and Golub–Kahan reduction methods applied to discrete ill-posed problems. Numerical Linear Algebra with Applications, 23(1): 187–204, January 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gavin:2018:KES

[GP18] Brendan Gavin and Eric Polizzi. Krylov eigenvalue strategy using the FEAST algorithm with inexact system solves. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Goossens:1999:RHR

[GR99] Serge Goossens and Dirk Roose. Ritz and harmonic Ritz values and the convergence of FOM and GMRES. Numerical Linear Algebra with Applications, 6(4):281–293, June 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/65500096/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=65500096&PLACEBO=IE.pdf.

Guo:2004:ELS

[GR04] Hongbin Guo and Rosemary A.

Renaut. Estimation of uT (A)v for large-scale unsymmetric matrices. Numerical Linear Algebra with Applications, 11(1): 75–89, February 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Guo:2005:PVD

[GR05] Hongbin Guo and Rosemary A.

Renaut. Parallel variable distribution for total least squares. Numerical Linear Algebra with Applications, 12(9): 859–876, November 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gross:2023:AOP

[GR23] Sven Gross and Arnold Reusken.

Analysis of optimal preconditioners for CutFEM. Numerical Linear Algebra with Applications, 30(5):e2486:1– e2486:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Grasedyck:2008:NMS

[Gra08]  L. Grasedyck.  Nonlinear

multigrid for the solution of large-scale Riccati equations in low-rank and H-matrix format. Numerical Linear Algebra with Applications, 15(9): 779–807, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Grosz:2000:PIB

[Gro00]  L. Grosz.  Preconditioning

by incomplete block elimination. Numerical Linear Algebra with Applications, 7(7–8): 527–541, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505481/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505481&PLACEBO=IE.pdf.

Govaerts:1997:RDM

[GS97] W. Govaerts and B. Sijnave. Rank-deficient matrices as a computational tool. Numerical Linear Algebra with Applications, 4(6): 443–458, November/December 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15048;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15048&PLACEBO=IE.pdf.

Genseberger:1999:ACE

[GS99] Menno Genseberger and Gerard L. G. Sleijpen. Alternative correction equations in the Jacobi–Davidson method. Numerical Linear Algebra with Applications, 6 (3):235–253, April/May 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.

com/cgi-bin/abstract?ID=
63003631;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=63003631&
PLACEBO=IE.pdf.

Gu:2005:NSP

[GS05] G.-D. Gu and V. Simoncini.

Numerical solution of parameterdependent linear systems. Numerical Linear Algebra with Applications, 12(9):923–940, November 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Greif:2007:PDT

[GS07] Chen Greif and Dominik Schötzau. Preconditioners for the discretized time-harmonic Maxwell equations in mixed form. Numerical Linear Algebra with Applications, 14(4): 281–297, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Golub:2001:PBS

[GSS01] Gene H. Golub, Ahmed H.

Sameh, and Vivek Sarin. A parallel balance scheme for banded linear systems. Numerical Linear Algebra with Applications, 8(5):297–316, July/August 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/82004028/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=82004028&PLACEBO=IE.pdf.

Gratton:2021:MCQ

[GSTPT21] Serge Gratton, Ehouarn Simon, David Titley-Peloquin, and Philippe L. Toint. Minimizing convex quadratics with variable precision conjugate gradients. Numerical Linear Algebra with Applications, 28 (1):e2337:1–e2337:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gopalakrishnan:2009:CMC

[GT09] Jayadeep Gopalakrishnan and Shuguang Tan. A convergent multigrid cycle for the hybridized mixed method. Numerical Linear Algebra with Applications, 16(9):689–714, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Grubisic:2019:DSM

[GT19] Luka Grubisić and Josip Tambaca. Direct solution method for the equilibrium problem for elastic stents. Numerical Linear Algebra with Applications, 26(3):e2231:1–e2231:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gratton:2016:SOE

[GTI16] S. Gratton and J. Tshimanga-Ilunga. On a second-order expansion of the truncated singular subspace decomposition. Numerical Linear Algebra with Applications, 23(3):

519–534, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Goreinov:1997:MFI

[GTY97] S. A. Goreinov, E. E. Tyrtyshnikov, and A. Yu. Yeremin. Matrix-free iterative solution strategies for large dense linear systems. Numerical Linear Algebra with Applications, 4(4): 273–294, July/August 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15039;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15039&PLACEBO=IE.pdf.

Gong:2018:ROP

[GTZ18] Wei Gong, Zhiyu Tan, and Shuo Zhang. A robust optimal preconditioner for the mixed finite element discretization of elliptic optimal control problems. Numerical Linear Algebra with Applications, 25(1): ??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gustafson:1997:OTI

[Gus97]  K. Gustafson.  Operator

trigonometry of iterative methods. Numerical Linear Algebra with Applications, 4(4): 333–347, July/August 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.

                                        com/cgi-bin/abstract?ID=
                                        15038;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15038&PLACEBO=IE.pdf.

Gustafson:1998:OTM

[Gus98]  Karl Gustafson.  Operator

trigonometry of the model problem. Numerical Linear Algebra with Applications, 5(5): 377–399, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62000046;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62000046&
                                        PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Gustafson:2003:OTP

[Gus03]  Karl Gustafson.  Opera-

tor trigonometry of preconditioning, domain decomposition, sparse approximate inverses, successive overrelaxation, minimum residual schemes. Numerical Linear Algebra with Applications, 10(4): 291–315, June 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gustafson:2004:IPL

[Gus04a] Karl Gustafson. An inner product lemma. Numerical Linear Algebra with Applications, 11(7):649–659, September 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gustafson:2004:ND

[Gus04b] Karl Gustafson. Normal degree. Numerical Linear Algebra with Applications, 11(7): 661–674, September 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Giraud:2003:GTO

[GVT03] L. Giraud, F. Guevara Vasquez, and R. S. Tuminaro. Grid transfer operators for highly variable coefficient problems in two-level non-overlapping domain decomposition methods. Numerical Linear Algebra with Applications, 10(5–6):467–484, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gulliksson:2000:PTG

[GW00] Mårten Gulliksson and Per-Åke Wedin. Perturbation theory for generalized and constrained linear least squares. Numerical Linear Algebra with Applications, 7(4):181–195, May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72507875/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72507875&PLACEBO=IE.pdf.

Guo:2014:NUB

[GX14] Cui Guo and Hua Xiang. A note on the upper bound in SA AMG convergence analysis. Numerical Linear Algebra with Applications, 21(3): 399–402, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gonzaga:2008:FNL

[GY08] Clovis Caesar Gonzaga and Jin Yun Yuan. Foz2006 numerical linear algebra and optimization. Numerical Linear Algebra with Applications, 15 (10):887–889, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Gillard:2016:WNS

[GZ16] J. W. Gillard and A. A. Zhigljavsky. Weighted norms in subspace-based methods for time series analysis. Numerical Linear Algebra with Applications, 23(5):947–967, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hackbusch:1992:PCG

[Hac92] Wolfgang Hackbusch. A parallel conjugate gradient method. Journal of Numerical Linear Algebra with Applications, 1 (2):133–147, 1992. CODEN ???? ISSN 0129-3281.

Hansen:2013:OPS

[Han13] Per Christian Hansen. Oblique projections and standard-form transformations for discrete inverse problems. Numerical Linear Algebra with Applications, 20(2):250–258, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hamilton:2010:NMS

[HBH10] Steven Hamilton, Michele Benzi, and Eldad Haber. New multigrid smoothers for the Oseen problem. Numerical Linear Algebra with Applications, 17 (2-3):557–576, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huckaby:2005:SPQ

[HC05] D. A. Huckaby and T. F. Chan.

Stewart’s pivoted QLP decomposition for low-rank matrices. Numerical Linear Algebra with Applications, 12(2–3):153–159, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2020:TGM

[HC20] Hanzhang Hu and Yanping Chen. A two-grid method for characteristic expanded mixed finite element solution of miscible displacement problem. Numerical Linear Algebra with Applications, 27(3):e2292:1– e2292:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hao:2015:SSP

[HCD15] C. L. Hao, C. F. Cui, and Y. H.

Dai. A sequential subspace projection method for extreme Zeigenvalues of supersymmetric tensors. Numerical Linear Algebra with Applications, 22(2): 283–298, March 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Heavner:2023:EAC

[HCGM23] Nathan Heavner, Chao Chen, Abinand Gopal, and Per-Gunnar Martinsson. Efficient algorithms for computing rankrevealing factorizations on a GPU. Numerical Linear Algebra with Applications, 30 (6):e2515:1–e2515:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2007:IBB

[HD07] Yi-Qing Hu and Yu-Hong Dai. Inexact Barzilai–Borwein method for saddle point problems. Numerical Linear Algebra with Applications, 14(4): 299–317, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hansen:2019:HEB

[HDA19] Per Christian Hansen, Yiqiu Dong, and Kuniyoshi Abe. Hybrid enriched bidiagonalization for discrete ill-posed problems. Numerical Linear Algebra with Applications, 26(3): e2230:1–e2230:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2019:UMC

[HDH19] Na Huang, Yu-Hong Dai, and QiYa Hu. Uzawa methods for a class of block three-by-three saddle-point problems. Numerical Linear Algebra with Applications, 26(6):e2265:1– e2265:??, December 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hollander:2018:ADM

[HDIS18] Gabriel Hollander, Philippe Dreesen, Mariya Ishteva, and Johan Schoukens. Approximate decoupling of multivariate polynomials using weighted tensor decomposition. Numerical Linear Algebra with Applications, 25(2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

He:2021:IPL

[He21]  Yunhui He.  Independence

of placement for local Fourier analysis. Numerical Linear Algebra with Applications, 28 (6):e2388:1–e2388:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

He:2023:VBP

[He23] Yunhui He. A Vanka-based parameter-robust multigrid relaxation for the Stokes–Darcy Brinkman problems. Numerical Linear Algebra with Applications, 30(6):e2514:1– e2514:??, December 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hemmingsson:1996:TPB

[Hem96] Lina Hemmingsson. Toeplitz preconditioners with block structure for first-order PDEs. Numerical Linear Algebra with Applications, 3(1):21–44, January/February 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000500.

Hezari:2015:PGI

[HES15] Davod Hezari, Vahid Edalatpour, and Davod Khojasteh Salkuyeh. Preconditioned GSOR iterative method for a class of complex symmetric system of linear equations. Numerical Linear Algebra with Applications, 22(4): 761–776, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hetmaniuk:2007:RQM

[Het07] U. Hetmaniuk. A Rayleigh quotient minimization algorithm based on algebraic multigrid. Numerical Linear Algebra with Applications, 14(7): 563–580, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huber:2022:FSI

[HFG^{+}22] Sarah Huber, Yasunori Futamura, Martin Galgon, Akira Imakura, Bruno Lang, and Tetsuya Sakurai. Flexible subspace iteration with moments for an effective contour integration-based eigensolver. Numerical Linear Algebra with Applications, 29 (6):e2447:1–e2447:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hemmingsson-Franden:2001:NOP

[HFW01] Lina Hemmingsson-Frändén and Andrew Wathen. A nearly optimal preconditioner for the Navier–Stokes equations. Numerical Linear Algebra with Applications, 8(4): 229–243, June 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76510167/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76510167&PLACEBO=IE.pdf.

He:2000:SBF

[HG00] J. W. He and R. Glowinski.

Steady Bingham fluid flow in cylindrical pipes: a time dependent approach to the iterative solution. Numerical Linear Algebra with Applications, 7(6):381–428, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516699/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516699&PLACEBO=IE.pdf.

Hakopian:2006:TLP

[HH06] Yuri R. Hakopian and Arsen S. Harutyunyan. Two-level preconditioners for serendipity finite element matrices. Numerical Linear Algebra with Applications, 13(10):847–864, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2016:NTP

[HHLL16] Tsung-Ming Huang, Wei-Qiang Huang, Ren-Cang Li, and Wen-Wei Lin. A new two-phase structure-preserving doubling algorithm for critically singular M-matrix algebraic Riccati equations. Numerical Linear Algebra with Applications, 23(2):291– 313, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Haber:2010:NOC

[HHM10] Eldad Haber, Raya Horesh, and Jan Modersitzki. Numerical optimization for constrained image registration. Numerical Linear Algebra with Applications, 17(2-3):343–359, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2013:FEE

[HHQ13] Shenglong Hu, Zheng-Hai Huang, and Liqun Qi. Finding the extreme Z-eigenvalues of tensors via a sequential semidefinite programming method. Numerical Linear Algebra with Applications, 20(6): 972–984, December 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hemker:2004:FTL

[HHvR04] P. W. Hemker, W. Hoffmann, and M. H. van Raalte. Fourier two-level analysis for discontinuous Galerkin discretization with linear elements. Numerical Linear Algebra with Applications, 11(5–6):473–491, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hached:2018:CKB

[HJ18] M. Hached and K. Jbilou.

Computational Krylov-based methods for large-scale differential Sylvester matrix problems. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Han:1997:NAP

[HJR97] Weimin Han, Søren Jensen, and B. Daya Reddy. Numerical approximations of problems in plasticity: error analysis and solution algorithms. Numerical Linear Algebra with Applications, 4(3):191–204, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15033;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15033&PLACEBO=IE.pdf.

Hackbusch:2002:BKA

[HK02] W. Hackbusch and B. N. Khoromskij. Blended kernel approximation in the Hmatrix techniques. Numerical Linear Algebra with Applications, 9(4):281–304, June 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/92012857/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=92012857&PLACEBO=IE.pdf.

Howle:2012:BPF

[HK12] Victoria E. Howle and Robert C. Kirby. Block preconditioners for finite element discretization of incompressible flow with thermal convection. Numerical Linear Algebra with Applications, 19(2):427–440, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hadjifotinou:2021:ATV

[HK21] Katerina G. Hadjifotinou and Nicholas P. Karampetakis. An algorithm for two-variable rational interpolation suitable for matrix manipulations with the evaluation-interpolation method. Numerical Linear Algebra with Applications, 28 (5):e2378:1–e2378:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Holm:2006:MDC

[HKH^{+}06] Randi Holm, Roland Kaufmann, Bjørn-Ove Heimsund, Erlend Øian, and Magne S. Espedal. Meshing of domains with complex internal geometries. Numerical Linear Algebra with Applications, 13(9): 717–731, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Haslinger:2007:PSC

[HKKP07] J. Haslinger, T. Kozubek, R. Kučera, and G. Peichl. Projected Schur complement method for solving non-symmetric systems arising from a smooth fictitious domain approach. Numerical Linear Algebra with Applications, 14(9):713–739, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hsu:2021:CMN

[HKL21] Fu-Shin Hsu, Yueh-Cheng Kuo, and Ching-Sung Liu. Continuation methods for nonnegative rank-1 approximation of nonnegative tensors. Numerical Linear Algebra with Applications, 28(6):e2398:1– e2398:??, December 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hong:2019:CDP

[HKLP19] Qingguo Hong, Johannes Kraus, Maria Lymbery, and Fadi Philo. Conservative discretizations and parameterrobust preconditioners for Biot and multiple-network fluxbased poroelasticity models. Numerical Linear Algebra with Applications, 26(4):e2242:1– e2242:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hackbusch:2012:UTF

[HKST12] Wolfgang Hackbusch, Boris N. Khoromskij, Stefan Sauter, and Eugene E. Tyrtyshnikov. Use of tensor formats in elliptic eigenvalue problems. Numerical Linear Algebra with Applications, 19(1):133–151, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hong:2016:MBM

[HL16] Jun-Tao Hong and Chen-Liang Li. Modulus-based matrix splitting iteration methods for a class of implicit complementarity problems. Numerical Linear Algebra with Applications, 23(4):629–641, August 2016. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2021:NSA

[HL21] Baohua Huang and Wen Li.

Numerical subspace algorithms for solving the tensor equations involving Einstein product. Numerical Linear Algebra with Applications, 28(2): e2351:1–e2351:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hlavacek:1999:RSS

[Hla99] Ivan Hlaváček. Reliable solution of a Signorini contact problem with friction. Numerical Linear Algebra with Applications, 6(6):411–434, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/67501476/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501476&PLACEBO=IE.pdf.

Huang:2013:SGA

[HLLL13] Wei-Qiang Huang, Tiexiang Li, Yung-Ta Li, and Wen-Wei Lin. A semiorthogonal generalized Arnoldi method and its variations for quadratic eigenvalue problems. Numerical Linear Algebra with Applications, 20 (2):259–280, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hwang:2005:JDM

[HLLW05] Tsung-Min Hwang, Wen-Wei Lin, Jinn-Liang Liu, and Weichung Wang. Jacobi–Davidson methods for cubic eigenvalue problems. Numerical Linear Algebra with Applications, 12 (7):605–624, September 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Haase:1992:DDP

[HLM92] Gundolf Haase, Ulrich Langer, and Arnd Meyer. Domain decomposition preconditioners with inexact subdomain solvers. Journal of Numerical Linear Algebra with Applications, 1(1):27–41, 1992. CO-DEN ???? ISSN 0129-3281.

Harizanov:2018:OSL

[HLM^{+}18] S. Harizanov, R. Lazarov, S. Margenov, P. Marinov, and Y. Vutov. Optimal solvers for linear systems with fractional powers of sparse SPD matrices. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hansen:1996:PPS

[HM96] Per Christian Hansen and Klaus Mosegaard. Piecewise polynomial solutions without a priori break points. Numerical Linear Algebra with Applications, 3(6):513–524, November/December 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001006.

Hanna:2003:CMB

[HM03] Magdy Tawfik Hanna and Sana Ahmed Mansoori. A centrosymmetric matrix based technique for the interpolation of a Hermitian signal. Numerical Linear Algebra with Applications, 10(8):701–720, December 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2014:SPC

[HM14] N. Huang and C. F. Ma.

Some predictor–corrector-type iterative schemes for solving nonsymmetric algebraic Riccati equations arising in transport theory. Numerical Linear Algebra with Applications, 21(6):761–780, December 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2016:MBM

[HM16] Na Huang and Changfeng Ma.

The modulus-based matrix splitting algorithms for a class of weakly nonlinear complementarity problems. Numerical Linear Algebra with Applications, 23(3):558–569, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

He:2018:LFA

[HM18] Yunhui He and Scott P. MacLachlan. Local Fourier analysis of block-structured multigrid relaxation schemes for the Stokes equations. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

He:2020:TLF

[HM20] Yunhui He and Scott MacLachlan. Two-level Fourier analysis of multigrid for higher-order finite-element discretizations of the Laplacian. Numerical Linear Algebra with Applications, 27(3):e2285:1–e2285:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hochstenbach:2019:SMT

[HMMP19] Michiel E. Hochstenbach, Karl Meerbergen, Emre Mengi, and Bor Plestenjak. Subspace methods for three-parameter eigenvalue problems. Numerical Linear Algebra with Applications, 26(4):e2240:1– e2240:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Heuer:1999:PMR

[HMS99] Norbert Heuer, Matthias Maischak, and Ernst P. Stephan. Preconditioned minimum residual iteration for the h-p version of the coupled FEM/BEM with quasiuniform meshes. Numerical Linear Algebra with Applications, 6(6):435–456, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/67501478/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501478&PLACEBO=IE.pdf.

Ho:2005:SIC

[HN05] Man-Kiu Ho and Michael K.

Ng. Splitting iterations for circulant-plus-diagonal systems. Numerical Linear Algebra with Applications, 12(8): 779–792, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hessenthaler:2018:CMR

[HNR^{+}18] A. Hessenthaler, D. Nordsletten, O. Röhrle, J. B. Schroder, and R. D. Falgout. Convergence of the multigrid reduction in time algorithm for the linear elasticity equations. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Homke:2006:MMA

[Höm06] Lars Hömke. A multigrid method for anisotropic PDEs in elastic image registration. Numerical Linear Algebra with Applications, 13(2–3):215–229, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Howell:2018:PSM

[How18] Jason S. Howell. Prestructuring sparse matrices with dense rows and columns via null space methods. Numerical Linear Algebra with Applications, 25(2):??, March 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hackbusch:1997:DGS

[HP97] Wolfgang Hackbusch and Thomas Probst. Downwind Gauss–Seidel smoothing for convection dominated problems. Numerical Linear Algebra with Applications, 4(2): 85–102, March/April 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15001017;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=15001017&
                                        PLACEBO=IE.pdf.

Hoppe:2004:PDN

[HP04] R. H. W. Hoppe and S. I. Petrova. Primal-dual Newton interior point methods in shape and topology optimization. Numerical Linear Algebra with Applications, 11(5– 6):413–429, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Harbrecht:2003:MPC

[HPPS03] Helmut Harbrecht, Freddy Paiva, Cristian Pérez, and Reinhold Schneider. Multiscale preconditioning for the coupling of FEM-BEM. Numerical Linear Algebra with Applications, 10(3):197–222, April/ May 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Harbrecht:2015:EAR

[HPS15] Helmut Harbrecht, Michael Peters, and Markus Siebenmorgen. Efficient approximation of random fields for numerical applications. Numerical Linear Algebra with Applications, 22 (4):596–617, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Heinig:2005:SAS

[HR05] G. Heinig and K. Rost.

Split algorithms for symmetric Toeplitz matrices with arbitrary rank profile. Numerical Linear Algebra with Applications, 12(2–3):141–151, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2005:OTS

[HS05] Yifan Hu and Jennifer Scott.

Ordering techniques for singly bordered block diagonal forms for unsymmetric parallel sparse direct solvers. Numerical Linear Algebra with Applications, 12(9):877–894, November 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hochstenbach:2008:HRR

[HS08] Michiel E. Hochstenbach and Gerard L. G. Sleijpen. Harmonic and refined Rayleigh– Ritz for the polynomial eigenvalue problem. Numerical Linear Algebra with Applications, 15(1):35–54, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hayami:2011:GVK

[HS11] Ken Hayami and Masaaki Sugihara. A geometric view of Krylov subspace methods on singular systems. Numerical Linear Algebra with Applications, 18(3):449–469, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See corrigenda [HS14, HS21b].

Hogg:2013:EAP

[HS13] Jonathan D. Hogg and Jennifer A. Scott. An efficient analyse phase for element problems. Numerical Linear Algebra with Applications, 20(3): 397–412, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hayami:2014:CGV

[HS14] Ken Hayami and Masaaki Sugihara. Corrigendum to: A geometric view of Krylov subspace methods on singular systems. Numerical Linear Algebra with Applications, 21(5): 701–702, October 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [HS11].

Hogg:2015:USM

[HS15] Jonathan Hogg and Jennifer Scott. On the use of suboptimal matchings for scaling and ordering sparse symmetric matrices. Numerical Linear Algebra with Applications, 22(4): 648–663, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Heidel:2018:RTR

[HS18] Gennadij Heidel and Volker Schulz. A Riemannian trustregion method for low-rank tensor completion. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hao:2021:SMW

[HS21a] Yue Hao and Valeria Simoncini. The Sherman– Morrison–Woodbury formula for generalized linear matrix equations and applications. Numerical Linear Algebra with Applications, 28(5):e2384:1– e2384:??, October 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hayami:2021:CGV

[HS21b] Ken Hayami and Kota Sugihara. Corrigendum 2 to: A geometric view of Krylov subspace methods on singular systems. Numerical Linear Algebra with Applications, 28 (4):e2368:1–e2368:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [HS11].

Huckle:2005:PSN

[HSCTP05] Thomas Huckle, Stefano Serra-Capizzano, and Cristina Tablino-Possio. Preconditioning strategies for non-Hermitian Toeplitz linear systems. Numerical Linear Algebra with Applications, 12(2–3):211–220, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2022:SAD

[HST22] Jonathan J. Hu, Christopher M. Siefert, and Raymond S. Tuminaro. Smoothed aggregation for difficult stretched mesh and coefficient variation problems. Numerical Linear Algebra with Applications, 29 (6):e2442:1–e2442:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2018:RIM

[HSY18] R. Huang, J. Sun, and C. Yang.

Recursive integral method with Cayley transformation. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2012:PTN

[Hua12] Rong Huang. On parametrization of totally nonpositive matrices and applications. Numerical Linear Algebra with Applications, 19(4):742–753, August 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huckle:1998:ECS

[Huc98] Thomas K. Huckle. Efficient computation of sparse approximate inverses. Numerical Linear Algebra with Applications, 5(1):57–71, January/February 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5962;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5962&PLACEBO=IE.pdf.

Hernandez:2021:GAC

[HVCY21] Taylor M. Hernandez, Roel Van Beeumen, Mark A. Caprio, and Chao Yang. A greedy algorithm for computing eigenvalues of a symmetric matrix with localized eigenvectors. Numerical Linear Algebra with Applications, 28(2): e2341:1–e2341:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2016:TGS

[HVX16] Xiaozhe Hu, Panayot S. Vassilevski, and Jinchao Xu. A two-grid SA-AMG convergence bound that improves when increasing the polynomial degree. Numerical Linear Algebra with Applications, 23(4): 746–771, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2018:OCF

[HW18] Zheng-Da Huang and Hui-Di Wang. On the optimal convergence factor of the accelerated parameterized inexact Uzawa method with three parameters for augmented systems. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Heidel:2019:PBC

[HW19] Gennadij Heidel and Andy Wathen. Preconditioning for boundary control problems in incompressible fluid dynamics. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2021:CFP

[HW21] Jun Huang and Gang Wu. Convergence of the fixed-point iteration for multilinear PageRank. Numerical Linear Algebra with Applications, 28 (5):e2379:1–e2379:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2022:EFI

[HW22] Wen Huang and Ke Wei.

An extension of fast iterative shrinkage-thresholding algorithm to Riemannian optimization for sparse principal component analysis. Numerical Linear Algebra with Applications, 29(1):e2409:1– e2409:??, January 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Huang:2019:KSM

[HXM19] Baohua Huang, Yajun Xie, and Changfeng Ma. Krylov subspace methods to solve a class of tensor equations via the Einstein product. Numerical Linear Algebra with Applications, 26(4):e2254:1–e2254:??, August 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hong:2022:ANS

[HY22] Tao Hong and Irad Yavneh.

On adapting Nesterov’s scheme to accelerate iterative methods for linear problems. Numerical Linear Algebra with Applications, 29(2):e2417:1– e2417:??, March 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hu:2023:CIC

[HZZC23] Qingjie Hu, Hongrun Zhang, Zhijuan Zhou, and Yu Chen. A class of improved conjugate gradient methods for nonconvex unconstrained optimization. Numerical Linear Algebra with Applications, 30 (4):e2482:1–e2482:??, August 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Iannazzo:2016:GMT

[Ian16] Bruno Iannazzo. The geometric mean of two matrices from a computational viewpoint. Numerical Linear Algebra with Applications, 23(2): 208–229, March 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ibraghimov:2002:ATW

[Ibr02] Ilghiz Ibraghimov. Application of the three-way decomposition for matrix compression. Numerical Linear Algebra with Applications, 9(6–7):551–565, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ixaru:1996:RP

[IDVV96] L. Gr. Ixaru, H. De Meyer, G. Vanden Berghe, and M. Van Daele. A regularization procedure for ∑_{i=1}^{n} f_{i}(z_{j})x_{i} = g(z_{j})(j = 1, 2, …, n). Numerical Linear Algebra with Applications, 3(1):81–90, January/February 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000499.

Iwashita:2023:CAP

[IIFM23] Takeshi Iwashita, Kota Ikehara, Takeshi Fukaya, and Takeshi Mifune. Convergence acceleration of preconditioned conjugate gradient solver based on error vector sampling for a sequence of linear systems. Numerical Linear Algebra with Applications, 30(6):e2512:1– e2512:??, December 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ikramov:2000:BGF

[IK00] Khakim D. Ikramov and Andrey B. Kucherov. Bounding the growth factor in Gaussian elimination for Buckley’s class of complex symmetric matrices. Numerical Linear Algebra with Applications, 7(5): 269–274, July/August 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72508409/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72508409&PLACEBO=IE.pdf.

Ibrahim:2022:ADF

[IKAA22] Abdulkarim Hassan Ibrahim, Poom Kumam, Auwal Bala Abubakar, and Abubakar Adamu. Accelerated derivativefree method for nonlinear monotone equations with an application. Numerical Linear Algebra with Applications, 29(3):e2424:1–e2424:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Iannazzo:2013:SST

[IP13] Bruno Iannazzo and Federico Poloni. A subspace shift technique for nonsymmetric algebraic Riccati equations associated with an M-matrix. Numerical Linear Algebra with Applications, 20(3):440–452, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Imakura:2009:IWS

[ISZ09] A. Imakura, T. Sogabe, and S.-L. Zhang. An implicit wavelet sparse approximate inverse preconditioner using block finger pattern. Numerical Linear Algebra with Applications, 16(11-12):915–928, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ilic:2005:KSA

[IT05] M. Ilić and I. W. Turner.

Krylov subspaces and the analytic grade. Numerical Linear Algebra with Applications, 12 (1):55–76, February 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ilic:2007:LSS

[ITS07] M. Ilić, I. W. Turner, and Y. Saad. Linear system solution by null-space approximation and projection (SNAP). Numerical Linear Algebra with Applications, 14(1):61–82, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Iontcheva:2004:MMM

[IV04] Ana H. Iontcheva and Panayot S.

Vassilevski. Monotone multigrid methods based on element agglomeration coarsening away from the contact boundary for the Signorini’s problem. Numerical Linear Algebra with Applications, 11(2–3):189–204, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jia:1996:IIO

[Jia96] Zhongxiao Jia. On IOM(q):

The incomplete orthogonalization method for large unsymmetric linear systems. Numerical Linear Algebra with Applications, 3(6):491–512, November/December 1996. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15001005.

Jian:2017:EEI

[Jia17] Yuan Jian. Extremal eigenvalue intervals of symmetric tridiagonal interval matrices. Numerical Linear Algebra with Applications, 24(2):??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Johnson:2009:GRC

[JK09] Charles Johnson and Lev Krukier. General resolution of a convergence question of L. Krukier. Numerical Linear Algebra with Applications, 16(11-12):949–950, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jia:2017:RBS

[JK17] Zhongxiao Jia and Wenjie Kang. A residual based sparse approximate inverse preconditioning procedure for large sparse linear systems. Numerical Linear Algebra with Applications, 24(2):??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jozi:2018:WSV

[JK18] Meisam Jozi and Saeed Karimi. A weighted singular value decomposition for the discrete inverse problems. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Johnson:2009:STC

[JL09] Charles R. Johnson and Joshua A. Link. Solution theory for complete bilinear systems of equations. Numerical Linear Algebra with Applications, 16(11-12):929–934, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jin:2005:CPS

[JLW05] Xiao-Qing Jin, Siu-Long Lei, and Yi-Min Wei. Circulant preconditioners for solving singular perturbation delay differential equations. Numerical Linear Algebra with Applications, 12(2–3):327–336, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jun:2010:SOT

[JM10] Younbae Jun and Tsun-Zee Mai. Second-order treatment of the interface of domain decomposition method for parabolic problems. Numerical Linear Algebra with Applications, 17(5):839–853, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jarlebring:2018:KML

[JMPR18] Elias Jarlebring, Giampaolo Mele, Davide Palitta, and Emil Ringh. Krylov methods for low-rank commuting generalized Sylvester equations. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:1992:JNL

[JNL92] Journal of numerical linear algebra with applications, 1992. ISSN 0129-3281. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA.

Jia:2019:RQM

[JNS19] Zhigang Jia, Michael K. Ng, and Guang-Jing Song. Robust quaternion matrix completion with applications to image inpainting. Numerical Linear Algebra with Applications, 26 (4):e2245:1–e2245:??, August 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Joubert:1994:IIC

[JO94] Wayne Joubert and Thomas Oppe. Improved SSOR and incomplete Cholesky solution of linear equations on shared memory and distributed memory parallel computers. Numerical Linear Algebra with Applications, 1(3):287–311, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Janovska:2001:NHT

[JO01] Dáša Janovská and Gerhard Opfer. A note on hyperbolic transformations. Numerical Linear Algebra with Applications, 8(2):127–146, March 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76506692/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76506692&PLACEBO=IE.pdf.

Joubert:1994:CBR

[Jou94]  Wayne Joubert.  On the

convergence behavior of the restarted GMRES algorithm for solving nonsymmetric linear systems. Numerical Linear Algebra with Applications, 1(5):427–447, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jagels:1994:FMR

[JR94] Carl F. Jagels and Lothar Reichel. A fast minimal residual algorithm for shifted unitary matrices. Numerical Linear Algebra with Applications, 1(6):555–570, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jones:1996:TSM

[JS96] Mark T. Jones and Daniel B.

Szyld. Two-stage multisplitting methods with overlapping blocks. Numerical Linear Algebra with Applications, 3(2): 113–124, March/April 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000983.

Jing:2017:SGA

[JYH17] Yan-Fei Jing, Pei Yuan, and Ting-Zhu Huang. A simpler GMRES and its adaptive variant for shifted linear systems. Numerical Linear Algebra with Applications, 24(1):??, January 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jiang:2017:TTC

[JYZ17] Bo Jiang, Fan Yang, and Shuzhong Zhang. Tensor and its Tucker core: The invariance relationships. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jia:2009:PSA

[JZ09] Zhongxiao Jia and Baochen Zhu. A power sparse approximate inverse preconditioning procedure for large sparse linear systems. Numerical Linear Algebra with Applications, 16(4):259–299, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Jiang:2011:DAS

[JZ11] Ying-Jun Jiang and Jin-Ping Zeng. Direct algorithm for the solution of two-sided obstacle problems with M-matrix. Numerical Linear Algebra with Applications, 18(1):167–173, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ke:2017:BTP

[KABH17] Guoyi Ke, Eugenio Aulisa, Giorgio Bornia, and Victoria Howle. Block triangular preconditioners for linearization schemes of the Rayleigh– Bénard convection problem. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kaporin:1994:NCR

[Kap94] I. E. Kaporin. New convergence results and preconditioning strategies for the conjugate gradient method. Numerical Linear Algebra with Applications, 1(2):179–210, 1994. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kaporin:1998:HQP

[Kap98] Igor E. Kaporin. High quality preconditioning of a general symmetric positive definite matrix based on its UTU +UTR+ RTU-decomposition. Numerical Linear Algebra with Applications, 5(6):483–509, November/December 1998. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002968;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002968&
PLACEBO=IE.pdf.

Kaporin:1999:PAF

[Kap99] Igor Kaporin. A practical algorithm for faster matrix multiplication. Numerical Linear Algebra with Applications, 6 (8):687–700, December 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/69000703/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=69000703&PLACEBO=IE.pdf.

Kaporin:2002:UMO

[Kap02] I. E. Kaporin. Using the modified 2nd order incomplete Cholesky decomposition as the conjugate gradient preconditioning. Numerical Linear Algebra with Applications, 9(6–7):401–408, September/ November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kaporin:2005:SCM

[Kap05] I. Kaporin. Superlinear convergence in minimum residual iterations. Numerical Linear Algebra with Applications, 12(5– 6):453–470, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kaufman:2007:RAF

[Kau07]  Linda Kaufman.  The re-

traction algorithm for factoring banded symmetric matrices. Numerical Linear Algebra with Applications, 14(3): 237–254, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kolberg:2015:EAS

[KBF15] Mariana Kolberg, Gerd Bohlender, and Luiz Gustavo Fernandes. An efficient approach to solve very large dense linear systems with verified computing on clusters. Numerical Linear Algebra with Applications, 22(2):299–316, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostic:2017:IBM

[KC17] V. R. Kostić and Lj. Cvetković.

On the inertia of the block H-matrices. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostic:2016:PLT

[KCC16] V. R. Kostić, Lj. Cvetković, and D. Lj. Cvetković. Pseudospectra localizations and their applications. Numerical Linear Algebra with Applications, 23(2):356–372, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kokiopoulou:2011:TOE

[KCS11] E. Kokiopoulou, J. Chen, and Y. Saad. Trace optimization and eigenproblems in dimension reduction methods. Numerical Linear Algebra with Applications, 18(3):565–602, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostic:2009:GTL

[KCV09] V. Kostić, L. J. Cvetković, and R. S. Varga. Geršgorin-type localizations of generalized eigenvalues. Numerical Linear Algebra with Applications, 16(11-12):883–898, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kemper:2012:LTB

[Kem12] Jens Kemper. Long-time behavior of the proper orthogonal decomposition method. Numerical Linear Algebra with Applications, 19(5):842–868, October 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kiskiras:2007:NCS

[KH07] J. Kiskiras and G. D. Halikias. A note on the complex semi-definite matrix Procrustes problem. Numerical Linear Algebra with Applications, 14(6): 485–502, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kalantzis:2023:EAS

[KH23] Vassilis Kalantzis and Lior Horesh. Enhanced algebraic substructuring for symmetric generalized eigenvalue problems. Numerical Linear Algebra with Applications, 30(2): e2473:1–e2473:??, March 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskij:1996:FCI

[Kho96] Boris N. Khoromskij. On fast computations with the inverse to harmonic potential operators via domain decomposition. Numerical Linear Algebra with Applications, 3 (2):91–111, March/April 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000986.

Kong:2012:SBR

[KJ12] Xu Kong and Yao-Lin Jiang.

Subtracting a best rank-1 approximation from p × p × 2(p ≥ 2) tensors. Numerical Linear Algebra with Applications, 19 (3):503–523, May 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kaporin:2002:PBO

[KK02] Igor E. Kaporin and Igor N. Konshin. A parallel block overlap preconditioning with inexact submatrix inversion for linear elasticity problems. Numerical Linear Algebra with Applications, 9(2): 141–162, March 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/89013913/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=89013913&PLACEBO=IE.pdf.

Kollmann:2013:PMS

[KK13] Markus Kollmann and Michael Kolmbauer. A preconditioned MinRes solver for time-periodic parabolic optimal control problems. Numerical Linear Algebra with Applications, 20(5): 761–784, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskaia:2016:FTM

[KK16] Venera Khoromskaia and Boris N. Khoromskij. Fast tensor method for summation of longrange potentials on 3D lattices with defects. Numerical Linear Algebra with Applications, 23 (2):249–271, March 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskij:2023:ETN

[KK23a] Boris N. Khoromskij and Venera Khoromskaia. Editorial: Tensor numerical methods and their application in scientific computing and data science. Numerical Linear Algebra with Applications, 30(3):e2493:1– e2493:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskij:2023:FST

[KK23b] Boris N. Khoromskij and Venera Khoromskaia. Fast solution of three-dimensional elliptic equations with randomly generated jumping coefficients by using tensor-structured preconditioners. Numerical Linear Algebra with Applications, 30(3):e2477:1–e2477:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kucera:2012:MPI

[KKMM12] R. Kučera, T. Kozubek, A. Markopoulos, and J. Machalová. On the Moore–Penrose inverse in solving saddle-point systems with singular diagonal blocks. Numerical Linear Algebra with Applications, 19(4): 677–699, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kharchenko:2001:RAT

[KKNY01] S. A. Kharchenko, L. Yu. Kolotilina, A. A. Nikishin, and A. Yu. Yeremin. A robust AINV-type method for constructing sparse approximate inverse preconditioners in factored form. Numerical Linear Algebra with Applications, 8(3):165–179, April/ May 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76506670/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76506670&PLACEBO=IE.pdf.

Khoromskaia:2020:NSS

[KKO20] Venera Khoromskaia, Boris N. Khoromskij, and Felix Otto. Numerical study in stochastic homogenization for elliptic partial differential equations: Convergence rate in the size of representative volume elements. Numerical Linear Algebra with Applications, 27(3): e2296:1–e2296:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kalantzis:2018:DDA

[KKPS18] Vassilis Kalantzis, James Kestyn, Eric Polizzi, and Yousef Saad. Domain decomposition approaches for accelerating contour integration eigenvalue solvers for symmetric eigenvalue problems. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Krukier:2014:GSH

[KKR14] Lev A. Krukier, Boris L.

Krukier, and Zhi-Ru Ren. Generalized skew-Hermitian triangular splitting iteration methods for saddle-point linear systems. Numerical Linear Algebra with Applications, 21(1): 152–170, January 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kandolf:2021:CLR

[KKRS21] Peter Kandolf, Antti Koskela, Samuel D. Relton, and Marcel Schweitzer. Computing low-rank approximations of the Fréchet derivative of a matrix function using Krylov subspace methods. Numerical Linear Algebra with Applications, 28 (6):e2401:1–e2401:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskij:2019:QCA

[KKS19] Boris N. Khoromskij, N. Kishore Kumar, and Jan Schneider. Quantized CP approximation and sparse tensor interpolation of function-generated data. Numerical Linear Algebra with Applications, 26(5):e2262:1– e2262:??, October 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kim:2006:FOS

[KLM^{+}06] Sang Dong Kim, Chang-Ock Lee, Thomas A. Manteuffel, Stephen F. McCormick, and Oliver Röhrle. First-order system least squares for the Oseen equations. Numerical Linear Algebra with Applications, 13(7):523–542, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2014:RMM

[KLM14] Johannes Kraus, Maria Lymbery, and Svetozar Margenov. Robust multilevel methods for quadratic finite element anisotropic elliptic problems. Numerical Linear Algebra with Applications, 21(3):375–398, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2015:ASM

[KLM15] J. Kraus, M. Lymbery, and S. Margenov. Auxiliary space multigrid method based on additive Schur complement approximation. Numerical Linear Algebra with Applications, 22(6):965–986, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kressner:2021:CRB

[KLMP21] Daniel Kressner, Kathryn Lund, Stefano Massei, and Davide Palitta. Compress-and-restart block Krylov subspace methods for Sylvester matrix equations. Numerical Linear Algebra with Applications, 28 (1):e2339:1–e2339:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Krizek:1999:PPG

[KLN99] Michal Křížek, Liping Liu, and Pekka Neittaanmäki. Postprocessing of Gauss–Seidel iterations. Numerical Linear Algebra with Applications, 6(2):147–156, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002995;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002995&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Kogler:2023:CAS

[KLS23] Lukas Kogler, Philip L. Lederer, and Joachim Schöberl. A conforming auxiliary space preconditioner for the mass conserving stress-yielding method. Numerical Linear Algebra with Applications, 30(5):e2503:1– e2503:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kutcherov:1992:AFM

[KM92] Andrew B. Kutcherov and Michail M. Makarov. An approximate factorization method for solving discrete elliptic problems on stretched domains. Journal of Numerical Linear Algebra with Applications, 1(1):1–26, 1992. CODEN ???? ISSN 0129-3281.

Kolev:1999:TLP

[KM99] Tzanio V. Kolev and Svetozar D. Margenov. Twolevel preconditioning of pure displacement non-conforming FEM systems. Numerical Linear Algebra with Applications, 6(7):533–555, October/November 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/68502740/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=68502740&PLACEBO=IE.pdf.

Kravvaritis:2009:GFH

[KM09] Christos Kravvaritis and Marilena Mitrouli. The growth factor of a Hadamard matrix of order 16 is 16. Numerical Linear Algebra with Applications, 16(9):715–743, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostic:2016:ACM

[KMC16] Vladimir R. Kostić, Agnieszka MiȨdlar, and Ljiljana Cvetković. An algorithm for computing minimal Geršgorin sets. Numerical Linear Algebra with Applications, 23(2): 272–290, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kalchev:2018:MLS

[KMM18] Delyan Z. Kalchev, Thomas A.

Manteuffel, and Steffen Münzenmaier. Mixed and least-squares finite element methods with application to linear hyperbolic problems. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kuchta:2019:SNP

[KMM19] Miroslav Kuchta, Kent-Andre Mardal, and Mikael Mortensen. On the singular Neumann problem in linear elasticity. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ketelsen:2010:FEM

[KMMR10] C. Ketelsen, T. Manteuffel, S. McCormick, and J. Ruge. Finite element methods for quantum electrodynamics using a Helmholtz decomposition of the gauge field. Numerical Linear Algebra with Applications, 17(2-3):539–556, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2008:MPC

[KMS08] J. Kraus, S. Margenov, and J. Synka. On the multilevel preconditioning of Crouzeix– Raviart elliptic problems. Numerical Linear Algebra with Applications, 15(5):395–416, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kilmer:2007:KPA

[KN07] M. E. Kilmer and J. G. Nagy. Kronecker product approximations for dense block Toeplitzplus-Hankel matrices. Numerical Linear Algebra with Applications, 14(8):581–602, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kindermann:2014:NAT

[KN14] Stefan Kindermann and Carmeliza Navasca. New algorithms for tensor decomposition based on a reduced functional. Numerical Linear Algebra with Applications, 21(3):340–374, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kwak:2003:DDM

[KNP03] D. Y. Kwak, S. V. Nepomnyaschikh, and H. C. Pyo. Domain decomposition for model heterogeneous anisotropic problem. Numerical Linear Algebra with Applications, 10(1–2): 129–157, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kirkland:2001:DCA

[KNX01] Stephen J. Kirkland, Michael Neumann, and Jianhong Xu. A divide and conquer approach to computing the mean first passage matrix for Markov chains via Perron complement reductions. Numerical Linear Algebra with Applications, 8(5): 287–295, July/August 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/80502039/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=80502039&PLACEBO=IE.pdf.

Kolotilina:1999:FSA

[KNY99] L. Yu. Kolotilina, A. A. Nikishin, and A. Yu. Yeremin. Factorized sparse approximate inverse preconditionings. IV: Simple approaches to rising efficiency. Numerical Linear Algebra with Applications, 6(7): 515–531, October/November 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/68502739/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=68502739&PLACEBO=IE.pdf.

Kolotilina:2000:IFA

[KNY00] L. Yu. Kolotilina, A. A. Nikishin, and A. Yu. Yeremin. An incomplete LU-factorization algorithm based on block bordering. Numerical Linear Algebra with Applications, 7(7–8): 543–567, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505471/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505471&PLACEBO=IE.pdf.

Kolesnikov:2018:CAP

[KO18] D. A. Kolesnikov and I. V. Oseledets. Convergence analysis of projected fixed-point iteration on a low-rank matrix manifold. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kolotilina:2005:BEB

[Kol05] L. Yu. Kolotilina. Bounds for the eigenvalues of block 2 × 2 Hermitian positive-definite matrices. Numerical Linear Algebra with Applications, 12(5– 6):393–417, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Konshin:2017:FPS

[KOV17] Igor Konshin, Maxim Olshanskii, and Yuri Vassilevski. LU factorizations and ILU preconditioning for stabilized discretizations of incompressible Navier–Stokes equations. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Klawonn:2000:COS

[KP00] Axel Klawonn and Luca F.

Pavarino. A comparison of overlapping Schwarz methods and block preconditioners for saddle point problems. Numerical Linear Algebra with Applications, 7(1): 1–25, January/February 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71001527/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71001527&PLACEBO=IE.pdf.

Kuznetsov:2010:NMA

[KP10] Yu Kuznetsov and A. Prokopenko.

A new multilevel algebraic preconditioner for the diffusion equation in heterogeneous media. Numerical Linear Algebra with Applications, 17(5): 759–769, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kressner:2014:PLR

[KPT14] Daniel Kressner, Martin Plešinger, and Christine Tobler. A preconditioned low-rank CG method for parameter-dependent Lyapunov matrix equations. Numerical Linear Algebra with Applications, 21(5):666–684, October 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kim:2006:MIC

[KPV06] T. Kim, J. E. Pasciak, and P. S. Vassilevski. Mesh-independent convergence of the modified inexact Newton method for a second order non-linear problem. Numerical Linear Algebra with Applications, 13(1): 23–47, February 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kolev:2008:CAM

[KPV08] Tzanio V. Kolev, Joseph E. Pasciak, and Panayot S. Vassilevski. H (curl) auxiliary mesh preconditioning. Numerical Linear Algebra with Applications, 15(5):455–471, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostler:2006:AMS

[KR06] H. Köstler and U. Rüde. An accurate multigrid solver for computing singular solutions of elliptic problems. Numerical Linear Algebra with Applications, 13(2–3):231–249, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kimmritz:2011:PMM

[KR11] Madlen Kimmritz and Thomas Richter. Parallel multigrid method for finite element simulations of complex flow problems on locally refined meshes. Numerical Linear Algebra with Applications, 18(4): 615–636, August 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kressner:2014:MEA

[KR14] Daniel Kressner and Jose E.

Roman. Memory-efficient Arnoldi algorithms for linearizations of matrix polynomials in Chebyshev basis. Numerical Linear Algebra with Applications, 21(4):569–588, August 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2002:APM

[Kra02]  J. K. Kraus.  An alge-

braic preconditioning method for M-matrices: linear versus non-linear multilevel iteration. Numerical Linear Algebra with Applications, 9(8): 599–618, December 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2006:AMP

[Kra06] J. K. Kraus. Algebraic multilevel preconditioning of finite element matrices using local Schur complements. Numerical Linear Algebra with Applications, 13(1):49–70, February 2006. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostler:2008:MSO

[KRW08] H. Köstler, K. Ruhnau, and R. Wienands. Multigrid solution of the optical flow system using a combined diffusionand curvature-based regularizer. Numerical Linear Algebra with Applications, 15(2-3): 201–218, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Krzyzanowski:2011:BPS

[Krz11] Piotr Krzyżanowski. On block preconditioners for saddle point problems with singular or indefinite (1, 1) block. Numerical Linear Algebra with Applications, 18(1):123–140, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Klawonn:2004:PEL

[KS04] Axel Klawonn and Gerhard Starke. A preconditioner for the equations of linear elasticity discretized by the PEERS element. Numerical Linear Algebra with Applications, 11(5– 6):493–510, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Knizhnerman:2010:NIE

[KS10] L. Knizhnerman and V. Simoncini. A new investigation of the extended Krylov subspace method for matrix function evaluations. Numerical Linear Algebra with Applications, 17 (4):615–638, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kressner:2015:TLR

[KS15] Daniel Kressner and Petar Sirković. Truncated low-rank methods for solving general linear matrix equations. Numerical Linear Algebra with Applications, 22(3):564–583, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kogler:2022:AMM

[KS22] Lukas Kogler and Joachim Schöberl. An algebraic multigrid method for elasticity based on an auxiliary topology with edge matrices. Numerical Linear Algebra with Applications, 29(1):e2408:1–e2408:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Keel:2013:RHS

[KSB13] L. H. Keel, T. F. Stratton, and S. P. Bhattacharyya. Robust Hurwitz and Schur stability via interval positivity. Numerical Linear Algebra with Applications, 20(2):281–290, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kraus:2008:MMD

[KT08] J. K. Kraus and S. K. Tomar. A multilevel method for discontinuous Galerkin approximation of three-dimensional anisotropic elliptic problems. Numerical Linear Algebra with Applications, 15(5):417– 438, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kublanovskaya:1992:RDA

[Kub92] V. N. Kublanovskaya. Rank division “algorithms and their applications”. Journal of Numerical Linear Algebra with Applications, 1(2):199–213, 1992. CODEN ???? ISSN 0129-3281.

Kuznetsov:1992:NPA

[Kuz92] Yu. A. Kuznetsov. A new parallel algebraic preconditioner. Journal of Numerical Linear Algebra with Applications, 1 (2):215–225, 1992. CODEN ???? ISSN 0129-3281.

Kaagstrom:1992:GSS

[KV92] Bo Kågström and Paul Van Dooren. A generalized statespace approach for the additive decomposition of a transfer matrix. Journal of Numerical Linear Algebra with Applications, 1(2):165–181, 1992. CODEN ???? ISSN 0129-3281.

Krizkova:1996:TLP

[KV96] Jitka Křížková and Petr Vaněk. Two-level preconditioner with small coarse grid appropriate for unstructured meshes. Numerical Linear Algebra with Applications, 3(4): 255–274, July/August 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000996.

Kolev:2006:AEA

[KV06] Tzanio V. Kolev and Panayot S.

Vassilevski. AMG by element agglomeration and constrained energy minimization interpolation. Numerical Linear Algebra with Applications, 13(9): 771–788, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kuzel:2015:EIS

[KV15] Roman Kužel and Petr Vaněk.

Exact interpolation scheme with approximation vector used as a column of the prolongator. Numerical Linear Algebra with Applications, 22(6): 950–964, December 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kostic:2012:LGE

[KVC12] V. Kostić, R. S. Varga, and L. Cvetković. Localization of generalized eigenvalues by Cartesian ovals. Numerical Linear Algebra with Applications, 19(4):728–741, August 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kincaid:2010:LTD

[KVW10] David R. Kincaid, Richard S. Varga, and Charles H. Warlick. The life and times of Dr David M. Young, Jr. Numerical Linear Algebra with Applications, 17(5):743–757, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Khoromskij:1999:RSC

[KW99] Boris N. Khoromskij and Gabriel Wittum. Robust Schur complement method for strongly anisotropic elliptic equations. Numerical Linear Algebra with Applications, 6 (8):621–653, December 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/69000701/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=69000701&PLACEBO=IE.pdf.

Kong:2018:FCT

[KWS^{+}18] Fande Kong, Yaqi Wang, Sebastian Schunert, John W. Peterson, Cody J. Permann, David Andrš, and Richard C. Martineau. A fully coupled two-level Schwarz preconditioner based on smoothed aggregation for the transient multigroup neutron diffusion equations. Numerical Linear Algebra with Applications, 25 (3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kim:2003:MMB

[KXZ03] HwanHo Kim, Jinchao Xu, and Ludmil Zikatanov. A multigrid method based on graph matching for convectiondiffusion equations. Numerical Linear Algebra with Applications, 10(1–2):181–195, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kharchenko:1995:ETB

[KY95] S. A. Kharchenko and A. Yu.

Yeremin. Eigenvalue translation based preconditioners for the GMRES(k) method. Numerical Linear Algebra with Applications, 2(1):51–77, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lai:1997:SCN

[Lai97] Chen-Yao G. Lai. Short communication: a note on optimal hybrid V -cycle multilevel algorithms for mixed finite element systems with penalty term. Numerical Linear Algebra with Applications, 4(6): 491–498, November/December 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15047;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15047&PLACEBO=IE.pdf.

Lambers:2012:EHO

[Lam12] James V. Lambers. Explicit high-order time stepping based on componentwise application of asymptotic block Lanczos iteration. Numerical Linear Algebra with Applications, 19(6): 970–991, December 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Langer:1997:E

[Lan97]  Ulrich Langer.  Editorial.

Numerical Linear Algebra with Applications, 4(3):131, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15034;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15034&PLACEBO=IE.pdf.

Layton:2005:MRC

[Lay05] W. Layton. Model reduction by constraints, discretization of flow problems and an induced pressure stabilization. Numerical Linear Algebra with Applications, 12(5–6):547–562, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lazzaro:2016:NAL

[Laz16] Damiana Lazzaro. A nonconvex approach to low-rank matrix completion using convex optimization. Numerical Linear Algebra with Applications, 23(5):801–824, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2008:CAN

[LB08] Yiqin Lin and Liang Bao.

Convergence analysis of the Newton–Shamanskii method for a nonsymmetric algebraic Riccati equation. Numerical Linear Algebra with Applications, 15(6):535–546, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2017:UAP

[LB17] Geunseop Lee and Jesse Barlow. Updating approximate principal components with applications to template tracking. Numerical Linear Algebra with Applications, 24(2): ??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2021:AMN

[LB21] Barry Lee and Enrique Pereira Batista. Algebraic multigrid for the nonlinear powerflow equations. Numerical Linear Algebra with Applications, 28(2):e2347:1–e2347:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Leon:2013:GSO

[LBG13] Steven J. Leon, Åke Björck, and Walter Gander. Gram– Schmidt orthogonalization: 100 years and more. Numerical Linear Algebra with Applications, 20(3):492–532, May 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2005:ITP

[LC05] Fu-Rong Lin and Wai-Ki Ching. Inverse Toeplitz preconditioners for Hermitian Toeplitz systems. Numerical Linear Algebra with Applications, 12(2–3):221–229, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2007:NBE

[LC07] Xinguo Liu and Chaoxian Chen. A note on backward errors for Toeplitz systems. Numerical Linear Algebra with Applications, 14(7): 547–562, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2013:PAL

[LC13] T. Li and E. K-W. Chu. Pole assignment for linear and quadratic systems with timedelay in control. Numerical Linear Algebra with Applications, 20(2):291–301, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2021:SPP

[LC21] Jingwei Li and Xiao-Chuan Cai. Summation pollution of principal component analysis and an improved algorithm for location sensitive data. Numerical Linear Algebra with Applications, 28 (5):e2370:1–e2370:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2018:CDS

[LCHH18] Junyuan Lin, Lenore J. Cowen, Benjamin Hescott, and Xiaozhe Hu. Computing the diffusion state distance on graphs via algebraic multigrid and random projections. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2013:PBP

[LCN13] Wen Li, Lu-Bin Cui, and Michael K. Ng. The perturbation bound for the Perron vector of a transition probability tensor. Numerical Linear Algebra with Applications, 20(6): 985–1000, December 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2021:GTI

[LCZZ21] Jinkui Liu, Haisong Cao, Yongxiang Zhao, and Liqiao Zhang. A gradient-type iterative method for impulse noise removal. Numerical Linear Algebra with Applications, 28 (4):e2358:1–e2358:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lukas:2007:OMP

[LD07] D. Lukáš and Z. Dostál. Optimal multigrid preconditioned semi-monotonic augmented Lagrangians applied to the Stokes problem. Numerical Linear Algebra with Applications, 14(9):741–750, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Luo:2008:CPS

[LD08] X. Y. Luo and C. Davies. Call for papers: Special issue on ‘Flow in Collapsible Tubes or Over Compliant Surfaces for Biomedical Applications’, Communications in Numerical Methods in Engineering (CNM). Numerical Linear Algebra with Applications, 15(4):391, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

LeBorne:2023:BCL

[Le 23] Sabine Le Borne. A block Cholesky–LU-based QR factorization for rectangular matrices. Numerical Linear Algebra with Applications, 30 (5):e2497:1–e2497:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Leblond:2002:SAM

[Leb02]  Michel Leblond.  H-self-

adjoint matrices. Application to radiosity. Numerical Linear Algebra with Applications, 9(2):181–193, March 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/89013908/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=89013908&PLACEBO=IE.pdf.

Lee:2010:USA

[Lee10] Young-Ju Lee. Uniform stability analysis of Austin, Manteuffel and McCormick finite elements and fast and robust iterative methods for the Stokeslike equations. Numerical Linear Algebra with Applications, 17(1):109–138, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2012:SAE

[Lee12] B. Lee. Space-angle-energy multigrid methods for S_{n} discretizations of the multienergetic Boltzmann equation. Numerical Linear Algebra with Applications, 19(4): 773–795, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2016:PPM

[Lee16] Barry Lee. Parallel preconditioners and multigrid solvers for stochastic polynomial chaos discretizations of the diffusion equation at the large scale. Numerical Linear Algebra with Applications, 23(1):5–36, January 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2018:MMR

[Lee18] B. Lee. Multigrid for model reduction of power grid networks. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2019:MSO

[Lee19] B. Lee. Multigrid for secondorder ADN elliptic systems. Numerical Linear Algebra with Applications, 26(5):e2256:1– e2256:??, October 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2021:AMS

[Lee21a] Barry Lee. Algebraic multigrid for systems of elliptic boundary-value problems. Numerical Linear Algebra with Applications, 28(3):e2303:1– e2303:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2021:SAU

[Lee21b] Barry Lee. Spectral analysis for uncertainty quantification. Numerical Linear Algebra with Applications, 28 (4):e2354:1–e2354:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2012:PNS

[LGS12] Dan Li, Chen Greif, and Dominik Schötzau. Parallel numerical solution of the timeharmonic Maxwell equations in mixed form. Numerical Linear Algebra with Applications, 19(3):525–539, May 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2008:ECN

[LH08] Zi-Cai Li and Hung-Tsai Huang. Effective condition number for numerical partial differential equations. Numerical Linear Algebra with Applications, 15(7):575–594, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ling:2017:AEE

[LH17] Yonghui Ling and Zhengda Huang. An analysis on the efficiency of Euler’s method for computing the matrix p th root. Numerical Linear Algebra with Applications, 24(6): ??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2007:SSM

[LHL07a] Hou-Biao Li, Ting-Zhu Huang, and Hong Li. On some subclasses of P -matrices. Numerical Linear Algebra with Applications, 14(5):391–405, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2007:MHS

[LHL07b] Liang Li, Ting-Zhu Huang, and Xing-Ping Liu. Modified Hermitian and skew-Hermitian splitting methods for non-Hermitian positive-definite linear systems. Numerical Linear Algebra with Applications, 14 (3):217–235, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See errata [SB12].

Li:2007:OGT

[LHLS07] Hou-Biao Li, Ting-Zhu Huang, Hong Li, and Shu-Qian Shen. Optimal Gerschgorin-type inclusion intervals of singular values. Numerical Linear Algebra with Applications, 14(2): 115–128, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2011:ICT

[LHW11] Zi-Cai Li, Hung-Tsai Huang, and Yimin Wei. Ill-conditioning of the truncated singular value decomposition, Tikhonov regularization and their applications to numerical partial differential equations. Numerical Linear Algebra with Applications, 18(2):205–221, March 2011. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2000:CBC

[Li00] Xiezhang Li. Comparison between the convergence rates of the Chebyshev method and the related (2, 2)-step methods. Numerical Linear Algebra with Applications, 7(4):169–180, May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72507876/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72507876&PLACEBO=IE.pdf.

Linsenmann:2012:CRT

[Lin12] Christopher Linsenmann. On the convergence of right transforming iterations for the numerical solution of PDEconstrained optimization problems. Numerical Linear Algebra with Applications, 19(4): 621–638, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2022:STD

[Liu22] Haixia Liu. Symmetric tensor decomposition by alternating gradient descent. Numerical Linear Algebra with Applications, 29(1):e2406:1– e2406:??, January 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Livne:2004:CCR

[Liv04a] O. E. Livne. Coarsening by compatible relaxation. Numerical Linear Algebra with Applications, 11(2–3):205–227, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Livshits:2004:AMW

[Liv04b] Irene Livshits. An algebraic multigrid wave-ray algorithm to solve eigenvalue problems for the Helmholtz operator. Numerical Linear Algebra with Applications, 11(2–3):229–239, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Livshits:2014:SMM

[Liv14]  Ira Livshits.  A scalable

multigrid method for solving indefinite Helmholtz equations with constant wave numbers. Numerical Linear Algebra with Applications, 21(2): 177–193, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lei:2004:BPS

[LJ04] Siu-Long Lei and Xiao-Qing Jin. BCCB preconditioners for systems of BVM-based numerical integrators. Numerical Linear Algebra with Applications, 11(1):25–40, February 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2014:HGA

[LJM14] Cao Lu, Xiangmin Jiao, and Nikolaos Missirlis. A hybrid geometric + algebraic multigrid method with semiiterative smoothers. Numerical Linear Algebra with Applications, 21(2):221–238, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Longley:1997:AGS

[LL97] James W. Longley and Roger D.

Longley. Accuracy of Gram– Schmidt orthogonalization and Householder transformation for the solution of linear least squares problems. Numerical Linear Algebra with Applications, 4(4):295–303, July/August 1997. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15036;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15036&PLACEBO=IE.pdf.

Li:2014:NEI

[LLK14] Chaoqian Li, Yaotang Li, and Xu Kong. New eigenvalue inclusion sets for tensors. Numerical Linear Algebra with Applications, 21(1):39–50, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lai:1997:III

[LLL97] Yu-Ling Lai, Kun-Yi Lin, and Wen-Wei Lin. An inexact inverse iteration for large sparse eigenvalue problems. Numerical Linear Algebra with Applications, 4(5): 425–437, September/October 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15044;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15044&PLACEBO=IE.pdf.

Li:2016:NFD

[LLLJ16] Shengguo Li, Xiangke Liao, Jie Liu, and Hao Jiang. New fast divide-and-conquer algorithms for the symmetric tridiagonal eigenvalue problem. Numerical Linear Algebra with Applications, 23(4):656–673, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2017:UMP

[LLNV17] Wen Li, Dongdong Liu, Michael K. Ng, and Seak-Weng Vong. The uniqueness of multilinear PageRank vectors. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lai:2023:AAF

[LLPC23] Fuqi Lai, Wen Li, Xiaofei Peng, and Yannan Chen. Anderson accelerated fixed-point iteration for multilinear PageRank. Numerical Linear Algebra with Applications, 30 (5):e2499:1–e2499:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lee:2012:FET

[LLS12] Spike T. Lee, Xin Liu, and Hai-Wei Sun. Fast exponential time integration scheme for option pricing with jumps. Numerical Linear Algebra with Applications, 19(1):87–101, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2019:RMS

[LLV19] Dongdong Liu, Wen Li, and Seak-Weng Vong. Relaxation methods for solving the tensor equation arising from the higher-order Markov chains. Numerical Linear Algebra with Applications, 26(5):e2260:1– e2260:??, October 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2009:ECN

[LLW09] Zi-Cai Li, Tzon-Tzer Lu, and Yimin Wei. Effective condition number of Treffitz methods for biharmonic equations with crack singularities. Numerical Linear Algebra with Applications, 16(2):145–171, ???? 2009. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-

1506 (electronic).

Limon:2006:MAS

[LM06] Alfonso Limon and Hedley Morris. A multilevel adaptive solver based on secondgeneration wavelet thresholding techniques. Numerical Linear Algebra with Applications, 13(2–3):251–273, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Loe:2022:TEP

[LM22] Jennifer A. Loe and Ronald B. Morgan. Toward efficient polynomial preconditioning for GMRES. Numerical Linear Algebra with Applications, 29 (4):e2427:1–e2427:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lube:2000:NNO

[LMM00] G. Lube, L. Müller, and H. Müller. A new nonoverlapping domain decomposition method for stabilized finite element methods applied to the non-stationary Navier– Stokes equations. Numerical Linear Algebra with Applications, 7(6):449–472, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516697/
                                        START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=72516697&PLACEBO=IE.pdf.

Litvinenko:2023:CDD

[LMM^{+}23] Alexander Litvinenko, Youssef Marzouk, Hermann G. Matthies, Marco Scavino, and Alessio Spantini. Computing fdivergences and distances of high-dimensional probability density functions. Numerical Linear Algebra with Applications, 30(3):e2467:1–e2467:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Luksan:2004:IPM

[LMV04] Ladislav Lukšan, Ctirad Matonoha, and Jan Vlček. Interiorpoint method for non-linear non-convex optimization. Numerical Linear Algebra with Applications, 11(5–6):431–453, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2012:CSD

[LNP12] Fang Li, Michael K. Ng, and Robert J. Plemmons. Coupled segmentation and denoising/deblurring models for hyperspectral material identification. Numerical Linear Algebra with Applications, 19(1): 153–173, January 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lim:2013:STT

[LNQ13] Lek-Heng Lim, Michael K. Ng, and Liqun Qi. The spectral theory of tensors and its applications. Numerical Linear Algebra with Applications, 20(6): 889–890, December 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2015:MFB

[LNY15] Xinxin Li, Michael K. Ng, and Xiaoming Yuan. Median filtering-based methods for static background extraction from surveillance video. Numerical Linear Algebra with Applications, 22(5):845–865, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

LeBorne:2013:RER

[LO13] S. Le Borne and J. S. Ovall. Rapid error reduction for block Gauss–Seidel based on p-hierarchical basis. Numerical Linear Algebra with Applications, 20(5):743–760, October 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Litsarev:2015:FLR

[LO15] M. S. Litsarev and I. V. Oseledets. Fast low-rank approximations of multidimensional integrals in ion-atomic collisions modelling. Numerical Linear Algebra with Applications, 22(6):1147–1160, December 2015. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Loring:2014:CLU

[Lor14] Terry A. Loring. Computing a logarithm of a unitary matrix with general spectrum. Numerical Linear Algebra with Applications, 21(6):744–760, December 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Leem:2004:AMA

[LOS04] K. H. Leem, S. Oliveira, and D. E. Stewart. Algebraic multigrid (AMG) for saddle point systems from meshfree discretizations. Numerical Linear Algebra with Applications, 11(2–3):293–308, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lotti:2007:NSB

[Lot07] G. Lotti. A note on the solution of not balanced banded Toeplitz systems. Numerical Linear Algebra with Applications, 14(8):645–657, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lottes:2023:OPS

[Lot23] James Lottes. Optimal polynomial smoothers for multigrid V -cycles. Numerical Linear Algebra with Applications, 30 (6):e2518:1–e2518:??, December 2023. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

LeBorne:2008:MPS

[LOY08] Sabine Le Borne, Suely Oliveira, and Fang Yang. H-matrix preconditioners for symmetric saddle-point systems from meshfree discretization. Numerical Linear Algebra with Applications, 15(10): 911–924, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2016:SPM

[LP16] Yinbing Lu and Jianyu Pan. Shifted power method for computing tensor H-eigenpairs. Numerical Linear Algebra with Applications, 23(3):410–426, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Leveque:2022:FIS

[LP22] Santolo Leveque and John W. Pearson. Fast iterative solver for the optimal control of timedependent PDEs with Crank– Nicolson discretization in time. Numerical Linear Algebra with Applications, 29(2):e2419:1– e2419:??, March 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lazarov:2006:OSB

[LPQ06] Raytcho Lazarov, Joseph Pasciak, and Guan Qin. On the occasion of the sixtieth birthday of Richard E. Ewing. Numerical Linear Algebra with Applications, 13(9):675–676, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2015:FAI

[LPS15] Xin Lu, Hong-Kui Pang, and Hai-Wei Sun. Fast approximate inversion of a block triangular Toeplitz matrix with applications to fractional subdiffusion equations. Numerical Linear Algebra with Applications, 22(5):866–882, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lamping:2016:SAK

[LPS16] Frank Lamping, Juan-Manuel Peña, and Tomas Sauer. Spline approximation, Kronecker products and multilinear forms. Numerical Linear Algebra with Applications, 23 (3):535–557, May 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2018:AIM

[LPSV18] Xin Lu, Hong-Kui Pang, Hai-Wei Sun, and Seak-Weng Vong. Approximate inversion method for time-fractional subdiffusion equations. Numerical Linear Algebra with Applications, 25 (2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lazarov:2001:ISC

[LPV01] Raytcho D. Lazarov, Joseph E. Pasciak, and Panayot S. Vassilevski. Iterative solution of a coupled mixed and standard Galerkin discretization method for elliptic problems. Numerical Linear Algebra with Applications, 8(1): 13–31, January/February 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501453/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501453&PLACEBO=IE.pdf.

Lyons:2006:FCF

[LPW06] S. L. Lyons, R. R. Parashkevov, and X. H. Wu. A family of H_{1}-conforming finite element spaces for calculations on 3D grids with pinch-outs. Numerical Linear Algebra with Applications, 13(9):789–799, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2013:EST

[LQY13] Guoyin Li, Liqun Qi, and Gaohang Yu. The Z-eigenvalues of a symmetric tensor and its application to spectral hypergraph theory. Numerical Linear Algebra with Applications, 20(6):1001–1029, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Layton:1995:PRP

[LR95] William J. Layton and Patrick J.

Rabier. Peaceman–Rachford procedure and domain decomposition for finite element problems. Numerical Linear Algebra with Applications, 2 (4):363–393, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Larin:2008:CSE

[LR08] Maxim Larin and Arnold Reusken. A comparative study of efficient iterative solvers for generalized Stokes equations. Numerical Linear Algebra with Applications, 15(1):13–34, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Luo:2017:USM

[LRGO17] P. Luo, C. Rodrigo, F. J. Gaspar, and C. W. Oosterlee. On an Uzawa smoother in multigrid for poroelasticity equations. Numerical Linear Algebra with Applications, 24(1): ??, January 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Langville:2004:KPA

[LS04] Amy N. Langville and William J.

Stewart. A Kronecker product approximate preconditioner for SANs. Numerical Linear Algebra with Applications, 11(8–9): 723–752, October/November 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2005:PBE

[LS05] Wen Li and Weiwei Sun. The perturbation bounds for eigenvalues of normal matrices. Numerical Linear Algebra with Applications, 12(2–3):89–94, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2006:SRP

[LS06] Wen Li and Weiwei Sun. Some remarks on the perturbation of polar decompositions for rectangular matrices. Numerical Linear Algebra with Applications, 13(4):327–338, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2015:NSI

[LS15] Yiding Lin and Valeria Simoncini. A new subspace iteration method for the algebraic Riccati equation. Numerical Linear Algebra with Applications, 22(1):26–47, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2022:NTM

[LS22] Tianyi Lu and Yangfeng Su. A Newton-type method for twodimensional eigenvalue problems. Numerical Linear Algebra with Applications, 29 (4):e2430:1–e2430:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2021:MST

[LSC21] Shishun Li, Xinping Shao, and Rongliang Chen. Multilevel space-time multiplicative Schwarz preconditioner for parabolic equations. Numerical Linear Algebra with Applications, 28(6):e2390:1–e2390:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2018:MGS

[LSJ18] Qiaohua Liu, Dongmei Shen, and Ziwei Jia. Making global simpler GMRES more stable. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2001:PMI

[LSL01] Wen Li, W. Sun, and K. Liu.

Parallel multisplitting iterative methods for singular Mmatrices. Numerical Linear Algebra with Applications, 8 (3):181–190, April/May 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76506671/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76506671&PLACEBO=IE.pdf.

Li:2003:PPV

[LSS03] Zhongze Li, Yousef Saad, and Masha Sosonkina. pARMS: a parallel version of the algebraic recursive multilevel solver. Numerical Linear Algebra with Applications, 10(5–6):485–509, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lungten:2018:PSP

[LSS18] Sangye Lungten, Wil H. A. Schilders, and Jennifer A. Scott. Preordering saddlepoint systems for sparse LDL^{T} factorization without pivoting. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2008:MER

[LT08] Yonghui Liu and Yongge Tian. More on extremal ranks of the matrix expressions A − BX ± X ∗ B∗ with statistical applications. Numerical Linear Algebra with Applications, 15(4): 307–325, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2009:CAB

[LT09] Jing Li and Xuemin Tu. Convergence analysis of a balancing domain decomposition method for solving a class of indefinite linear systems. Numerical Linear Algebra with Applications, 16(9):745–773, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2011:SDM

[LT11] Yonghui Liu and Yongge Tian.

A simultaneous decomposition of a matrix triplet with applications. Numerical Linear Algebra with Applications, 18 (1):69–85, January 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2013:HTG

[LT13] Yonghui Liu and Yongge Tian. Hermitian-type generalized singular value decomposition with applications. Numerical Linear Algebra with Applications, 20(1):60–73, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2005:NIN

[Lu05] Lin-Zhang Lu. Newton iterations for a non-symmetric algebraic Riccati equation. Numerical Linear Algebra with Applications, 12(2–3):191–200, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lund:2020:TFD

[Lun20] Kathryn Lund. The tensor t-function: A definition for functions of third-order tensors. Numerical Linear Algebra with Applications, 27(3): e2288:1–e2288:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Luksan:1998:IPI

[LV98] Ladislav Lukšan and Jan Vlček. Indefinitely preconditioned inexact Newton method for large sparse equality constrained non-linear programming problems. Numerical Linear Algebra with Applications, 5(3):219–247, May/June 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
61002410;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=61002410&
PLACEBO=IE.pdf.

LeTallec:1999:ESM

[LV99] Patrick Le Tallec and Marina Vidrascu. Efficient solution of mechanical and biomechanical problems by domain decomposition. Numerical Linear Algebra with Applications, 6(7): 599–616, October/November 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/68502743/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=68502743&PLACEBO=IE.pdf.

Li:2004:EAS

[LV04] C. Li and C. Vuik. Eigenvalue analysis of the SIMPLE preconditioning for incompressible flow. Numerical Linear Algebra with Applications, 11(5– 6):511–523, June/August 2004.

CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lashuk:2008:SVE

[LV08] Ilya Lashuk and Panayot S.

Vassilevski. On some versions of the element agglomeration AMGe method. Numerical Linear Algebra with Applications, 15(7):595–620, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lashuk:2012:EAC

[LV12] I. V. Lashuk and P. S. Vassilevski. Element agglomeration coarse Raviart–Thomas spaces with improved approximation properties. Numerical Linear Algebra with Applications, 19(2):414–426, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lemmerling:2002:STL

[LVD02] P. Lemmerling, S. Van Huffel, and B. De Moor. The structured total least-squares approach for non-linearly structured matrices. Numerical Linear Algebra with Applications, 9(4):321–332, June 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/92012856/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=92012856&PLACEBO=IE.pdf.

Lacroix:2001:DPI

[LVW01] Sébastien Lacroix, Yuri V. Vassilevski, and Mary F. Wheeler. Decoupling preconditioners in the implicit parallel accurate reservoir simulator (IPARS). Numerical Linear Algebra with Applications, 8(8):537–549, December 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88010596/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88010596&PLACEBO=IE.pdf.

Leyk:1998:ELE

[LW98] Zbigniew Leyk and Henryk Woźniakowski. Estimating a largest eigenvector by Lanczos and polynomial algorithms with a random start. Numerical Linear Algebra with Applications, 5 (3):147–164, May/June 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
61002412;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=61002412&
PLACEBO=IE.pdf.

Loghin:2003:SCP

[LW03] D. Loghin and A. J. Wathen. Schur complement preconditioning for elliptic systems of partial differential equations. Numerical Linear Algebra with Applications, 10(5–6):423–443, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2004:PPA

[LW04] Wen-Wei Lin and Jenn-Nan Wang. Partial pole assignment for the vibrating system with aerodynamic effect. Numerical Linear Algebra with Applications, 11(1):41–58, February 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2005:PPA

[LW05] Wen-Wei Lin and Jenn-Nan Wang. Partial pole assignment for the quadratic pencil by output feedback control with feedback designs. Numerical Linear Algebra with Applications, 12(10):967–979, December 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lin:2007:NCP

[LW07] Yiqin Lin and Yimin Wei. A note on constraint preconditioners for nonsymmetric saddle point problems. Numerical Linear Algebra with Applications, 14(8):659–664, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2015:IRP

[LW15] Hanyu Li and Yimin Wei. Improved rigorous perturbation bounds for the LU and QR factorizations. Numerical Linear Algebra with Applications, 22 (6):1115–1130, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2016:CIM

[LW16] Xin-Guo Liu and Wei-Guo Wang. On convergence of iterative methods for maximal correlation problems. Numerical Linear Algebra with Applications, 23(4):588–606, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2017:WMM

[LW17] Qiaohua Liu and Minghui Wang. On the weighting method for mixed least squares–total least squares problems. Numerical Linear Algebra with Applications, 24 (5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2021:RGL

[LW21] Ke Li and Gang Wu. A randomized generalized low rank approximations of matrices algorithm for high dimensionality reduction and image compression. Numerical Linear Algebra with Applications, 28 (1):e2338:1–e2338:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2023:SPQ

[LW23] Tao Li and Qing-Wen Wang.

Structure preserving quaternion full orthogonalization method with applications. Numerical Linear Algebra with Applications, 30(5):e2495:1– e2495:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2016:FCP

[LWC16] Hui Liu, Kun Wang, and Zhangxin Chen. A family of constrained pressure residual preconditioners for parallel reservoir simulations. Numerical Linear Algebra with Applications, 23(1):120–146, January 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2023:TAH

[LWS^{+}23] Pengjie Liu, Xiaoyu Wu, Hu Shao, Yan Zhang, and Shuhan Cao. Three adaptive hybrid derivative-free projection methods for constrained monotone nonlinear equations and their applications. Numerical Linear Algebra with Applications, 30(2):e2471:1– e2471:??, March 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2009:GBD

[LWW09] Xin-Guo Liu, Wei-Guo Wang, and Yi-Min Wei. A generalization of the Bott–Duffin inverse and its applications. Numerical Linear Algebra with Applications, 16(3):173–196, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2022:GBI

[LWZ22] Tao Li, Qing-Wen Wang, and Xin-Fang Zhang. Gradient based iterative methods for solving symmetric tensor equations. Numerical Linear Algebra with Applications, 29(2): e2414:1–e2414:??, March 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Laub:2008:ASC

[LX08] A. J. Laub and J. Xia. Applications of statistical condition estimation to the solution of linear systems. Numerical Linear Algebra with Applications, 15(6):489–513, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2016:SCB

[LXS16] Ruipeng Li, Yuanzhe Xi, and Yousef Saad. Schur complement-based domain decomposition preconditioners with low-rank corrections. Numerical Linear Algebra with Applications, 23(4):706–729, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2013:NSP

[LXW13] Zhao Li, Qingxiang Xu, and Yimin Wei. A note on stable perturbations of Moore– Penrose inverses. Numerical Linear Algebra with Applications, 20(1):18–26, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2017:SMT

[LXX17] Dong-Hui Li, Shuilian Xie, and Hong-Ru Xu. Splitting methods for tensor equations. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Levi:2015:SLQ

[LY15] Yoni Levi and Irad Yavneh.

Speed limit quasi splines and their application to interpolation with bounded first order derivative. Numerical Linear Algebra with Applications, 22 (3):433–452, May 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Lu:2015:NLD

[LYL15] Linzhang Lu, Fei Yuan, and Ren-Cang Li. A new look at the doubling algorithm for a structured palindromic quadratic eigenvalue problem. Numerical Linear Algebra with Applications, 22(3):393–409, May 2015. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Laszkiewicz:2009:PFI

[LZ09] Beata Laszkiewicz and Krystyna Ziátak. A Padé family of iterations for the matrix sector function and the matrix p th root. Numerical Linear Algebra with Applications, 16(11-12):951–970, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2012:LEB

[LZ12] Xin-Guo Liu and Na Zhao. Linearization estimates of the backward errors for least squares problems. Numerical Linear Algebra with Applications, 19(6):954–969, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2022:AAM

[LZ22] Zhizhi Li and Huai Zhang. Anderson acceleration of the modulus-based matrix splitting algorithms for horizontal nonlinear complementarity systems. Numerical Linear Algebra with Applications, 29 (5):e2438:1–e2438:??, October 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2023:EOP

[LZ23] Zhizhi Li and Huai Zhang. On estimation of the optimal parameter of the modulusbased matrix splitting algorithm for linear complementarity problems on second-order cones. Numerical Linear Algebra with Applications, 30 (4):e2480:1–e2480:??, August 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ling:2012:SRA

[LZQ12] Chen Ling, Xinzhen Zhang, and Liqun Qi. Semidefinite relaxation approximation for multivariate bi-quadratic optimization with quadratic constraints. Numerical Linear Algebra with Applications, 19(1): 113–131, January 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2011:NIM

[LZY11] Xiaoji Liu, Guangping Zhou, and Yaoming Yu. Note on the iterative methods for computing the generalized inverse over Banach spaces. Numerical Linear Algebra with Applications, 18(4):775–787, August 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liu:2020:IAD

[LZZ20] Zhongyun Liu, Yang Zhou, and Yulin Zhang. On inexact alternating direction implicit iteration for continuous Sylvester equations. Numerical Linear Algebra with Applications, 27 (5):e2320:1–e2320:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ma:2022:TSU

[Ma22]  Wei Ma.  Two-step Ulm–

Chebyshev-like method for inverse singular value problems. Numerical Linear Algebra with Applications, 29(5):e2440:1– e2440:??, October 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mailybaev:2006:CME

[Mai06] Alexei A. Mailybaev. Computation of multiple eigenvalues and generalized eigenvectors for matrices dependent on parameters. Numerical Linear Algebra with Applications, 13 (5):419–436, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Margenov:1994:UBC

[Mar94] S. D. Margenov. Upper bound of the constant in the strengthened C.B.S. inequality for FEM 2D elasticity equations. Numerical Linear Algebra with Applications, 1(1):65–74, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Marek:1995:SRB

[Mar95]  I. Marek.  Some remarks

on the barrier lemma and K-monotonicity. Numerical Linear Algebra with Applications, 2(5):431–445, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Special issue dedicated to David M. Young, Jr.

Margenov:1998:SCA

[Mar98] Svetozar D. Margenov. Semicoarsening AMLI algorithms for elasticity problems. Numerical Linear Algebra with Applications, 5(5):347–362, September/October 1998. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62000044;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62000044&
PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Marek:2000:GEN

[Mar00]  Ivo Marek.  Guest Edito-

rial: Numerical linear algebra methods for computational fluid flow problems. Numerical Linear Algebra with Applications, 7(6):361, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516696/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516696&PLACEBO=IE.pdf.

Martinez:2016:TPE

[Mar16] Ángeles Martínez. Tuned preconditioners for the eigensolution of large SPD matrices arising in engineering problems. Numerical Linear Algebra with Applications, 23(3): 427–443, May 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mathias:1996:AAO

[Mat96] Roy Mathias. Analysis of algorithms for orthogonalizing products of unitary matrices. Numerical Linear Algebra with Applications, 3(2): 125–145, March/April 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000985.

Mavriplis:2001:MAN

[Mav01] Dimitri J. Mavriplis. Multigrid approaches to non-linear diffusion problems on unstructured meshes. Numerical Linear Algebra with Applications, 8(8):499–512, December 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88010578/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88010578&PLACEBO=IE.pdf.

Mayer:2005:ICI

[May05] Jan Mayer. ILUCP: a Crout ILU preconditioner with pivoting. Numerical Linear Algebra with Applications, 12(9):

941–955, November 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mayer:2007:MCI

[May07]  Jan Mayer.  A multilevel

Crout ILU preconditioner with pivoting and row permutation. Numerical Linear Algebra with Applications, 14(10): 771–789, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ma:2021:RIN

[MB21] Ru-Ru Ma and Zheng-Jian Bai. A Riemannian inexact Newton–CG method for stochastic inverse singular value problems. Numerical Linear Algebra with Applications, 28(1):e2336:1– e2336:??, January 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meynen:1997:APA

[MBW97] S. Meynen, A. Boersma, and P. Wriggers. Application of a parallel algebraic multigrid method for the solution of elastoplastic shell problems. Numerical Linear Algebra with Applications, 4 (3):223–238, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=

                                        15032;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15032&PLACEBO=IE.pdf.

Murillo:2004:FIP

[MC04] Maria Murillo and Xiao-Chuan Cai. A fully implicit parallel algorithm for simulating the non-linear electrical activity of the heart. Numerical Linear Algebra with Applications, 11(2–3):261–277, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Michelini:2008:SAA

[MC08] Pablo Navarrete Michelini and Edward J. Coyle. A semialgebraic approach that enables the design of inter-grid operators to optimize multigrid convergence. Numerical Linear Algebra with Applications, 15 (2-3):219–247, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meerbergen:2009:CCB

[MC09] Karl Meerbergen and Jean-Pierre Coyette. Connection and comparison between frequency shift time integration and a spectral transformation preconditioner. Numerical Linear Algebra with Applications, 16(1):1–17, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Michael:2012:SMT

[MCC^{+}12] Ng Michael, Raymond Chan, Xiaojun Chen, Liqun Qi, Franklin Luk, Michael Ng, and Eugene Tyrtyshnikov. Structured matrices and tensors. Numerical Linear Algebra with Applications, 19(1):1–2, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Miao:2020:UJD

[MCLM20] Jifei Miao, Guanghui Cheng, Wenrui Li, and Eric Moreau. A unitary joint diagonalization algorithm for nonsymmetric higher-order tensors based on Givens-like rotations. Numerical Linear Algebra with Applications, 27(3):e2291:1– e2291:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mastronardi:2001:FST

[MCV01] Nicola Mastronardi, Shivkumar Chandrasekaran, and Sabine Van Huffel. Fast and stable two-way algorithm for diagonal plus semiseparable systems of linear equations. Numerical Linear Algebra with Applications, 8(1):7–12, January/February 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501454/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501454&PLACEBO=IE.pdf.

Mandel:2003:CBD

[MD03] Jan Mandel and Clark R. Dohrmann. Convergence of a balancing domain decomposition by constraints and energy minimization. Numerical Linear Algebra with Applications, 10(7):639–659, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meng:2021:CNT

[MDB21] Qing-Le Meng, Huai-An Diao, and Zheng-Jian Bai. Condition numbers for the truncated total least squares problem and their estimations. Numerical Linear Algebra with Applications, 28(5):e2369:1–e2369:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mazza:2023:MBS

[MDMS23] Mariarosa Mazza, Marco Donatelli, Carla Manni, and Hendrik Speleers. On the matrices in B-spline collocation methods for Riesz fractional equations and their spectral properties. Numerical Linear Algebra with Applications, 30 (1):e2462:1–e2462:??, January 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meerbergen:2001:CPR

[Mee01]  K. Meerbergen.  Changing

poles in the rational Lanczos method for the Hermitian eigenvalue problem. Numerical Linear Algebra with Applications, 8(1):33–52, January/February 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501452/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501452&PLACEBO=IE.pdf.

Meyer:1994:CSI

[Mey94] Arnd Meyer. The concept of special inner products for deriving new conjugate gradientlike solvers for nonsymmetric sparse linear systems. Numerical Linear Algebra with Applications, 1(2):129–139, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mezzadri:2020:MBS

[Mez20] Francesco Mezzadri. Modulusbased synchronous multisplitting methods for solving horizontal linear complementarity problems on parallel computers. Numerical Linear Algebra with Applications, 27 (5):e2319:1–e2319:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Magri:2018:MAF

[MFFJ18] Victor A. P. Magri, Andrea Franceschini, Massimiliano Ferronato, and Carlo Janna. Multilevel approaches for FSAI preconditioning. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Matioli:2008:NFP

[MG08] L. C. Matioli and C. C. Gonzaga. A new family of penalties for augmented Lagrangian methods. Numerical Linear Algebra with Applications, 15 (10):925–944, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Montero:2002:AIC

[MGF^{+}02] G. Montero, L. González, E. Flórez, M. D. García, and A. Suárez. Approximate inverse computation using Frobenius inner product. Numerical Linear Algebra with Applications, 9(3):239–247, April/ May 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/90512121/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=90512121&PLACEBO=IE.pdf.

Mitchell:2010:HMM

[Mit10] William F. Mitchell. The hpmultigrid method applied to hp-adaptive refinement of triangular grids. Numerical Linear Algebra with Applications, 17(2-3):211–228, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Miyajima:2015:FEM

[Miy15] Shinya Miyajima. Fast enclosure for the minimum norm least squares solution of the matrix equation AXB = C. Numerical Linear Algebra with Applications, 22(3):548–563, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Miyajima:2017:FVC

[Miy17] Shinya Miyajima. Fast verified computation for solutions of algebraic Riccati equations arising in transport theory. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ma:1994:NCI

[MK94] Fu Ming Ma and Tassilo Küpper. Numerical calculation of invariant manifolds for maps. Numerical Linear Algebra with Applications, 1 (2):141–150, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mickelin:2020:ACT

[MK20] Oscar Mickelin and Sertac Karaman. On algorithms for and computing with the tensor ring decomposition. Numerical Linear Algebra with Applications, 27(3):e2289:1–e2289:??, May 2020. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Mickelin:2023:ROT

[MK23] Oscar Mickelin and Sertac Karaman. Recovering orthogonal tensors under arbitrarily strong, but locally correlated, noise. Numerical Linear Algebra with Applications, 30 (4):e2479:1–e2479:??, August 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mastronardi:2005:FRS

[MLV05] N. Mastronardi, P. Lemmerling, and S. Van Huffel. Fast regularized structured total least squares algorithm for solving the basic deconvolution problem. Numerical Linear Algebra with Applications, 12(2–3):201–209, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Margenov:1995:OAM

[MM95] S. Margenov and J. Maubach. Optimal algebraic multilevel preconditioning for local refinement along a line. Numerical Linear Algebra with Applications, 2(4):347–361, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meyer:1997:MAS

[MM97] Arnd Meyer and Detlef Michael. A modern approach to the solution of problems of classic elastoplasticity on parallel computers. Numerical Linear Algebra with Applications, 4 (3):205–221, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15030;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15030&PLACEBO=IE.pdf.

Marek:1998:CAI

[MM98] Ivo Marek and Petr Mayer.

Convergence analysis of an iterative aggregation/disaggregation method for computing stationary probability vectors of stochastic matrices. Numerical Linear Algebra with Applications, 5(4):253–274, July/August 1998. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
10005753;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=10005753&
PLACEBO=IE.pdf.

Mihajlovi:2002:CDP

[MM02] M. D. Mihajlovi and S. Mijalkovi. A component decomposition preconditioning for 3D stress analysis problems. Numerical Linear Algebra with Applications, 9(6–7):567–583, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

mongaMade:2004:PPI

[mM04] M. Magolu monga Made. Performance of parallel incomplete LDL t factorizations for solving acoustic wave propagation problems from industry. Numerical Linear Algebra with Applications, 11(8–9):813–830, October/November 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mastroianni:2009:WCM

[MM09] Giuseppe Mastroianni and Gradimir V. Milovanović. Well-conditioned matrices for numerical treatment of Fredholm integral equations of the second kind. Numerical Linear Algebra with Applications, 16 (11-12):995–1011, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mehrmann:2011:ACS

[MM11] Volker Mehrmann and Agnieszka Miedlar. Adaptive computation of smallest eigenvalues of self-adjoint elliptic partial differential equations. Numerical Linear Algebra with Applications, 18(3):387–409, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mitchell:2018:AIT

[MM18] Wayne Mitchell and Tom Manteuffel. Advances in implementation, theoretical motivation, and numerical results for the nested iteration with range decomposition algorithm. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

MacLachlan:2012:RAM

[MMC12] S. P. MacLachlan, J. D. Moulton, and T. P. Chartier. Robust and adaptive multigrid methods: comparing structured and algebraic approaches. Numerical Linear Algebra with Applications, 19(2):389–413, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

MacLachlan:2006:ARB

[MMM06] Scott MacLachlan, Tom Manteuffel, and Steve McCormick. Adaptive reduction-based AMG. Numerical Linear Algebra with Applications, 13(8):599–620, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mackey:2009:NMP

[MMMM09] D. Steven Mackey, Niloufer Mackey, Christian Mehl, and Volker Mehrmann. Numerical methods for palindromic eigenvalue problems: Computing the anti-triangular Schur form. Numerical Linear Algebra with Applications, 16(1): 63–86, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Manteuffel:2010:FRE

[MMN^{+}10] Thomas Manteuffel, Steven McCormick, Joshua Nolting, John Ruge, and Geoff Sanders. Further results on error estimators for local refinement with first-order system least squares (FOSLS). Numerical Linear Algebra with Applications, 17 (2-3):387–413, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

mongaMade:1999:EPL

[mMP99] Magolu monga Made and Ben Polman. Efficient planewiselike preconditioners for solving 3D problems. Numerical Linear Algebra with Applications, 6(5):379–406, July/August 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/65500101/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=65500101&PLACEBO=IE.pdf.

Manteuffel:2010:OBI

[MMPR10] T. Manteuffel, S. McCormick, M. Park, and J. Ruge. Operator-based interpolation for bootstrap algebraic multigrid. Numerical Linear Algebra with Applications, 17(2-3):

519–537, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

mongaMade:2002:SAP

[mMvdV02] M. Magolu monga Made and Henk A. van der Vorst. Spectral analysis of parallel incomplete factorizations with implicit pseudo-overlap. Numerical Linear Algebra with Applications, 9(1):45–64, January/February 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013649/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013649&PLACEBO=IE.pdf.

Moriya:2000:DGM

[MN00] Kentaro Moriya and Takashi Nodera. The DEFLATED-GMRES(m, k) method with switching the restart frequency dynamically. Numerical Linear Algebra with Applications, 7(7–8):569–584, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505474/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505474&PLACEBO=IE.pdf.

Moret:2005:IFM

[MN05] I. Moret and P. Novati. Interpolating functions of matrices on zeros of quasi-kernel polynomials. Numerical Linear Algebra with Applications, 12(4): 337–353, May 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mardal:2007:OOS

[MNCT07] Kent-Andre Mardal, Bjørn Fredrik Nielsen, Xing Cai, and Aslak Tveito. An order optimal solver for the discretized bidomain equations. Numerical Linear Algebra with Applications, 14 (2):83–98, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Manteuffel:1994:ROP

[MO94] Thomas A. Manteuffel and James S. Otto. On the roots of the orthogonal polynomials and residual polynomials associated with a conjugate gradient method. Numerical Linear Algebra with Applications, 1(5):449–475, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

MacLachlan:2011:LFA

[MO11] Scott P. MacLachlan and C. W. Oosterlee. Local Fourier analysis for multigrid with overlapping smoothers applied to systems of PDEs. Numerical Linear Algebra with Applications, 18(4):751–774, August 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

MacLachlan:2014:TBA

[MO14] Scott P. MacLachlan and Luke N. Olson. Theoretical bounds for algebraic multigrid performance: review and analysis. Numerical Linear Algebra with Applications, 21(2): 194–220, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mikhalev:2016:IRS

[MO16] A. Yu. Mikhalev and I. V. Oseledets. Iterative representing set selection for nested cross approximation. Numerical Linear Algebra with Applications, 23(2):230–248, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Moret:2007:RRK

[Mor07] I. Moret. On RD-rational Krylov approximations to the core-functions of exponential integrators. Numerical Linear Algebra with Applications, 14 (5):445–457, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Moret:2009:RLA

[Mor09] Igor Moret. Rational Lanczos approximations to the matrix square root and related functions. Numerical Linear Algebra with Applications, 16(6): 431–445, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mehrmann:2013:GSD

[MP13] Volker Mehrmann and Federico Poloni. A generalized structured doubling algorithm for the numerical solution of linear quadratic optimal control problems. Numerical Linear Algebra with Applications, 20(1):112–137, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Moret:2014:RSI

[MP14] Igor Moret and Marina Popolizio. The restarted shift-and-invert Krylov method for matrix functions. Numerical Linear Algebra with Applications, 21(1): 68–80, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meerbergen:2015:SAT

[MP15] Karl Meerbergen and Bor Plestenjak. A Sylvester– Arnoldi type method for the generalized eigenvalue problem with two-by-two operator determinants. Numerical Linear Algebra with Applications, 22(6):1131–1146, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mehrmann:2016:IFA

[MP16] Volker Mehrmann and Federico Poloni. An inverse-free ADI algorithm for computing Lagrangian invariant subspaces. Numerical Linear Algebra with Applications, 23(1):

147–168, January 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mainar:2018:ACC

[MP18a] E. Mainar and J. M. Peña. Accurate computations with collocation matrices of a general class of bases. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meini:2018:PBA

[MP18b] Beatrice Meini and Federico Poloni. Perron-based algorithms for the multilinear PageRank. Numerical Linear Algebra with Applications, 25 (6):??, December 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mainar:2020:ABD

[MPR20] Esmeralda Mainar, Juan Manuel Peña, and Beatriz Rubio. Accurate bidiagonal decomposition of collocation matrices of weighted φ-transformed systems. Numerical Linear Algebra with Applications, 27(3): e2295:1–e2295:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mainar:2022:AEC

[MPR22] Esmeralda Mainar, Juan M.

Peña, and Beatriz Rubio. Accurate and efficient computations with Wronskian matrices of Bernstein and related bases. Numerical Linear Algebra with Applications, 29(3): e2423:1–e2423:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mainar:2023:TPA

[MPR23] E. Mainar, J. M. Peña, and B. Rubio. Total positivity and accurate computations with Gram matrices of Said– Ball bases. Numerical Linear Algebra with Applications, 30 (6):e2521:1–e2521:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Migallon:1996:BTS

[MPS96] Violeta Migallón, José Penadés, and Daniel B. Szyld. Block two-stage methods for singular systems and Markov chains. Numerical Linear Algebra with Applications, 3(5): 413–426, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001000.

Marques:2006:CES

[MPV06] Osni A. Marques, Beresford N. Parlett, and Christof Vömel. Computations of eigenpair subsets with the MRRR algorithm. Numerical Linear Algebra with Applications, 13(8):

643–653, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Myllykoski:2014:PRB

[MR14] M. Myllykoski and T. Rossi.

A parallel radix-4 block cyclic reduction algorithm. Numerical Linear Algebra with Applications, 21(4):540–556, August 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Massei:2022:HAL

[MRK22] Stefano Massei, Leonardo Robol, and Daniel Kressner. Hierarchical adaptive low-rank format with applications to discretized partial differential equations. Numerical Linear Algebra with Applications, 29 (6):e2448:1–e2448:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Maryska:1996:PFF

[MRT96] Jiří Maryška, Miroslav Rozložník, and Miroslav Tůma. The potential fluid flow problem and the convergence rate of the minimal residual method. Numerical Linear Algebra with Applications, 3(6): 525–542, November/December 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001007.

Mendoza:1998:CND

[MRT98] María Mendoza, Marcos Raydan, and Pablo Tarazaga. Computing the nearest diagonally dominant matrix. Numerical Linear Algebra with Applications, 5(6):461–474, November/December 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002966;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002966&
                                        PLACEBO=IE.pdf.

Martikainen:2002:FDS

[MRT02] Janne Martikainen, Tuomo Rossi, and Jari Toivanen. A fast direct solver for elliptic problems with a divergence constraint. Numerical Linear Algebra with Applications, 9 (8):629–652, December 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Marcato:2007:IEC

[MS07] Rachel Marques Marcato and Claudia Sagastizábal. Introducing environmental constraints in generation expansion problems. Numerical Linear Algebra with Applications, 14(4):351–368, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Malyshev:2014:FSU

[MS14] A. N. Malyshev and M. Sadkane. Fast solution of unsymmetric banded Toeplitz systems by means of spectral factorizations and Woodbury’s formula. Numerical Linear Algebra with Applications, 21(1): 13–23, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ma:2022:AAL

[MS22] Linjian Ma and Edgar Solomonik.

Accelerating alternating least squares for tensor decomposition by pairwise perturbation. Numerical Linear Algebra with Applications, 29(4):e2431:1– e2431:??, August 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mohamed:2018:ISG

[MSB18] M Shadi Mohamed, Mohammed Seaid, and Abderrahman Bouhamidi. Iterative solvers for generalized finite element solution of boundaryvalue problems. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mitchell:2021:PPA

[MSF21] Wayne B. Mitchell, Robert Strzodka, and Robert D. Falgout. Parallel performance of algebraic multigrid domain decomposition. Numerical Linear Algebra with Applications, 28(3): e2342:1–e2342:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mathew:2007:ABM

[MSS07] T. P. Mathew, M. Sarkis, and C. E. Schaerer. Analysis of block matrix preconditioners for elliptic optimal control problems. Numerical Linear Algebra with Applications, 14(4):257–279, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Morini:2016:SEU

[MST16] Benedetta Morini, Valeria Simoncini, and Mattia Tani. Spectral estimates for unreduced symmetric KKT systems arising from Interior Point methods. Numerical Linear Algebra with Applications, 23(5): 776–800, October 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Meerbergen:2013:JDM

[MSV13] Karl Meerbergen, Christian Schröder, and Heinrich Voss. A Jacobi–Davidson method for two-real-parameter nonlinear eigenvalue problems arising from delay-differential equations. Numerical Linear Algebra with Applications, 20(5): 852–868, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mastronardi:2005:NLA

[MV05] Nicola Mastronardi and Sabine Van Huffel. Numerical linear algebra and its applications. Numerical Linear Algebra with Applications, 12(8):683, October 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mancini:2013:ISP

[MV13] R. Mancini and S. Volkwein.

An inverse scattering problem for the time-dependent Maxwell equations: nonlinear optimization and model-order reduction. Numerical Linear Algebra with Applications, 20 (4):689–711, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mohammadi:2019:DRE

[MV19] Seyyed Abbas Mohammadi and Heinrich Voss. On the distribution of real eigenvalues in linear viscoelastic oscillators. Numerical Linear Algebra with Applications, 26(2): e2228:1–e2228:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Markovsky:2004:CMS

[MVK04] Ivan Markovsky, Sabine Van Huffel, and Alexander Kukush. On the computation of the multivariate structured total least squares estimator. Numerical Linear Algebra with Applications, 11(5–6):591–608, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Minster:2023:CDT

[MVLB23] Rachel Minster, Irina Viviano, Xiaotian Liu, and Grey Ballard. CP decomposition for tensors via alternating least squares with QR decomposition. Numerical Linear Algebra with Applications, 30 (6):e2511:1–e2511:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mastronardi:2008:FAC

[MVV08] N. Mastronardi, M. Van Barel, and R. Vandebril. A fast algorithm for computing the smallest eigenvalue of a symmetric positive-definite Toeplitz matrix. Numerical Linear Algebra with Applications, 15(4): 327–337, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mardal:2011:PDS

[MW11] Kent-Andre Mardal and Ragnar Winther. Preconditioning discretizations of systems of partial differential equations. Numerical Linear Algebra with Applications, 18(1):1–40, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Maurer:2016:SPF

[MW16] Daniel Maurer and Christian Wieners. A scalable parallel factorization of finite element matrices with distributed Schur complements. Numerical Linear Algebra with Applications, 23(5):848–864, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Murray:2021:SAF

[MW21] Charles D. Murray and Tobias Weinzierl. Stabilized asynchronous fast adaptive composite multigrid using additive damping. Numerical Linear Algebra with Applications, 28(3): e2328:1–e2328:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mehl:2006:COS

[MWZ06] M. Mehl, T. Weinzierl, and Chr. Zenger. A cache-oblivious self-adaptive full multigrid method. Numerical Linear Algebra with Applications, 13(2– 3):275–291, March/April 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Mitchell:2020:NAA

[MYD20] Drew Mitchell, Nan Ye, and Hans De Sterck. Nesterov acceleration of alternating least squares for canonical tensor decomposition: Momentum step size selection and restart mechanisms. Numerical Linear Algebra with Applications, 27 (4):e2297:1–e2297:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Morgan:2016:PBD

[MYZ16] Ronald B. Morgan, Zhao Yang, and Baojiang Zhong. Pseudoeigenvector bases and deflated GMRES for highly nonnormal matrices. Numerical Linear Algebra with Applications, 23(6):1032–1045, December 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Morgan:1998:HPM

[MZ98] Ronald B. Morgan and Min Zeng. Harmonic projection methods for large nonsymmetric eigenvalue problems. Numerical Linear Algebra with Applications, 5(1): 33–55, January/February 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5961;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5961&PLACEBO=IE.pdf.

Ma:2015:CUM

[MZ15] C. F. Ma and Q. Q. Zheng. The corrected Uzawa method for solving saddle point problems. Numerical Linear Algebra with Applications, 22(4): 717–730, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Marioni:2017:NAS

[MZHB17] L. Marioni, J. R. Alves Z., E. Hachem, and F. Bay. A new approach to solve complex valued systems arising from the solution of Maxwell equations in the frequency domain through real-equivalent formulations. Numerical Linear Algebra with Applications, 24(2): ??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:1997:NOP

[NA97] Yvan Notay and Zakaria Ould Amar. A nearly optimal preconditioning based on recursive red-black orderings. Numerical Linear Algebra with Applications, 4(5):369– 391, September/October 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15042;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15042&PLACEBO=IE.pdf.

Nabben:1997:BRM

[Nab97] Reinhard Nabben. Book review: Matrices of Sign-Solvable Linear Systems by R. A. Brualdi and B. L. Shader. Cambridge University Press, 1995, Cambridge, ISBN: 0-521-48296-8 (hardback). Price UK£30.00 (US$49.95). Numerical Linear Algebra with Applications, 4(5):439–440, September/October 1997. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15045;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15045&PLACEBO=IE.pdf.

Nazareth:1995:TRB

[Naz95] J. L. Nazareth. Trust regions based on conic functions in linear and nonlinear programming. Numerical Linear Algebra with Applications, 2 (3):235–241, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nedoma:1999:NMC

[NBKS99] J. Nedoma, M. Bartoš, Z. Kestřánek, Jr., and J. Stehlík. Numerical methods for constrained optimization in 2D and 3D biomechanics. Numerical Linear Algebra with Applications, 6(7): 557–586, October/November 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/68502741/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=68502741&PLACEBO=IE.pdf.

Noutsos:2005:PPI

[NCV05] D. Noutsos, S. Serra Capizzano, and P. Vassalos. A preconditioning proposal for ill-conditioned Hermitian twolevel Toeplitz systems. Numerical Linear Algebra with Applications, 12(2–3):231–239, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Neymeyr:2002:PEE

[Ney02] Klaus Neymeyr. A posteriori error estimation for elliptic eigenproblems. Numerical Linear Algebra with Applications, 9(4):263–279, June 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/92012854/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=92012854&PLACEBO=IE.pdf.

Neymeyr:2005:NII

[Ney05] Klaus Neymeyr. A note on Inverse Iteration. Numerical Linear Algebra with Applications, 12(1):1–8, February 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Naumovich:2010:AMW

[NFD10] Anna Naumovich, Malte Förster, and Richard Dwight. Algebraic multigrid within defect correction for the linearized Euler equations. Numerical Linear Algebra with Applications, 17 (2-3):307–324, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

NGuessan:2015:CAM

[NG15] Assi N’Guessan and Issa Cherif Geraldo. A cyclic algorithm for maximum likelihood estimation using Schur complement. Numerical Linear Algebra with Applications, 22 (6):1161–1179, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noordmans:1998:CRS

[NH98] J. Noordmans and P. W. Hemker. Convergence results for 3D sparse grid approaches. Numerical Linear Algebra with Applications, 5(5): 363–376, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62000045;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62000045&
                                        PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Niessner:2006:MSM

[NH06] J. Niessner and R. Helmig. Multi-scale modelling of twophase–two-component processes in heterogeneous porous media. Numerical Linear Algebra with Applications, 13(9): 699–715, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nithiarasu:2009:JCN

[NL09] Perumal Nithiarasu and Rainald Löhner. The journal Communications in Numerical Methods in Engineering with Biomedical Applications becomes the International Journal for Numerical Methods in Biomedical Engineering (IJNMBE) from 1st January 2010. Numerical Linear Algebra with Applications, 16(11-12):1013– 1014, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Napov:2016:AMP

[NL16] Artem Napov and Xiaoye S. Li.

An algebraic multifrontal preconditioner that exploits the low-rank property. Numerical Linear Algebra with Applications, 23(1):61–82, January 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Anonymous:1994:NLA

[NLA94] Numerical Linear Algebra with Applications, 1994. ISSN 1070-5325 (print), 1099-1506 (electronic). John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia.

Ng:2011:SLD

[NLZ11] Michael K. Ng, Li-Zhi Liao, and Leihong Zhang. On sparse linear discriminant analysis algorithm for high-dimensional data classification. Numerical Linear Algebra with Applications, 18(2):223–235, March 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Napov:2010:WDT

[NN10] Artem Napov and Yvan Notay. When does two-grid optimality carry over to the V - cycle? Numerical Linear Algebra with Applications, 17(2-3):273–290, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Napov:2011:AAA

[NN11] Artem Napov and Yvan Notay. Algebraic analysis of aggregation-based multigrid. Numerical Linear Algebra with Applications, 18(3):539–564, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nagy:2015:DRJ

[NN15] James Nagy and Michael Ng. Dedication to Robert J. Plemmons. Numerical Linear Algebra with Applications, 22(5): 793–794, October 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nakajima:2004:PIS

[NO04] Kengo Nakajima and Hiroshi Okuda. Parallel iterative solvers with selective blocking preconditioning for simulations of fault-zone contact. Numerical Linear Algebra with Applications, 11(8–9):831–852, October/November 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:1994:DDV

[Not94] Yvan Notay. DRIC: a dynamic version of the RIC method. Numerical Linear Algebra with Applications, 1 (6):511–532, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:1998:OCA

[Not98] Yvan Notay. Optimal V -cycle algebraic multilevel preconditioning. Numerical Linear Algebra with Applications, 5(5): 441–459, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62000048;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62000048&
PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Notay:2002:CJD

[Not02a] Y. Notay. Combination of Jacobi–Davidson and conjugate gradients for the partial symmetric eigenproblem. Numerical Linear Algebra with Applications, 9(1): 21–44, January/February 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.

                                        com/cgi-bin/abstract/88013650/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=88013650&PLACEBO=IE.pdf.

Notay:2002:RPF

[Not02b] Y. Notay. Robust parameterfree algebraic multilevel preconditioning. Numerical Linear Algebra with Applications, 9(6–7):409–428, September/ November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notarnicola:2005:MRE

[Not05a] Filippo Notarnicola. Matrix recursive expressions of the DFT of even and odd complex sequences. Numerical Linear Algebra with Applications, 12(8): 793–808, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:2005:AMA

[Not05b] Y. Notay. Algebraic multigrid and algebraic multilevel methods: a theoretical comparison. Numerical Linear Algebra with Applications, 12(5–6):419–451, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:2010:AAT

[Not10] Yvan Notay. Algebraic analysis of two-grid methods: The nonsymmetric case. Numerical Linear Algebra with Applications, 17(1):73–96, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Novati:2003:SLI

[Nov03] P. Novati. Solving linear initial value problems by Faber polynomials. Numerical Linear Algebra with Applications, 10(3): 247–270, April/May 2003. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2013:TTM

[NPR13] Silvia Noschese, Lionello Pasquini, and Lothar Reichel. Tridiagonal Toeplitz matrices: properties and novel applications. Numerical Linear Algebra with Applications, 20(2): 302–326, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nazareth:1996:GNM

[NQ96] J. L. Nazareth and Liqun Qi. Globalization of Newton’s method for solving nonlinear equations. Numerical Linear Algebra with Applications, 3(3):239–249, May/June 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000989.

Noschese:2011:SDN

[NR11] Silvia Noschese and Lothar Reichel. The structured distance to normality of Toeplitz matrices with application to preconditioning. Numerical Linear Algebra with Applications, 18(3):429–447, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2012:GCS

[NR12] Silvia Noschese and Lothar Reichel. Generalized circulant Strang-type preconditioners. Numerical Linear Algebra with Applications, 19(1):3– 17, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2014:ISP

[NR14a] Silvia Noschese and Lothar Reichel. Inverse subspace problems with applications. Numerical Linear Algebra with Applications, 21(5):589–603, October 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2014:MTS

[NR14b] Silvia Noschese and Lothar Reichel. A modified truncated singular value decomposition method for discrete illposed problems. Numerical Linear Algebra with Applications, 21(6):813–822, December 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2017:ASP

[NR17] Silvia Noschese and Lothar Reichel. Approximated structured pseudospectra. Numerical Linear Algebra with Applications, 24(2):??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2019:ESU

[NR19] Silvia Noschese and Lothar Reichel. Eigenvector sensitivity under general and structured perturbations of tridiagonal Toeplitz-type matrices. Numerical Linear Algebra with Applications, 26(3):e2232:1– e2232:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Noschese:2022:EIS

[NR22] Silvia Noschese and Lothar Reichel. Estimating and increasing the structural robustness of a network. Numerical Linear Algebra with Applications, 29(2):e2418:1–e2418:??, March 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ng:2005:NBM

[NSCTP05] Michael K. Ng, Stefano Serra-Capizzano, and Cristina Tablino-Possio. Numerical behaviour of multigrid methods for symmetric Sinc–Galerkin systems. Numerical Linear Algebra with Applications, 12(2– 3):261–269, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nguyen:2022:SAF

[NSCTPW22] Quoc Khanh Nguyen, Stefano Serra-Capizzano, Cristina Tablino-Possio, and Eddie Wadbro. Spectral analysis of the finite element matrices approximating 2D linearly elastic structures and multigrid proposals. Numerical Linear Algebra with Applications, 29 (4):e2433:1–e2433:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ng:2003:E

[NT03] Esmond Ng and Wei-Pai Tang. Editorial. Numerical Linear Algebra with Applications, 10(5–6):383, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ng:2004:PEC

[NT04] Esmond Ng and Wei-Pai Tang. Preface 2 Editorial comments. Numerical Linear Algebra with Applications, 11(8–9):693, October/November 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nabben:2008:CAV

[NV08a] R. Nabben and C. Vuik. A comparison of abstract versions of deflation, balancing and additive coarse grid correction preconditioners. Numerical Linear Algebra with Applications, 15(4):355–372, ????

2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Notay:2008:RKB

[NV08b] Yvan Notay and Panayot S.

Vassilevski. Recursive Krylovbased multigrid cycles. Numerical Linear Algebra with Applications, 15(5):473–487, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nobile:2023:PRO

[NV23] Fabio Nobile and Tommaso Vanzan. Preconditioners for robust optimal control problems under uncertainty. Numerical Linear Algebra with Applications, 30(2):e2472:1– e2472:??, March 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nielsen:2015:CNR

[NW15] A. J. N. Nielsen and M. Weber. Computing the nearest reversible Markov chain. Numerical Linear Algebra with Applications, 22(3):483–499, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ng:2017:VAR

[NWZ17] Michael K. Ng, Wei Wang, and Xile Zhao. A variational approach for restoring images corrupted by noisy blur kernels and additive noise. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Neumann:2003:SCS

[NX03] Michael Neumann and Jianhong Xu. On the stability of the computation of the stationary probabilities of Markov chains using Perron complements. Numerical Linear Algebra with Applications, 10(7): 603–618, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Nikishin:2003:PTA

[NY03] A. A. Nikishin and A. Yu. Yeremin. Prefiltration technique via aggregation for constructing low-density highquality factorized sparse approximate inverse preconditionings. Numerical Linear Algebra with Applications, 10(3): 235–246, April/May 2003. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Neymeyr:2014:IMR

[NZ14] Klaus Neymeyr and Ming Zhou. Iterative minimization of the Rayleigh quotient by block steepest descent iterations. Numerical Linear Algebra with Applications, 21(5): 604–617, October 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ovtchinnikov:2004:OLN

[OC04] Serguei Ovtchinnikov and Xiao-Chuan Cai. One-level Newton–Krylov–Schwarz algorithm for unsteady non-linear radiation diffusion problem. Numerical Linear Algebra with Applications, 11(10):867–881, December 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oktay:2022:MMP

[OC22] Eda Oktay and Erin Carson. Multistage mixed precision iterative refinement. Numerical Linear Algebra with Applications, 29(4):e2434:1– e2434:??, August 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oishi:2008:SNS

[OCYM08] C. M. Oishi, J. A. Cuminato, J. Y. Yuan, and S. McKee. Stability of numerical schemes on staggered grids. Numerical Linear Algebra with Applications, 15(10):945–967, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oviedo:2021:HGM

[ODH21] Harry Oviedo, Oscar Dalmau, and Rafael Herrera. A hybrid gradient method for strictly convex quadratic programming. Numerical Linear Algebra with Applications, 28 (4):e2360:1–e2360:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

OHara:2014:PRO

[O’H14] Michael J. O’Hara. On the perturbation of rank-one symmetric tensors. Numerical Linear Algebra with Applications, 21 (1):1–12, January 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Olshanskii:1999:ISO

[Ols99] Maxim A. Olshanskii. An iterative solver for the Oseen problem and numerical solution of incompressible Navier–Stokes equations. Numerical Linear Algebra with Applications, 6(5):353–378, July/August 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/65500100/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=65500100&PLACEBO=IE.pdf.

Ozaki:2011:TEE

[OOO11] Katsuhisa Ozaki, Takeshi Ogita, and Shin’ichi Oishi. Tight and efficient enclosure of matrix multiplication by using optimized BLAS. Numerical Linear Algebra with Applications, 18(2):237–248, March 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ozaki:2016:EFT

[OOO16] Katsuhisa Ozaki, Takeshi Ogita, and Shin’ichi Oishi. Error-free transformation of matrix multiplication with a posteriori validation. Numerical Linear Algebra with Applications, 23(5):931–946, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oseledets:2023:LCA

[ORU23] Ivan V. Oseledets, Maxim V.

Rakhuba, and André Uschmajew. Local convergence of alternating low-rank optimization methods with overrelaxation. Numerical Linear Algebra with Applications, 30(3): e2459:1–e2459:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Olshevsky:2001:SFH

[OS01] Vadim Olshevsky and Michael Stewart. Stable factorization for Hankel and Hankellike matrices. Numerical Linear Algebra with Applications, 8(6–7):401–434, September/November 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007283/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85007283&PLACEBO=IE.pdf.

Olson:2010:SAH

[OS10] Luke N. Olson and Jacob B. Schroder. Smoothed aggregation for Helmholtz problems. Numerical Linear Algebra with Applications, 17(2-3):361–386, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Olson:2010:NPS

[OST10a] Luke N. Olson, Jacob Schroder, and Raymond S. Tuminaro. A new perspective on strength measures in algebraic multigrid. Numerical Linear Algebra with Applications, 17(4): 713–733, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oseledets:2010:CAT

[OST10b] I. V. Oseledets, D. V. Savostyanov, and E. E. Tyrtyshnikov. Cross approximation in tensor electron density computations. Numerical Linear Algebra with Applications, 17(6):935–952, December 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Oswald:1995:MPD

[Osw95]  Peter Oswald.  Multilevel

preconditioners for discretizations of the biharmonic equation by rectangular finite elements. Numerical Linear Algebra with Applications, 2 (6):487–505, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Olshanskii:2022:RAL

[OZ22] Maxim A. Olshanskii and Alexander Zhiliakov. Recycling augmented Lagrangian preconditioner in an incompressible fluid solver. Numerical Linear Algebra with Applications, 29(2):e2415:1–e2415:??, March 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ozbay:2013:CCI

[Özb13] H. Özbay. Computation of H_{∞} controllers for infinite dimensional plants using numerical linear algebra. Numerical Linear Algebra with Applications, 20(2):327–335, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Osborn:2018:SHP

[OZB^{+}18] Sarah Osborn, Patrick Zulian, Thomas Benson, Umberto Villa, Rolf Krause, and Panayot S. Vassilevski. Scalable hierarchical PDE sampler for generating spatially correlated random fields using nonmatching meshes. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Padiy:1999:PMS

[Pad99] Alexander Padiy. On a parallel multilevel solver for linear elasticity problems. Numerical Linear Algebra with Applications, 6(3):171–188, April/ May 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
63003628;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=63003628&
PLACEBO=IE.pdf.

Parlett:1992:RMS

[Par92] Beresford N. Parlett. The rewards for maintaining semiorthogonality among Lanczos vectors. Journal of Numerical Linear Algebra with Applications, 1(2):243–267, 1992. CO-DEN ???? ISSN 0129-3281.

Parlett:2003:SDF

[Par03] Beresford N. Parlett. The spectral diameter as a function of the diagonal entries. Numerical Linear Algebra with Applications, 10(7):595–602, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pasini:2019:CAA

[Pas19] Massimiliano Lupo Pasini. Convergence analysis of Andersontype acceleration of Richardson’s iteration. Numerical Linear Algebra with Applications, 26(4):e2241:1–e2241:??, August 2019. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Pan:2005:TSW

[PBN05] Jian-Yu Pan, Zhong-Zhi Bai, and Michael K. Ng. Two-step waveform relaxation methods for implicit linear initial value problems. Numerical Linear Algebra with Applications, 12(2–3):293–304, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Papy:2005:EDF

[PDV05] J. M. Papy, L. De Lathauwer, and S. Van Huffel. Exponential data fitting using multilinear algebra: the single-channel and multi-channel case. Numerical Linear Algebra with Applications, 12(8):809–826, October 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pena:2003:SBS

[Peñ03] J. M. Peña. Simultaneous backward stability of Gauss and Gauss–Jordan elimination. Numerical Linear Algebra with Applications, 10(4): 317–321, June 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pena:2007:RGD

[Peñ07] J. M. Peña. Refining Gerschgorin disks through new criteria for nonsingularity. Numerical Linear Algebra with Applications, 14(8):665–671,

????  2007.  CODEN

NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Peng:2008:SSC

[Pen08] Zhen-Yun Peng. Solutions of symmetry-constrained leastsquares problems. Numerical Linear Algebra with Applications, 15(4):373–389, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pena:2009:EBS

[Peñ09] J. M. Peña. Eigenvalue bounds for some classes of P -matrices. Numerical Linear Algebra with Applications, 16(11-12):871– 882, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Perrone:2006:KPA

[Per06] Lisa Perrone. Kronecker product approximations for image restoration with anti-reflective boundary conditions. Numerical Linear Algebra with Applications, 13(1):1–22, February 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pflaum:1999:AAM

[Pfl99]  Christoph Pflaum.  Alge-

braic analysis of multigrid algorithms. Numerical Linear Algebra with Applications, 6 (8):701–728, December 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/69000704/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=69000704&PLACEBO=IE.pdf.

Pestman:2014:CSW

[PGT14] Wiebe R. Pestman, Rolf H. H.

Groenwold, and Steven Teerenstra. Comparison of strategies when building linear prediction models. Numerical Linear Algebra with Applications, 21(5): 618–628, October 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pitton:2019:AIS

[PH19] G. Pitton and L. Heltai. Accelerating the iterative solution of convection-diffusion problems using singular value decomposition. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pultarova:2021:TSG

[PL21] Ivana Pultarová and Martin Ladecký. Two-sided guaranteed bounds to individual eigenvalues of preconditioned finite element and finite difference problems. Numerical Linear Algebra with Applications, 28(5):e2382:1–e2382:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Liao:2007:LSS

[pLL07] An ping Liao and Yuan Lei. Least-squares solutions of matrix inverse problem for bisymmetric matrices with a submatrix constraint. Numerical Linear Algebra with Applications, 14(5):425–444, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ponce:2023:MWT

[PLMV23] Colin Ponce, Ruipeng Li, Christina Mao, and Panayot Vassilevski. Multilevel-in-width training for deep neural network regression. Numerical Linear Algebra with Applications, 30(5):e2501:1– e2501:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Payer:1997:ISS

[PM97] H.-J. Payer and H. A. Mang. Iterative strategies for solving systems of linear, algebraic equations arising in 3D BE-FE analyses of tunnel drivings. Numerical Linear Algebra with Applications, 4 (3):239–268, May/June 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        15028;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=15028&PLACEBO=IE.pdf.

Pultarova:2011:PPI

[PM11] I. Pultarová and I. Marek.

Physiology and pathology of iterative aggregation–disaggregation methods. Numerical Linear Algebra with Applications, 18 (6):1051–1065, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pan:2018:SOA

[PN18] Junjun Pan and Michael K. Ng.

Symmetric orthogonal approximation to symmetric tensors with applications to image reconstruction. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Poirier:2000:EPS

[Poi00] Bill Poirier. Efficient preconditioning scheme for block partitioned matrices with structured sparsity. Numerical Linear Algebra with Applications, 7(7–8):715–726, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505484/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505484&PLACEBO=IE.pdf.

Pearson:2020:IPM

[PPS20] John W. Pearson, Margherita Porcelli, and Martin Stoll.

Interior-point methods and preconditioning for PDEconstrained optimization problems involving sparsity terms. Numerical Linear Algebra with Applications, 27(2):e2276:1– e2276:??, March 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Paige:1995:ASE

[PPv95] Chris C. Paige, Beresford N. Parlett, and Henk A. van der Vorst. Approximate solutions and eigenvalue bounds from Krylov subspaces. Numerical Linear Algebra with Applications, 2(2):115–133, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pester:1995:PVP

[PR95] Matthias Pester and Sergej Rjasanow. A parallel version of the preconditioned conjugate gradient method for boundary element equations. Numerical Linear Algebra with Applications, 2(1):1–16, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pester:1996:PPI

[PR96] Matthias Pester and Sergej Rjasanow. A parallel preconditioned iterative realization of the panel method in 3D. Numerical Linear Algebra with Applications, 3(1): 65–80, January/February 1996.

CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000498.

Pflaum:2011:ISF

[PR11] Christoph Pflaum and Zhabiz Rahimi. An iterative solver for the finite-difference frequencydomain (FDFD) method for the simulation of materials with negative permittivity. Numerical Linear Algebra with Applications, 18(4): 653–670, August 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Poloni:2016:SPD

[PR16] Federico Poloni and Timo Reis.

A structure-preserving doubling algorithm for Luré equations. Numerical Linear Algebra with Applications, 23(1): 169–186, January 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Petkovic:2009:CLR

[PRPI09] M. S. Petković, L. Rančić, L. D. Petković, and S. Ilić. Chebyshev-like root-finding methods with accelerated convergence. Numerical Linear Algebra with Applications, 16(11-12):971–994, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pranic:2016:RAP

[PRR^{+}16] Miroslav S. Pranić, Lothar Reichel, Giuseppe Rodriguez, Zhengsheng Wang, and Xuebo Yu. A rational Arnoldi process with applications. Numerical Linear Algebra with Applications, 23(6):1007–1022, December 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pantazis:1995:RCR

[PS95] Ricardo D. Pantazis and Daniel B. Szyld. Regions of convergence of the Rayleigh quotient iteration method. Numerical Linear Algebra with Applications, 2(3):251–269, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Perugia:2000:BDI

[PS00] I. Perugia and V. Simoncini. Block-diagonal and indefinite symmetric preconditioners for mixed finite element formulations. Numerical Linear Algebra with Applications, 7(7–8): 585–616, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505478/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505478&PLACEBO=IE.pdf.

Pang:2011:SIL

[PS11] Hong-Kui Pang and Hai-Wei Sun. Shift-invert Lanczos method for the symmetric positive semidefinite Toeplitz matrix exponential. Numerical Linear Algebra with Applications, 18(3):603–614, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pironkov:2008:PDC

[PSK08] P. Pironkov, M. Schäfer, and K. Krastev. Properties of discrete Chebyshev collocation differential operators in curvilinear geometries. Numerical Linear Algebra with Applications, 15(8):701–716, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pearson:2014:PSC

[PSW14] John W. Pearson, Martin Stoll, and Andrew J. Wathen. Preconditioners for stateconstrained optimal control problems with Moreau–Yosida penalty function. Numerical Linear Algebra with Applications, 21(1):81–97, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Prokopenko:2017:AMM

[PT17] Andrey Prokopenko and Raymond S. Tuminaro. An algebraic multigrid method for Q_{2} − Q_{1} mixed discretizations of the Navier–Stokes equations. Numerical Linear Algebra with Applications, 24(6):??, December 2017. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Pultarova:2008:NSL

[Pul08]  Ivana Pultarová.  Neces-

sary and sufficient local convergence condition of one class of iterative aggregation– disaggregation methods. Numerical Linear Algebra with Applications, 15(4):339–354, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pultarova:2009:PPE

[Pul09] I. Pultarová. Preconditioning and a posteriori error estimates using h- and p-hierarchical finite elements with rectangular supports. Numerical Linear Algebra with Applications, 16(5):415–430, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pultarova:2016:CTE

[Pul16] I. Pultarová. Convergence theory of exact interpolation scheme for computing several eigenvectors. Numerical Linear Algebra with Applications, 23 (2):373–390, March 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Paige:1999:SAL

[PV99] Christopher C. Paige and Paul Van Dooren. Sensitivity analysis of the Lanczos reduction. Numerical Linear Algebra with Applications, 6(1): 29–50, January/February 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002985;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002985&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part I (Milovy, 1997).

Pearson:2012:NAS

[PW12] John W. Pearson and Andrew J. Wathen. A new approximation of the Schur complement in preconditioners for PDE-constrained optimization. Numerical Linear Algebra with Applications, 19(5): 816–829, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pestana:2013:CPS

[PW13] J. Pestana and A. J. Wathen.

Combination preconditioning of saddle point systems for positive definiteness. Numerical Linear Algebra with Applications, 20(5):785–808, October 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pencheva:2003:BDD

[PY03] Gergina Pencheva and Ivan Yotov. Balancing domain decomposition for mortar mixed finite element methods. Numerical Linear Algebra with Applications, 10(1–2):159–180, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Pascual:2022:LSR

[PY22] Hector Pascual and Xin C. Yee. Least squares regression principal component analysis: a supervised dimensionality reduction method. Numerical Linear Algebra with Applications, 29 (1):e2411:1–e2411:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Qian:2018:MSU

[QACT18] Jiang Qian, Alan L. Andrew, Delin Chu, and Roger C. E. Tan. Methods for solving underdetermined systems. Numerical Linear Algebra with Applications, 25(1):??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Quaife:2015:PLD

[QB15] Bryan Quaife and George Biros. On preconditioners for the Laplace double-layer in 2D. Numerical Linear Algebra with Applications, 22(1): 101–122, January 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Qin:2023:STB

[QPS23] Hai-Hua Qin, Hong-Kui Pang, and Hai-Wei Sun. Sine transform based preconditioning techniques for space fractional diffusion equations. Numerical Linear Algebra with Applications, 30(4):e2474:1– e2474:??, August 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Quiring:2021:MGE

[QV21] Benjamin Quiring and Panayot S.

Vassilevski. Multilevel graph embedding. Numerical Linear Algebra with Applications, 28(3):e2326:1–e2326:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Qiu:2021:PNS

[QvGvW^{+}21] Yue Qiu, Martin B. van Gijzen, Jan-Willem van Wingerden, Michel Verhaegen, and Cornelis Vuik. Preconditioning Navier–Stokes control using multilevel sequentially semiseparable matrix computations. Numerical Linear Algebra with Applications, 28(2): e2349:1–e2349:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Qian:2009:SLR

[QXB09] Jiang Qian, Shu-Fang Xu, and Feng-Shan Bai. Symmetric low-rank corrections to quadratic models. Numerical Linear Algebra with Applications, 16(5):397–413, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ragni:2014:RKM

[Rag14] S. Ragni. Rational Krylov methods in exponential integrators for European option pricing. Numerical Linear Algebra with Applications, 21(4): 494–512, August 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rakowsky:1999:SCM

[Rak99] Natalja Rakowsky. The Schur complement method as a fast parallel solver for elliptic partial differential equations in oceanography. Numerical Linear Algebra with Applications, 6(6):497–510, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/67501477/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501477&PLACEBO=IE.pdf.

Rosseel:2008:AMS

[RBV08] E. Rosseel, T. Boonen, and S. Vandewalle. Algebraic multigrid for stationary and time-dependent partial differential equations with stochastic coefficients. Numerical Linear Algebra with Applications, 15 (2-3):141–163, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Renaut:1998:PMS

[Ren98]  R. A. Renaut.  A par-

allel multisplitting solution of the least squares problem. Numerical Linear Algebra with Applications, 5(1): 11–31, January/February 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
5960;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5960&PLACEBO=IE.pdf.

Reusken:1996:MMB

[Reu96] Arnold Reusken. A multigrid method based on incomplete Gaussian elimination. Numerical Linear Algebra with Applications, 3(5): 369–390, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000998.

Fang:2009:TCM

[rFS09] Haw ren Fang and Yousef Saad. Two classes of multisecant methods for nonlinear acceleration. Numerical Linear Algebra with Applications, 16 (3):197–221, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ren:2007:ELP

[RGG07] Jihong Ren, Chen Greif, and Mark R. Greenstreet. An efficient linear programming solver for optimal filter synthesis. Numerical Linear Algebra with Applications, 14(9): 695–712, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ruthotto:2017:SMS

[RGM17] Lars Ruthotto, Chen Greif, and Jan Modersitzki. A stabilized multigrid solver for hyperelastic image registration. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rietsch:2009:CIE

[Rie09] Eike Rietsch. Consistent interpolation of equidistantly sampled data. Numerical Linear Algebra with Applications, 16 (4):319–344, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Riseth:2019:OAU

[Ris19] Asbjørn Nilsen Riseth. Objective acceleration for unconstrained optimization. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Rjasanow:1998:SBE

[Rja98] Sergej Rjasanow. The structure of the boundary element matrix for the threedimensional Dirichlet problem in elasticity. Numerical Linear Algebra with Applications, 5(3):203–217, May/June 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
61002409;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=61002409&
PLACEBO=IE.pdf.

Rao:2018:HPA

[RK18] S. Chandra Sekhara Rao and Rabia Kamra. A hybrid parallel algorithm for large sparse linear systems. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Renaut:2012:MRL

[RLG12] Rosemary A. Renaut, Youzuo Lin, and Hongbin Guo. Multisplitting for regularized least squares with Krylov subspace recycling. Numerical Linear Algebra with Applications, 19 (4):655–676, August 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ruymbeek:2019:CMM

[RMM19] Koen Ruymbeek, Karl Meerbergen, and Wim Michiels. Calculating the minimal/ maximal eigenvalue of symmetric parameterized matrices using projection. Numerical Linear Algebra with Applications, 26(5):e2263:1–e2263:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ruymbeek:2022:SMM

[RMM22] Koen Ruymbeek, Karl Meerbergen, and Wim Michiels. Subspace method for multiparametereigenvalue problems based on tensor-train representations. Numerical Linear Algebra with Applications, 29(5):e2439:1– e2439:??, October 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Robbe:2021:EMI

[RNV21] Pieterjan Robbe, Dirk Nuyens, and Stefan Vandewalle. Enhanced multi-index Monte Carlo by means of multiple semicoarsened multigrid for anisotropic diffusion problems. Numerical Linear Algebra with Applications, 28(3):e2281:1– e2281:??, May 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ruijter:2013:FCS

[ROA13] M. J. Ruijter, C. W. Oosterlee, and R. F. T. Aalbers. On the Fourier cosine series expansion method for stochastic control problems. Numerical Linear Algebra with Applications, 20 (4):598–625, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rohn:1992:AFS

[Roh92] Jiří Rohn. An algorithm for finding a singular matrix in an interval matrix. Journal of Numerical Linear Algebra with Applications, 1(1):43–47, 1992. CODEN ???? ISSN 0129-3281.

Ritter:2012:EAF

[RR12] Daniel Ritter and Ulrich Rüde.

Experimental analysis of a Fast Adaptive Compositebased grid expansion scheme for open boundary problems. Numerical Linear Algebra with Applications, 19(2): 268–278, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Reid:2001:RRO

[RS01] J. K. Reid and J. A. Scott.

Reversing the row order for the row-by-row frontal method. Numerical Linear Algebra with Applications, 8(1):1–6, January/February 2001. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76501451/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76501451&PLACEBO=IE.pdf.

Reitzinger:2002:AMM

[RS02] S. Reitzinger and J. Schöberl. An algebraic multigrid method for finite element discretizations with edge elements. Numerical Linear Algebra with Applications, 9(3):223– 238, April/May 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/90512119/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=90512119&PLACEBO=IE.pdf.

Routray:2007:FAD

[RS07] Aurobinda Routray and Smarak Swain. Fast algorithms for designing variable FIR notch filters. Numerical Linear Algebra with Applications, 14(9): 673–694, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rees:2010:BTP

[RS10] Tyrone Rees and Martin Stoll. Block-triangular preconditioners for PDE-constrained optimization. Numerical Linear Algebra with Applications, 17(6): 977–996, December 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rees:2018:CSN

[RS18] Tyrone Rees and Jennifer Scott. A comparative study of null-space factorizations for sparse symmetric saddle point systems. Numerical Linear Algebra with Applications, 25(1): ??, January 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Russo:2015:QOP

[RSCTP15] Alessandro Russo, Stefano Serra-Capizzano, and Cristina Tablino-Possio. Quasi-optimal preconditioners for finite element approximations of diffusion dominated convection– diffusion equations on (nearly) equilateral triangle meshes. Numerical Linear Algebra with Applications, 22(1):123–144, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rahla:2020:SAP

[RSCTP20] Ryma Imene Rahla, Stefano Serra-Capizzano, and Cristina Tablino-Possio. Spectral analysis of P_{k} finite element matrices in the case of Friedrichs– Keller triangulations via Generalized Locally Toeplitz technology. Numerical Linear Algebra with Applications, 27 (4):e2302:1–e2302:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ritter:2010:FAC

[RSR10] Daniel Ritter, Markus Stürmer, and Ulrich Rüde. A fastadaptive composite grid algorithm for solving the free-space Poisson problem on the Cell Broadband Engine. Numerical Linear Algebra with Applications, 17(2-3):291–305, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rossi:1999:PFD

[RT99] Tuomo Rossi and Jari Toivanen. Parallel fictitious domain method for a non-linear elliptic Neumann boundary value problem. Numerical Linear Algebra with Applications, 6 (1):51–60, January/February 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62002986;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62002986&
                                        PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part I (Milovy, 1997).

Raydan:2002:PPA

[RT02] Marcos Raydan and Pablo Tarazaga. Primal and polar approach for computing the symmetric diagonally dominant projection. Numerical Linear Algebra with Applications, 9(5):333–345, July/August 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Raghavan:2003:LTH

[RTN03] Padma Raghavan, Keita Teranishi, and Esmond G. Ng.

A latency tolerant hybrid sparse solver using incomplete Cholesky factorization. Numerical Linear Algebra with Applications, 10(5–6):541–560, July/September 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Reichel:2022:TGK

[RU22] Lothar Reichel and Ugochukwu O.

Ugwu. The tensor Golub– Kahan–Tikhonov method applied to the solution of ill-posed problems with a t-product structure. Numerical Linear Algebra with Applications, 29 (1):e2412:1–e2412:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Reps:2012:AWN

[RV12] B. Reps and W. Vanroose. Analyzing the wave number dependency of the convergence rate of a multigrid preconditioned Krylov method for the Helmholtz equation with an absorbing layer. Numerical Linear Algebra with Applications, 19(2):232–252, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rusten:1998:DEP

[RVW98] Torgeir Rusten, Panayot S.

Vassilevski, and Ragnar Winther. Domain embedding preconditioners for mixed systems. Numerical Linear Algebra with Applications, 5(5):321– 345, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62000043;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62000043&
                                        PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Reichel:2008:GLA

[RY08] Lothar Reichel and Qiang Ye. A generalized LSQR algorithm. Numerical Linear Algebra with Applications, 15(7): 643–660, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saad:1994:IDT

[Saa94] Yousef Saad. ILUT: a dual threshold incomplete LU factorization. Numerical Linear Algebra with Applications, 1 (4):387–402, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saad:2000:E

[Saa00a] Yousef Saad. Editorial. Numerical Linear Algebra with Applications, 7(7–8):489–490, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505473/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505473&PLACEBO=IE.pdf.

Saad:2000:FAM

[Saa00b] Yousef Saad. Further analysis of minimum residual iterations. Numerical Linear Algebra with Applications, 7(2):67–93, March 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71008526/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71008526&PLACEBO=IE.pdf.

Sacksteder:2005:ADS

[Sac05] V. E. Sacksteder. O(N) algorithms for disordered systems. Numerical Linear Algebra with Applications, 12(8): 827–838, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sauter:1995:SIC

[Sau95] Stefan A. Sauter. On the stability of the incomplete Cholesky decomposition for a singular perturbed problem, where the coefficient matrix is not an M-matrix. Numerical Linear Algebra with Applications, 2(1):17–28, 1995. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Solak:2003:NBN

[SB03] Süleyman Solak and Durmu Bozkurt. A note on bound for norms of Cauchy–Hankel matrices. Numerical Linear Algebra with Applications, 10(4): 377–382, June 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Salkuyeh:2012:EMH

[SB12] Davod Khojasteh Salkuyeh and Shiva Behnejad. Errata: ‘Modified Hermitian and skew-Hermitian splitting methods for non-Hermitian positive-definite linear systems’ [Numer. Linear Algebra Appl. 14 (2007) 217– 235]. Numerical Linear Algebra with Applications, 19(5): 885–890, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [LHL07b].

Sistek:2015:BMH

[ŠBS15] Jakub Šístek, Jan Březina, and Bedřich Sousedík. BDDC for mixed-hybrid formulation of flow in porous media with combined mesh dimensions. Numerical Linear Algebra with Applications, 22(6):903–929, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Spedicato:1994:CDA

[SCD94] E. Spedicato, Z. Chen, and N. Deng. A class of difference ABS-type algorithms for a nonlinear system of equations. Numerical Linear Algebra with Applications, 1 (3):313–329, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schrader:1999:CAJ

[Sch99] Uwe Schrader. Convergence of asynchronous Jacobi-Newton-iterations. Numerical Linear Algebra with Applications, 6(2):157–165, March 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002996;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002996&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part 2 (Milovy, 1997).

Schroder:2012:SAS

[Sch12] Jacob B. Schroder. Smoothed aggregation solvers for anisotropic diffusion. Numerical Linear Algebra with Applications, 19(2): 296–312, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Scott:1999:NRO

[Sco99] Jennifer A. Scott. A new row ordering strategy for frontal solvers. Numerical Linear Algebra with Applications, 6 (3):189–211, April/May 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
63003629;   http://www3.

                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=63003629&
                                        PLACEBO=IE.pdf.

Saibaba:2020:EKS

[SCP20] Arvind K. Saibaba, Julianne Chung, and Katrina Petroske. Efficient Krylov subspace methods for uncertainty quantification in large Bayesian linear inverse problems. Numerical Linear Algebra with Applications, 27(5):e2325:1– e2325:??, October 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Seibold:2010:PAM

[Sei10]  B. Seibold.  Performance

of algebraic multigrid methods for non-symmetric matrices arising in particle methods. Numerical Linear Algebra with Applications, 17(2-3): 433–451, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sousedik:2014:HSC

[SGP14] Bedrich Sousedík, Roger G. Ghanem, and Eric T. Phipps. Hierarchical Schur complement preconditioner for the stochastic Galerkin finite element methods: Dedicated to Professor Ivo Marek on the occasion of his 80th birthday. Numerical Linear Algebra with Applications, 21(1):136–151, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Salas:2015:SRS

[SGSM15] Pablo Salas, Luc Giraud, Yousef Saad, and Stéphane Moreau. Spectral recycling strategies for the solution of nonlinear eigenproblems in thermoacoustics. Numerical Linear Algebra with Applications, 22(6):1039–1058, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Scott:2014:LBH

[SH14] Jennifer Scott and Yifan Hu.

Level-based heuristics and hill climbing for the antibandwidth maximization problem. Numerical Linear Algebra with Applications, 21(1):51–67, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Skalna:2019:DIM

[SH19] Iwona Skalna and Milan Hladík. Direct and iterative methods for interval parametric algebraic systems producing parametric solutions. Numerical Linear Algebra with Applications, 26(3):e2229:1– e2229:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Shapira:1998:AMD

[Sha98]  Yair Shapira.  Analysis

of matrix-dependent multigrid algorithms. Numerical Linear Algebra with Applications, 5(3):165–202 (or 165–201??), May/June 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
61002411;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=61002411&
PLACEBO=IE.pdf.

Shapira:1999:MCA

[Sha99] Yair Shapira. Model case analysis of an algebraic multilevel method. Numerical Linear Algebra with Applications, 6(8):655–685, December 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/69000702/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=69000702&PLACEBO=IE.pdf.

Shi:2002:CPR

[Shi02] Yixun Shi. A combination of potential reduction steps and steepest descent steps for solving convex programming problems. Numerical Linear Algebra with Applications, 9 (3):195–203, April/May 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/90512113/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=90512113&PLACEBO=IE.pdf.

Shi:2004:PSD

[Shi04]  Yixun Shi.  A projected-

steepest-descent potential-reduction algorithm for convex programming problems. Numerical Linear Algebra with Applications, 11(10):883–893, December 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sun:2018:BGM

[SHJC18] Dong-Lin Sun, Ting-Zhu Huang, Yan-Fei Jing, and Bruno Carpentieri. A block GMRES method with deflated restarting for solving linear systems with multiple shifts and multiple right-hand sides. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sun:2011:SNI

[SHT11] Jian-Qing Sun, Xing-Biao Hu, and Hon-Wah Tam. Short note: An integrable numerical algorithm for computing eigenvalues of a specially structured matrix. Numerical Linear Algebra with Applications, 18(2): 261–274, March 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saibaba:2021:RAG

[SHvBW21] Arvind K. Saibaba, Joseph Hart, and Bart van Bloemen Waanders. Randomized algorithms for generalized singular value decomposition with application to sensitivity analysis. Numerical Linear Algebra with Applications, 28 (4):e2364:1–e2364:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sugihara:2020:RPM

[SHZ20] Kota Sugihara, Ken Hayami, and Ning Zheng. Right preconditioned MINRES for singular systems. Numerical Linear Algebra with Applications, 27(3): e2277:1–e2277:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sidje:1997:APK

[Sid97] Roger B. Sidje. Alternatives for parallel Krylov subspace basis computation. Numerical Linear Algebra with Applications, 4(4):305–331, July/August 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15037;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15037&PLACEBO=IE.pdf.

Sidje:2011:IUG

[Sid11] Roger B. Sidje. Inexact uniformization and GMRES methods for large Markov chains. Numerical Linear Algebra with Applications, 18(6):947–960, November 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Simoncini:1999:NVR

[Sim99] V. Simoncini. A new variant of restarted GMRES. Numerical Linear Algebra with Applications, 6(1):61–77, January/February 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62002987;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62002987&
PLACEBO=IE.pdf. Czech-US

Workshop in Iterative Methods and Parallel Computing, Part I (Milovy, 1997).

Simoncini:2003:AFS

[Sim03] V. Simoncini. Algebraic formulations for the solution of the nullspace-free eigenvalue problem using the inexact Shift-and-Invert Lanczos method. Numerical Linear Algebra with Applications, 10(4): 357–375, June 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sirkovic:2019:RBA

[Sir19] Petar Sirković. A reduced basis approach to large-scale pseudospectra computation. Numerical Linear Algebra with Applications, 26(2):e2222:1– e2222:??, March 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Shen:2014:EDP

[SJBH14] Shu-Qian Shen, Ling Jian, Wen-Di Bao, and Ting-Zhu Huang. On the eigenvalue distribution of preconditioned nonsymmetric saddle point matrices. Numerical Linear Algebra with Applications, 21(4): 557–568, August 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sayed:2001:SSF

[SK01] A. H. Sayed and T. Kailath. A survey of spectral factorization methods. Numerical Linear Algebra with Applications, 8(6–7):467–496, September/November 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/85007288/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85007288&PLACEBO=IE.pdf.

Susnjara:2021:FSD

[SK21] Ana Susnjara and Daniel Kressner. A fast spectral divide-and-conquer method for banded matrices. Numerical Linear Algebra with Applications, 28(4):e2365:1– e2365:??, August 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schmitt:2022:TMO

[SKKS22] Britta Schmitt, Boris N.

Khoromskij, Venera Khoromskaia, and Volker Schulz. Tensor method for optimal control problems constrained by fractional three-dimensional elliptic operator with variable coefficients. Numerical Linear Algebra with Applications, 29 (1):e2404:1–e2404:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sturmer:2008:FFM

[SKR08] M. Stürmer, H. Köstler, and U. Rüde. A fast full multigrid solver for applications in image processing. Numerical Linear Algebra with Applications, 15 (2-3):187–200, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sun:2010:CPN

[SL10] Li-Ying Sun and Jun Liu. Constraint preconditioning for nonsymmetric indefinite linear systems. Numerical Linear Algebra with Applications, 17(4): 677–689, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schreiber:2019:PTI

[SL19] Martin Schreiber and Richard Loft. A parallel time integrator for solving the linearized shallow water equations on the rotating sphere. Numerical Linear Algebra with Applications, 26(2):e2220:1–e2220:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Swirydowicz:2021:LSG

[ŚLA^{+}21] Katarzyna Świrydowicz, Julien Langou, Shreyas Ananthan, Ulrike Yang, and Stephen Thomas. Low synchronization Gram–Schmidt and generalized minimal residual algorithms. Numerical Linear Algebra with Applications, 28(2): e2343:1–e2343:??, March 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saibaba:2016:RAG

[SLK16] Arvind K. Saibaba, Jonghyun Lee, and Peter K. Kitanidis. Randomized algorithms for generalized Hermitian eigenvalue problems with application to computing Karhunen– Loève expansion. Numerical Linear Algebra with Applications, 23(2):314–339, March 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schuermans:2004:SWL

[SLV04] M. Schuermans, P. Lemmerling, and S. Van Huffel. Structured weighted low rank approximation. Numerical Linear Algebra with Applications, 11(5–6):609–618, June/August 2004. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-1506 (electronic).

Schuermans:2006:BRH

[SLV06] M. Schuermans, P. Lemmerling, and S. Van Huffel. Blockrow Hankel weighted low rank approximation. Numerical Linear Algebra with Applications, 13(4):293–302, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sheikh:2013:CSL

[SLV13] A. H. Sheikh, D. Lahaye, and C. Vuik. On the convergence of shifted Laplace preconditioner combined with multilevel deflation. Numerical Linear Algebra with Applications, 20(4): 645–662, August 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Smietanski:2019:NEI

[Śmi19] Marek J. Śmietański. On a new exponential iterative method for solving nonsmooth equations. Numerical Linear Algebra with Applications, 26 (5):e2255:1–e2255:??, October 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sosonkina:2000:PSL

[SMSW00] Maria Sosonkina, John T.

Melson, Yousef Saad, and Layne T. Watson. Preconditioning strategies for linear systems arising in tire design. Numerical Linear Algebra with Applications, 7(7–8): 743–757, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505476/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505476&PLACEBO=IE.pdf.

Song:2020:RTC

[SNZ20] Guangjing Song, Michael K. Ng, and Xiongjun Zhang. Robust tensor completion using transformed tensor singular value decomposition. Numerical Linear Algebra with Applications, 27(3):e2299:1– e2299:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Soleymani:2014:FCI

[Sol14] F. Soleymani. A fast convergent iterative solver for approximate inverse of matrices. Numerical Linear Algebra with Applications, 21(3): 439–452, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Soto:2013:FRC

[Sot13] Ricardo L. Soto. A family of realizability criteria for the real and symmetric nonnegative inverse eigenvalue problem. Numerical Linear Algebra with Applications, 20(2): 336, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Salapaka:2006:ANP

[SP06] S. Salapaka and A. Peirce.

Analysis of a novel preconditioner for a class of plevel lower rank extracted systems. Numerical Linear Algebra with Applications, 13(6): 437–472, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schneider:2018:IGF

[SP18] F. S. Schneider and M. Pisarenco. Inverse generating function approach for the preconditioning of Toeplitz-block systems. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Salapaka:2005:ACB

[SPD05] S. Salapaka, A. Peirce, and M. Dahleh. Analysis of a circulant based preconditioner for a class of lower rank extracted systems. Numerical Linear Algebra with Applications, 12(1): 9–32, February 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Spigler:2021:NSI

[Spi21] Renato Spigler. On the numerical solution of ill-conditioned linear systems by regularization and iteration. Numerical Linear Algebra with Applications, 28(1):e2335:1– e2335:??, January 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Salinas:2013:MMC

[SRGL13] P. Salinas, C. Rodrigo, F. J. Gaspar, and F. J. Lisbona. Multigrid methods for cellcentered discretizations on triangular meshes. Numerical Linear Algebra with Applications, 20(4):626–644, August 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Simoncini:1997:AAS

[SS97] V. Simoncini and E. Sjöström. An algorithm for approximating the singular triplets of complex symmetric matrices. Numerical Linear Algebra with Applications, 4(6): 469–489, November/December 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15050;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=15050&PLACEBO=IE.pdf.

Saad:2002:AAR

[SS02] Y. Saad and B. Suchomel. ARMS: an algebraic recursive multilevel solver for general sparse linear systems. Numerical Linear Algebra with Applications, 9(5):359–378, July/August 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Simoncini:2007:RCD

[SS07] Valeria Simoncini and Daniel B.

Szyld. Recent computational developments in Krylov subspace methods for linear systems. Numerical Linear Algebra with Applications, 14(1): 1–59, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Stoyan:2004:GDA

[SSB04] G. Stoyan, G. Strauber, and Á. Baran. Generalizations to discrete and analytical Crouzeix-Velte decompositions. Numerical Linear Algebra with Applications, 11(5– 6):565–590, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sundar:2015:CMA

[SSB15] Hari Sundar, Georg Stadler, and George Biros. Comparison of multigrid algorithms for high-order continuous finite element discretizations. Numerical Linear Algebra with Applications, 22(4):664–680, August 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Shishkin:2019:FAT

[SSB19] Serge L. Shishkin, Arkadi Shalaginov, and Shaunak D. Bopardikar. Fast approximate truncated SVD. Numerical Linear Algebra with Applications, 26(4):e2246:1– e2246:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sugiyama:2023:WRM

[SSSF23] Masumi Sugiyama, Jacob B. Schroder, Ben S. Southworth, and Stephanie Friedhoff. Weighted relaxation for multigrid reduction in time. Numerical Linear Algebra with Applications, 30(1):e2465:1– e2465:??, January 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Stotland:2018:HPD

[SST18] Vadim Stotland, Oded Schwartz, and Sivan Toledo. Highperformance direct algorithms for computing the sign function of triangular matrices. Numerical Linear Algebra with Applications, 25(2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sadkane:2017:CAC

[ST17a] Miloud Sadkane and Ahmed Touhami. Convergence analysis of the ChebFilterCG algorithm. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Scott:2017:ISR

[ST17b] Jennifer Scott and Miroslav Tůma. Improving the stability and robustness of incomplete symmetric indefinite factorization preconditioners. Numerical Linear Algebra with Applications, 24(5):??, October 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sadkane:2019:CBV

[ST19] Miloud Sadkane and Ahmed Touhami. ChebStaBlkCG: A block variant of ChebFilterCG. Numerical Linear Algebra with Applications, 26(2): e2227:1–e2227:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sarabandi:2023:SMN

[ST23] Soheil Sarabandi and Federico Thomas. Solution methods to the nearest rotation matrix problem in R^{3} : a comparative survey. Numerical Linear Algebra with Applications, 30(5):e2492:1–e2492:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Starke:1996:MMR

[Sta96] Gerhard Starke. Multilevel minimal residual methods for nonsymmetric elliptic problems. Numerical Linear Algebra with Applications, 3(5): 351–367, September/October 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001002.

Stewart:1995:SBH

[Ste95]  G. W. Stewart.  On the

solution of block Hessenberg systems. Numerical Linear Algebra with Applications, 2 (3):287–296, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Stefanovski:1999:GEA

[Ste99]  Jovan Stefanovski.  Gen-

erating equations approach for quadratic matrix equations. Numerical Linear Algebra with Applications, 6(4): 295–326, June 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/65500097/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=65500097&PLACEBO=IE.pdf.

Stoer:1992:DAS

[Sto92] Josef Stoer. A dual algorithm for solving degenerate linearly constrained linear least squares problems. Journal of Numerical Linear Algebra with Applications, 1(2):103–131, 1992. CODEN ???? ISSN 0129-3281.

Savostyanov:2012:FTM

[STZ12] D. V. Savostyanov, E. E. Tyrtyshnikov, and N. L. Zamarashkin. Fast truncation of mode ranks for bilinear tensor operations. Numerical Linear Algebra with Applications, 19(1):103–111, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sun:2005:NBE

[Sun05] Ji-Guang Sun. A note on backward errors for structured linear systems. Numerical Linear Algebra with Applications, 12 (7):585–603, September 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sun:2006:SEI

[Sun06] Li-Ying Sun. Some extensions of the improved modified Gauss–Seidel iterative method for H-matrices. Numerical Linear Algebra with Applications, 13(10):869–876, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Stammberger:2011:AML

[SV11] Markus Stammberger and Heinrich Voss. Automated multi-level sub-structuring for fluid-solid interaction problems. Numerical Linear Algebra with Applications, 18(3): 411–427, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Swijsen:2022:TCU

[SVV22] Lars Swijsen, Joeri Van der Veken, and Nick Vannieuwenhoven. Tensor completion using geodesics on Segrè manifolds. Numerical Linear Algebra with Applications, 29 (6):e2446:1–e2446:??, December 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saad:1996:DDQ

[SW96] Yousef Saad and Kesheng Wu. DQGMRES: a direct quasiminimal residual algorithm based on incomplete orthogonalization. Numerical Linear Algebra with Applications, 3(4):329–343, July/August 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000994.

Stoll:2012:PPD

[SW12] Martin Stoll and Andy Wathen. Preconditioning for partial differential equation constrained optimization with control constraints. Numerical Linear Algebra with Applications, 19(1):53–71, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Sosonkina:1998:NAG

[SWKW98] Maria Sosonkina, Layne T. Watson, Rakesh K. Kapania, and Homer F. Walker. A new adaptive GMRES algorithm for achieving high accuracy. Numerical Linear Algebra with Applications, 5(4): 275–297, July/August 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
10005754;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=10005754&
PLACEBO=IE.pdf.

Schneck:2021:IMP

[SWW21] Jakob Schneck, Martin Weiser, and Florian Wende. Impact of mixed precision and storage layout on additive Schwarz smoothers. Numerical Linear Algebra with Applications, 28 (4):e2366:1–e2366:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Schleicher:2007:MMM

[SWY07] Jörg Schleicher, Lei Wang, and Jin Yun Yuan. Mathematical modelling and mathematical methods in energy. Numerical Linear Algebra with Applications, 14(4):255, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Shen:2015:UAV

[SX15] Hailun Shen and Hua Xiang.

Uzawa algorithms with variable relaxation for nonsymmetric generalized saddle point problems. Numerical Linear Algebra with Applications, 22 (6):1020–1038, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Song:2018:TPS

[SY18a] Xiao-Liang Song and Bo Yu. A two-phase strategy for control constrained elliptic optimal control problems. Numerical Linear Algebra with Applications, 25(4):??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Steinbach:2018:CAM

[SY18b] Olaf Steinbach and Huidong Yang. Comparison of algebraic multigrid methods for an adaptive space–time finite-element discretization of the heat equation in 3D and 4D. Numerical Linear Algebra with Applications, 25(3):??, May 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Saad:1999:DTT

[SZ99] Yousef Saad and Jun Zhang. Diagonal threshold techniques in robust multi-level ILU preconditioners for general sparse linear systems. Numerical Linear Algebra with Applications, 6(4):257–280, June 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.

com/cgi-bin/abstract/65500095/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=65500095&PLACEBO=IE.pdf.

Shi:2011:NEM

[SZ11] Lei Shi and Ding-Xuan Zhou.

Normal estimation on manifolds by gradient learning. Numerical Linear Algebra with Applications, 18(2):249– 259, March 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Szularz:2014:ISA

[Szu14] M. Szularz. Imposing symmetry in augmented linear systems. Numerical Linear Algebra with Applications, 21(6): 781–795, December 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Szyld:1994:ECC

[Szy94] Daniel B. Szyld. Equivalence of conditions for convergence of iterative methods for singular equations. Numerical Linear Algebra with Applications, 1(2):151–154, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Thum:2010:TPO

[TC10] Peter Thum and Tanja Clees.

Towards physics-oriented smoothing in algebraic multigrid for systems of partial differential equations arising in multi-ion transport and reaction models. Numerical Linear Algebra with Applications, 17(2-3): 253–271, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tabeart:2018:CLP

[TDH^{+}18] Jemima M. Tabeart, Sarah L. Dance, Stephen A. Haben, Amos S. Lawless, Nancy K. Nichols, and Joanne A. Waller. The conditioning of leastsquares problems in variational data assimilation. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tabeart:2022:NBC

[TDL^{+}22] Jemima M. Tabeart, Sarah L. Dance, Amos S. Lawless, Nancy K. Nichols, and Joanne A. Waller. New bounds on the condition number of the Hessian of the preconditioned variational data assimilation problem. Numerical Linear Algebra with Applications, 29 (1):e2405:1–e2405:??, January 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Thekale:2010:ONM

[TGKR10] A. Thekale, T. Gradl, K. Klamroth, and U. Rüde. Optimizing the number of multigrid cycles in the full multigrid algorithm. Numerical Linear Algebra with Applications, 17(2-3):199–210, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Treister:2019:MSH

[TH19] Eran Treister and Eldad Haber. A multigrid solver to the Helmholtz equation with a point source based on travel time and amplitude. Numerical Linear Algebra with Applications, 26(1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tian:2009:DII

[THC09] Gui-Xian Tian, Ting-Zhu Huang, and Shu-Yu Cui. The digraphs and inclusion intervals of matrix singular values. Numerical Linear Algebra with Applications, 16(8): 677–687, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Thess:1998:PMP

[The98] Michael Thess. Parallel multilevel preconditioners for thin smooth shell finite element analysis. Numerical Linear Algebra with Applications, 5(5): 401–440, September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
62000047;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=62000047&
PLACEBO=IE.pdf. Special Is-

sue: PRISM 97 (Nijmegen).

Tian:2013:LSS

[Tia13] Yongge Tian. Least-squares solutions and least-rank solutions of the matrix equation AXA^{∗} = B and their relations. Numerical Linear Algebra with Applications, 20(5): 713–722, October 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Thao:2021:CAC

[TR21] Nguyen T. Thao and Dominik Rzepka. Coarse-arithmetic coordinate descent for the resolution of semidefinite linear systems. Numerical Linear Algebra with Applications, 28 (5):e2377:1–e2377:??, October 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Trench:2005:ARB

[Tre05] W. F. Trench. Asymptotic relationships between singular values of structured matrices similarly generated by different formal expansions of a rational function. Numerical Linear Algebra with Applications, 12(2–3):111–116, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Trench:2013:IPU

[Tre13] W. Trench. Inverse problems for unilevel block αcirculants. Numerical Linear Algebra with Applications, 20 (2):349–356, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tang:2012:PMC

[TS12] Jok M. Tang and Yousef Saad.

A probing method for computing the diagonal of a matrix inverse. Numerical Linear Algebra with Applications, 19(3): 485–501, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tavakolipour:2020:AEE

[TS20] Hanieh Tavakolipour and Fatemeh Shakeri. Asymptotics of the eigenvalues for exponentially parameterized pentadiagonal matrices. Numerical Linear Algebra with Applications, 27(6):e2330:1–e2330:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tavernier:2021:TLP

[TSMM21] Joris Tavernier, Jaak Simm, Karl Meerbergen, and Yves Moreau. Two-level preconditioning for ridge regression. Numerical Linear Algebra with Applications, 28(4):e2371:1– e2371:??, August 2021. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Trias:2006:DSF

[TSPSO06] F. X. Trias, M. Soria, C. D.

Pérez-Segarra, and A. Oliva. A direct Schur–Fourier decomposition for the efficient solution of high-order Poisson equations on loosely coupled parallel computers. Numerical Linear Algebra with Applications, 13(4):303–326, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tebbens:2010:PUS

[TT10] Jurjen Duintjer Tebbens and Miroslav Tůma. Preconditioner updates for solving sequences of linear systems in matrix-free environment. Numerical Linear Algebra with Applications, 17(6):997–1019, December 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tsuji:2015:AAS

[TT15] P. Tsuji and R. Tuminaro. Augmented AMG-shifted Laplacian preconditioners for indefinite Helmholtz problems. Numerical Linear Algebra with Applications, 22(6):1077–1101, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Turek:2000:CIF

[Tur00] S. Turek. CFD for incompressible flow: numerical efficiency versus gigaflops. Numerical Linear Algebra with Applications, 7(6):473–482, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516701/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516701&PLACEBO=IE.pdf.

Tomljanovic:2020:SAD

[TV20] Zoran Tomljanović and Matthias Voigt. Semi-active H_{∞} damping optimization by adaptive interpolation. Numerical Linear Algebra with Applications, 27(4):e2300:1–e2300:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Toivanen:2020:FFT

[TW20] Jari Toivanen and Monika Wolfmayr*. A Fast Fourier transform based direct solver for the Helmholtz problem. Numerical Linear Algebra with Applications, 27(3):e2283:1– e2283:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Treister:2010:SSM

[TY10] Eran Treister and Irad Yavneh.

Square and stretch multigrid for stochastic matrix eigenproblems. Numerical Linear Algebra with Applications, 17 (2-3):229–251, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tian:2023:MVS

[TYF23] Yingjie Tian, Xiaotong Yu, and Saiji Fu. Multi-view side information-incorporated tensor completion. Numerical Linear Algebra with Applications, 30(5):e2485:1–e2485:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tyrtyshnikov:1992:SPL

[Tyr92] E. E. Tyrtyshnikov. On symmetrizing preconditioners with low rank updates. Journal of Numerical Linear Algebra with Applications, 1(2): 227–235, 1992. CODEN ???? ISSN 0129-3281.

Tyrtyshnikov:2005:SPO

[Tyr05] Eugene Tyrtyshnikov. Structured preconditioners for operator equations. Numerical Linear Algebra with Applications, 12(2–3):251–259, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Tang:2023:SPM

[TYZL23] Ling Tang, Yajie Yu, Yanjun Zhang, and Hanyu Li. Sketchand-project methods for tensor linear systems. Numerical Linear Algebra with Applications, 30(2):e2470:1–e2470:??, March 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Uhlig:2023:CFV

[Uhl23] Frank Uhlig. Constructing the field of values of decomposable and general matrices using the ZNN based path following method. Numerical Linear Algebra with Applications, 30 (6):e2513:1–e2513:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Umetani:2009:MBS

[UMO09] N. Umetani, S. P. MacLachlan, and C. W. Oosterlee. A multigrid-based shifted Laplacian preconditioner for a fourth-order Helmholtz discretization. Numerical Linear Algebra with Applications, 16 (8):603–626, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ubaru:2019:SMC

[US19] Shashanka Ubaru and Yousef Saad. Sampling and multilevel coarsening algorithms for fast matrix approximations. Numerical Linear Algebra with Applications, 26(3): e2234:1–e2234:??, May 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ullah:2021:DFS

[USS21] Najib Ullah, Jamilu Sabi’u, and Abdullah Shah. A derivative-free scaling memoryless Broyden–Fletcher–Goldfarb– Shanno method for solving a system of monotone nonlinear equations. Numerical Linear Algebra with Applications, 28(5):e2374:1–e2374:??, October 2021. CODEN NLAAEM.

ISSN 1070-5325 (print), 1099-

1506 (electronic).

Vabishchevich:2020:AFP

[Vab20] P. N. Vabishchevich. Approximation of a fractional power of an elliptic operator. Numerical Linear Algebra with Applications, 27(3):e2287:1–e2287:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vabishchevich:2023:NAF

[Vab23] Petr N. Vabishchevich. Nonlinear approximation of functions based on nonnegative least squares solver. Numerical Linear Algebra with Applications, 30(6):e2522:1–e2522:??, December 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vanderstraeten:2000:APB

[Van00] Denis Vanderstraeten. An accurate parallel block Gram– Schmidt algorithm without reorthogonalization. Numerical Linear Algebra with Applications, 7(4):219–236, May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72507877/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72507877&PLACEBO=IE.pdf.

Varga:2008:SPR

[Var08] A. Varga. On solving periodic Riccati equations. Numerical Linear Algebra with Applications, 15(9):809–835, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:1992:PNI

[Vas92] P. S. Vassilevski. Preconditioning nonsymmetric and indefinite finite element matrices. Journal of Numerical Linear Algebra with Applications, 1(1):59–76, 1992. CODEN ???? ISSN 0129-3281.

Vassilevski:2002:SME

[Vas02] Panayot S. Vassilevski. Sparse matrix element topology with application to AMG(e) and preconditioning. Numerical Linear Algebra with Applications, 9(6–7):429–444, September/November 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:2003:OBR

[Vas03] Panayot S. Vassilevski. On the occasion of the 60th birthday of Raytcho Lazarov. Numerical Linear Algebra with Applications, 10(1–2):1, January/March 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:2005:SIN

[Vas05] Panayot S. Vassilevski. Special issue of NLA on the occasion of the 70th birthday of Owe Axelsson. Numerical Linear Algebra with Applications, 12(5–6):391–392, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vandenEshof:2002:CJD

[vdE02] Jasper van den Eshof. The convergence of Jacobi–Davidson iterations for Hermitian eigenproblems. Numerical Linear Algebra with Applications, 9(2):163–179, March 2002. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/89013909/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=89013909&PLACEBO=IE.pdf.

Verbeek:2000:RNS

[Ver00] Menno E. Verbeek. Repairing near-singularity for dense EMC problems by adaptive basis techniques. Numerical Linear Algebra with Applications, 7(7–8):617–634, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505482/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505482&PLACEBO=IE.pdf.

Vakili:2023:PES

[VEV23] Seryas Vakili, Ghodrat Ebadi, and Cornelis Vuik. A parameterized extended shift-splitting preconditioner for nonsymmetric saddle point problems. Numerical Linear Algebra with Applications, 30(4):e2478:1– e2478:??, August 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vargas:2013:SMC

[VFdV13] Alessandro N. Vargas, Walter Furloni, and João B. R. do Val. Second moment constraints and the control problem of Markov jump linear systems. Numerical Linear Algebra with Applications, 20(2): 357–368, March 2013. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanGijzen:2015:FMS

[vGSZ15] Martin B. van Gijzen, Gerard L. G. Sleijpen, and Jens-Peter M. Zemke. Flexible and multi-shift induced dimension reduction algorithms for solving large sparse linear systems. Numerical Linear Algebra with Applications, 22(1):1–25, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Voronin:2022:LOP

[VHM^{+}22] Alexey Voronin, Yunhui He, Scott MacLachlan, Luke N. Olson, and Raymond Tuminaro. Low-order preconditioning of the Stokes equations. Numerical Linear Algebra with Applications, 29(3): e2426:1–e2426:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

VanBeeumen:2016:REI

[VJM16] Roel Van Beeumen, Elias Jarlebring, and Wim Michiels. A rank-exploiting infinite Arnoldi algorithm for nonlinear eigenvalue problems. Numerical Linear Algebra with Applications, 23(4):607–628, August 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanKan:2000:FPC

[vKVW00] J. van Kan, C. Vuik, and P. Wesseling. Fast pressure calculation for 2D and 3D time dependent incompressible flow. Numerical Linear Algebra with Applications, 7 (6):429–447, September 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516698/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516698&PLACEBO=IE.pdf.

Vassilevski:1996:PMF

[VL96] Panayot S. Vassilevski and Raytcho D. Lazarov. Preconditioning mixed finite element saddle-point elliptic problems. Numerical Linear Algebra with Applications, 3(1): 1–20, January/February 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000501.

Vecharynski:2011:ACC

[VL11] Eugene Vecharynski and Julien Langou. Any admissible cycleconvergence behavior is possible for restarted GMRES at its initial cycles. Numerical Linear Algebra with Applications, 18(3):499–511, May 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vlachkova:2000:NTA

[Vla00]  Krassimira Vlachkova.  A

Newton-type algorithm for solving an extremal constrained interpolation problem. Numerical Linear Algebra with Applications, 7 (3):133–146, April/May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72001234/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72001234&PLACEBO=IE.pdf.

vanNotay:2000:RAM

[vN00] Y. van Notay. A robust algebraic multilevel preconditioner for non-symmetric Mmatrices. Numerical Linear Algebra with Applications, 7(5): 243–267, July/August 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72508405/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72508405&PLACEBO=IE.pdf.

vanNoorden:2007:CPG

[vNR07] Tycho van Noorden and Joost Rommes. Computing a partial generalized real Schur form using the Jacobi–Davidson method. Numerical Linear Algebra with Applications, 14(3): 197–215, ???? 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Voevodin:1992:PSS

[Voe92]  V. V. Voevodin.  Parallel

software from the standpoint of a mathematician. Journal of Numerical Linear Algebra with Applications, 1(2): 237–242, 1992. CODEN ???? ISSN 0129-3281.

Vomel:2010:NHR

[Vöm10] Christof Vömel. A note on harmonic Ritz values and their reciprocals. Numerical Linear Algebra with Applications, 17 (1):97–108, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vomel:2012:NGF

[Vöm12]  Christof Vömel.  A note

on generating finer-grain parallelism in a representation tree. Numerical Linear Algebra with Applications, 19(5): 869–879, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Voss:2009:MPN

[Vos09] Heinrich Voss. A minmax principle for nonlinear eigenproblems depending continuously on the eigenparameter. Numerical Linear Algebra with Applications, 16(11-12):899– 913, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

VanHuffel:1995:ERA

[VP95] Sabine Van Huffel and Haesun Park. Efficient reduction algorithms for bordered band matrices. Numerical Linear Algebra with Applications, 2(2):95– 113, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vysotsky:2023:TRB

[VR23] Lev Vysotsky and Maxim Rakhuba. Tensor rank bounds and explicit QTT representations for the inverses of circulant matrices. Numerical Linear Algebra with Applications, 30(3):e2461:1–e2461:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanRaalte:2005:TLM

[vRH05] M. H. van Raalte and P. W.

Hemker. Two-level multigrid analysis for the convectiondiffusion equation discretized by a discontinuous Galerkin method. Numerical Linear Algebra with Applications, 12(5– 6):563–584, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vo:2017:ALS

[VS17] Huy D. Vo and Roger B. Sidje. Approximating the large sparse matrix exponential using incomplete orthogonalization and Krylov subspaces of variable dimension. Numerical Linear Algebra with Applications, 24(3):??, May 2017. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vanlent:2009:EMC

[VSG09] Jan Van lent, Robert Scheichl, and Ivan G. Graham. Energyminimizing coarse spaces for two-level Schwarz methods for multiscale PDEs. Numerical Linear Algebra with Applications, 16(10):775–799, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanderVorst:1994:GFN

[vV94] H. A. van der Vorst and C. Vuik. GMRESR: a family of nested GMRES methods. Numerical Linear Algebra with Applications, 1(4):369– 386, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

VanBarel:2005:OST

[VVM05a] M. Van Barel, E. Van Camp, and N. Mastronardi. Orthogonal similarity transformation into block-semiseparable matrices of semiseparability rank k. Numerical Linear Algebra with Applications, 12(10):981– 1000, December 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vandebril:2005:IQA

[VVM05b] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. An implicit QR algorithm for symmetric semiseparable matrices. Numerical Linear Algebra with Applications, 12(7): 625–658, September 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vandebril:2005:NRD

[VVM05c] Raf Vandebril, Marc Van Barel, and Nicola Mastronardi. A note on the representation and definition of semiseparable matrices. Numerical Linear Algebra with Applications, 12(8): 839–858, October 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanVenetie:2023:EST

[vVW23] Raymond van Venetië and Jan Westerdiep. Efficient spacetime adaptivity for parabolic evolution equations using wavelets in time and finite elements in space. Numerical Linear Algebra with Applications, 30(1):e2457:1–e2457:??, January 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:1997:SHB

[VW97] Panayot S. Vassilevski and Junping Wang. Stabilizing the hierarchical basis by approximate wavelets. I. Theory. Numerical Linear Algebra with Applications, 4(2): 103–126, March/April 1997. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15001015;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=15001015&
PLACEBO=IE.pdf.

Vassilevski:2001:SIS

[VW01] Panayot Vassilevski and Carol S. Woodward. Special issue on Solution Methods for Large-Scale Non-linear Problems. Numerical Linear Algebra with Applications, 8(8):497, December 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88010576/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88010576&PLACEBO=IE.pdf.

Vasquez:2015:PIR

[VW15] Fernando Guevara Vasquez and Benjamin Z. Webb. Pseudospectra of isospectrally reduced matrices. Numerical Linear Algebra with Applications, 22(1):145–174, January 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:2014:RCA

[VY14] Panayot S. Vassilevski and Ulrike Meier Yang. Reducing communication in algebraic multigrid using additive variants. Numerical Linear Algebra with Applications, 21(2): 275–296, March 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:2008:AMM

[VZ08] Panayot S. Vassilevski and Ludmil T. Zikatanov. Advanced multigrid methods for systems of PDEs. Numerical Linear Algebra with Applications, 15(5):393, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Vassilevski:2014:CPG

[VZ14] Panayot S. Vassilevski and Ludmil T. Zikatanov. Commuting projections on graphs. Numerical Linear Algebra with Applications, 21(3):297–315, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wan:2000:IPC

[Wan00] Wing Lok Wan. Interface preserving coarsening multigrid for elliptic problems with highly discontinuous coefficients. Numerical Linear Algebra with Applications, 7(7–8): 727–742, October/December 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73505480/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73505480&PLACEBO=IE.pdf.

Wang:2018:NBA

[Wan18a] Zeng-Qi Wang. A note on the block alternating splitting implicit iteration method for complex saddle-point problems. Numerical Linear Algebra with Applications, 25(6): ??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2018:CAS

[Wan18b] Zeng-Qi Wang. On a Chebyshev accelerated splitting iteration method with application to two-by-two block linear systems. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2014:NED

[WBL14] Shi-Liang Wu, Luca Bergamaschi, and Cui-Xia Li. A note on eigenvalue distribution of constraint-preconditioned symmetric saddle point matrices. Numerical Linear Algebra with Applications, 21(1): 171–174, January 2014. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Warsa:2004:PMD

[WBWM04] James S. Warsa, Michele Benzi, Todd A. Wareing, and Jim E. Morel. Preconditioning a mixed discontinuous finite element method for radiation diffusion. Numerical Linear Algebra with Applications, 11(8–9): 795–811, October/November 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2020:PTS

[WCW20] Xuezhong Wang, Maolin Che, and Yimin Wei. Preconditioned tensor splitting AOR iterative methods for h -tensor equations. Numerical Linear Algebra with Applications, 27 (6):e2329:1–e2329:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2015:CAB

[WCZ15] Yiju Wang, Louis Caccetta, and Guanglu Zhou. Convergence analysis of a block improvement method for polynomial optimization over unit spheres. Numerical Linear Algebra with Applications, 22 (6):1059–1076, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2008:LCG

[WD08] Li-Ping Wang and Yu-Hong Dai. Left conjugate gradient method for non-Hermitian linear systems. Numerical Linear Algebra with Applications, 15 (10):891–909, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2009:PIA

[WDS09] Yingxi Wang, Kui Du, and Weiwei Sun. Preconditioning iterative algorithm for the electromagnetic scattering from a large cavity. Numerical Linear Algebra with Applications, 16(5):345–363, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Webster:2010:AMO

[Web10a] R. Webster. Algebraic multigrid and 4th-order discretedifference equations of incompressible fluid flow. Numerical Linear Algebra with Applications, 17(4):655–676, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Webster:2010:AMM

[Web10b] R. Webster. Algebraic multigrid, mixed-order interpolation, and incompressible fluid flow. Numerical Linear Algebra with Applications, 17(1): 17–42, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Webster:2018:CAS

[Web18] Ronald Webster. CLC in AMG solvers for saddle-point problems. Numerical Linear Algebra with Applications, 25(2): ??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Weiss:1994:PGC

[Wei94] Rüdiger Weiss. Properties of generalized conjugate gradient methods. Numerical Linear Algebra with Applications, 1(1):45–63, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2015:TCF

[WF15] Gang Wu and Ting-Ting Feng.

A theoretical contribution to the fast implementation of null linear discriminant analysis with random matrix multiplication. Numerical Linear Algebra with Applications, 22 (6):1180–1188, December 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Washio:1994:PBP

[WH94] Takumi Washio and Ken Hayami. Parallel block preconditioning based on SSOR and MILU. Numerical Linear Algebra with Applications, 1(6):533–553, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wieners:1999:MMP

[Wie99]  Christian Wieners.  Multi-

grid methods for Prandtl-Reuss plasticity. Numerical Linear Algebra with Applications, 6(6):457–478, September 1999. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/67501479/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=67501479&PLACEBO=IE.pdf.

Walden:1995:OBP

[WKS95] Bertil Waldén, Rune Karlson, and Ji Guang Sun. Optimal backward perturbation bounds for the linear least squares problem. Numerical Linear Algebra with Applications, 2 (3):271–286, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wei:2003:IPB

[WL03] Yimin Wei and Xiezhang Li. An improvement on perturbation bounds for the Drazin inverse. Numerical Linear Algebra with Applications, 10(7): 563–575, October/November 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wei:2008:GFM

[WL08] Musheng Wei and Qiaohua Liu. On growth factors of the modified Gram–Schmidt algorithm. Numerical Linear Algebra with Applications, 15(7): 621–636, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2012:LBS

[WLBH12] Lei Wang, Heng Liang, Fengshan Bai, and Yan Huo. A load balancing strategy for parallel computation of sparse permanents. Numerical Linear Algebra with Applications, 19 (6):1017–1030, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2021:LRT

[WLC21] Ping-Ping Wang, Liang Li, and Guang-Hui Cheng. Low-rank tensor completion with sparse regularization in a transformed domain. Numerical Linear Algebra with Applications, 28 (6):e2387:1–e2387:??, December 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2012:PTT

[WM12] Chuan-Long Wang and Guo-Yan Meng. A practical twoterm acceleration algorithm for linear systems. Numerical Linear Algebra with Applications, 19(3):571–583, May 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wei:2005:DSG

[WN05] Yimin Wei and Michael K. Ng. Displacement structure of group inverses. Numerical Linear Algebra with Applications, 12(2–3):103–110, March/April 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2018:LRA

[WN18] Xiaofei Wang and Carmeliza Navasca. Low-rank approximation of tensors via sparse optimization. Numerical Linear Algebra with Applications, 25 (2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2007:SSR

[WQ07] Yiju Wang and Liqun Qi. On the successive supersymmetric rank-1 decomposition of higher-order supersymmetric tensors. Numerical Linear Algebra with Applications, 14 (6):503–519, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2009:PMC

[WQZ09] Yiju Wang, Liqun Qi, and Xinzhen Zhang. A practical method for computing the largest M-eigenvalue of a fourth-order partially symmetric tensor. Numerical Linear Algebra with Applications, 16 (7):589–601, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2018:SPC

[WRW18] Jun-Gang Wang, Yu-Hong Ran, and Dong-Ling Wang. On structure preserving and circulant preconditioners for the space fractional coupled nonlinear Schrödinger equations. Numerical Linear Algebra with Applications, 25(4):??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wen:2019:PDM

[WSN19] You-Wei Wen, Hai-Wei Sun, and Michael K. Ng. A primaldual method for the Meyer model of cartoon and texture decomposition. Numerical Linear Algebra with Applications, 26(2):e2224:1–e2224:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2015:ISI

[WtFW15] Gang Wu, Ting ting Feng, and Yimin Wei. An inexact shiftand-invert Arnoldi algorithm for Toeplitz matrix exponential. Numerical Linear Algebra with Applications, 22(4): 777–792, August 2015. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wiesner:2014:MTN

[WTWG14] T. A. Wiesner, R. S. Tuminaro, W. A. Wall, and M. W. Gee. Multigrid transfers for nonsymmetric systems based on Schur complements and Galerkin projections. Numerical Linear Algebra with Applications, 21(3): 415–438, May 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2010:CEI

[WTZD10] Ai-Guo Wu, Lin Tong, Ying Zhang, and Guang-Ren Duan. Comments on ‘An efficient iterative method for solving the matrix equation AXB + CY D = E’. Numerical Linear Algebra with Applications, 17(4):735–739, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See [yPxP06].

Wu:2015:SVH

[Wu15] Shi-Liang Wu. Several variants of the Hermitian and skew-Hermitian splitting method for a class of complex symmetric linear systems. Numerical Linear Algebra with Applications, 22(2):338–356, March 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2007:PAA

[WW07] Gang Wu and Yimin Wei. A Power–Arnoldi algorithm for computing PageRank. Numerical Linear Algebra with Applications, 14(7):521–546, ????

2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wei:2008:NME

[WW08a] Guo-Wei Wei and Ge Wang.

Numerical methods in engineering (CNM). Numerical Linear Algebra with Applications, 15(1):83, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wei:2008:PAG

[WW08b] Musheng Wei and Minghui Wang. Perturbation analysis for the generalized Schur complement of a positive semidefinite matrix. Numerical Linear Algebra with Applications, 15(1):1–11, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2011:NIR

[WW11] Xinyuan Wu and Zhengyu Wang. A new iterative refinement with roundoff error analysis. Numerical Linear Algebra with Applications, 18(2): 275–282, March 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wang:2020:NSD

[WW20] Yunjie Wang and Gang Wu.

New strategies for determining backward perturbation bound of approximate twosided Krylov subspaces. Numerical Linear Algebra with Applications, 27(6):e2324:1– e2324:??, December 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wen:2015:SSN

[WWC^{+}15] You-Wei Wen, Man Wang, Zhiying Cao, Xiaoqing Cheng, Wai-Ki Ching, and Vassilios S. Vassiliadis. Sparse solution of nonnegative least squares problems with applications in the construction of probabilistic Boolean networks. Numerical Linear Algebra with Applications, 22(5):883–899, October 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wheeler:2010:EAM

[WWX10] Mary F. Wheeler, Tim Wildey, and Guangri Xue. Efficient algorithms for multiscale modeling in porous media. Numerical Linear Algebra with Applications, 17(5):771–785, October 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2015:MMB

[wX15] Wei wei Xu. Modified modulusbased matrix splitting iteration methods for linear complementarity problems. Numerical Linear Algebra with Applications, 22(4):748–760, August 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2021:SAR

[WX21] Nianci Wu and Hua Xiang.

Semiconvergence analysis of the randomized row iterative method and its extended variants. Numerical Linear Algebra with Applications, 28 (1):e2334:1–e2334:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Walker:1994:SG

[WZ94] Homer F. Walker and Lu Zhou.

A simpler GMRES. Numerical Linear Algebra with Applications, 1(6):571–581, 1994. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wu:2018:STP

[WZZ18] Shu-Lin Wu, Hui Zhang, and Tao Zhou. Solving timeperiodic fractional diffusion equations via diagonalization technique and multigrid. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xie:2013:ESL

[XC13] Jinshan Xie and An Chang.

On the Z-eigenvalues of the signless Laplacian tensor for an even uniform hypergraph. Numerical Linear Algebra with Applications, 20(6):1030–1045, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2016:PPI

[XCG16] Wei-Ru Xu, Guoliang Chen, and Yi Gong. Procrustes problems and inverse eigenproblems for multilevel block αcirculants. Numerical Linear Algebra with Applications, 23(5):906–930, October 2016. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xia:2010:FAH

[XCGL10] Jianlin Xia, Shivkumar Chandrasekaran, Ming Gu, and Xiaoye S. Li. Fast algorithms for hierarchically semiseparable matrices. Numerical Linear Algebra with Applications, 17(6):953–976, December 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xiang:2010:KPA

[XG10] Hua Xiang and Laura Grigori. Kronecker product approximation preconditioners for convection–diffusion model problems. Numerical Linear Algebra with Applications, 17 (4):691–712, ???? 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xie:2003:SCI

[XHZ03] Dongxiu Xie, Xiyan Hu, and Lei Zhang. The solvability conditions for inverse eigenproblem of symmetric and antipersymmetric matrices and its approximation. Numerical Linear Algebra with Applications, 10(3):223–234, April/ May 2003. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xia:2012:RIO

[Xia12] J. Xia. A robust inner–outer hierarchically semi-separable preconditioner. Numerical Linear Algebra with Applications, 19(6):992–1016, December 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xie:2011:IBM

[Xie11] Huiqing Xie. IRAM-based method for eigenpairs and their derivatives of large matrixvalued functions. Numerical Linear Algebra with Applications, 18(3):513–538, May 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xie:2021:PBM

[Xie21] Huiqing Xie. A perturbationbased method for a parameterdependent nonlinear eigenvalue problem. Numerical Linear Algebra with Applications, 28 (4):e2355:1–e2355:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xiao:2012:ADM

[XJ12] Yun-Hai Xiao and Zheng-Fen Jin. An alternating direction method for linear-constrained matrix nuclear norm minimization. Numerical Linear Algebra with Applications, 19(3): 541–554, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2017:AIB

[XM17] Xiaowen Xu and Zeyao Mo. Algebraic interface-based coarsening AMG preconditioner for multi-scale sparse matrices with applications to radiation hydrodynamics computation. Numerical Linear Algebra with Applications, 24 (2):??, March 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2009:TFM

[XQ09] Wei Xu and Sanzheng Qiao. A twisted factorization method for symmetric SVD of a complex symmetric tridiagonal matrix. Numerical Linear Algebra with Applications, 16(10): 801–815, ???? 2009. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xiao:2011:RPH

[XS11] Yingxiong Xiao and Shi Shu. A robust preconditioner for higher order finite element discretizations in linear elasticity. Numerical Linear Algebra with Applications, 18(5): 827–842, October 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xiao:2009:GBA

[XSZ09] Yingxiong Xiao, Shi Shu, and Tuyan Zhao. A geometricbased algebraic multigrid method for higher-order finite element equations in two-dimensional linear elasticity. Numerical Linear Algebra with Applications, 16(7): 535–559, ???? 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2022:IRK

[XX22] Shengjie Xu and Fei Xue. Inexact rational Krylov subspace method for eigenvalue problems. Numerical Linear Algebra with Applications, 29 (5):e2437:1–e2437:??, October 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xin:2020:ERR

[XXCB20] Zixing Xin, Jianlin Xia, Stephen Cauley, and Venkataramanan Balakrishnan. Effectiveness and robustness revisited for a preconditioning technique based on structured incomplete factorization. Numerical Linear Algebra with Applications, 27(3):e2294:1– e2294:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xie:2019:RAT

[XXW19] Pengpeng Xie, Hua Xiang, and Yimin Wei. Randomized algorithms for total least squares problems. Numerical Linear Algebra with Applications, 26 (1):??, January 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xu:2010:TLA

[XZS10] Hongru Xu, Jinping Zeng, and Zhe Sun. Two-level additive Schwarz algorithms for nonlinear complementarity problem with an M-function. Numerical Linear Algebra with Applications, 17(4):599–613, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Xiao:2015:EAM

[XZS15] Yingxiong Xiao, Zhiyang Zhou, and Shi Shu. An efficient algebraic multigrid method for quadratic discretizations of linear elasticity problems on some typical anisotropic meshes in three dimensions. Numerical Linear Algebra with Applications, 22(3):465–482, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2004:URP

[Yan04] Ulrike Meier Yang. On the use of relaxation parameters in hybrid smoothers. Numerical Linear Algebra with Applications, 11(2–3):155–172, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2010:LRI

[Yan10]  Ulrike Meier Yang.  On

long-range interpolation operators for aggressive coarsening. Numerical Linear Algebra with Applications, 17(2-3): 453–472, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2018:SNM

[Yan18] Ai-Li Yang. Scaled norm minimization method for computing the parameters of the HSS and the two-parameter HSS preconditioners. Numerical Linear Algebra with Applications, 25(4):??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2023:FAA

[Yan23]  Zhao Yang.  A fast and

accurate algorithm for solving linear systems associated with a class of negative matrix. Numerical Linear Algebra with Applications, 30 (5):e2483:1–e2483:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yavneh:2004:E

[Yav04]  Irad Yavneh.  Editorial.

Numerical Linear Algebra with Applications, 11(2–3):91, March/April 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yu:2023:NOT

[YB23] Quan Yu and Minru Bai. Nonconvex optimization for thirdorder tensor completion under wavelet transform. Numerical Linear Algebra with Applications, 30(5):e2489:1– e2489:??, October 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yao:2019:RVF

[YBZ19] Teng-Teng Yao, Zheng-Jian Bai, and Zhi Zhao. A Riemannian variant of the Fletcher– Reeves conjugate gradient method for stochastic inverse eigenvalue problems with partial eigendata. Numerical Linear Algebra with Applications, 26(2):e2221:1–e2221:??, March 2019. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yin:2017:FAO

[YCY17] Guojian Yin, Raymond H. Chan, and Man-Chung Yeung. A FEAST algorithm with oblique projection for generalized eigenvalue problems. Numerical Linear Algebra with Applications, 24(4):??, August 2017. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2011:CNB

[YDH11] Xingdong Yang, Hua Dai, and Qingquan He. Condition numbers and backward perturbation bound for linear matrix equations. Numerical Linear Algebra with Applications, 18(1):155–165, January 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ye:2020:PAS

[Ye20] Qiang Ye. Preconditioning for accurate solutions of illconditioned linear systems. Numerical Linear Algebra with Applications, 27(4):e2315:1– e2315:??, August 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yuan:2020:CMG

[YHAG20] Xinru Yuan, Wen Huang, P.- A. Absil, and Kyle A. Gallivan. Computing the matrix geometric mean: Riemannian versus Euclidean conditioning, implementation techniques, and a Riemannian BFGS method. Numerical Linear Algebra with Applications, 27(5):e2321:1–e2321:??, October 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2018:MIM

[YHS18] Xue Yang, Yu-Mei Huang, and Li Sun. A modulus iteration method for retinex problem. Numerical Linear Algebra with Applications, 25(6):??, December 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2008:PBW

[YL08] Hu Yang and Hanyu Li. Perturbation bounds for weighted polar decomposition in the weighted unitarily invariant norm. Numerical Linear Algebra with Applications, 15(8): 685–700, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yuan:2011:PEB

[YLH11] Quan Yuan, Huinan Leng, and Zhiqing He. A property of eigenvalue bounds for a class of symmetric tridiagonal interval matrices. Numerical Linear Algebra with Applications, 18 (4):707–717, August 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2022:CEQ

[YM22] Zhao Yang and Xiao-Xiao Ma. Computing eigenvalues of quasi-rational Bernstein– Vandermonde matrices to high relative accuracy. Numerical Linear Algebra with Applications, 29(3):e2421:1–e2421:??, May 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yaniv:2023:SSR

[YMS^{+}23] Yotam Yaniv, Jacob D. Moorman, William Swartworth, Thomas Tu, Daji Landis, and Deanna Needell. Selectable Set Randomized Kaczmarz. Numerical Linear Algebra with Applications, 30(1):e2458:1– e2458:??, January 2023. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yang:2004:MFC

[YNP04] Chao Yang, Esmond G. Ng, and Pawel A. Penczek. Matrixfree constructions of circulant and block circulant preconditioners. Numerical Linear Algebra with Applications, 11(8–9):773–793, October/November 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yong:1996:SCT

[Yon96] Xue-Rong Yong. Short communication: Two properties of diagonally dominant matrices. Numerical Linear Algebra with Applications, 3(2): 173–177, March/April 1996. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
15000984.

Yotov:2001:MNK

[Yot01]  Ivan Yotov.  A multi-

level Newton–Krylov interface solver for multiphysics couplings of flow in porous media. Numerical Linear Algebra with Applications, 8 (8):551–570, December 2001. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/88010592/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=88010592&PLACEBO=IE.pdf.

Yeung:2020:ART

[YPC20] Yu-Hong Yeung, Alex Pothen, and Jessica Crouch. AMPS: Real-time mesh cutting with augmented matrices for surgical simulations. Numerical Linear Algebra with Applications, 27(6):e2323:1–e2323:??, December 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Peng:2007:PDS

[yPES07] Zhen yun Peng and Salah M. El-Sayed. On positive definite solution of a nonlinear matrix equation. Numerical Linear Algebra with Applications, 14(2):99–113, ???? 2007. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ypma:1995:SFP

[Ypm95] Tjalling J. Ypma. A SAXPY formulation for plane rotations. Numerical Linear Algebra with Applications, 2 (6):533–541, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Peng:2006:EIM

[yPxP06] Zhen yun Peng and Ya xin Peng. An efficient iterative method for solving the matrix equation AXB + CY D = E. Numerical Linear Algebra with Applications, 13(6): 473–485, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). See comments [WTZD10].

Peng:2004:IPB

[yPyHZ04] Zhen yun Peng, Xi yan Hu, and Lei Zhang. The inverse problem of bisymmetric matrices with a submatrix constraint. Numerical Linear Algebra with Applications, 11(1): 59–73, February 2004. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yavneh:2012:NBB

[YW12] Irad Yavneh and Marion Weinzierl. Nonsymmetric Black Box multigrid with coarsening by three. Numerical Linear Algebra with Applications, 19(2):194–209, March 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yu:2013:ATL

[YXZ13] Guozhu Yu, Jinchao Xu, and Ludmil T. Zikatanov. Analysis of a two-level method for anisotropic diffusion equations on aligned and nonaligned grids. Numerical Linear Algebra with Applications, 20(5):832–851, October 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yin:2012:AAA

[YYN12] Jun-Feng Yin, Guo-Jian Yin, and Michael Ng. On adaptively accelerated Arnoldi method for computing PageRank. Numerical Linear Algebra with Applications, 19(1):73–85, January 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Ye:2013:AAA

[YZ13] Qiang Ye and Weifeng Zhi. Analysis of alignment algorithms with mixed dimensions for dimensionality reduction. Numerical Linear Algebra with Applications, 20(2): 369–384, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Yu:2023:LTR

[YZCQ23] Quan Yu, Xinzhen Zhang, Yannan Chen, and Liqun Qi. Low Tucker rank tensor completion using a symmetric block coordinate descent method. Numerical Linear Algebra with Applications, 30(3):e2464:1– e2464:??, May 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2023:CAB

[ZBCN23] Ming Zhou, Zhaojun Bai, Yunfeng Cai, and Klaus Neymeyr. Convergence analysis of a block preconditioned steepest descent eigensolver with implicit deflation. Numerical Linear Algebra with Applications, 30 (5):e2498:1–e2498:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2011:STP

[ZCW11] Chengjian Zhang, Hao Chen, and Leiming Wang. Strangtype preconditioners applied to ordinary and neutral differentialalgebraic equations. Numerical Linear Algebra with Applications, 18(5):843–855, October 2011. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zvonarev:2022:FSM

[ZG22] Nikita Zvonarev and Nina Golyandina. Fast and stable modification of the Gauss– Newton method for lowrank signal estimation. Numerical Linear Algebra with Applications, 29(4):e2428:1– e2428:??, August 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zha:1992:TWC

[Zha92] Hong Yuan Zha. A two-way chasing scheme for reducing a symmetric arrowhead matrix to tridiagonal form. Journal of Numerical Linear Algebra with Applications, 1(1):49–57, 1992. CODEN ???? ISSN 0129-3281.

Zhang:2018:EVH

[Zha18] Ju-Li Zhang. An efficient variant of HSS preconditioner for generalized saddle point problems. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2012:FIC

[ZHJL12] Yong Zhang, Ting-Zhu Huang, Yan-Fei Jing, and Liang Li. Flexible incomplete Cholesky factorization with multi-parameters to control the number of nonzero elements in preconditioners. Numerical Linear Algebra with Applications, 19(3):555–569, May 2012. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2006:SJD

[Zho06]  Yunkai Zhou.  Studies on

Jacobi–Davidson, Rayleigh quotient iteration, inverse iteration generalized Davidson and Newton updates. Numerical Linear Algebra with Applications, 13(8):621–642, ???? 2006. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2016:TTE

[Zho16] Weiqi Zhou. Triangular truncation and its extremal matrices. Numerical Linear Algebra with Applications, 23(4): 642–655, August 2016. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2018:CEN

[Zho18] Ming Zhou. Convergence estimates of nonrestarted and restarted block-Lanczos methods. Numerical Linear Algebra with Applications, 25(5): ??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhu:2008:DDP

[Zhu08] Yunrong Zhu. Domain decomposition preconditioners for elliptic equations with jump coefficients. Numerical Linear Algebra with Applications, 15(2-3):271–289, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhu:2014:AMP

[Zhu14] Yunrong Zhu. Analysis of a multigrid preconditioner for Crouzeix–Raviart discretization of elliptic partial differential equation with jump coefficients. Numerical Linear Algebra with Applications, 21(1): 24–38, January 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhao:2010:PLS

[ZHZ10] Lijun Zhao, Xiyan Hu, and Lei Zhang. Properties and leastsquares problems for row extended matrices. Numerical Linear Algebra with Applications, 17(1):57–71, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zikatanov:2008:TSB

[Zik08] Ludmil T. Zikatanov. Twosided bounds on the convergence rate of two-level methods. Numerical Linear Algebra with Applications, 15(5): 439–454, ???? 2008. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zitko:2000:GCC

[Zít00] Jan Zítko. Generalization of convergence conditions for a restarted GMRES. Numerical Linear Algebra with Applications, 7(3):117–131, April/ May 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72001233/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72001233&PLACEBO=IE.pdf.

Zitko:2005:CCR

[Zít05] Jan Zítko. Convergence conditions for a restarted GM-RES method augmented with eigenspaces. Numerical Linear Algebra with Applications, 12 (4):373–390, May 2005. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zeng:2006:DAS

[ZJ06] J. P. Zeng and Y. J. Jiang. Direct algorithms to solve the two-sided obstacle problem for an M-matrix. Numerical Linear Algebra with Applications, 13(7):543–551, ???? 2006. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2022:GMK

[ZL22] Yanjun Zhang and Hanyu Li. Greedy Motzkin–Kaczmarz methods for solving linear systems. Numerical Linear Algebra with Applications, 29 (4):e2429:1–e2429:??, August 2022. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2023:LUB

[ZLLH23] Li-Ping Zhang, Zi-Cai Li, Ming-Gong Lee, and Hung-Tsai Huang. Lower and upper bounds of condition number for Vandermonde-wise matrices and method of fundamental solutions using pseudo radial-lines. Numerical Linear Algebra with Applications, 30(2):e2466:1–e2466:??, March 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhong:2008:CCR

[ZM08] Baojiang Zhong and Ronald B.

Morgan. Complementary cycles of restarted GMRES. Numerical Linear Algebra with Applications, 15(6):559–571, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zou:2020:PEH

[ZM20] Qinmeng Zou and Frédéric Magoulès. Parameter estimation in the Hermitian and skew-Hermitian splitting method using gradient iterations. Numerical Linear Algebra with Applications, 27(4):e2304:1– e2304:??, August 2020. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zubair:2010:GMM

[ZMO10] H. Bin Zubair, S. P. MacLachlan, and C. W. Oosterlee. A geometric multigrid method based on L-shaped coarsening for PDEs on stretched grids. Numerical Linear Algebra with Applications, 17(6): 871–894, December 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2018:ADM

[ZN18] Jianjun Zhang and James G. Nagy. An alternating direction method of multipliers for the solution of matrix equations arising in inverse problems. Numerical Linear Algebra with Applications, 25(4): ??, August 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zeng:2020:DTO

[ZN20] Chao Zeng and Michael K. Ng. Decompositions of thirdorder tensors: HOSVD, T-SVD, and beyond. Numerical Linear Algebra with Applications, 27(3):e2290:1–e2290:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2022:RTT

[ZN22] Xiongjun Zhang and Michael K.

Ng. Robust tensor train component analysis. Numerical Linear Algebra with Applications, 29(1):e2403:1– e2403:??, January 2022. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zou:2023:FBC

[Zou23]  Qinmeng Zou.  A flexible

block classical Gram–Schmidt skeleton with reorthogonalization. Numerical Linear Algebra with Applications, 30 (5):e2491:1–e2491:??, October 2023. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2012:LCA

[ZQ12] Liping Zhang and Liqun Qi.

Linear convergence of an algorithm for computing the largest eigenvalue of a nonnegative tensor. Numerical Linear Algebra with Applications, 19(5): 830–841, October 2012. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2013:DEE

[ZQLX13] L. P. Zhang, L. Q. Qi, Z. Y.

Luo, and Y. Xu. The dominant eigenvalue of an essentially nonnegative tensor. Numerical Linear Algebra with Applications, 20(6):929–941, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2013:LES

[ZQW13] Guanglu Zhou, Liqun Qi, and Soon-Yi Wu. On the largest eigenvalue of a symmetric nonnegative tensor. Numerical Linear Algebra with Applications, 20(6):913–928, December 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2008:AIS

[ZS08] Yunkai Zhou and D. C. Sorensen. Approximate implicit subspace iteration with alternating directions for LTI system model reduction. Numerical Linear Algebra with Applications, 15(9):873–886, ???? 2008. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhong:2010:CAE

[ZSCX10] Liuqiang Zhong, Shi Shu, Long Chen, and Jinchao Xu. Convergence of adaptive edge finite element methods for H(curl)- elliptic problems. Numerical Linear Algebra with Applications, 17(2-3):415–432, ???? 2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2018:RTS

[ZSKA18] Jiani Zhang, Arvind K.

Saibaba, Misha E. Kilmer, and Shuchin Aeron. A randomized tensor singular value decomposition based on the t-product. Numerical Linear Algebra with Applications, 25(5):??, October 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhong:2013:TGM

[ZSWX13] Liuqiang Zhong, Shi Shu, Junxian Wang, and J. Xu. Two-grid methods for timeharmonic Maxwell equations. Numerical Linear Algebra with Applications, 20(1):93– 111, January 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhao:2014:MFS

[ZVO14] Jing Zhao, Edwin A. H.

Vollebregt, and Cornelis W. Oosterlee. Multigrid with FFT smoother for a simplified 2D frictional contact problem. Numerical Linear Algebra with Applications, 21(2): 256–274, March 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2010:CGS

[ZW10] Naimin Zhang and Yi-Min Wei.

On the convergence of general stationary iterative methods for range-Hermitian singular linear systems. Numerical Linear Algebra with Applications, 17(1):139–154, ????

2010. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhou:2018:FAS

[ZWQA18] Guanglu Zhou, Gang Wang, Liqun Qi, and Mohammed Alqahtani. A fast algorithm for the spectral radii of weakly reducible nonnegative tensors. Numerical Linear Algebra with Applications, 25(2):??, March 2018. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zheng:2020:MLR

[ZXS20] Qingqing Zheng, Yuanzhe Xi, and Yousef Saad. Multicolor low-rank preconditioner for general sparse linear systems. Numerical Linear Algebra with Applications, 27 (4):e2316:1–e2316:??, August 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zheng:2019:PAM

[ZY19] Bing Zheng and Zhanshan Yang. Perturbation analysis for mixed least squares-total least squares problems. Numerical Linear Algebra with Applications, 26(4):e2239:1– e2239:??, August 2019. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2011:PCF

[ZYFG11] Jiansong Zhang, Danping Yang, Hongfei Fu, and Hui Guo. Parallel characteristic finite element method for timedependent convection-diffusion problem. Numerical Linear Algebra with Applications, 18(4): 695–705, August 2011. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhao:2013:IIF

[ZYL13] Haifeng Zhao, Liang Yan, and Jijun Liu. On the interface identification of free boundary problem by method of fundamental solution. Numerical Linear Algebra with Applications, 20(2):385–396, March 2013. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2015:MQN

[ZZ15] Benxin Zhang and Zhibin Zhu.

A modified quasi-Newton diagonal update algorithm for total variation denoising problems and nonlinear monotone equations with applications in compressive sensing. Numerical Linear Algebra with Applications, 22(3):500–522, May 2015. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhang:2021:PID

[ZZ21] Chengjian Zhang and Yongtao Zhou. A preconditioned implicit difference scheme for semilinear two-dimensional time-space fractional Fokker– Planck equations. Numerical Linear Algebra with Applications, 28(4):e2357:1–e2357:??, August 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Zhao:2020:LCC

[ZZLX20] Ruijuan Zhao, Bing Zheng, Maolin Liang, and Yangyang Xu. A locally and cubically convergent algorithm for computing Z-eigenpairs of symmetric tensors. Numerical Linear Algebra with Applications, 27(3):e2284:1–e2284:??, May 2020. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).